A Complete Bibliography of Publications in Statistics and Computing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 May 2025 Version 1.36

Title word cross-reference

1 [1297]. 2 [177, 1311]. 2 × 2 [68]. 3 [1548]. ^{+} [2162]. ^{2} [1865, 1419]. _{2} [565]. A [1144, 1793, 1561]. A + tB [1781]. α [1787, 1325, 1170]. β [1603]. c [1895]. D [2081, 1054, 73, 169, 807, 1576]. ℓ_{1} [1401, 1420]. ℓ_{2,0} [2204]. F [669, 423, 1379, 43, 151, 1984, 1745]. g [428, 2213]. GARCH(p, q) [1777]. h [749, 428, 2213]. J [834]. K [2194, 779, 2163, 1766, 1043, 1501, 428, 170, 376]. L^{p} [2201]. L_{1} [989, 507]. L_{2} [1281, 549, 1519]. λ [510]. log GARCH [2144]. M [1276, 1942, 2021, 269, 294]. O(n) [1469]. m × n [729]. n [756]. N^{−1/2} [1297]. P [251, 236, 1132, 898, 591, 417, 1787, 644, 756, 37, 2215]. π [1769]. Q_{B} [1790]. r [1330]. S^{2} [569]. σ [1068, 1067]. T [1464, 798, 893, 1913, 1598, 1147, 525, 726, 1466, 842, 1000, 992, 1193, 613, 753, 1017, 825, 375, 1056]. U [1356]. G_{m}^{0} [1910]. W [1232, 35]. Y [972].

-and- [428]. -Boosting [1519]. -dimensional [1548]. -distributions [893, 992, 1193, 2213, 73, 169]. -divergences [1984]. -ensembles [1603]. -expectiles [2194]. -factor [798]. -fold [779, 2163]. -GAN [1745]. -graph [1232]. -likelihood [749]. -link [1056]. -linked [972]. -means [1043]. -measure [1379]. -median [644, 591]. -monotonicity [1766]. -nearest [376]. -norm [1281]. -normal [842, 1017]. -optimal [1895, 1144, 2081, 1790, 1793, 807, 1464, 1561]. -penalized [1420, 1401, 2204, 989]. -permanents [1170]. -quantile [1276, 1942, 2021]. -shaped [1356]. -sphere [1576]. -splines [1132, 898]. -stable [1068, 1787, 1325, 1067]. -statistic [43]. -tempered [1787]. -test [1000, 834, 35]. -th [170]. -value [2215, 236]. -values [417, 37, 251]. -vines [1054].

21st [85]. 25th [1114]. 2D [824].

3D [824, 359].

4 [230]. 4D [1738]. 4D-Var [1738].

90 [8].

A. [791]. ABC [1075, 907, 1537, 908, 913, 1166, 1076, 1317]. abrupt [776]. absence [1891]. absolute [449]. abstract [1829]. Accelerated [2107, 1983, 1019, 1272, 2041, 1873, 1696, 1439, 1602, 1991, 2001, 1751]. Accelerating [1658, 186, 1915, 1559, 1622]. Acceleration [52, 615, 684, 1332, 788]. acceptance [1201, 1139, 891, 1297]. acceptance-rejection [1201, 1297]. Access [501, 502]. account [1968]. Accuracy [642, 1898]. Accurate [1355, 163, 569, 2126, 1881, 1392, 1026]. Achieving [517, 1750]. acquisition [28]. across [785, 1883, 367, 626]. activity [1874]. actual [418]. acyclic [1431, 1085, 1208, 1247]. Adaptation [1822]. adaptative [1788]. adapted [974, 1491]. adapting [1865, 2092]. Adaptive [419, 1181, 678, 1497, 871, 679, 1002, 1372, 686, 1202, 675, 1757, 1860, 582, 673, 1075, 1076, 1179, 1055, 1828, 2117, 1561, 1608, 672, 2141, 1319, 1022, 676, 1096, 1403, 892, 1886, 1668, 671, 804, 923, 1183, 599, 1198, 1718, 2033, 2209, 1203, 2167, 1218, 1931, 1879, 895, 818, 1280, 1939, 1612, 2090, 1960, 917, 646, 1098, 1564, 1169, 160, 2149, 1257, 891, 1602, 2112, 1315, 874, 2064]. Adaptive-Ridge [2064]. adaptive-to-model [1315]. adaptively [1731]. ADASYN [2090]. adding [431]. Additive [1160, 1399, 2047, 1607, 1679, 2211, 1066, 1680, 648, 2125, 879, 1955, 742, 620, 1357, 666, 414, 172, 2086, 1111, 2132, 1211, 920, 995, 1258, 2160, 2077, 1930, 1616, 1104, 1916, 180, 565, 2193, 1065, 1550, 1429, 1221, 1324, 1742]. additivity [524]. ADE [230]. ADE-4 [230]. adequacy [926]. Adjusted [1060, 1059, 1397, 2210, 572, 2062]. adjustment [1330, 440, 1551]. Admissible [31]. adulteration [69]. advanced [498]. advancement [1033]. adversarial [2094, 1896]. Affine [1770]. Affine-mapping [1770]. after [1264]. aftershock [1706]. against [1124, 661, 394]. age [159]. agents [2080]. Aggregate [132]. aggregated [1409]. aggregation [1582, 1441]. agreement [1717]. AIC [251, 1810]. aided [110]. Air [2209]. Air-HOLP [2209]. Aitkin [253, 252]. Akaike [271]. Albert [690]. algebra [380, 1023, 387]. Algebraic [452, 772]. Algorithm [404, 1144, 1661, 1201, 189, 1576, 1243, 72, 48, 2165, 684, 1216, 1938, 487, 1271, 590, 324, 1268, 1459, 1967, 1606, 634, 615, 1156, 20, 1731, 1285, 1859, 1412, 1053, 1868, 327, 1634, 244, 944, 287, 1059, 1779, 1474, 1861, 2081, 154, 2156, 1980, 2134, 339, 1640, 340, 945, 1592, 2185, 1218, 773, 248, 616, 1029, 962, 886, 1009, 1845, 530, 1695, 1879, 1720, 122, 1832, 1705, 818, 1228, 1295, 1977, 2037, 465, 1343, 269, 913, 237, 974, 449, 1956, 673, 722, 1132, 200, 300, 623, 1235, 1564, 1973, 2049, 1206, 1317, 2046, 1637, 476, 891, 2093, 2034, 518, 1016, 1411, 2104]. algorithm [103, 729, 835, 1145, 1305, 807, 535, 438, 619, 1834, 2064, 1060]. algorithm-based [1980]. Algorithmic [899]. Algorithms [47, 474, 1817, 431, 1081, 663, 227, 1986, 1180, 676, 928, 1133, 265, 2203, 578, 1435, 106, 521, 68, 726, 506, 1379, 165, 79, 842, 959, 46, 2131, 796, 1299, 225, 1489, 2171, 1274, 1044, 10, 432, 856, 1572, 1696, 1205, 1797, 1776, 758, 1710, 202, 1790, 1170, 1895, 520, 1482, 1410, 788]. aliasing [1112, 1891]. aligned [1611, 2029]. allocation [1181, 1534, 618]. allocations [1765]. allowing [1683]. along [706]. alpha [572]. alternating [463]. alternative [576, 825, 467]. alternatives [771, 964, 654]. ALVar [1860]. American [1600]. among [1591]. amounts [1047]. amplitude [484]. analyses [397, 1816, 930]. analysing [332]. Analysis [291, 238, 1635, 208, 758, 438, 874, 1623, 1654, 69, 54, 206, 437, 1604, 591, 893, 865, 2066, 1273, 223, 2095, 353, 488, 1146, 355, 564, 13, 1147, 38, 1088, 1314, 1275, 1307, 29, 2089, 66, 53, 2, 549, 1605, 704, 1995, 299, 1902, 1950, 785, 1061, 1354, 548, 889, 685, 373, 1103, 1082, 644, 848, 1212, 2175, 1127, 1217, 2094, 1861, 2132, 2226, 2131, 627, 875, 1858, 781, 1926, 1501, 232, 357, 698, 1413, 2028, 1964, 2208, 516, 1381, 1341, 2217, 122, 261, 1747, 395, 876, 1062, 354, 1101, 559, 1390, 2042, 276, 1289, 689, 974, 1976, 770]. analysis [1797, 1186, 956, 5, 751, 1340, 502, 1454, 2190, 1169, 637, 581, 271, 1317, 1948, 1311, 230, 368, 1904, 2197, 12, 721, 209, 241, 193, 229, 1188, 450, 1991, 1422, 1482, 1554, 1869, 2019, 2192, 651, 1798, 2070]. analytic [185, 1529, 1707]. Analytical [40]. analyzer [1741]. analyzers [798, 1020, 1084]. Analyzing [759, 1990]. anchor [1719]. angle [2090]. angular [1388]. animation [536]. Anisotropic [2189, 1511]. Anisotropy [1415]. Annealed [392, 1681]. Annealing [799, 1150, 48, 634, 298, 244, 154, 46, 1299, 18, 248, 96]. Anniversary [1114]. announcement [61]. anomalies [1548]. anomalous [2216]. Anomaly [1583, 1728]. ANOVA [478]. Answer [156]. Answer-perturbation [156]. Antithetic [1487, 615, 1832]. any [1690]. Anytime [1667]. Aperture [357]. apparent [1595]. APPLE [1036]. applicable [1293, 1131]. Application [573, 456, 164, 565, 554, 241, 862, 1556, 283, 2144, 1874, 1660, 861, 679, 217, 1060, 1701, 1047, 1883, 244, 1357, 1059, 1631, 1212, 1484, 761, 580, 1097, 2208, 1052, 1721, 646, 1186, 2133, 1838, 1415, 1448, 2102, 626, 1762, 1311, 36, 43, 1856, 450, 835, 1168, 2109, 1394]. Applications [20, 445, 376, 935, 587, 1074, 1192, 878, 704, 1032, 1950, 1073, 1989, 692, 1421, 1159, 1446, 1449, 1844, 1148, 1691, 605, 2052, 1226, 1857, 632, 1318, 1943, 673, 1754, 1709]. applicative [1301]. Applied [123, 601, 1600, 158, 196, 836, 2007]. Applying [227]. Approach [458, 403, 1144, 1150, 1898, 1532, 1843, 1573, 601, 1330, 1310, 1773, 2082, 243, 266, 265, 733, 1479, 167, 1374, 1371, 446, 176, 701, 1517, 373, 1631, 1627, 1321, 1160, 2086, 1645, 1801, 2167, 1211, 1331, 442, 617, 2139, 1947, 1917, 895, 716, 1557, 1992, 1308, 765, 185, 1366, 929, 2090, 2164, 2091, 1396, 1819, 2169, 1348, 1940, 1189, 2024, 1723, 120, 434, 209, 1818, 1411, 193, 1614, 2113, 2145, 2001, 1742]. approaches [907, 2100, 954, 358, 1929, 268]. Approximate [1304, 1320, 730, 905, 1755, 906, 1199, 732, 1089, 1150, 1157, 982, 1204, 380, 1278, 1245, 1246, 1284, 892, 2162, 1758, 1061, 1244, 2135, 1160, 1825, 910, 474, 1447, 1062, 909, 1396, 912, 1800, 1560, 1469, 2149, 1036, 594]. Approximating [595, 1213, 35, 1181, 1015, 388, 74, 288, 182]. Approximation [564, 1536, 836, 1015, 1733, 1163, 1530, 1621, 1466, 1006, 1299, 1592, 472, 1694, 1853, 799, 1695, 1735, 856, 247, 1960, 2044, 1842, 1748, 1361, 1568, 512, 1736, 1358, 1813, 538, 438, 619, 2137, 1834, 1725, 1408, 2131]. Approximations [459, 1334, 1215, 295, 1455, 40, 1808, 2201, 1809, 304, 533, 1939, 1366, 990, 1121, 1135, 1152]. AR-ADASYN [2090]. arbitrary [1855, 1958, 471, 1591, 274]. ARCH [548]. Archimedean [1476]. architectures [1626]. arcs [1576]. areal [1647]. areas [82, 90]. arising [959]. ARL [569]. arm [689]. ARMA [695, 457]. arrival [1475]. arrivals [1567]. art [55]. artificial [485, 1905, 1772, 67]. ascent [1552]. aspects [102, 668, 1103, 1655, 1095, 930]. assess [363, 1262]. Assessing [330, 1124, 1717, 215]. assessment [284, 279, 498, 484, 926, 376]. assimilation [2108]. associated [2063, 981, 2146]. association [1655, 1137, 1037, 656]. assortativity [1764]. assumptions [1177, 768]. astronomy [627]. asymmetric [674, 2068, 1101, 1154, 538]. Asymmetry [772, 610]. Asymptotic [379, 2182, 1425, 2054, 37, 974, 870]. asymptotics [1355, 385, 1710, 1736]. attained [161]. attribution [2073, 1906]. attrition [370]. Augmentation [1239, 567, 1092]. Augmented [1687, 838, 1143]. autocorrelation [546, 1217]. autoencoder [2127]. autoencoders [2088]. automata [4]. Automated [1330, 235, 2033, 219, 1594]. Automatic [111, 707, 1799, 1780, 112, 455, 468, 1357, 472, 534, 1784, 1762, 1831, 1737, 1493, 1218, 160, 1309]. Automatically [1865]. Automating [718]. automotive [1943]. Autoregressive [456, 1506, 2015, 1489, 2118, 1632, 929, 918, 1820, 1676, 302, 2214]. Auxiliary [1417, 723, 226, 974]. available [26]. average [115, 2025, 267, 694, 1195]. averaging [939, 1399, 1232, 2206, 2163, 1907, 2083, 1900]. Avoiding [1109]. aware [1938, 2079]. AXIOM [381]. axis [1611, 2029]. axis-aligned [1611, 2029].

B [1426, 2128, 839]. B-spline [1426, 2128, 839]. back [139, 517, 141, 119, 145, 138, 143, 144, 142]. background [1738, 2102]. backward [1868]. backwards [1125]. balanced [1681, 1143]. bands [605, 2025]. Bandwidth [1222, 860, 557, 305]. bandwidths [1326].

Barber [2138]. Barnard [236]. Barthelmé [1065]. barycenters [1555]. Based [406, 543, 1595, 188, 1679, 1997, 893, 257, 2143, 1310, 2157, 1333, 1841, 223, 748, 1128, 1216, 1713, 80, 1537, 1268, 1773, 1147, 1998, 1133, 593, 1272, 243, 961, 1083, 600, 417, 715, 2100, 1284, 433, 1053, 1161, 1683, 832, 526, 1468, 492, 123, 963, 1354, 815, 1574, 1685, 1979, 1020, 1518, 1846, 820, 1108, 1212, 976, 838, 926, 2068, 1717, 1571, 2156, 1980, 628, 2072, 970, 1839, 1858, 1882, 687, 2167, 769, 1226, 1012, 1905, 1478, 442, 1009, 1647, 1673, 1873, 1896, 2057, 2163, 968, 1288, 1175, 1306, 1746, 1001, 793, 936, 869, 1308]. based [1975, 1143, 1949, 1167, 765, 1366, 1367, 1039, 917, 1976, 1903, 510, 1819, 1348, 1956, 2212, 2215, 2205, 2024, 1776, 170, 2025, 1105, 1811, 1723, 870, 623, 758, 764, 1564, 1037, 1897, 160, 636, 811, 2059, 2098, 1948, 1121, 2197, 705, 209, 1262, 1818, 1256, 1630, 969, 1770, 1614, 1302, 1699, 739, 418, 844, 1441, 973, 1866, 2096, 1732, 1430]. baseline [589]. baselines [1876]. bases [871, 382, 1332]. Basic [102, 142]. basics [119, 138, 139, 141, 145, 143, 144]. basis [413, 495, 496, 839]. batch [1592, 1543, 1257]. Bayes [54, 146, 188, 251, 1150, 806, 1941, 238, 1382, 900, 1804, 1462, 171, 373, 1198, 1645, 1232, 571, 711, 1405, 1870, 382, 1838, 1698, 1677, 2093, 2034, 2214, 260]. Bayes-error [1870]. Bayesian [1063, 1073, 1061, 1329, 1089, 1304, 1623, 558, 699, 1739, 1898, 1532, 1523, 1242, 709, 1320, 1659, 1184, 2189, 2063, 2157, 1333, 1117, 1210, 2199, 730, 515, 1278, 1713, 587, 794, 652, 590, 372, 397, 2216, 1760, 1074, 13, 2015, 905, 1816, 1383, 1382, 707, 787, 349, 346, 690, 882, 595, 1953, 1988, 205, 2043, 600, 919, 2221, 1479, 1335, 892, 53, 2, 425, 285, 396, 1638, 1682, 167, 784, 362, 2124, 1337, 1032, 1426, 1404, 1701, 1755, 1267, 742, 1178, 1902, 2162, 446, 586, 785, 950, 2058, 1758, 1883, 1094, 1293, 641, 1808, 1049, 1303, 1248, 674].

Bayesian [939, 1244, 548, 603, 598, 1915, 1688, 972, 1125, 1183, 1476, 1593, 1833, 1703, 743, 2097, 1399, 1600, 509, 948, 1111, 2068, 347, 1493, 1718, 1825, 910, 978, 2002, 277, 2009, 1640, 2130, 643, 1499, 1494, 1801, 288, 1211, 915, 348, 829, 1704, 1905, 846, 1872, 225, 1663, 1413, 2060, 1213, 1873, 539, 1757, 2139, 1947, 1261, 516, 1052, 1778, 1695, 2217, 1894, 2223, 1558, 374, 1653, 942, 1802, 1447, 1705, 906, 2040, 1228, 577, 614, 1001, 1610, 965, 1769, 1706, 1131, 1062, 922, 750, 1101, 1977, 2037, 911, 1092, 624, 2042, 276, 618, 547, 2047, 951, 1939, 689, 1366]. Bayesian [1851, 1529, 1707, 2220, 1570, 1708, 1417, 909, 1864, 2164, 1286, 1064, 2091, 1987, 1616, 1396, 1837, 1524, 1908, 1916, 1416, 912, 1549, 2212, 1350, 1800, 268, 980, 1199, 1811, 1723, 1550, 1710, 1323, 1485, 1564, 1454, 667, 1118, 637, 581, 1578, 898, 754, 626, 1263, 1948, 1711, 554, 1498, 1619, 523, 1786, 1328, 12, 1907, 216, 721, 1313, 1793, 209, 786, 1336, 1602, 1651, 1188, 450, 835, 1699, 1991, 1850, 1422, 633, 1055, 1822, 1937, 568, 973, 1187, 1866, 2192, 2137, 1433]. BayesProject [1590]. be [444]. been [421]. bees [249]. before [919]. behavior [1177]. belief [1386, 577]. Bernoulli [1009, 583]. Bernstein [1909]. Best [1423, 1999]. Beta [728, 669, 587, 1994]. better [421]. Between [385, 65, 1045, 74, 46, 1984, 2101]. between-unit [1045]. Beyond [1253, 1826, 1006, 1689, 866, 1297]. bi [1628]. bi-directional [1628]. Bias [2054, 975, 1737, 1349, 1504, 1852, 2056, 758]. Bias-reduced [2054]. biased [1272]. Bichon [1824]. Biclustering [1966, 1726, 1009]. big [1600, 1485, 1856, 2112, 1422]. Bilinear [741]. bimonotonicity [740]. bin [1843, 1774]. bin-marginal [1843]. Binary [456, 282, 370, 1287, 1320, 1216, 168, 1914, 826, 1643, 1684, 2130, 204, 1312, 1009, 2187, 1885, 1443, 1917, 951, 1056, 1220, 927, 1650, 1613, 512, 302, 1798]. Bingham [1636, 1178, 508, 697, 1397, 1091]. binned [1771]. Binomial [1034, 479]. bins [297]. bioassay [243]. bioequivalence [634]. biplots [150]. birds [249]. Birnbaum [2065, 1842]. birth [994]. bisection [2069]. bit [1285]. Bitcoin [2052]. Bivariate [1135, 2076, 1035, 608, 80, 1260, 525, 643, 1054, 1516, 1909, 1240, 160]. Blackwellized [1907]. Block [78, 850, 1878, 1345, 1455, 1461, 1684, 235, 234, 1149, 1312, 1418, 1721, 1837, 332, 1730]. blocked [1197, 368]. blocking [1664, 1793, 450]. blocking-MCMC [450]. blockmodels [1570, 1708]. blocks [465]. BLUP [749]. BMDP [8]. BOJA [152]. Boltzmann [45]. Book [791, 61, 1192, 690, 108, 98, 197, 128, 116, 790, 60, 221]. Boosted [880]. Boosting [1298, 736, 396, 1519, 2186, 1821, 549, 1161, 1510, 796, 1221, 1387, 2147]. Bootstrap [2148, 1349, 454, 191, 1344, 418, 201, 593, 2158, 163, 40, 3, 1124, 192, 442, 1291, 1288, 1240, 271, 376, 36, 447, 1171, 2007]. bootstrap-based [1288]. Bootstrapping [152, 1000, 573, 1968, 19, 445, 555]. Both [401]. bottleneck [1672]. bouncy [1664]. bound [1661, 2069, 1861, 1297]. boundary [76, 1334, 682]. Bounded [2062, 1998, 903]. Bounded-memory [2062]. Bounding [387]. bounds [1060, 1059, 1261, 1744, 518]. Box [825, 1762]. boxes [203]. brain [2014]. branch [1661]. branch-and-bound [1661]. branching [972]. brand [2208]. break [931, 2191]. breakdown [968]. breaking [1609]. breakpoint [1514]. breaks [845]. Bregman [1212]. bridge [863, 1954, 1296, 1547, 1745, 1937]. bridges [1633, 2055]. Brownian [605, 1802]. Bucy [1722]. buffet [1316]. BUGS [466]. Builder [2097]. building [1279]. Bump [308]. bundle [1535]. business [492]. buttons [517]. Byzantine [2082]. Byzantine-robust [2082].

C [1061, 1077]. Caimo [1073]. Calculating [423, 513, 1091, 1863, 635, 526, 1198, 435]. Calculation [125, 1355, 1636, 1474, 236]. calculations [386, 59]. calculator [7]. calculus [381]. calibrated [1582, 1645, 1939]. Calibrating [1407, 1107]. calibration [251, 1775, 1371, 1652, 1801, 1143]. candidate [509]. canonical [1191, 1190, 1193, 5]. capture [2171]. capture-recapture [2171]. carbon [956]. Carlo [186, 773, 1071, 762, 1588, 1157, 813, 1122, 424, 406, 1117, 1181, 1633, 1658, 823, 590, 181, 284, 279, 13, 1759, 873, 849, 1015, 1606, 2043, 1002, 265, 1625, 892, 1731, 1597, 1072, 362, 429, 2124, 1492, 1539, 671, 479, 1292, 1530, 1875, 1408, 607, 1251, 585, 1593, 551, 506, 1238, 1534, 1402, 1589, 172, 1230, 1458, 226, 1375, 605, 1395, 1097, 1353, 304, 472, 1693, 182, 1362, 1646, 1873, 799, 1757, 1880, 1487, 1228, 1546, 921, 965, 1318, 1961, 247, 1572, 1696, 1205, 909, 834, 1524, 673, 927, 1003, 1194, 1086, 1177, 1256, 1170, 1972, 520, 1699, 1422, 1410, 1552, 1332]. Carlo [438, 1725]. CARMA [1095]. CART [346]. cascade [154]. cascading [879]. case [840, 1019, 635, 2203, 1861, 1131, 567, 200, 1439]. case-cohort [1019, 1861]. Cases [868]. CAT [1501]. categorical [1273, 1919, 2196, 701, 1020, 1137, 1149, 436, 1442, 1486, 770, 902, 768]. categorisation [1728]. Cauchy [1950, 1711, 2197]. causal [1938, 1440, 2094, 1876]. causation [30]. cautionary [72]. cell [462, 1883]. cellular [4]. censored [1209, 1569, 1164, 961, 1932, 2185, 920, 1048, 2217, 559, 2047, 1338, 739, 1927, 201, 2083, 1300]. censoring [2195]. centered [2136]. centering [567]. centers [984]. Central [1540, 769, 151]. centred [1834]. centring [1333]. century [85]. Certified [2058]. cGAN [1896]. Chain [590, 950, 585, 506, 1059, 1375, 677, 762, 813, 1122, 935, 181, 284, 279, 849, 1606, 2043, 186, 265, 1072, 429, 1060, 479, 416, 1408, 1238, 1458, 226, 773, 1362, 1757, 1546, 965, 276, 1572, 1071, 1205, 520, 1410, 438, 1725]. chains [1157, 424, 1556, 1680, 212, 517, 1804, 1484, 387, 636, 677]. Change [185, 1281, 1877, 888, 800, 1172, 1406, 298, 536, 776, 1090, 855, 1987, 2224, 2177, 1940, 883, 1050, 1323]. change-over [298]. Change-point [185, 1877, 1172, 1090, 2224, 2177, 883, 1323]. change-points [855]. Changepoint [1783, 861, 2221, 586, 785, 1590, 1321, 1274, 1363, 1486, 1712, 2023, 2039]. changepoints [1319, 2153, 2086, 998]. changes [1906, 1418]. channel [94]. Chaos [402, 541, 401]. Chaotic [407, 404, 1553]. characteristic [1882]. characteristics [938]. Characterization [1571, 1415]. chart [727]. charts [45, 569, 2195]. checking [2041]. checkpoints [1712]. chi [1215]. chi-squared [1215]. Choice [443, 146, 83, 557, 1785, 472, 846, 2187, 305, 465, 1075, 2024, 1076, 1257]. choices [219]. Cholesky [878, 1691, 1839, 1395]. Cholesky-GARCH [878]. Christoffel [1446]. chromosomes [110]. chronic [1631]. cipher [849]. ciphers [168, 235]. circular [1997, 2095, 264, 1746, 1240, 351]. circumference [297]. claims [1943]. class [2140, 2122, 2196, 212, 1068, 1203, 773, 530, 1219, 1675, 1067, 1417, 1188]. classes [678]. classical [849, 1121]. Classification [802, 1051, 1929, 1567, 735, 893, 1941, 227, 27, 110, 1933, 1752, 919, 549, 784, 336, 1370, 1703, 1934, 204, 1993, 245, 1363, 744, 1870, 1207, 1838, 2024, 623, 1037, 341, 833, 67, 31, 376, 78, 259, 1562, 553, 1254]. classified [1595, 2196]. Classifier [2075, 1595, 1879, 2115, 1900]. Classifier-dependent [2075]. classifiers [1838, 376, 1224, 710]. classifying [810]. climate [692, 1906, 1186]. clinical [435, 689]. cloning [862]. closed [1919, 1492, 1643, 2171, 1037]. closed-form [1919, 1643]. closed-population [2171]. closure [936]. Cluster [627, 1255, 1584, 223, 1354, 1858, 763, 984, 1416, 1795]. clustered [1255, 2065, 1964, 2166, 1451]. Clustering [1460, 2005, 611, 2119, 1601, 1481, 203, 1529, 1628, 1582, 1523, 893, 1843, 1046, 1216, 1146, 1345, 1988, 1406, 2138, 2194, 1374, 1043, 1468, 1047, 963, 815, 1574, 1685, 1979, 1020, 1518, 1212, 1124, 1227, 1148, 1690, 1640, 687, 1226, 2139, 530, 1175, 1306, 1721, 793, 1033, 2079, 2168, 1367, 1708, 1903, 1819, 1956, 1520, 2024, 1282, 170, 764, 1741, 664, 1971, 350, 2059, 351, 1196, 2053, 306, 850, 2117, 1433, 1430, 639]. clusterings [1582]. clusters [1582, 1124, 1996, 1309]. Clusterwise [1970]. Co [1721, 1628, 1004, 1552, 1730]. co-block [1730]. Co-clustering [1721, 1628]. co-information [1004]. co-ordinate [1552]. Coaching [259]. cocktail [807]. CoDa [2040]. code [1173]. codes [126]. coefficient [370, 2225, 961, 949, 1038, 191, 880, 960, 1927]. coefficients [1434, 694]. coerced [891]. coherence [1130]. cohort [1019, 1861]. cohorts [1765]. collapsed [1188, 850]. collapsibility [395]. colour [1311]. colourability [59]. coloured [299]. column [332]. COM [1383]. Combatting [541]. combination [1083]. combination-based [1083]. combinations [413, 1113]. Combinatorial [1044, 1169, 733, 1895]. combined [727, 987]. combining [1571, 1472]. COMBSS [1999]. comment [496, 839]. Commentary [253]. Comments [139, 141, 143, 144, 497, 145, 142]. commercially [26]. commissioning [1400]. common [2184]. Communication [1922, 541, 2039]. Communication-efficient [1922, 2039]. communities [1571, 1570]. community [2069, 2221, 1959, 1400, 1418]. Comparative [591, 2066, 975, 644, 411]. Comparing [1582, 1128, 1763, 1172, 1773]. Comparison [244, 1286, 146, 1333, 648, 1215, 800, 349, 532, 1715, 1267, 674, 113, 74, 19, 46, 2099, 841, 951, 332, 1709, 1378, 1313, 1411]. comparisons [47, 187, 899]. compartmental [416]. competing [1943, 2025]. competitive [2080, 1577]. Complement [644]. complete [1306]. complete-data [1306]. completely [1595, 1659]. completion [1654]. complex [1310, 654, 1004, 1875, 277, 875, 508, 1450, 1051, 1910, 1414, 1903, 1229, 1871]. complex-valued [1450]. Complexity [1720, 1898, 1285, 804, 1672, 1485, 31]. component [589, 2095, 488, 1088, 1314, 1043, 1950, 2226, 781, 2028, 1341, 816, 1288, 158, 196, 1390, 2042, 1976, 1186, 930, 2192, 1798]. component-wise [1043]. components [345, 487, 932, 575, 167, 704, 505, 1779, 1241, 195, 698, 1099, 614, 158, 196, 922, 618, 1529, 1342, 555, 193]. Composable [1413]. Composite [1579, 748, 1128, 2082, 1967, 1004, 1848, 2164, 2169, 2205, 1199]. compositional [715, 592, 1993, 1301, 991, 2017, 1889]. compound [1381, 922, 973]. Comprehensive [1941, 2099]. Computation [190, 730, 596, 952, 544, 911, 1214, 178, 1162, 669, 1333, 1278, 905, 690, 2027, 1403, 892, 546, 1758, 1061, 388, 525, 1244, 171, 373, 1125, 1590, 761, 1825, 910, 569, 398, 101, 107, 477, 1062, 909, 1396, 912, 1199, 120, 554, 391, 151, 731, 971, 1356, 1586, 1089]. Computational [6, 323, 654, 1103, 806, 601, 114, 1074, 500, 1073, 668, 1655, 232, 906, 477, 1485, 1469, 1206, 1095, 1818, 1736, 790]. Computationally [1245, 1246, 1476, 1576, 1688, 1321, 19, 559]. computations [1150, 1320, 1122, 1066, 1867, 780, 1559, 1065, 901]. compute [361, 962]. Computer [110, 1392, 1308, 865, 872, 380, 1446, 1517, 868, 1645, 1801, 387, 876, 1567, 1515, 866]. Computer-aided [110]. computers [104]. Computing [1209, 334, 1176, 448, 981, 1026, 37, 790, 1740, 273, 1119, 283, 1893, 590, 1598, 696, 1115, 1295, 470, 1151, 200, 55, 1918, 626, 1170, 2104]. concave [1209, 2033, 1040, 1720, 2008]. concentrated [1701]. concentration [1138, 991]. Concepts [374, 171]. concerning [260]. concomitants [560]. concurrent [2188, 1840]. Conditional [403, 1620, 1368, 1364, 456, 267, 450, 1569, 380, 115, 1714, 1015, 934, 1290, 963, 2094, 1211, 474, 398, 1810, 1291, 2128, 160, 161, 302, 1358, 538]. Conditionally [1568, 594]. conditioned [444, 2155, 1488]. conditioning [1455, 2153, 1135]. Conditions [407]. conductors [682]. Confidence [605, 870, 1648, 982, 1204, 40, 112, 1634, 544, 1026, 2101, 5, 2025, 925, 732, 694, 572]. confidential [1818]. confidentiality [500, 1289]. configuration [277]. configurations [269]. confirmatory [1147]. conflict [2199]. conflicting [1674]. conformal [745]. congruential [177]. conjugate [1779, 2031]. connection [821]. connectivity [1444, 2014]. conquer [2218]. consensus [2038]. consequences [59]. Considerate [907]. Consistency [1913, 1483, 2155, 518]. Consistent [1938, 1672, 499, 1449, 1591]. consisting [455, 468]. constant [1636, 1347, 1091, 2023, 1195]. constant-per-iteration [2023]. Constants [459, 1722]. Constrained [1685, 726, 1671, 2173, 1201, 2011, 1988, 1473, 1084, 1155, 2152, 331, 1104, 1549, 1923]. constraint [1766, 1574, 1250]. constraints [740, 1799, 1294, 1266, 951]. Constructing [329, 1790, 907, 1268, 710]. construction [163, 1462, 985, 1288, 917, 732, 2053]. constructs [1296]. Contacts [9, 16, 23, 34, 42, 50, 63, 71, 77, 93, 99, 109, 117, 129, 137, 153, 162, 173, 184, 199, 213, 222, 231, 239, 246, 256, 262, 270, 281, 293, 303, 321, 333, 343, 352, 360, 369, 377, 389, 399, 408, 420, 430, 439, 451, 460, 469, 481, 490, 503, 511, 519, 528, 537, 545, 552, 561, 570, 579, 588, 597, 604, 612, 622, 630, 640, 650, 660, 670, 680, 691, 702, 713, 724, 734, 746, 757, 767, 778, 792, 805, 822, 843, 864, 877, 890, 904, 914]. Contacts [924, 937, 947, 957, 967, 977, 988, 999, 1013, 1027, 1042, 1057, 1126, 1140]. contamination [1043]. content [1571]. contiguity [1988]. contingency [68, 772, 1012, 290, 203, 161, 209, 729]. Continuity [2137]. Continuous [1319, 536, 2094, 1575, 1635, 2070, 1653, 1432, 1999, 267, 1762, 695, 148]. continuous-time [1653, 695]. contoured [687]. contours [215, 477]. contribute [1463]. Control [1456, 192, 1646, 2195, 1841, 727, 1886, 462, 45, 569, 1478, 1112, 909]. controlled [2129, 874]. Controlling [712, 814, 2117]. Convergence [1608, 284, 458, 1230, 1596, 1744, 1482, 1862, 1157, 257, 72, 279, 186, 265, 149, 1605, 1060, 1252, 1644, 1059, 2072, 387, 1496, 921, 390, 583, 300, 1751]. convergent [945]. convex [1848, 1920, 1675, 2168, 2031, 682]. Conway [1864]. cookbook [2027]. Coordinate [1016, 1542, 1271, 1268, 1156, 2058, 1040, 1899, 1295]. coordinate-exchange [1271, 1156]. COPD [1631]. COPICA [1088]. COPICA-independent [1088]. coping [1358]. Copula [1247, 1220, 667, 1322, 1088, 1440, 1270, 1717, 1054, 1211, 1976, 1358]. copula-based [1717, 1976]. copulas [1476, 1801, 1365, 769, 1226, 1909, 1219, 1437, 1369]. COR [2103, 2087]. Core [2114]. Core-elements [2114]. Corrected [417, 2054, 1557]. Correction [1942, 2140, 1420, 1682, 2103, 1471, 1807, 2070, 2037, 2077, 2224, 2093, 11, 76, 2056, 758, 2184]. corrections [1349]. correlated [444, 327, 2036, 2209, 915, 1341, 1207, 1056, 512, 1895]. Correlation [1283, 1839, 65, 1035, 854, 66, 57, 2014, 916, 191, 1364, 2220, 638, 367, 705, 2016]. Correlation-based [1839, 705]. correspondence [1273]. Cost [1907, 1469]. count [589, 2179, 1647, 1721, 531, 1945, 1577]. counting [1194]. counts [783, 53, 741]. coupled [813, 424, 1725]. Coupling [1229, 1373, 1744]. course [1141]. covariance [1580, 709, 1184, 748, 1508, 993, 1763, 1598, 1185, 2051, 1468, 641, 74, 2067, 562, 968, 2040, 922, 210, 2024, 1723, 367, 837, 2004, 1016, 1182, 1377, 1376, 1187, 1394]. covariances [1738]. covariate [1716, 1928, 1977, 2037, 744, 1407, 1039]. covariates [1102, 1225, 1210, 2107, 1266, 846, 1502, 1511, 902, 808, 2078, 1324, 1871]. covariation [1979]. covariogram [514]. Coverage [215, 694]. covers [1194]. Cox [1538, 1666, 825, 750, 1762]. credible [1673]. credit [1861]. criteria [1268, 349, 95, 1049, 559, 883, 341, 78]. Criterion [271, 2010, 1824, 1293, 2103, 2087, 364, 1288, 1131, 1854, 1142]. critical [1156]. Cross [65, 846, 817, 894, 2196, 2043, 1513, 779, 363, 364, 195, 1213, 1896, 2163, 1001, 158, 196, 1368, 350, 1329, 1328, 1900]. cross-classified [2196]. Cross-correlation [65]. cross-entropy [894]. cross-modality [1896]. cross-sectional [1001]. cross-validated [350]. Cross-validation [846, 817, 2043, 779, 363, 364, 195, 2163, 158, 196, 1368, 1329, 1328, 1900]. cross-validatory [1213]. crossed [1649]. crossing [1164, 1334, 2026]. crossover [330, 1650]. crossproducts [1502]. cryptocurrency [1990, 2052]. cubature [1023, 1493, 1494]. Cucala [1075]. cumulant [386]. cumulants [379, 696]. cumulative [186, 795]. cumulative-link [186]. cure [2066, 2217]. current [105, 1125, 2132]. Curve [664, 707, 224, 824, 706, 217, 970, 946, 808]. curves [602, 896, 566, 455, 468, 2072, 448, 1340, 44]. cusum [280]. cusums [279]. cut [2167, 1695]. cut-off [2167]. Cuts [1063, 1064]. cutting [1816]. cyclical [1051, 175].

D [1077, 1311, 1360]. D-robust [1360]. Dahlin [1069]. Dantzig [2104]. Data [1654, 457, 1531, 291, 238, 2179, 1165, 499, 435, 632, 869, 456, 1700, 543, 762, 589, 1966, 1255, 2005, 1997, 1877, 1209, 817, 862, 1164, 1035, 1216, 2199, 1611, 1319, 1893, 1022, 1130, 80, 2216, 1660, 1788, 861, 1096, 219, 1389, 1967, 1019, 961, 2196, 715, 1440, 1479, 1666, 2021, 2089, 500, 784, 1412, 1548, 774, 336, 327, 1158, 832, 1950, 526, 1409, 1883, 286, 455, 468, 308, 1684, 723, 776, 2065, 159, 939, 1517, 1915, 1020, 1431, 1082, 1833, 1783, 1115, 979, 1138, 79, 2159, 1227, 1600, 838, 1217, 1888, 1655, 2108, 1861, 2132]. data [1571, 1137, 1932, 1690, 2209, 2020, 544, 847, 1858, 643, 1265, 1149, 1203, 2052, 1926, 2167, 136, 1501, 662, 2185, 1312, 1931, 920, 436, 478, 698, 1413, 183, 756, 1400, 1647, 1810, 2084, 88, 2187, 987, 1996, 1048, 1106, 1873, 2057, 1964, 1947, 1341, 2217, 1026, 1993, 855, 916, 445, 1274, 533, 1306, 1746, 1001, 1301, 1929, 1555, 1910, 567, 1092, 2106, 559, 1949, 185, 951, 2048, 1339, 1367, 1529, 2090, 1960, 1567, 1207, 1436, 1486, 531, 770, 738, 1943, 1056, 1903, 759, 1819, 1186, 2224, 2177, 1837, 132, 1189, 2133, 1916, 1876, 565, 751, 1282, 747, 301, 991, 2017]. data [1662, 1485, 1741, 1971, 67, 1794, 1587, 2059, 2149, 1451, 554, 1551, 2004, 990, 1338, 1904, 1637, 1889, 2197, 2142, 1970, 810, 1618, 1818, 1259, 1856, 2112, 229, 302, 2214, 306, 1302, 1994, 2113, 1991, 1182, 1376, 739, 1921, 1422, 1927, 2145, 2001, 1768, 1792, 1586, 2083, 1869, 2178, 1798, 2181, 2062, 1433, 207, 1577, 1377]. data-adaptive [1022, 1931, 1960]. data-based [1997]. Data-driven [1531, 869, 1517, 1810, 1873, 1794]. database [130]. databases [156, 133, 134, 135]. dataset [1578]. datasets [1843, 692, 627, 2160, 471, 1447, 1747, 632, 1398, 2183]. dating [956]. David [1192]. De-noising [1821, 1491]. death [994]. Debiased [2219, 2218]. decentralized [2169]. Decision [1465, 446]. Decision-making [1465]. decomposability [395]. decomposable [709, 1250, 1753]. decomposed [993]. Decomposing [225]. decomposition [1536, 455, 468, 1691, 18, 869, 2102, 1187, 1726]. decompositions [848, 46, 1326]. deconvolution [860, 1766, 635, 703, 248]. Decrease [758]. Decrypting [849]. Deep [2118, 1424, 1618, 1678, 2211, 1941, 1743, 1807, 1749, 1769, 2026, 2109]. Default [1749, 1382]. defined [1513]. definition [2196]. definitive [1782]. deformable [415, 275]. deformation [722]. degree [2184]. degree-correction [2184]. degrees [1168]. Delaunay [968]. Delaunay-triangulation-based [968]. Delayed [1139]. Delete [294]. Delete- [294]. deleting [1264]. Deletion [638, 521]. demand [1357, 19]. demodulation [1051]. Dempster [252]. Deng [504]. dense [1941]. densities [1355, 2141, 426, 563, 647, 1701, 125, 1984, 828, 836, 1740, 160, 2008]. Density [80, 1766, 403, 793, 1384, 2193, 43, 1921, 860, 611, 933, 372, 1537, 423, 1133, 595, 635, 556, 549, 1426, 286, 1890, 703, 1718, 76, 215, 632, 1033, 247, 1339, 1364, 1987, 1186, 700, 55, 160, 1818, 1093, 714]. Density-based [793, 1537]. departure [526]. departures [1035]. Dependence [486, 1112, 1512, 706, 785, 2014, 1038, 1516, 1001, 1100, 1724, 1378, 1358]. dependency [1460]. dependent [1823, 2075, 1773, 774, 1148, 2167, 1218, 1746, 1738, 1943, 1282, 1179]. depth [663, 1829, 1212, 981, 477, 273]. derivative [1660, 168, 506, 1716, 1302]. derivatives [896, 1133, 236, 839]. derived [28]. descent [1081, 1040, 1635, 2070, 1899, 1295, 432, 124, 1016]. Design [323, 889, 866, 1623, 1144, 1661, 574, 867, 2157, 872, 2216, 871, 1156, 2124, 1641, 298, 327, 1755, 1703, 1669, 25, 1228, 1812, 1908, 1564, 1360]. Designs [943, 1162, 2123, 1351, 1268, 966, 1134, 2080, 325, 326, 1053, 1372, 1507, 1765, 1294, 1462, 985, 2135, 1671, 2081, 328, 329, 1392, 1261, 330, 1639, 331, 1366, 1173, 332, 1782, 1897, 1650, 1962, 1231, 1790, 1793, 1262, 1895, 969, 1360, 2053, 807, 1464]. desired [747]. destroyed [275]. Detecting [602, 1548, 1804, 1981, 1674]. Detection [2086, 1028, 1877, 56, 1514, 1583, 2069, 861, 2221, 1406, 1638, 1682, 1959, 1728, 1094, 642, 1590, 1783, 1321, 847, 1012, 1931, 1906, 1052, 289, 1418, 855, 1486, 1987, 2224, 2177, 1940, 883, 1050, 1100, 1662, 1323, 1712, 1309, 2191, 2183, 2023, 2039, 1394]. detection/localization [855]. determinant [1781, 1508, 1574, 2004]. determinant-and-shape [1574]. determinantal [190]. Determination [403]. determining [167]. Deterministic [845, 594, 1633, 1297]. development [626]. developments [232]. deviates [65]. deviations [449]. DEXPERT [25]. diagnosing [1981]. diagnostic [636, 150, 280]. Diagnostics [501, 733, 793, 638]. Diagonal [1312, 2120]. diagram [1761]. diagrams [521, 78]. Dickman [1952]. difference [766, 2101, 1350, 732]. different [928, 1929]. Differential [590, 677, 1573, 2063, 851, 1497, 266, 1307, 1284, 1244, 1530, 1471, 1467, 1905, 1496, 1559, 1802, 1696, 1393, 836, 2121, 1098, 1498, 1141, 986, 1651, 1296, 1742]. differentially [1739, 1477]. differentiation [1831, 1762]. Difficulties [226]. diffuse [1620]. diffusion [2198, 1946, 1633, 1537, 2055, 1687, 1920, 121]. diffusions [1700, 2140, 2122, 1421, 598, 1525, 863, 2012]. Diffusive [819]. Diggle [791]. digraphs [1085]. dilution [69]. Dimension [1944, 764, 257, 1933, 1925, 2100, 1975, 1870, 1960, 1570, 1214, 1533, 795, 1297]. dimension-based [257].

Dimension-independent [1944]. dimensional [1877, 953, 997, 1760, 1455, 1598, 1314, 2010, 1275, 1925, 1756, 1624, 1666, 1955, 1914, 1129, 500, 784, 1548, 1755, 1950, 2058, 308, 685, 1476, 1227, 1137, 2036, 2209, 1691, 847, 1858, 1926, 1395, 1743, 1807, 1704, 2084, 2057, 1928, 1757, 1327, 855, 2204, 233, 1977, 2037, 1891, 1581, 1393, 1207, 738, 1396, 2224, 2177, 1520, 1918, 636, 67, 1676, 2061, 1887, 1541, 1182, 2003, 2111, 2207, 1394, 788, 1300]. dimensions [1958, 1365, 1437, 2044, 273, 1591, 1723, 1346]. DINDSCAL [852]. Direct [550, 1805, 250, 1656, 852]. directed [1431, 466, 1208, 1247, 1565, 1764]. direction [1590]. Directional [2197, 1628, 715, 578, 592, 1949, 874]. Directions [401, 105, 286, 795]. Dirichlet [797, 1270, 1138, 1160, 2040, 571, 1301, 2193, 1813]. Disclosure [498, 462, 1478]. discontinuities [602]. discoveries [1794]. discovery [2156, 1892, 2117]. discrepancy [1758, 2096, 1297, 1797]. Discrete [38, 1966, 487, 1986, 1978, 53, 546, 550, 1431, 1476, 1458, 1250, 544, 1785, 1585, 1026, 382, 1709, 267, 1803, 1257, 274, 694]. discretely [1700, 652, 1089]. Discretisation [483]. discretization [514, 624, 2091]. discretized [1502, 816]. discriminant [893, 1146, 38, 1275, 1870, 1196, 1554]. discriminating [1231]. discrimination [371, 240, 446, 1703]. Discriminative [1690, 1640, 2079]. Discussion [318, 313, 310, 319, 309, 320, 311, 316, 312, 315, 412, 252, 317, 314, 253]. disease [159, 603, 1631, 1887]. dispersed [1584]. dispersion [1237, 563, 647, 2071, 1393]. display [230]. distance [1881, 1060, 802, 1059, 2134, 1858, 636, 1262, 1441]. distance-based [1858, 636, 1262, 1441]. distances [847, 1365]. distinctions [142]. distinguishing [1001]. Distributed [2132, 2085, 1879, 855, 1856, 2111, 1665, 1922, 2082, 1955, 706, 2103, 2087, 1265, 1785, 1830, 2172, 278, 2112, 1302, 2039, 1586]. Distribution [459, 188, 251, 669, 1913, 2119, 423, 780, 727, 1147, 1998, 1636, 797, 996, 167, 529, 1178, 228, 668, 1230, 541, 2175, 2068, 970, 1325, 508, 697, 516, 613, 1017, 825, 17, 1852, 1113, 2040, 1384, 884, 1101, 461, 1388, 375, 1940, 1838, 720, 870, 1091, 554, 2197, 1269, 1089, 2181, 2030]. Distributional [1719, 780, 1854, 1387]. distributions [189, 699, 1936, 2005, 893, 888, 953, 997, 94, 1681, 928, 728, 789, 1535, 264, 75, 47, 1859, 1536, 1270, 718, 550, 517, 1047, 125, 388, 1875, 1952, 2201, 113, 726, 2067, 842, 338, 2033, 1882, 687, 1620, 2052, 915, 348, 1334, 829, 592, 1397, 2185, 73, 398, 962, 992, 1193, 1291, 753, 1720, 816, 1445, 194, 908, 831, 1982, 1910, 218, 1438, 1240, 242, 480, 510, 428, 883, 1151, 1776, 555, 667, 523, 1815, 810, 1014, 169, 351, 694, 2145, 2213, 2096, 148, 2136]. divergence [2143, 926, 2092, 1822]. divergences [1212, 1984]. diverging [1324]. diverse [2115]. divide [2218]. divide-and-conquer [2218]. Divisive [1227, 2069]. DNA [1103]. Do [2007]. Does [1158]. domain [386]. Domains [72, 867, 1929]. dose [1765, 985, 391]. dose-escalation [1765, 985]. dose-response [391]. DOSTM [229]. Double [950, 2121, 1735, 1422]. Double-loop [2121]. Double-Parallel [1422]. Doubly [2020, 1963, 423, 1074, 1073, 1888, 2217]. doubly-sketching [1888]. Douc [1192, 1071]. Dr. [412]. DRAM [599]. drift [1700]. driven [1531, 1958, 1517, 1471, 1467, 1810, 1873, 1802, 869, 1794, 28]. driving [472]. dropping [330]. Drovandi [1061]. dual [1414]. dual-tree [1414]. dually [976]. durations [2191]. Durmus [1059]. dyadic [1826]. Dynamic [1474, 1778, 1418, 1367, 1406, 1461, 983, 954, 685, 551, 2188, 339, 1721, 1992, 624, 2071, 1581, 1340, 2190, 1923, 786]. dynamics [1951, 261, 175, 1751].

each [1793]. early [124]. earthquake [2098]. easy [590]. ECME [1834]. economical [1201]. ECOPICA [1976]. EDF [37]. Edge [976, 642]. edges [1345, 1461]. edition [1490]. Editor [1114, 1242, 1154]. Editorial [24, 51, 64, 118, 323, 671, 1490, 1, 100, 174, 322, 422, 470, 735, 400, 353, 344, 409, 378, 491]. eds [791]. effect [324, 325, 2188, 441, 576, 1870, 2025]. effective [297, 1068, 1067]. effectiveness [87]. effects [1010, 1427, 2011, 1298, 693, 1509, 2080, 524, 1045, 1163, 159, 1241, 1029, 749, 1228, 831, 1977, 2037, 1031, 828, 1039, 1454, 1836, 1762, 1613, 216, 986, 572, 1482, 619]. efficiency [431, 1470, 149, 47, 1252, 946, 1626]. Efficient [762, 2073, 1665, 823, 1074, 1133, 1383, 2225, 1849, 1881, 1625, 1886, 2124, 546, 785, 983, 950, 1079, 1073, 1989, 1462, 1419, 1488, 1787, 68, 1767, 599, 1241, 580, 195, 962, 1261, 331, 2056, 196, 477, 365, 911, 547, 658, 1549, 1080, 14, 971, 214, 1945, 2178, 1642, 1576, 1355, 601, 115, 1215, 2055, 1636, 1922, 2082, 882, 163, 586, 1634, 1465, 1664, 1688, 1449, 1476, 2222, 1631, 1589, 1321, 192, 19, 1801, 1123, 158, 269, 2169, 1105, 1485, 1962, 2061, 2039]. efficiently [1863, 1893]. eigen [993]. eigen-decomposed [993]. Eigenfunction [1700, 1767]. eigenvalue [780]. eigenvalues [74]. elastic [1855, 804, 1203, 833, 2003]. elastic-net [2003]. electrical [601, 682]. element [1384, 2091]. elements [2114]. elicitation [1935]. Eliciting [1270]. elimination [1863, 165]. elitist [1268]. ellipsoid [440]. ellipsoidal [658]. elliptic [1117, 1873]. elliptical [1438, 1662, 1815]. elliptically [687, 1388]. EM-based [510]. EM-type [842, 237, 2049]. EM-variational [1642]. EM/Gauss [189]. Embedded [1878, 205, 2]. embedding [1993, 1795]. emerging [1309]. emission [358, 1014]. Empirical [2186, 238, 1501, 1048, 1927, 1281, 806, 2144, 532, 900, 820, 949, 2072, 970, 1645, 1857, 2171, 571, 1143, 1976, 1774, 28, 1324, 1541, 1963]. empirical-likelihood [949]. Emulation [1696, 1806]. Emulation-accelerated [1696]. emulator [1645]. emulators [1153]. endogenous [783]. energetic [1630]. energy [940, 882, 1671, 2150]. engineering [349, 912]. Enhancing [2066, 2174, 747]. Enmsp [2003]. Ensemble [406, 1624, 1605, 1657, 1362, 1699, 1846, 1349, 2108, 913, 2224, 2177, 1105, 1722, 1770, 710]. ensembles [1611, 1603, 2115]. Entropic [1814]. entropy [894, 1191, 1190, 820, 1910, 493]. enumeration [1782]. envelope [1868]. envelopes [1892, 658]. environment [55, 150, 241]. environmental [159]. Envisioning [15]. epidemic [1457, 596, 1915, 1400, 1706, 464]. epidemics [1052, 567, 911]. Epidemiologic [1505]. epigraph [1819]. epilepsy [589]. equal [2096]. equality [166, 367]. equation [851, 1497, 266, 1284, 1402, 1905, 1696, 2121, 2146, 986]. equations [2063, 1307, 1244, 1530, 19, 1467, 2002, 987, 1496, 1559, 1802, 1289, 1738, 638, 836, 1350, 1098, 1498, 1141, 1651, 1296, 1742, 1471]. equi [940]. equi-energy [940]. equilibrium [621]. equispaced [574]. equivalence [1200]. Erdos [1689]. ergodic [1060, 1059]. Eric [1192]. Erlang [1692]. Erratum [1204, 62, 198, 1246, 1191, 1359, 1329, 1377]. Error [407, 1595, 437, 1210, 608, 1821, 2107, 779, 1627, 172, 1947, 1112, 1870, 1342, 2091, 1908, 1391, 376, 1423, 260, 2181]. error-prone [1821, 2107]. errors [857, 2116, 1223, 1785, 1906, 1928, 1511, 211, 1454, 811]. errors-in-variables [1906, 1928]. escalation [1765, 985]. estimate [2049, 1619, 512, 1411, 2113]. estimated [417, 215, 943, 418]. Estimates [573, 1403, 1230, 441, 1968]. Estimating [938, 2151, 896, 879, 776, 1264, 1007, 473, 366, 1369, 990, 986, 1346, 1565, 1700, 1823, 1244, 1844, 339, 987, 711, 1289, 638, 161]. Estimation [1936, 1025, 608, 932, 824, 1512, 779, 454, 1149, 2012, 1528, 831, 1982, 411, 1632, 737, 124, 626, 482, 2213, 538, 1700, 860, 699, 1739, 419, 574, 1679, 1997, 857, 1527, 709, 1010, 72, 997, 611, 1164, 1310, 872, 740, 659, 115, 1319, 935, 587, 372, 993, 1986, 994, 861, 873, 693, 593, 894, 708, 1185, 1636, 1922, 1821, 2225, 2158, 934, 1272, 821, 1538, 961, 2196, 212, 1955, 2089, 1886, 1859, 217, 1412, 429, 2124, 1824, 777, 1426, 1755, 1959, 950, 1409, 1030, 641, 631, 1266, 1421, 416, 603, 1465, 1737, 1686, 1084, 989, 1431, 1476]. estimation [2103, 2087, 2222, 1108, 1779, 1241, 2148, 1474, 1127, 415, 347, 1718, 1347, 1617, 931, 489, 645, 629, 970, 76, 1325, 2020, 781, 1054, 881, 1331, 1450, 903, 1809, 945, 304, 1029, 442, 1673, 1106, 1489, 2114, 2057, 1928, 471, 1381, 2171, 2218, 1418, 1639, 816, 916, 2152, 1472, 1852, 1860, 632, 2056, 1033, 1384, 921, 1368, 984, 1830, 1208, 1364, 1442, 1570, 1696, 1039, 390, 449, 1987, 510, 1549, 428, 1774, 722, 1415, 268, 755, 1448, 1740, 1391, 1722, 409, 1748, 2026, 1771, 202, 1206, 1229, 1587, 837, 376, 1622, 1423, 2008, 1338, 1121, 1637, 859, 786, 1818]. estimation [1256, 1093, 1249, 214, 1182, 1377, 1376, 2110, 1921, 1963, 1927, 2145, 714, 2219, 418, 1277, 1871, 1730, 1742, 651, 2062, 874]. Estimator [457, 454, 1569, 1913, 854, 1508, 1919, 712, 440, 2051, 774, 1589, 968, 1860, 946, 1339, 160, 2004, 1483, 1141, 2083]. estimators [1255, 1010, 1773, 728, 800, 2155, 224, 696, 557, 514, 234, 703, 959, 761, 1693, 441, 981, 1487, 384, 965, 1480, 1342, 1416, 208, 1378, 1177, 1745, 1195, 1036]. ETAS [1475, 1706, 2098]. euMMD [1893]. Evaluating [1639, 718, 1379]. Evaluation [727, 217, 647, 1167, 1037, 1895, 94, 563, 1325, 1777, 1213, 17, 1774, 1709, 2149, 1329, 1328, 1965]. Evaluations [814, 506, 489]. even [421]. event [935, 873, 1959, 1353, 1557, 1567, 1396, 1916, 1918, 1803, 1948, 874]. events [1528, 2146]. evidence [21, 1030, 1474, 834]. Evolution [590, 677, 102]. Evolutionary [105, 101, 1986, 1606, 607, 799, 107, 673]. evolving [1506, 1721]. EVPPI [1465]. EWMA [727, 569]. EWMA- [569]. Exact [1210, 780, 1294, 1178, 2202, 586, 1688, 643, 290, 1092, 1208, 573, 883, 1754, 1323, 161, 2149, 391, 694, 1477, 1661, 1015, 797, 1513, 1507, 479, 1139, 68, 544, 772, 1397, 365]. example [1835, 387]. examples [1192]. exceed [136]. exceedances [1025]. exchange [1271, 1268, 1156]. exchangeable [1783, 1996]. excitable [4]. excitatory [1600]. exciting [1515, 1622]. excursion [1824, 750]. exogenous [175]. expansion [2076, 616]. expansions [379, 1767, 836]. expectation [777, 1644, 1973, 619]. expectations [1823, 1886]. Expected [560, 1595, 1285, 2124, 1755, 952, 376]. expected-posterior [952]. Expectile [2021, 1990, 2027, 1737, 2057, 925, 1336]. expectiles [2194, 2054, 1259]. experiment [1392]. Experimental [323, 1623, 1144, 1661, 2157, 1268, 2124, 1641, 1294, 1812, 1895]. experiments [865, 872, 2123, 1271, 1629, 1156, 325, 1507, 298, 327, 1004, 889, 1446, 1517, 868, 1669, 2135, 25, 331, 1308, 866]. expert [21, 20, 25, 269, 26]. experts [2022, 1002, 2078, 1866]. Explainable [1930, 2077]. Explaining [1177]. explanation [410]. explanatory [1919]. explicit [2161, 1683, 177]. Exploiting [1598, 1395, 1693]. exploits [1494]. exploration [80, 882, 1669]. Exploratory [2190, 930]. Exploring [286, 815, 247]. exponential [54, 146, 1874, 1176, 556, 563, 647, 1989, 2060, 2206, 2223, 1101, 1543, 1154, 1637, 2023]. exponential-family [2060]. exponentials [595]. Extended [1393, 1176, 1809, 395, 1143, 531]. Extending [798, 2135, 808]. extensibility [374]. extensible [1425]. extension [1938, 837, 302, 1577]. Extensions [1225, 2064, 643]. external [1978]. extraction [1022, 1749]. extrapolation [1670]. extremal [1777, 1516]. Extreme [1906, 1936, 2105, 2054, 157, 1737, 2128, 1918]. extremes [115, 1906, 2061, 1378].

F [1069, 1079, 1071]. face [704, 610]. facial [536, 610]. factor [1654, 54, 188, 1604, 798, 1659, 1913, 324, 1147, 1084, 1844, 2067, 1127, 1649, 1219, 1992, 1529, 1707, 1340, 1741, 930, 721, 1793, 1464, 651, 1394]. Factored [704]. factorial [325, 327]. factorisation [382]. factorised [1089]. factorization [1839, 885]. factors [146, 251, 2123, 1382, 1804, 2081, 1198]. failure [872, 2107, 1019, 1272, 2041, 759, 1338, 1602, 1991, 2145, 2001]. fairness [1750]. False [1892, 2117]. familial [159]. Families [341, 146, 1176, 388]. family [181, 112, 1047, 2060, 1112, 1982, 1543, 1984, 1637, 1815, 2023]. family-wise [1112]. FANOVA [869]. FANOVA-decomposition [869]. FAST [484, 1066, 2141, 1019, 789, 21, 1513, 1403, 1644, 1603, 1244, 1590, 1493, 1325, 1809, 1704, 1996, 2223, 1939, 1656, 2220, 1342, 1837, 1132, 1065, 1182, 1376, 2126, 696, 1808, 944, 1631, 931, 1839, 1645, 1395, 245, 1864, 1819, 1698, 1089, 1145, 1377]. Faster [1502, 1387]. FastHCS [1235]. Favaro [1067]. favor [1611]. FDR [814]. FDR-controlling [814]. Feasible [755]. Feature [620, 736, 2075, 2209, 1749, 1993, 2079, 1207, 1748, 1041, 1300]. features [2011, 2203]. feedback [1816]. Ferguson [1243]. few [435]. few-stage [435]. fewer [677]. fibre [1548]. field [1605, 1711]. fields [1576, 1580, 1287, 1320, 1128, 335, 1513, 1558, 1656, 1415, 1075, 1076]. filling [866, 1262]. filter [1521, 1859, 1643, 1846, 913, 1581, 974, 1105, 1373, 1770, 1699]. filtered [1700]. Filtering [526, 910, 1788, 362, 123, 954, 1575, 1691, 1567, 918, 2035, 1098, 1498]. filters [584, 1668, 505, 1108, 1347, 1471, 1467, 1575, 1620, 1596, 1222, 1860, 1722]. final [596]. finance [878]. financial [1151, 2191]. find [1634]. Finding [345, 52, 1053, 1372, 165, 277, 1772, 269, 1366]. Finetti [1122]. Finite [1276, 992, 1193, 810, 653, 783, 2005, 587, 397, 932, 1644, 1159, 58, 1599, 1964, 1445, 1175, 1543, 618, 2091, 1943, 1151, 555, 1385, 1541, 1360, 306, 2213, 1942]. finite-sample [1541]. finite-sum [1644]. first [1144, 1459, 1842]. first-order [1144]. first-passage-time [1842]. Fisher [1640, 2030, 307, 318, 1636, 1176, 313, 1621, 310, 319, 309, 697, 1397, 320, 1649, 311, 316, 312, 315, 317, 1551, 314]. Fishers [1212]. fit [1958, 414, 1882, 2052, 769, 2099, 887, 884, 1240, 1709, 2181]. Fitting [1427, 463, 1831, 576, 1735, 991, 216, 1607, 543, 621, 648, 707, 1683, 692, 329, 2185, 1894, 465, 646, 1697, 731, 2001]. fixed [1514, 52, 487, 1045, 441, 510]. fixed-effect [441]. fixed-point [52]. flat [976]. Flattening [1601]. flexibility [431, 2175]. Flexible [475, 1124, 2068, 2130, 1054, 1017, 825, 1803, 1110, 1889, 601, 1298, 1833, 1516, 2040, 1219, 1546, 1429]. flow [1738]. flow-dependent [1738]. flows [1786]. focus [1547]. focused [1142]. fold [779, 2163]. Föllmer [1786]. forces [1670]. forecast [1660, 1563]. forecaster [1005]. Forecasting [406, 803, 1775, 1357, 539]. forecasts [760, 1820, 1945]. forest [1933, 2100]. forest-based [2100]. forests [1683, 1283, 979]. ForLion [2081]. form [1919, 1643, 1693]. formation [1826]. forms [43]. formula [2076, 2165, 509, 1105, 554]. formulæ [1023]. formulation [1846, 971]. Forthcoming [33]. Forward [2006, 334, 1868, 448]. forward-backward [1868]. forwards [1125]. four [2123, 324, 59]. four-colourability [59]. four-level [2123]. Fourier [2141, 524, 647, 484, 2012, 1740]. fractional [794, 994, 845, 1802, 1469, 1676]. frailty [2065, 945]. framework [806, 114, 1155, 1934, 1717, 374, 808, 1869, 2147]. Fréchet [1858]. Fredholm [1459]. Free [1912, 882, 1588, 1951, 955, 1816, 1165, 2089, 1715, 1370, 1670, 2180, 1674, 911, 1560, 1533, 2098, 1907, 1887, 1792, 1300]. freedom [1168]. frequencies [797, 1232]. frequency [47, 1947, 386]. Frequent [635, 1037]. Frequentist [2206, 2163]. Friedman [307, 318, 313, 310, 319, 309, 320, 311, 316, 312, 315, 317, 314]. Frost [464]. Frugal [1843]. fruit [69]. Fuentes [791]. full [272]. fully [1366, 974]. funBIalign [2156]. Function [454, 2188, 487, 111, 423, 1509, 2158, 777, 1702, 2074, 1716, 489, 1325, 1882, 2152, 1136, 427, 1031, 1987, 1805, 1911, 1549, 200, 409, 859, 1378, 1356, 2181]. Function-on-function [2188, 1911]. function-on-scalar [2158]. Functional [291, 1946, 2022, 1082, 781, 1341, 1545, 2127, 1840, 1531, 1066, 1298, 851, 1314, 2089, 1702, 1779, 1217, 1716, 2156, 1858, 304, 1931, 2014, 2139, 1610, 1929, 946, 1308, 1390, 1960, 1436, 1903, 1819, 2115, 1282, 1065, 1344, 1741, 664, 2078, 809, 1901, 1182, 1377, 1376, 2192, 1798, 2181]. functionals [1680, 1823, 1133]. functions [1700, 1580, 257, 748, 595, 635, 1217, 970, 288, 413, 1232, 533, 1366, 839, 1838, 1199, 737, 1119, 2190, 1744, 1429, 1269, 971, 1411, 193, 1332, 538]. fundamental [1193]. Further [74]. Fused [2044, 2048, 1174]. fusion [1654]. future [81, 82, 83, 421, 91]. Fuzzy [2138, 2194, 530, 2038].

G [1079]. gain [2124, 1755]. GAM [2045]. gamma [242, 523, 708]. GAN [1745]. gap [1944]. gappy [1512]. GARCH [878, 761, 1980, 695, 1969]. GARCH-type [761]. gas [1357]. Gauss [189, 1446]. Gaussian [1453, 1576, 1523, 1843, 953, 997, 2063, 2211, 1333, 1847, 1867, 2141, 1035, 688, 1128, 527, 1046, 854, 1146, 2032, 1133, 335, 2155, 1290, 1538, 789, 882, 1629, 1752, 1512, 1513, 1440, 2054, 1129, 996, 532, 900, 1270, 123, 723, 963, 1686, 1767, 1621, 1466, 976, 1018, 1717, 1640, 1386, 1839, 1596, 1743, 1807, 1672, 1694, 1853, 398, 248, 1639, 2152, 1175, 1746, 665, 750, 1830, 1167, 1339, 1364, 1388, 1656, 1827, 1806, 573, 456, 480, 1153, 1837, 2133, 1549, 1415, 1214, 1800, 211, 700, 1098, 1697, 540, 664, 581, 1578, 1526, 1469, 1923, 1236, 1361]. Gaussian [1568, 1095, 1562, 721, 1424, 1152, 351, 302, 520, 450, 901, 1736, 1107, 535, 2178, 1678, 594]. Gaussian-process [1839]. Gelfand [791]. gene [912]. General [1237, 1857, 206, 437, 2198, 678, 2041, 20, 1264, 2081, 617, 1048, 1985, 1511, 2044, 1393, 1812, 1194, 1346, 1822]. generalisation [2079]. generalisations [685]. Generalised [1974, 1577, 596, 830, 1686, 2086, 1957, 1448, 2093, 2034, 1277]. generality [47]. generalizable [1849]. generalization [257, 1681, 818, 1348]. generalizations [518]. Generalized [1877, 1888, 1663, 2028, 816, 2048, 1056, 1429, 1008, 1607, 206, 437, 862, 782, 648, 823, 1919, 879, 186, 2051, 1955, 556, 75, 996, 1855, 484, 1045, 228, 620, 668, 1688, 989, 1082, 1018, 2068, 1040, 364, 1504, 1258, 1502, 1884, 2160, 1445, 191, 2195, 411, 2077, 1930, 1031, 828, 1173, 510, 428, 2215, 565, 755, 1132, 1723, 1973, 1897, 410, 2129, 1174, 1963, 1168, 2016, 438, 2062]. generalizers [260]. generated [65, 708, 1571, 750]. Generating [1018, 461, 720, 169, 1764]. Generation [14, 1201, 75, 996, 2033, 1085, 73, 1996, 445, 1656, 2090, 658, 719, 1815, 768]. generative [2094, 1896]. generator [1021, 66, 1761]. generators [504]. Generic [717, 1023, 1358]. genesis [1115]. Genetic [106, 104, 10, 431, 590, 244, 159, 521, 2222, 225, 103]. genetics [1631]. genomic [861]. genotype [150]. Gentle [790]. geoadditive [1152, 1336]. geodesic [1501]. geodesics [1176]. Geographically [2179, 1034, 2176]. geometric [1553, 1979, 373, 1393, 1075, 1076]. geometrically [1060, 1059]. Geometry [1751, 976, 477, 2079]. geometry-aware [2079]. Geometry-informed [1751]. geospatial [1055]. geostatistical [1747, 633]. German [2052]. Gibbs [1075, 1197, 459, 587, 149, 113, 2180, 1559, 2223, 2050, 1560, 208, 300, 1076, 1439, 368, 12, 1188, 260, 1542, 1482]. Gibbs-like [1542]. gift [366]. Gillespie [1077]. GIPSCAL [434]. gives [366]. GKnock [2129]. GLM [1683]. GLMs [1393]. Global [875, 364, 876, 1994, 1731, 1371, 1299, 799, 2223, 921, 170, 1792]. Global-local [1994]. Globally [1851, 1681]. Gneiting [1580]. Gneiting-type [1580]. Goodness [1882, 887, 884, 1958, 2052, 769, 2099, 1240, 1709, 2181]. Goodness-of-fit [1882, 887, 884, 1958, 2052, 769, 2099, 1240, 1709, 2181]. GP [1706]. GP-ETAS [1706]. GParareal [1806]. GPGPUs [1325]. GPU [2131, 1439]. GPU-accelerated [1439]. Gradient [1387, 1588, 1069, 1456, 487, 1849, 1070, 1731, 1161, 1435, 1848, 1466, 1779, 1617, 1397, 1635, 2070, 1472, 1832, 1899, 2031, 1091, 124, 2149, 434, 1983, 2109]. gradient-free [1588]. Graph [1689, 1776, 1351, 1914, 225, 1232, 1647, 1708]. Graph-based [1776, 1647]. Graphical [1709, 1453, 1523, 709, 1569, 1063, 812, 1290, 900, 1079, 373, 2182, 976, 466, 1250, 2084, 1778, 717, 1830, 2071, 1142, 1064, 2133, 211, 1323, 383, 1080, 230, 1016, 1648, 214, 2113, 901, 1187]. graphics [249, 373, 209]. graphics-based [209]. graphon [1232]. graphs [1420, 1401, 655, 1431, 1689, 1501, 1028, 2060, 763, 1208, 1753, 1247, 1724, 1194, 1648]. Gray [126]. great [421]. great-and [421]. Greater [83]. Greedy [2069, 2075, 1797, 1485]. grid [1372, 1669, 1425]. Gröbner [382]. Group [1081, 2046, 2207, 2204, 2168, 1145, 1305]. group-lasso [1145]. grouped [1081, 1666, 804, 1964, 2048, 1748, 2149]. groups [815, 367]. growing [2057]. growing-dimensional [2057]. growth [2198, 719]. growth-interaction [719]. Guest [353, 323, 344, 409, 378, 735, 491, 400]. guided [287, 1581]. Gumbel [1113]. gunfire [1600]. Guttorp [791].

H [307, 609]. H-likelihood [609]. Haar [821]. Hamiltonian [1395, 1537, 1759, 1873, 1696, 1524, 1003, 1972, 1332]. Handbook [791]. Hara [901]. Hastings [1069, 1789, 1234, 676, 1070, 578, 1687, 923, 959, 1233, 1359, 1179]. Having [470]. Hawkes [1874, 2009, 1771, 1622, 1642]. haystacks [1772]. hazard [414, 182, 445, 1987, 1348, 1356]. hazards [1666, 1861, 2132, 1324, 201]. healing [1252]. heaps [79]. heavy [1936, 728, 1047, 1006, 1980, 831]. heavy-tailed [1936, 728, 1047, 1006, 1980, 831]. Help [9, 16, 23, 34, 42, 50, 63, 71, 77, 93, 99, 109, 117, 129, 137, 153, 162, 173, 184, 199, 213, 222, 231, 239, 246, 256, 262, 270, 281, 293, 303, 321, 333, 343, 352, 360, 369, 377, 389, 399, 408, 420, 430, 439, 451, 460, 469, 481, 490, 503, 511, 519, 528, 537, 545, 552, 561, 570, 579, 588, 597, 604, 612, 622, 630, 640, 650, 660, 670, 680, 691, 702, 713, 724, 734, 746, 757, 767, 778, 792, 805, 822, 843, 864, 877, 890, 904, 914]. Help [924, 937, 947, 957, 967, 977, 988, 999, 1013, 1027, 1042, 1057, 1126, 1140]. herded [1482]. herded-Gibbs-type [1482]. herding [1814]. Hessian [1069, 1070, 795, 1561]. Heston [1263]. heterogeneity [701, 180, 2172, 2183, 2019]. Heterogeneous [1964, 1514, 1948]. heteroscedastic [817, 1466, 617, 661, 856, 1562, 482]. heteroskedastic [674]. Heuristic [165, 1790]. heuristics [853, 644]. HGLMs [886]. Hidden [1159, 1148, 2106, 1014, 1102, 1978, 1191, 1190, 1032, 2202, 1990, 1052, 1558, 1653, 716, 711, 803, 1734, 268, 1075, 2035, 1076, 1483, 302, 1840]. Hierarchical [1455, 346, 458, 53, 1337, 1691, 540, 2069, 296, 641, 723, 1212, 2156, 441, 2208, 194, 1368, 1735, 885, 1908, 1956, 971, 1813]. Hierarchical-block [1455]. hierarchical-likelihood [441]. hierarchies [2038, 1585]. High [2010, 1925, 1129, 2084, 1676, 1887, 1541, 1877, 953, 997, 1298, 1760, 1455, 1598, 1933, 1756, 1666, 1955, 500, 784, 1755, 1950, 2058, 286, 308, 1476, 1227, 2036, 2209, 1691, 847, 1365, 1926, 1395, 1743, 1807, 1704, 357, 1928, 1757, 1947, 968, 1327, 855, 2204, 1437, 1977, 2037, 1581, 1393, 1207, 1396, 2224, 2177, 1520, 719, 1214, 1723, 67, 2061, 1346, 1182, 2003, 2207, 1394, 788]. high-dimension [1933]. High-dimensional [2010, 1925, 1129, 2084, 1676, 1887, 1541, 1877, 1760, 1455, 1598, 1756, 1666, 1955, 500, 1755, 1950, 2058, 308, 1476, 2036, 1691, 1926, 1395, 1743, 1807, 1928, 1757, 1327, 855, 2204, 1977, 2037, 1581, 1393, 1207, 1396, 2224, 2177, 1520, 67, 2061, 1182, 2003, 2207, 1394, 788]. high-frequency [1947]. high-resolution [357]. higher [1133, 1566, 2118, 385, 273]. higher-order [1566, 2118, 385]. highly [1538, 1026, 450]. Hilbert [2010, 1800, 1119, 1526]. Hill [66]. histogram [297, 1579]. histogram-valued [1579]. historical [1298, 1610]. hit [243]. HMC [2120]. Hölder [1415]. holonomic [1397, 1091]. HOLP [2209]. homo [1562]. homo- [1562]. homogeneity [1124]. homogeneous [1028, 803]. homophilic [1826]. homoscedastic [661]. horizontally [939]. horse [366]. Horseshoe [1439]. Hougaard [1842]. household [911]. HPC [2162]. Huber [1748, 2150]. Huber-energy [2150]. Huberised [2068]. Huberised-type [2068]. huge [1843]. human [110, 610, 835]. hunting [308]. Hybrid [1005, 435, 1935, 189, 862, 1268, 595, 442, 2195, 624, 276, 2109, 2137]. hyper [1907]. hyper-parameter [1907]. hyperbolic [75, 1957, 1445, 510]. hypercube [867, 966]. Hypergeometric [797, 529, 962]. hypergraph [395]. hypergraph-theoretic [395]. Hyperparameter [2203, 1811]. hyperparameters [258]. hyperuniformity [1844]. hypograph [1819]. hypograph-based [1819]. hypotheses [1007, 367]. Hypothesis [1238, 558, 2143, 2119, 1833, 1534, 814, 1007, 1112, 2111].

I-robust [1360]. idea [280]. Ideal [376]. identically [1665, 278]. Identifiability [1978, 1730, 1707]. Identifiable [306]. Identification [275, 954, 2108, 610]. identified [1851]. identify [383]. identifying [568]. IDGM [2113]. if [1595]. II [478]. IID [260]. III [478]. illustration [1537]. illustrative [1032]. Image [353, 354, 653, 355, 219, 1752, 1548, 244, 1061, 685, 1212, 154, 1599, 1896, 1062, 722, 1415]. imagery [642]. images [1995, 1375, 357, 289]. imaginary [2142]. Imaging [682, 1883]. imbalanced [1843, 2090]. impact [470, 1826]. Implementation [295, 2131, 853, 1032, 171, 1105, 1485, 1086]. implementations [1138]. Implementing [155]. implication [194]. implications [770]. implicit [1581]. Implicitly [1612]. Implied [888, 760]. Importance [1457, 1714, 556, 407, 725, 1589, 1608, 1281, 1823, 1841, 1665, 935, 678, 1015, 679, 1886, 826, 1755, 1380, 1283, 979, 192, 1670, 442, 187, 1026, 473, 1280, 392, 1612, 2121, 2146, 1568]. Important [82, 84, 90]. imprecise [489]. imprecision [2092]. improper [1154]. improve [1158]. Improved [659, 894, 484, 723, 172, 1647, 1876, 1224, 1296, 1255, 736, 1474, 1029, 1387, 901]. improvement [1004, 1626]. Improvements [2009]. Improving [431, 2153, 1470, 1083, 1030, 340, 908, 2024, 2110, 1745]. impulsive [2151]. Impurity [758]. Imputation [1587, 1273, 2186, 2084, 841, 2045, 1614, 1871]. IMS [1058]. incidence [1915]. including [859]. inclusions [682]. Incomplete [1551, 762, 1412, 161, 554]. Incomplete-data [1551]. incorporating [94, 1908, 1034]. increasing [777]. incremental [1380, 1644, 465]. increments [1251]. Independence [2116, 1035, 2010, 963, 675, 1693, 1591, 1177, 1352]. independent [938, 1823, 1665, 2126, 1088, 1172, 1251, 187, 1113, 2077, 1930, 1976, 1944, 278, 1862, 568]. index [1389, 2010, 2124, 1402, 948, 949, 2052, 756, 1106, 1880, 1832, 929, 1545, 1262, 960, 1927, 1277, 1315]. Indexes [255, 1582]. Indian [1316]. indices [361, 1391]. individual [1765, 2132, 1826]. individual-level [2132]. Individualized [2094]. INDSCAL [852]. induced [1629, 2056, 808]. induction [29]. inductive [28]. inequality [951]. inexpensive [165, 559]. Inference [2198, 1584, 223, 1874, 264, 1159, 2176, 1475, 1353, 624, 1581, 1519, 1304, 806, 1951, 1505, 380, 782, 621, 1791, 1938, 608, 515, 652, 483, 2153, 1147, 1816, 1383, 2158, 1756, 882, 1165, 1881, 2054, 1284, 433, 1178, 2162, 586, 1409, 479, 1808, 954, 1303, 674, 1664, 598, 1139, 1915, 1688, 666, 972, 68, 1183, 2182, 1370, 172, 842, 1600, 2180, 509, 1980, 1825, 628, 978, 2002, 2009, 1839, 1265, 1801, 1211, 1743, 1807, 474, 829, 1397, 1905, 1232, 2014, 1647, 1048, 1052, 1695, 2218, 1653, 1802, 1447, 1674, 936, 1706, 750, 1092, 1830, 1078, 1827, 1417, 2164]. inference [1837, 1524, 2212, 1077, 1098, 1323, 1118, 1533, 2098, 1263, 1923, 1236, 1257, 1698, 523, 721, 1324, 1630, 2112, 2183, 1089, 2214, 1699, 1736, 1969, 960, 2219, 1552, 1822, 2207, 1937, 2109, 2181, 1642, 594]. inferences [949, 749]. inferential [1301]. Inferring [1401, 812, 1420]. infinite [1624, 1053, 1918]. infinite-dimensional [1624, 1918]. infinitesimal [2071]. inflated [2164]. inflection [933]. influence [190, 521]. Influential [207]. Information [1102, 494, 1478, 271, 589, 1069, 1573, 1070, 1638, 1682, 2124, 774, 1755, 2116, 2202, 1004, 1293, 642, 2174, 1049, 1522, 1038, 233, 1131, 1854, 2079, 1142, 332, 15]. informative [2135, 1932]. Informed [1447, 1757, 1407, 1751]. INGARCH [1094, 2164]. inhibition [1874]. Inhomogeneous [1566, 984]. Initial [407, 2033, 1620, 1428]. Initializing [527]. INLA [2162]. inner [546]. innovations [1632]. input [1978, 1901]. input-output [1978]. inputs [1308, 175, 1229]. insights [1121]. instead [478]. instrumental [2020]. Insufficient [2050]. Integer [1507, 1021, 1980, 1173, 1790]. integer-valued [1021, 1980]. integral [1686, 1740]. integrals [1459, 635]. integrated [1808, 1213, 1306, 1735, 1408]. integrating [1705]. integration [1553, 2161, 2066, 2141, 845, 205, 2, 1492, 123, 1198, 1646, 1472]. integration-based [123]. Integrative [1926]. integrators [1496]. integro [1350]. integro-difference [1350]. intensities [708, 1692]. intensity [1480, 719]. inter [526, 332]. inter-block [332]. inter-departure [526]. Interacting [928, 515, 1299, 2210]. interaction [1287, 719]. interactions [565, 1100, 1826]. Interactive [305, 249]. interchange [324]. interesting [656]. intermediate [1581, 941]. international [87]. Interpolating [1781]. Interpolation [258, 867, 2128]. interpolatory [1023]. Interpretable [1436]. interpretation [821, 28]. interpretive [373]. interrupted [1908]. intersection [1200, 383]. interval [1209, 94, 1932, 2185, 920, 2217, 1481, 2101, 732, 2004, 1338, 241, 2113, 1850]. interval-censored [1932, 2185, 2217, 1338]. interval-valued [2004, 2113, 1850]. intervals [982, 1204, 1514, 40, 112, 1634, 544, 1784, 1123, 2025, 925, 694, 572]. interventional [1876]. intractable [1074, 1073, 1825, 1078, 1077, 1974]. intrinsic [952, 1804, 1870]. introducing [1307]. Introduction [1069, 1063, 1079, 131, 1073, 1061, 1059, 1067, 1071, 1075, 1065, 1077, 491, 400, 105, 1058]. invariant [1641, 1530]. Inverse [688, 1517, 2063, 1117, 1624, 996, 1446, 1686, 2175, 1018, 1617, 1839, 1704, 1872, 1663, 1873, 837, 732, 1907, 1249, 958]. inversion [1531, 1605, 1337, 647, 1701, 2058, 1704, 1727, 1711, 1269]. inversive [177]. Investigation [1293, 657]. investigations [2186]. involving [367]. ion [94]. Irregularly [456, 1752]. Irreversible [1432, 1751]. ISBA [1058]. island [1086]. Isometric [958]. isotonic [2044]. isotropic [1576, 514]. Issue [1114, 671, 2176, 1058]. Issues [323, 149, 500, 1138, 1301, 1207]. Itô [381]. item [1184, 1973]. Iterated [272, 1181, 40, 1343]. iteration [2023]. iterations [593]. Iterative [953, 975, 13, 1105, 1744, 858, 1561]. iteratively [572]. Ito [836].

J [791, 1069, 1071, 1075, 1077]. J. [139, 141, 143, 144]. Jackknife [294, 1857, 2007]. jacknifing [152]. Jaeger [1686]. James [790, 208]. Jelinski [36]. Jerome [307]. Jim [690]. Jittering [2053]. John [142, 889]. Johnson [1240]. Joint [1557, 1058, 744, 1742, 1898, 524, 876, 2079, 1916, 1723, 1741, 1177, 1008, 1577]. Journal [1924, 2170, 2200, 2227]. juice [69]. jump [857, 824, 718, 830, 1303, 1488, 2108, 717, 1432, 535]. junction [1753, 901].

Kalman [1605, 123, 1846, 913, 1105, 1722, 1770, 1699]. Kent [1993, 991]. Kernel [1146, 549, 58, 860, 574, 1997, 867, 372, 1933, 728, 1403, 784, 703, 76, 1785, 1487, 1987, 1186, 1119, 1662, 2197, 193, 1352]. kernel-based [2197]. kernels [1157, 980]. kind [1459]. kinetic [652, 1139, 1419, 1097, 1052, 936]. Kingman [1067, 1068]. KL [1162, 2092]. KL-optimum [1162]. Klass [1243]. knockoffs [2129]. knots [2089]. knowledge [28]. known [938, 487, 278]. Kolmogorov [1334]. Kriging [869, 1398]. Krishnamurty [496]. Kronecker [752]. Kryging [1747]. Krylov [1747]. Kullback [2143, 926]. Kumaraswamy [1998].

L [1853, 1075]. L-RVGA [1853]. label [754, 1254]. labeled [1690]. labeling [844]. labelled [956]. lack [414]. lack-of-fit [414]. landmark [1883]. Langevin [1060, 1059, 1961, 1625, 1868, 1380, 1421, 895, 1751]. languages [135]. Laplace [1408, 2157, 1284, 1755, 2058, 1808, 1466, 1905, 1735, 720, 1564, 1269, 1736]. Laplace-based [2157, 1564]. Large [1995, 2014, 2152, 456, 1352, 1665, 1096, 1420, 1401, 961, 1629, 1156, 692, 1264, 1138, 1241, 1085, 756, 1489, 1502, 2114, 2160, 1274, 1447, 1747, 1837, 1398, 927, 1710, 1971, 1257, 214, 450, 2111, 1828, 2062, 1433]. large-sample [1710]. Large-scale [1995, 2014, 2152, 1352, 961, 1629, 1156, 1264, 1489, 2114, 2160, 1747, 1971, 1257, 1828]. largest [1330, 780]. LARS [1393, 1305]. LARS-type [1305]. Lasso [1631, 2218, 2048, 897, 1569, 743, 942, 2044, 1174, 1016, 1145, 1305, 1010, 1788, 1772, 1437]. LASSO-type [1010]. last [1169]. late [476]. Latent [1389, 436, 1885, 1708, 902, 1451, 621, 1986, 2196, 1479, 1129, 1191, 1190, 832, 1684, 1020, 2067, 645, 1149, 1809, 1312, 1213, 530, 1721, 1442, 1570, 1416, 1973, 1741, 2190, 1724, 1177, 1152, 520, 1188, 850, 1441]. latent-scale [1441]. Latin [966]. latinized [328]. lattice [1493]. lattices [433]. law [1817, 1940]. laws [659]. Lawson [889]. layer [485]. Layered [1280]. layout [572].

Lazy [1166]. lead [1775]. learn [45]. learners [1611]. Learning [1841, 27, 1440, 1914, 1684, 2088, 1250, 1485, 535, 2066, 2119, 1583, 2082, 600, 607, 1449, 1621, 2222, 2097, 1690, 1522, 261, 1917, 577, 1899, 1769, 1543, 2168, 2031, 2090, 918, 2169, 2115, 1876, 1578, 1676, 1786, 1856, 1965, 2127, 1145, 1828, 1983, 2219, 958, 2147, 2137]. Learning-based [1841]. Least [740, 1929, 449, 2105, 334, 2073, 155, 1702, 2074, 1670, 2114, 2029, 2173, 943, 2001]. least-squares [2105, 2073, 2029]. leave [1329, 1328]. leave-one-out [1329, 1328]. left [2126, 242]. legal [277]. Leibler [2143, 926]. length [727, 1272, 1804, 274]. length-biased [1272]. less [2153, 2007]. lesser [83]. Letter [1154]. level [2123, 1337, 1530, 2132, 1639, 1962, 1790]. levels [161, 418]. Lévy [1471, 1452, 1467, 1957, 1961]. Lévy-driven [1471, 1467]. lifting [681, 1331, 582]. like [1542]. Likelihood [1951, 955, 1147, 416, 1370, 1975, 1560, 973, 206, 189, 251, 437, 558, 1532, 782, 748, 1128, 1986, 1816, 1401, 708, 1636, 212, 250, 1412, 1715, 1004, 1634, 1518, 1084, 820, 1108, 1127, 842, 466, 2180, 949, 580, 1617, 2072, 645, 1981, 304, 472, 1397, 1029, 441, 609, 749, 1009, 1048, 1489, 1857, 2171, 661, 1472, 1852, 1674, 1306, 1136, 384, 1982, 1992, 911, 1143, 1405, 1442, 1935, 2164, 510, 1189, 428, 2205, 268, 1723, 1897, 1533, 202, 350, 120, 1836, 2098, 2149, 1676, 2008, 1338, 1483, 731, 1324, 1356, 2023, 1974, 1579, 1963, 1927, 1792, 1036, 619, 1420]. Likelihood-based [1147, 1975, 973, 748, 1723]. Likelihood-free [1951, 955, 1370, 1560, 1816, 1715, 2180, 1674, 911, 1533, 2098]. likelihoods [1658, 163, 532, 1267, 526, 473, 1078, 1740, 1077, 410, 1974]. Limit [1777, 2076, 2165, 1605, 769]. Limitations [1768]. limited [1853]. limited-memory [1853]. limiting [517]. limits [140, 1026]. Lin [504]. Lindsten [1069]. line [861, 1323].

Linear [458, 1011, 453, 139, 146, 206, 437, 1460, 862, 2063, 1495, 782, 730, 2125, 1081, 1919, 224, 1756, 186, 1566, 706, 1641, 1045, 830, 641, 1223, 228, 631, 1357, 171, 373, 141, 989, 1264, 1108, 1888, 1111, 1575, 2188, 1040, 875, 413, 1504, 1029, 1810, 617, 1106, 1502, 1884, 1928, 2206, 2163, 1327, 140, 122, 194, 395, 1113, 1649, 1528, 887, 1610, 411, 1977, 2037, 119, 145, 138, 856, 1289, 2048, 1031, 828, 1173, 646, 1056, 1911, 2215, 1448, 143, 747, 1098, 144, 1897, 124, 202, 811, 1423, 2129, 1221, 142, 1008, 2093, 2034, 1141, 193, 450, 1963, 1041, 1168]. linear [1277, 1871, 973, 438, 2062]. linearity [1136]. LINEX [629]. link [186, 1056, 859, 1411]. linkage [498, 1379, 368]. linked [972]. Local [860, 6, 566, 486, 646, 1533, 1035, 1791, 1271, 372, 224, 1925, 1733, 1079, 642, 1109, 373, 1326, 1599, 398, 2223, 305, 1364, 1080, 1994]. local-seasonal-global [2223]. localization [855]. Localized [553, 1559]. Localizing [1190, 1191]. Locally [1629, 2074, 1681, 1702, 686, 881, 1414, 1939, 1339, 1891, 1311]. Location [440, 1607, 72, 371, 601, 824, 1172, 920, 968, 477, 1984, 273, 306]. location-scale [920]. loess [455, 468]. log [146, 1781, 1209, 1538, 2033, 1617, 1720, 194, 395, 1151, 2008, 1991]. log-concave [2033, 1720, 2008]. log-determinant [1781]. log-likelihood [1617]. log-linear [146, 194, 395]. log-logistic [1991]. log-stable [1151]. logistic [1081, 240, 1902, 479, 2099, 2040, 1836, 1602, 1991, 1798, 207]. logistic-normal [2040]. Logit [2187, 1994]. logit-beta [1994]. Lomeli [1067]. Long [66, 1450, 1363, 1512, 1331, 2218]. long-memory [1331]. Long-range [66, 1512]. longitudinal [282, 370, 589, 1966, 2005, 1184, 296, 832, 2130, 987, 1106, 916, 716, 1557, 1916, 1454, 1008]. look [1156, 1864, 1041]. look-up [1864]. Looking [280, 366]. looks [421]. loop [2121]. loss [728, 1006, 629, 2206, 1368, 765, 1366, 200, 538]. lossless [1955]. Low [1701, 1137, 1654, 1598, 1933, 1914, 1934, 1265, 471, 1960, 738, 1587]. Low-dimensional [1137, 738]. Low-rank [1701, 1654, 1598, 1934, 1265, 1960, 1587]. low-storage [471].

M [791, 1063, 1061, 761, 1067, 412]. M-estimators [761]. machine [2066, 2222, 1917, 2112]. machines [1510, 45]. Macintosh [41]. MAD [1309]. magnetic [1896]. magnitudes [1499]. main [324, 325]. main-effect [325]. mainshock [1475]. maintaining [1898]. Maire [1071]. Majorization [1040]. making [1465]. MALA [1681]. Mallows [1881, 2166, 2083]. Mallows-type [2083]. management [130, 136]. Manifold [1395, 1923, 1972, 993, 1535, 1779, 958]. Manifold-constrained [1923]. manifold-restricted [1535]. manifolds [1492, 1578, 1677]. Manipulating [625]. many [846, 547]. Map [666, 917, 2155]. mapping [1887, 1770]. maps [1946, 486]. Marchetti [1079]. Marginal [429, 1522, 445, 1843, 1047, 125, 1687, 509, 1472, 233, 1228, 1368, 247, 638, 1740, 667, 2149, 1676, 1724, 1177]. marginalized [1068, 1067]. margins [1476]. mark [884]. marked [2140, 2122]. marking [750]. Markov [1191, 1059, 1071, 1077, 762, 1157, 857, 1287, 813, 1320, 1122, 424, 1102, 1556, 1680, 935, 590, 181, 284, 279, 849, 1606, 2043, 186, 265, 212, 1978, 1072, 429, 1190, 1060, 1032, 336, 832, 2202, 517, 950, 1990, 479, 1159, 1303, 416, 1488, 1408, 1688, 585, 506, 1238, 1458, 1484, 1148, 226, 1375, 387, 1353, 1258, 773, 1362, 1757, 1052, 1558, 1443, 1653, 1432, 1705, 716, 803, 1546, 1734, 965, 2106, 276, 1078, 1572, 1205, 268, 2035, 636, 1711, 990, 1483, 1014, 768, 520, 1089, 1410, 280, 1840, 438, 677, 1725]. Markov-switching [1258]. Markovian [1335]. marks [884]. MARS [285, 547]. martingale [1700, 1680]. mass [1894]. massive [2159, 1888, 627, 1265, 2057, 471, 632, 2001, 1792, 1869]. Massively [283, 241]. matched [512]. Matching [557, 274, 1243, 787, 1883, 1689, 359, 1472, 856, 700, 2017]. material [356]. Matérn [1831]. Mathematical [355, 1144]. matrices [1468, 74, 57, 1009, 1721, 210, 2024, 367, 383, 1361, 2120]. Matrix [2019, 1654, 2005, 1781, 709, 1066, 1102, 1763, 2015, 693, 774, 546, 1902, 2202, 1979, 1846, 2067, 1934, 2185, 1502, 916, 922, 1830, 885, 1065, 810, 1394, 1726]. matrix-variate [2005, 1902, 2185]. Max [1231, 1025, 1650, 518]. Max-min [1231]. max-product [518]. max-stable [1025]. Maximal [1217, 1038, 1037]. Maximin [867]. maximization [1644, 619]. maximizers [1109]. Maximum [1986, 1636, 1412, 1084, 842, 1852, 493, 1836, 2008, 731, 619, 206, 189, 437, 797, 429, 580, 645, 304, 1397, 1029, 1489, 448, 921, 1992, 1173, 510, 428, 268, 202, 1338, 1483, 1619, 1356, 2096, 1797]. Maxwell [1864]. may [1595]. MCD [1913, 712]. MCEN [1520]. McKean [1402, 2121, 2146]. MCMC [1739, 672, 1456, 1333, 1556, 296, 600, 1849, 458, 1733, 425, 1731, 532, 1032, 718, 654, 1239, 983, 1139, 599, 1729, 2088, 717, 1447, 895, 1769, 390, 1098, 1602, 450, 1547, 1542, 633, 1477, 535]. MCMSki [1058]. MDE [1768]. MDL [583]. Mean [1504, 384, 1797, 1221, 817, 1514, 1605, 2156, 1864, 638, 1423, 1152, 2111, 2096]. mean-field [1605]. mean-parameterised [1864]. meaningless [847]. means [1002, 1043, 104, 366]. measure [1281, 1537, 1379, 979, 2150]. measure-preserving [1537]. Measurement [407, 437, 857, 1210, 2116, 1627, 354, 1908]. measurements [682]. measures [1527, 1659, 2095, 2155, 1307, 1758, 1530, 1251, 2088, 1381, 432, 758, 1764, 2016]. Measuring [706, 39, 235]. mechanical [261]. mechanisms [770]. mechanistic [1304]. median [591, 584, 1979, 644, 513, 1504]. mediation [2094, 1948, 2207, 1869]. medical [722]. meeting [1058]. melding [1705]. memory [1331, 1450, 903, 1853, 1363, 2062]. Mengersen [1061]. Mervyn [253]. mesh [1443]. messy [88]. meta [1310, 1181]. meta-model [1310]. meta-models [1181]. metadata [176, 136]. metamodels [1639, 876]. Method [14, 1997, 862, 2126, 1633, 823, 13, 693, 894, 2055, 1967, 892, 696, 159, 1183, 1449, 19, 2108, 277, 2030, 662, 1397, 1931, 1873, 1928, 1880, 1894, 1001, 1546, 2031, 1738, 658, 1864, 2224, 2177, 1520, 1774, 1626, 170, 1091, 1697, 1897, 1771, 1169, 1551, 1724, 1309, 1269, 1602, 151, 1850, 858, 2039, 1464, 568, 710, 1358, 438]. methodology [462, 931, 808]. Methods [238, 6, 452, 337, 735, 1553, 2140, 2122, 953, 1573, 424, 1117, 1863, 1128, 648, 52, 181, 1763, 2075, 1192, 905, 1134, 1821, 882, 679, 524, 529, 1072, 362, 1715, 1652, 654, 671, 830, 642, 416, 1530, 113, 1767, 923, 1251, 551, 2222, 235, 74, 1402, 1349, 226, 347, 2009, 645, 435, 1494, 504, 182, 1777, 442, 749, 1291, 1747, 906, 793, 107, 921, 477, 495, 841, 1071, 1286, 496, 836, 332, 700, 1918, 1385, 1344, 581, 1526, 36, 1086, 1651, 520, 1579, 1983, 1041, 973, 1352]. metric [1848, 1829, 2096]. Metrics [1555, 1501]. Metropolis [1069, 1059, 1359, 1789, 813, 1234, 676, 1606, 615, 1070, 1060, 578, 1687, 923, 287, 959, 1559, 1233, 1675, 1710, 1206, 2210, 1179, 891, 1862]. Metropolis-adjusted [2210]. Metropolis-class [1675]. Metropolis-coupled [813]. Metropolis-within-Gibbs [1559]. Metropolized [187]. micro [1818, 49]. micro-data [1818]. microarray [692]. microdata [498, 492, 493, 502]. MIMCA [1273]. Min [1650, 1231]. Min-max [1650]. Mini [1543, 1592]. Mini-batch [1543, 1592]. minimal [1148, 1012]. Minimax [1134, 629, 1053]. minimising [1797]. minimization [1040, 649, 1423]. Minimum [593, 1444, 1508, 797, 440, 1671, 2004]. mining [656]. Mira [1073]. misclassification [1902, 1338]. Mises [2030, 668, 351]. Missing [770, 837, 2186, 1440, 1479, 1684, 979, 881, 2185, 2084, 2163, 2059, 2214, 1963, 1871, 1587]. mixed [1010, 371, 862, 1310, 1237, 782, 1427, 693, 708, 1509, 1440, 1479, 2080, 1955, 1045, 830, 1435, 641, 1163, 228, 1011, 989, 1476, 1627, 1160, 1241, 2081, 1029, 183, 749, 1810, 2000, 2040, 1649, 831, 1977, 2037, 1031, 828, 1056, 1448, 1454, 1897, 1962, 1836, 2149, 1637, 1970, 1790, 1008, 2093, 2034, 986, 306, 941, 1168, 973, 2109, 438, 899, 1507]. mixed-effects [1010, 1509, 2080, 1241, 831, 1977, 2037, 986]. mixed-integer [1790]. mixed-level [1962]. mixed-mode [183, 306]. mixed-type [1970]. mixing [1859, 1138, 1206]. Mixture [1752, 263, 1020, 183, 1279, 1189, 271, 189, 188, 653, 783, 2005, 1197, 709, 587, 483, 993, 678, 2032, 882, 1002, 655, 167, 532, 1072, 1404, 1380, 505, 550, 983, 723, 1109, 1103, 1518, 1212, 1138, 46, 339, 1599, 775, 1692, 1743, 1807, 773, 1964, 516, 613, 753, 1017, 2217, 661, 825, 816, 1445, 331, 571, 665, 1301, 1543, 1544, 237, 375, 1071, 1943, 980, 623, 1697, 811, 754, 1609, 1451, 2078, 1179, 1424, 731, 971, 2145, 844, 535, 1866]. mixture-of-experts [1866]. Mixtures [1601, 453, 998, 1276, 1942, 798, 893, 1116, 527, 487, 397, 1714, 2022, 932, 595, 1881, 1129, 575, 1223, 1266, 726, 1084, 1251, 1160, 842, 338, 1148, 340, 829, 1029, 992, 1193, 1175, 614, 465, 618, 1529, 242, 2154, 2193, 555, 540, 664, 2049, 1904, 810, 1618, 1014, 1093, 306, 1921, 2213, 1834, 2136]. mixtures-of-experts [2022]. ML [651]. MLE [870]. MLMC [1465]. MMD [1893]. MML [345, 336, 351]. mobility [835]. Modal [1601, 2225]. modalities [1883]. modality [2175, 1896]. Mode [1601, 183, 306]. Model [893, 621, 1216, 950, 1468, 1574, 743, 1212, 2097, 687, 1365, 1226, 1472, 1175, 917, 1956, 501, 350, 2059, 844, 1792, 1300, 1900, 54, 146, 1005, 2198, 653, 783, 1523, 371, 1310, 907, 223, 2144, 1514, 1878, 1345, 652, 372, 1919, 426, 1660, 1760, 1096, 349, 1272, 2041, 1799, 2196, 600, 1461, 240, 29, 1666, 655, 1335, 326, 425, 1638, 157, 563, 647, 1683, 1473, 1267, 484, 492, 1684, 692, 963, 1354, 815, 1685, 944, 674, 804, 939, 548, 1979, 1020, 1518, 414, 1846, 1703, 1466, 1627, 1484, 1399, 1127, 2086, 1148, 1670, 926, 2108, 949]. model [1861, 415, 1645, 1801, 1149, 2006, 1312, 1232, 1400, 2084, 2060, 1502, 1873, 2118, 2206, 2163, 2208, 530, 1894, 2223, 467, 1306, 1721, 1746, 571, 711, 1301, 1706, 946, 1992, 559, 1414, 276, 765, 1544, 1367, 1851, 885, 929, 1342, 2166, 1286, 1616, 1943, 638, 1903, 1774, 883, 180, 1075, 1151, 2024, 208, 623, 1100, 657, 1428, 764, 1973, 1741, 667, 1968, 1469, 2098, 1076, 1110, 2046, 2078, 1439, 1724, 150, 36, 1329, 1328, 1086, 1313, 1008, 512, 786, 1648, 214, 302, 969, 2113, 1991, 960, 1107, 2083, 2213, 535, 1730, 1866, 1840, 1315]. Model-based [893, 1216, 1468, 1574, 1212, 687, 1226, 1175, 917, 1956, 2059, 223, 1683, 492, 963, 1354, 815, 1685, 1979, 1020, 1518, 1846, 1873, 1306, 1367, 1903, 764, 969]. model-checking [2041]. Model-free [1792, 1300]. Modeling [865, 536, 1947, 453, 1184, 1019, 1479, 1109, 1796, 868, 1518, 2130, 643, 1692, 829, 357, 1872, 1906, 613, 753, 825, 1445, 354, 2091, 1350, 980, 1451, 835, 2145, 2178]. Modelling [1354, 456, 146, 485, 1434, 123, 1655, 915, 1516, 516, 1017, 374, 427, 375, 1916, 2102, 1095, 2061, 302, 1586]. Models [458, 452, 1159, 1014, 1453, 1304, 1607, 543, 69, 139, 206, 282, 437, 370, 589, 1276, 1942, 1255, 1898, 419, 1679, 1197, 709, 1010, 2066, 1951, 1320, 1659, 1025, 862, 1310, 1237, 1066, 1120, 1102, 1333, 1506, 782, 1181, 1210, 1046, 648, 2199, 1457, 730, 2125, 1427, 1146, 1081, 483, 1298, 993, 1986, 1322, 851, 879, 1063, 2032, 2015, 693, 1383, 708, 800, 1389, 2107, 1967, 1509, 1756, 296, 1290, 2041, 243, 2196, 679, 186, 1881, 463, 1440, 2080, 1955, 1284, 1978, 878, 900, 1072, 1053, 1641, 1190, 1032, 1404, 1270, 1755, 2202, 2162, 1068, 505, 785, 475]. models [1045, 983, 1409, 1990, 479, 830, 1435, 952, 1079, 641, 723, 1163, 1223, 2065, 228, 620, 631, 1357, 1049, 1011, 674, 416, 603, 1462, 1664, 373, 1139, 1419, 1915, 666, 972, 141, 1183, 989, 1090, 2182, 1593, 1264, 1138, 172, 2067, 1458, 1160, 1241, 976, 466, 1148, 1888, 2180, 948, 1111, 2081, 1934, 761, 1250, 1825, 356, 2188, 1040, 1932, 931, 645, 329, 875, 1599, 775, 1265, 1211, 1097, 1743, 1807, 1475, 1809, 1504, 772, 1785, 945, 1905, 920, 1258, 1232, 773, 1029, 1413, 576, 749, 1810, 2187, 1048, 1106, 1213, 1489, 1502, 1884, 2160, 1964, 2163, 1052, 1327]. models [140, 1778, 1695, 2099, 2217, 2171, 1594, 1418, 717, 1443, 1653, 258, 2000, 194, 395, 1705, 473, 1279, 876, 1087, 2040, 716, 741, 1557, 1649, 1228, 665, 1727, 887, 803, 1001, 1734, 1610, 831, 1368, 936, 1735, 869, 1067, 411, 1632, 1101, 1977, 2037, 2106, 119, 145, 138, 1830, 1543, 1949, 2071, 1363, 547, 856, 464, 1854, 951, 2048, 237, 175, 1985, 1407, 1511, 1031, 828, 1442, 1707, 1581, 1696, 1071, 1864, 2164, 1142, 1064, 531, 902, 918, 1056, 1515, 1104, 1837, 1220, 132, 1908, 1189, 2133, 1916, 1416, 2215, 268, 565, 755, 1448, 2193]. models [143, 2035, 211, 1154, 1065, 1820, 1545, 1550, 991, 737, 1323, 144, 1697, 880, 1973, 1454, 664, 1897, 124, 31, 581, 811, 898, 1229, 754, 1429, 1803, 1609, 626, 1080, 1263, 1762, 2149, 2129, 1676, 554, 1257, 1221, 990, 1231, 1483, 1889, 142, 216, 1424, 1177, 2093, 2034, 786, 1602, 731, 986, 2023, 1974, 214, 520, 1089, 2214, 450, 306, 941, 850, 901, 1850, 1969, 1963, 1927, 633, 1055, 2001, 844, 1168, 1277, 1441, 2207, 1464, 201, 1871, 973, 1813, 2109, 1394, 438, 619, 1315, 2062, 594, 1834, 207, 1191]. moderately [2224, 2177]. Modern [1911, 591, 1500]. Modifications [886].

Modified [584, 177, 1395, 1524, 2049]. modularity [2069, 2138]. modularity-bound [2069]. Modularized [1816, 1695]. modulated [1688]. modulation [741]. modulus [177]. Moment [936, 1818, 1243, 1151, 700]. Moment-based [1818]. moment-matching [1243]. Moments [1815, 379, 560, 1842]. monitoring [1319, 1130, 435, 2195, 390, 2039]. monotone [1427, 726, 946, 1411]. Monotonic [2211, 964]. monotonicity [1766]. Monte [1625, 1597, 1375, 1395, 1873, 1961, 1696, 1071, 1003, 1972, 1422, 762, 1588, 1157, 813, 1122, 424, 406, 1117, 1181, 1633, 1658, 823, 590, 181, 284, 279, 13, 1759, 873, 849, 1015, 1606, 2043, 1002, 186, 265, 892, 1731, 1072, 362, 429, 2124, 1492, 1539, 671, 479, 1292, 1530, 1875, 1408, 607, 1251, 585, 1593, 551, 506, 1238, 1534, 1402, 1589, 172, 1230, 1458, 226, 605, 1097, 1353, 304, 472, 1693, 773, 182, 1362, 1646, 799, 1757, 1880, 1487, 1228, 1546, 921, 965, 1318, 247, 1572, 1205, 909, 834, 1524, 673, 927, 1194, 1086, 1177, 1256, 1170, 520, 1699, 1410, 1552, 1332]. Monte [438, 1725]. Moores [1061]. moral [225]. Moranda [36]. morphology [355]. Morrison [1105]. most [1463, 90, 269]. motif [2156, 1232]. motion [605, 1802]. Moulines [1192, 1059]. mouth [366]. moves [940, 1477]. moving [115, 1022, 1375, 267]. MRCD [1662]. Müller [412]. Multi [1314, 1530, 1402, 1832, 2168, 649, 1586, 485, 1604, 1271, 1268, 243, 871, 2124, 298, 2132, 1858, 1926, 1880, 2208, 1417, 751, 1962, 351, 2003, 1107, 1254, 1464]. multi- [1962]. multi-brand [2208]. multi-class [1417]. multi-criteria [1268]. Multi-dimensional [1314, 1858]. multi-factor [1464]. multi-hit [243]. Multi-index [1402, 1832, 2124, 1880]. multi-label [1254]. multi-layer [485]. Multi-level [1530]. multi-objective [1271, 298]. multi-rater [751]. multi-reference [1604]. multi-resolution [871]. Multi-scale [1586]. multi-site [2132]. multi-source [1926]. multi-state [351]. Multi-step [649, 2003]. multi-stratum [1271]. multi-target [1107]. Multi-task [2168]. multidimensional [2189, 2143, 1132, 1811, 737, 1973]. multifacet [2208]. multifactor [1669]. multilayer [2184]. Multilevel [2196, 1402, 1347, 1467, 1575, 995, 2146, 1722, 2140, 2122, 1589, 2060, 1471]. Multilocus [368]. multimodal [218]. multimodality [39]. Multinomial [797, 762, 1255, 1412, 1270, 125, 479, 962, 276, 1851, 1520, 2219]. Multinomial/Dirichlet [797]. Multiobjective [1153]. multiparameter [769]. Multiple [1406, 2013, 1318, 929, 1486, 1249, 1877, 888, 1273, 1164, 928, 1096, 812, 1606, 1886, 586, 785, 1292, 1090, 1534, 814, 1007, 1375, 931, 1112, 258, 1917, 1274, 1557, 632, 1208, 744, 951, 689, 1943, 2224, 2177, 1956, 883, 2215, 1709, 1550, 2026, 1968, 581, 626, 2045, 1309, 216, 1994, 1862, 615]. multiple-arm [689]. Multiple-index [929]. Multiple-output [2013, 1550]. Multiple-population [1249]. multiple-try [1606, 581, 1862, 615]. multiplicative [463, 2085]. Multiplier [1291, 769]. Multiscale [821, 1326, 1609, 1700, 2176, 354]. Multivariate [797, 575, 562, 752, 1949, 828, 456, 2164, 811, 1311, 783, 1966, 1877, 798, 893, 56, 2165, 114, 964, 840, 1216, 1874, 1130, 564, 1455, 1147, 712, 1133, 708, 1538, 1752, 1083, 524, 1859, 157, 1536, 1045, 950, 1047, 286, 598, 1952, 726, 1833, 1590, 1980, 2009, 363, 1882, 769, 232, 962, 992, 1885, 516, 753, 1017, 825, 533, 1445, 1087, 614, 1734, 1929, 2106, 1984, 237, 1339, 1529, 510, 1741, 1771, 230, 801, 1135, 1637, 523, 1970, 1378, 1887, 241, 809, 1093, 302, 1614, 1994, 535, 2136]. Muralidhar [496]. Murray [253]. mutated [2080]. mutual [2116, 2079, 1591].

naïve [1838]. Natural [2109, 1357, 1466, 79, 104]. navigation [1924, 2170, 2200, 2227]. Nearest [1837, 802, 1879, 170, 376, 1224, 1732]. nearly [2044]. nearly-isotonic [2044]. need [87]. needles [1772]. Negative [1034, 2092, 43, 1563]. negativity [944, 695]. Neighbor [1912, 1837, 1879, 376, 1224]. Neighborhood [763, 1732]. neighbour [170]. neighbourhood [1287, 1757, 637]. neighbours [802]. Nelder [139, 141, 143, 144, 142]. Nested [1408, 1398, 819, 1760, 1470, 1808, 1163, 1474, 1735, 1342, 1180]. net [1855, 804, 1203, 833, 2003]. network [1505, 2221, 1959, 2013, 1571, 2167, 2014, 1400, 1749, 855, 1528, 2077, 1930, 1567, 1515, 759, 1826]. network-based [2167]. Networks [1005, 485, 1835, 1936, 1841, 1941, 1345, 1074, 595, 1406, 1461, 1566, 1073, 1589, 2094, 277, 18, 348, 1905, 225, 2060, 1489, 1896, 2118, 539, 1261, 1772, 577, 624, 2077, 1930, 1417, 2169, 912, 1956, 1485, 2190, 2026, 637, 67, 1712, 1565, 1764, 1901, 2039, 2184, 2137]. Neumann [1285]. Neural [1005, 1901, 485, 1936, 1941, 2013, 2167, 1905, 2118, 539, 1749, 1772, 2077, 1930, 2026, 67, 1712, 1786, 1565]. neuronal [1874]. Newton [189, 1697, 788]. next [86, 89, 90, 92]. Neyman [984]. NGG [1197]. NGG-mixture [1197]. Nicholas [307]. no [1670]. Node [521, 758, 1724]. node-wise [1724]. nodes [1023]. noise [938, 2151, 541, 354, 1891, 1469]. noiseless [1038]. noising [1821, 1491]. Noisy [1157, 289, 1639, 1233, 1359]. nomination [1835]. Non [371, 1234, 730, 777, 1599, 847, 204, 107, 1338, 447, 574, 867, 1164, 1335, 706, 2162, 1848, 492, 123, 723, 1223, 1159, 944, 1783, 1575, 1920, 687, 288, 1475, 248, 1029, 122, 816, 1746, 803, 1675, 884, 1977, 2037, 567, 2168, 2031, 1827, 974, 1911, 1350, 808, 1448, 35, 1098, 1369, 1662, 1454, 2026, 2049, 202, 581, 811, 1578, 1762, 1889, 695, 43, 721, 274, 151, 1856, 193, 1563, 1542, 858, 568, 651]. Non-binary [204]. non-centering [567]. non-central [151]. non-convex [1848, 1920, 1675, 2168, 2031]. non-crossing [1164, 2026]. non-elliptical [1662]. non-elliptically [687]. non-EM [651]. non-equispaced [574]. non-exchangeable [1783]. non-Gaussian [123, 723, 248, 581, 1578, 721]. non-homogeneous [803]. non-iterative [858]. Non-linear [730, 706, 1575, 122, 1977, 2037, 1911, 1098, 202, 193]. Non-local [1599]. non-negative [43, 1563]. non-negativity [944, 695]. non-normal [1223, 288, 1454, 811, 1762]. non-normality [35]. Non-parametric [371, 847, 1338, 447, 816, 1350, 808, 2049, 1889]. non-Poissonian [1475]. non-random [568]. non-randomly [1856]. Non-reversible [1234, 1542]. non-simplified [1369]. non-sparse [2162]. non-spatial [492]. Non-standard [107, 1448]. non-stationary [1335, 1746, 884, 1827, 274]. nonasymptotic [1644]. noncausal [2051]. noncentral [669, 423]. nonconcave [1127]. nonconjugate [13]. Nonconvex [2104, 1588, 1081, 1983]. Noncrossing [1550]. noncyclical [1387]. nondecimated [681]. nonignorable [2163]. Nonlinear [1509, 1507, 1186, 1010, 488, 851, 693, 2107, 2155, 2080, 1053, 1755, 954, 598, 551, 848, 2188, 539, 1772, 1099, 831, 946, 581, 1498, 1651, 958, 1192]. Nonlinearity [405]. Nonnegative [1659]. Nonparametric [1527, 954, 1108, 703, 970, 413, 920, 1381, 571, 573, 1987, 1515, 859, 1093, 1356, 419, 1679, 1659, 611, 608, 602, 1967, 934, 824, 524, 1335, 1859, 217, 1426, 388, 514, 1183, 1321, 1241, 926, 915, 2118, 2139, 1947, 1558, 1033, 841, 765, 582, 828, 980, 971, 1411, 482, 1642, 1433]. Nonparametrics [1242]. nonseparable [1580]. nonstationary [1867, 1718, 931, 698, 427, 1614]. norm [1281, 1308]. norm-based [1308]. Normal [2165, 458, 188, 2076, 483, 1455, 1598, 1773, 952, 1223, 525, 2175, 842, 46, 288, 2185, 1017, 661, 17, 1445, 2040, 465, 1366, 147, 2193, 1754, 1454, 811, 1762, 1135, 43, 810, 1834, 2136]. normal-based [1366]. normalised [1267]. Normalising [459, 1091]. normality [1425, 2167, 662, 35, 809]. normalization [1722]. normalized [1251]. normalizing [1636, 1347]. normals [575]. note [1114, 1242, 1021, 1379, 667, 59, 695]. novel [1005, 1479, 931, 1774, 2145, 710, 2030]. Novelty [1583, 1638, 1682]. NPMLE [1209]. Nudging [1521]. nugget [868]. nuisance [2157, 417, 418]. null [558, 1007]. number [487, 1865, 1298, 297, 66, 575, 167, 1668, 505, 815, 1124, 1007, 504, 1218, 1653, 614, 465, 618, 1529, 1398, 555, 1815, 1324]. number-of-bins [297]. numbers [1582, 1322, 177, 461, 720]. Numerical [94, 525, 234, 489, 428, 1553, 1355, 953, 2126, 1307, 10, 1656, 2092, 1428, 151, 241, 1651, 1985]. numerically [1449, 1705]. numerics [1490, 1500]. NuZZ [1985].

obfuscation [494]. object [601, 359]. Objective [1382, 1271, 635, 298]. objects [1375, 625, 275]. oblique [507]. observational [2025]. observations [1225, 2155, 1264, 1525, 881, 716, 1398, 1895, 201]. observed [1700, 2140, 2122, 857, 94, 1457, 652, 1537, 1687, 978, 2002, 1802, 1442, 1089]. obstructive [1631]. obtain [241]. occurrence [917]. occurrences [2098]. ODA [229]. ODE [1596, 1806, 1619]. Off [1965, 2167, 1323, 260]. off-line [1323]. Off-policy [1965]. off-training-set [260]. Oja [1845]. Olsson [1071]. Omega [1717]. omission [94]. On-line [861]. One [1919, 1604, 243, 388, 1670, 1219, 1744, 476, 1329, 1328, 572]. one-factor [1219]. one-hit [243]. one-shot [1744]. One-step [1919]. one-step-late [476]. one-way [572]. ongoing [381]. Online [1022, 2158, 2221, 1959, 2128, 2191, 1130, 1571, 580, 1672, 1413, 1400, 1845, 2208, 1860, 1987, 623, 2023, 2039]. online-surveillance [1130]. onset [159]. OpenCL [1325]. operator [1491]. Optimal [1281, 1615, 2123, 1351, 634, 1156, 2080, 1507, 1641, 1765, 1492, 1534, 1703, 606, 1669, 18, 2160, 2101, 974, 1416, 747, 1710, 2172, 1897, 1562, 1563, 1623, 1144, 1661, 813, 2157, 1841, 1268, 1134, 1733, 326, 1053, 1372, 298, 1294, 557, 1462, 2103, 2087, 985, 2135, 2081, 329, 2020, 435, 1274, 1366, 1231, 1790, 1793, 1196, 1895, 229, 1171, 260, 807, 1464, 1561, 1430]. optimality [1845, 1131]. Optimally [1718]. optimisation [1625, 1371, 1299, 1647, 1895]. optimised [1608]. Optimization [1912, 2051, 1588, 591, 111, 993, 1134, 1731, 2203, 1004, 1644, 1848, 1250, 489, 1920, 364, 799, 10, 921, 1999, 1153, 673, 170, 1811, 1037, 1856, 2110, 858, 1828, 788]. optimized [1355]. optimizer [2080]. Optimizing [966, 635]. optimum [1162]. Option [426]. Order [403, 278, 1144, 560, 1330, 938, 1133, 1566, 2118, 2056, 1343, 385, 1887]. order-free [1887]. ordered [1982]. ordinal [2005, 1216, 1389, 1260, 1473, 2130, 885, 902, 1189, 1613]. ordinary [1905, 1496, 467, 1498, 1141, 1742]. ordinate [1552]. Orthant [1438, 1167, 1214]. Orthodox [749]. Orthogonal [993, 2123, 324, 266, 1926, 1167, 1438, 646]. orthogonalization [839]. oscillatory [455, 468]. other [202]. outcome [2186, 596, 1557, 2047, 1948]. outcomes [1557, 1220, 1008, 1826]. Outlier [1012, 1662, 712, 1094, 1931, 12]. outliers [56, 2148]. Outlyingness [1463]. Output [663, 1978, 2013, 466, 1550, 1901]. Output-sensitive [663]. outputs [1308]. over- [1584]. over-dispersion [2071]. Over-relaxation [424]. overdamped [1868]. overdispersion [206, 1255, 1034]. overlapped [1730]. overlapping [345, 1689]. overview [853]. Owen [1077].

P [791, 766, 1075, 1302, 2189]. P-splines [766, 1302, 2189]. package [380, 342, 121]. packet [1610, 427]. pair [1054]. pair-copula [1054]. paired [1000, 2133]. pairs [1128, 1442, 512]. Pairwise [708, 515, 1442, 1189]. panel [2021, 2202, 1001, 2106, 990, 1970, 1259]. paper [318, 253, 313, 310, 319, 309, 320, 311, 316, 312, 315, 412, 317, 314]. papers [33, 252]. Parabolic [684]. paradox [1595]. Parallel [1588, 940, 1299, 1265, 801, 520, 178, 955, 283, 181, 1606, 2043, 600, 1325, 1218, 1663, 122, 1806, 1637, 1086, 241, 2104, 1667, 1422]. parallelism [431]. parallelizable [1410]. parallelization [217]. Parallelized [1808]. Parallelizing [1729, 633]. Parameter [997, 1310, 994, 415, 1453, 1997, 590, 993, 426, 393, 879, 861, 219, 1409, 1517, 1084, 1138, 1474, 347, 978, 2002, 1325, 616, 1784, 1418, 942, 1802, 1279, 936, 946, 1696, 1393, 918, 510, 1132, 1771, 1229, 1907, 2145, 1822, 1730]. parameterisations [1333]. parameterised [1864]. parameterization [899]. Parameters [457, 417, 832, 125, 1831, 1244, 711, 638, 722, 1448, 1050, 211, 1154, 1098, 1923, 1141, 418, 2184]. Parametric [964, 972, 454, 1240, 1532, 371, 760, 896, 157, 1260, 777, 1159, 1124, 2081, 847, 662, 816, 445, 531, 1940, 1350, 180, 808, 751, 2049, 1110, 990, 1338, 1889, 36, 447, 1411, 2001, 1315, 1834]. parametrization [1763]. parametrizations [210]. parametrized [1935, 1850]. pareto [1268, 1962]. pareto-based [1268]. Pareto-efficient [1962]. Parsimonious [1506, 2038, 2032, 665, 1904, 1046, 1146, 641, 963, 1685]. Partial [1912, 462, 1871, 2105, 2063, 155, 1702, 2074, 2029, 1487, 411, 909, 1651, 1929]. Partially [328, 1595, 2140, 2122, 857, 1457, 2125, 1537, 1129, 631, 1687, 1111, 978, 2002, 1106, 1884, 1802, 956, 2129, 1221, 1277, 1871, 1315]. partially-latent [1129]. Partially-latinized [328]. Particle [1069, 1070, 505, 2180, 645, 1630, 1521, 1181, 966, 1134, 1668, 1239, 954, 1664, 1139, 1347, 1471, 1467, 1575, 1620, 1222, 1860, 1546, 974, 2035, 1373, 1385, 1169, 2210, 1086, 1725]. Particle-based [1630]. particles [1865, 1668]. particular [289]. partition [1407]. Partitioned [202, 939]. Partitioning [1087, 2038, 1729, 606]. partitions [272, 2038, 1606, 1796, 203]. partly [1932, 275]. pass [471]. passage [1842]. path [2006, 836, 280, 1036]. pattern [1184, 299, 1555, 367, 835, 1871]. patterns [337, 45, 1012, 719, 1317, 568]. PC [1938, 8]. PC-90 [8]. PCA [1890, 1845, 2224, 2177, 1235]. PCA-uCPD [2224, 2177]. PDAEs [2161]. PDMP [1791]. pedigrees [59]. Penalised [1277, 1448]. penalize [1145]. Penalized [1666, 1431, 2226, 2171, 1903, 705, 1324, 714, 1453, 574, 1081, 896, 1420, 1401, 1435, 989, 1127, 1040, 781, 1054, 1009, 1784, 1749, 2204, 1295, 1960, 1369, 1897, 1836, 1637, 1259, 1856, 858, 1036]. penalties [2168, 1302, 1983]. penalty [766, 1031, 1750, 705, 1577]. percentage [52]. percentile [79]. Percentiles [1006, 151]. perceptron [485]. perceptrons [261]. percolation [603]. Perfect [444, 464, 480, 729, 338]. perfectly [1689]. perfectly-overlapping [1689]. Performance [17, 1797, 2073, 532, 1668, 234, 363, 1774, 518]. periodic [1506, 2144, 1567, 2178]. permanents [1170]. Permutation [1045, 166, 771, 964, 1083, 701, 1670, 1200, 367, 394]. Permutational [57]. permutations [1727]. permuting [1000]. perpetuities [1952]. persistence [1761]. perspective [1072, 742, 1498]. Perturbation [499, 2159, 156, 495, 496]. perturbations [1751]. Perturbed [1289, 74]. pharmacokinetic [329]. pharmacovigilance [1721]. phase [1271, 1772, 1776, 990, 1793]. phase-type [1776, 990]. phylogenetic [1863]. Phylourny [1863]. physical [461]. pieces [1940]. piecewise [1633, 1357, 1916, 1089]. Pitman [1714]. Pitman-Yor [1714]. Pivotal [1394, 1404]. placement [673]. plans [324]. plant [603]. plot [56]. plots [763, 215, 280]. PLS [1288]. Plummer [1063]. Point [1686, 1957, 1256, 558, 2140, 2122, 1877, 1584, 888, 938, 515, 52, 111, 1172, 1566, 1683, 299, 1959, 1090, 1108, 79, 1844, 968, 1894, 884, 984, 1318, 1555, 185, 1987, 2224, 2177, 1940, 883, 870, 1323, 637, 1622]. points [52, 933, 800, 1406, 2033, 1920, 855, 200, 148, 207]. Pointwise [304, 1740]. Poisson [1067, 146, 1237, 1120, 444, 1383, 708, 2221, 53, 1068, 1688, 1449, 562, 643, 1692, 2057, 1381, 614, 1864, 531, 555, 1650, 2197, 351, 973, 1813]. Poissonian [1475, 1100]. polar [1944]. Polettini [497]. policy [1965]. pollution [565]. PolSAR [1910]. Polya [1732, 1245, 1246, 962, 1949]. polychotomous [1260]. polygenic [1631]. Polynomial [452, 533, 1925, 1326, 534, 1464]. polynomials [1427, 794, 1909]. polytope [2135]. population [1303, 628, 1097, 2171, 1249]. population-based [628]. populations [1773, 603, 58, 2060]. portable [122]. portmanteau [2000]. Positive [2, 444, 480, 2147]. positive-unlabeled [2147]. positivity [1499]. positron [358]. possibly [1877]. post [1573, 2153, 2182]. post-selection [2153, 2182]. Posterior [54, 1098, 146, 188, 251, 558, 1470, 1284, 952, 1030, 228, 113, 625, 472, 182, 883, 1740, 1698, 1089]. posteriori [429, 921, 1619]. posteriors [882, 1758, 1578, 1725]. potentials [106]. Potts [711]. Power [2174, 1940, 564, 1660, 2153, 1083, 556, 177, 1030, 1981, 191, 1101, 646, 1154, 391, 394]. Power-law [1940]. power-scaling [1981]. powers [1781]. Practical [587, 1800, 1151, 1328, 1162, 842, 1329]. practice [227, 1780, 486, 1541]. Pre [1061, 1062, 1222, 164]. Pre-processing [1061, 1062, 164]. pre-smoothed [1222]. precipitation [1096]. precision [1830, 1361]. preconditioning [1362]. predetermined [1764]. predicting [1917]. Prediction [1024, 407, 838, 1794, 1763, 736, 1158, 779, 363, 1392, 2128, 1327, 1749, 737, 664, 376, 1541]. Prediction-based [838]. predictions [1398]. predictive [426, 919, 263, 167, 1049, 606, 1825, 348, 1213, 1368, 1286, 554, 897, 207]. predictor [742, 1594]. predictors [1081, 1666, 1473, 745, 2036, 1087, 547, 1224]. prefer [2007]. prescribed [693, 461]. Presence [407, 188, 2151, 2116, 1050]. Preserving [500, 1537, 494, 1955, 179, 1478, 1503]. Presto [121]. pricing [426]. primary [443]. Principal [698, 44, 148, 2095, 488, 1314, 704, 566, 1950, 1779, 2226, 781, 195, 2028, 1341, 1099, 158, 196, 1390, 2042, 1186, 200, 930, 193, 2192, 1798]. principle [1875]. Principles [919]. Prior [1287, 1443, 2199, 1638, 1682, 1270, 2058, 952, 2067, 1981, 846, 1769, 1174]. prior-data [2199]. priori [1197]. Priors [1122, 346, 1426, 1251, 1484, 1558, 1549, 1154, 1711, 2142, 1994]. Privacy [1955, 1573]. Privacy-preserving [1955]. private [1739, 1477]. Probabilistic [2161, 257, 1495, 32, 1820, 754, 1498, 1553, 1527, 1497, 21, 29, 20, 1652, 1490, 18, 1496, 269, 1500, 1800, 1428, 350, 1578, 1945]. probabilities [1863, 115, 1455, 1598, 1015, 596, 525, 1334, 962, 1167, 1438, 2092, 366, 1214, 1918, 1135, 151, 241, 694, 874]. Probability [146, 11, 1281, 2076, 2165, 872, 935, 595, 1307, 1536, 2174, 1484, 629, 442, 1123, 194, 382, 59, 1269, 2110]. probable [269]. probit [846, 276, 1851, 512, 302, 1439]. problem [1144, 1661, 591, 72, 249, 266, 644, 1575, 754, 507]. problematic [532]. problems [888, 2063, 1117, 1210, 1791, 663, 1074, 295, 961, 84, 1624, 733, 635, 586, 1073, 692, 1617, 1704, 1872, 1663, 609, 1873, 744, 2133, 673, 883, 1428, 31, 12, 1907, 1145]. procedure [1130, 2041, 900, 1683, 245, 1363, 2215, 367, 1206, 901, 2003]. Procedures [454, 126, 492, 814]. Process [1847, 456, 1333, 1146, 1462, 1686, 1621, 1466, 1138, 1844, 1160, 1980, 2009, 2072, 1839, 1692, 1957, 1672, 182, 1381, 1639, 571, 1769, 884, 984, 1827, 1842, 1806, 1153, 531, 1549, 2193, 540, 664, 1971, 1526, 1923, 1236, 1256, 302, 1586, 1813, 2178, 1642]. process-based [1333, 1256]. processed [1165]. processes [2198, 2140, 2122, 857, 1584, 2063, 938, 2211, 1867, 2141, 621, 115, 515, 1874, 1537, 994, 1538, 1629, 264, 1752, 2221, 1566, 2051, 433, 1316, 1959, 1024, 1303, 1686, 1488, 1687, 1688, 1452, 1787, 1767, 1108, 1600, 1718, 978, 1920, 1957, 1353, 1777, 1432, 2152, 1746, 1001, 1318, 750, 1078, 1891, 719, 1800, 1077, 1754, 267, 1771, 637, 1622, 121, 1311, 1562, 695, 1678]. processing [178, 1061, 176, 1062, 354, 164]. processor [181]. Procrustes [266, 299, 507]. Product [1693, 2189, 546, 1407, 518, 941, 1302]. Product-form [1693]. products [1165, 752]. Profile [466, 949, 1634, 410]. profiles [1191, 1190, 1974]. programming [1144, 1053, 1372, 105, 685, 104, 122, 1173, 1800, 1790, 1041, 1507]. programs [171]. progressive [2195]. projected [1949, 1870, 434]. Projection [488, 1953, 715, 95, 1652, 1590, 1444, 756, 981]. projection-based [715]. projections [1535, 784, 1248, 2029, 1167, 1438, 747, 1118]. prone [1821, 2107]. propagation [865, 20, 1386, 2212]. proper [112]. Properties [1592, 1547, 1162, 527, 854, 1777, 2046]. proportion [2148]. proportional [1666, 1861, 1812, 201]. proportions [982, 1204, 397, 339, 2101]. proposal [1681, 928, 718, 1894, 700]. proposals [1681, 676, 1572, 1179]. protect [233]. protection [498, 492, 156, 1289, 493]. Proximal [1760, 1205, 1435, 1848, 2134]. proximal-gradient [1435]. PRticle [1859]. Pseudo [212, 1724, 65, 2186, 1278, 1687, 1228, 1676]. Pseudo-likelihood [212, 1676]. Pseudo-marginal [1724, 1687, 1228]. pseudo-outcome [2186]. pseudo-random [65]. pseudo-sample [1278]. pseudodistance [1773]. pseudorandom [177]. Pudlo [1075]. pulmonary [1631]. pursuit [787, 1953, 95, 1444, 756, 856, 274]. PXEM [476].

quadratic [1861, 1870, 870, 43].

quadrature [1355, 1446, 1499, 288, 2149]. quadrilateral [682]. quality [1262]. quantal [243]. quantification [1553, 1696, 2212]. Quantifying [1660, 1458]. Quantile [832, 2018, 454, 1734, 1838, 1748, 2016, 1276, 1942, 699, 1164, 1998, 728, 707, 2082, 1967, 2179, 1540, 2021, 1011, 2013, 948, 1111, 2068, 2036, 2072, 987, 471, 908, 1295, 2195, 1939, 1039, 1935, 2169, 1545, 1550, 2026, 833, 1110, 1221, 1152, 2183]. quantile-based [2072]. Quantile-distribution [1838]. quantile-parametrized [1935]. quantiles [1936, 1082, 2128, 632, 538]. Quantitative [279, 1060, 1059, 358]. quantization [1531, 2150]. Quasi [1322, 3, 533, 1992, 247, 1897, 2178, 788]. quasi-likelihood [1897]. quasi-maximum [1992]. quasi-Newton [788]. quasi-periodic [2178]. Quasi-random [1322, 3, 533]. quaternary [1173]. quaternary-code [1173]. Query [135, 1835, 176]. queueing [1417]. queues [526]. quick [1683, 1292]. QuickMMCTest [1292]. quotient [43].

R [1192, 1071, 1694, 2093, 2034, 690]. R-VGA [1694]. R-VGAL [2093, 2034]. Radar [357]. radial [1163]. radially [193]. radius [2090]. radius-adaptive [2090]. Randal [1192]. Random [1553, 370, 1933, 149, 996, 1248, 2159, 356, 73, 133, 1031, 14, 1201, 1576, 1580, 65, 1287, 1320, 1659, 1021, 1823, 1665, 2126, 1128, 1215, 444, 1322, 1513, 655, 66, 326, 706, 75, 3, 2203, 1266, 1303, 1283, 1251, 506, 979, 1006, 1018, 1689, 625, 1085, 504, 576, 749, 2060, 1757, 1585, 2029, 1558, 445, 533, 1113, 1228, 461, 1656, 828, 1761, 658, 1774, 1415, 720, 1075, 1591, 1710, 1119, 1744, 1206, 1229, 2059, 1076, 1762, 1711, 2142, 1815, 216, 169, 1579, 568, 1332, 619, 1587]. random-effect [576]. random-scan [1206]. randomization [1573, 126]. Randomized [1489, 1971, 1972, 1817, 2203, 1880]. randomly [1856, 2083]. randomness [168, 394, 2007]. range [797, 1512, 66, 1855, 1371, 4]. range-restricted [1371]. Rank [1604, 1272, 739, 1441, 1654, 854, 1598, 1185, 1275, 961, 1701, 1767, 2159, 1934, 1265, 1926, 2204, 1917, 1960, 1526, 1587, 941, 1254]. Rank-based [1272, 739, 961]. Rank-one [1604]. ranking [1541]. rankings [1487]. Rao [1907]. rare [935, 873, 1353, 1396, 2146, 874]. Rate [857, 1733, 1595, 2066, 182, 448, 1112, 1892, 1348, 1744, 891, 1862]. Rate-optimal [1733]. rater [751]. Rates [458, 1608, 265, 1596, 387, 1496]. Rathindra [496]. Ratio [14, 558, 982, 1204, 2018, 820, 1690, 661, 445, 1136, 1982, 658, 2023]. Ratio-of-Uniforms [14, 658]. RBF [1005]. RDELA [968]. reaction [1841, 1589]. Real [1728, 88, 2183, 590, 1945]. Real-time [2183]. real-valued [1945]. recapture [2171]. recognition [704, 359]. recognize [45]. Recognizing [1035]. reconciliation [1563, 1945]. reconstruction [1701, 244, 154]. record [498, 1379]. recovery [1517, 1449, 2046]. rectangle [1855]. rectangle-range [1855]. Rectangular [1998, 525, 962]. recursions [526]. Recursive [1727, 2165, 2069, 1133, 1694, 1853, 1195]. recycling [447, 1410]. Reduced [1254, 1185, 2054, 1767, 1926, 1917, 1852, 1526]. Reduced-rank [1254, 1185, 1767, 1926, 1917, 1526]. reducing [1898]. reduction [1789, 1556, 1680, 1925, 975, 263, 2100, 1848, 1737, 1517, 959, 1504, 1218, 1975, 1960, 1153, 764, 1533, 1346, 2110, 795]. Reed [464]. reference [1604, 289]. refinement [1733]. regime [1914, 2012]. regime-switching [2012]. regimes [2020]. region [1697]. regions [1096, 1995, 5, 870]. Regression [282, 1847, 688, 6, 1399, 2047, 501, 956, 735, 1034, 54, 1276, 1942, 419, 574, 334, 760, 1164, 1237, 114, 190, 2073, 2199, 730, 608, 1022, 1713, 602, 1081, 1298, 1998, 1383, 707, 800, 1185, 2082, 2010, 2179, 2225, 2158, 1925, 934, 824, 871, 1953, 2021, 524, 1129, 1260, 1473, 742, 1902, 832, 2018, 1990, 479, 952, 1223, 2065, 1248, 766, 1446, 1517, 414, 1702, 1449, 1082, 2074, 1631, 743, 2013, 1466, 1627, 2159, 2148, 1155, 2176, 838, 1716, 948, 1111, 2068, 949, 1861, 2132, 1934, 2188, 2036, 562, 1926, 1211, 413, 1719, 1672, 846, 995, 1478, 534, 987, 617, 2085, 2057]. regression [1928, 2029, 2099, 467, 1954, 1594, 2204, 1087, 1136, 158, 196, 1295, 1949, 765, 582, 1854, 1511, 1827, 1039, 1864, 646, 738, 1616, 2169, 1220, 1911, 1520, 2193, 273, 1723, 1545, 1550, 55, 737, 1748, 540, 664, 2026, 833, 522, 925, 811, 1526, 1803, 837, 1836, 1110, 1236, 2078, 1387, 259, 1889, 2173, 705, 1613, 1970, 1152, 1336, 1602, 1249, 1324, 1346, 1541, 2183, 943, 1850, 2003, 858, 2219, 958, 482, 1464, 1937, 2083, 2019, 1300, 1732, 1834, 1577]. regression-based [1478]. Regression-incorporating [1034]. regression-SARIMA [1594]. Regressions [453, 463, 1266, 1264, 2049, 998, 795]. regressors [956]. regular [1351, 433]. Regularisation [735, 742, 1336]. regularised [1992]. regularization [219, 1290, 2138, 2089, 1855, 838, 737, 1041]. Regularized [1628, 688, 1538, 1386, 2199, 1508, 38, 2182, 1934, 2209, 1905, 863, 2168, 1726]. Regularizing [1611]. regulatory [912]. reinforcement [1965]. Rejection [1912, 1452, 1201, 933, 1285, 2033, 187, 818, 909, 1626, 1297]. Rejoinder [145, 497]. Relabelling [1404, 754]. related [2198, 663, 504, 208]. relatedness [2222]. relationship [1038]. relative [1528, 946, 1423, 572]. relaxation [424, 1928, 1464]. relaxed [1756, 733]. release [492]. reliability [1623, 36]. Reliable [1391, 1997]. relying [1216]. REML [886]. Remote [501, 502, 1289]. Renewable [1967]. renewal [1024, 1804, 1622]. Rényi [1773, 1689]. repairable [1024]. Reparameterization [268]. repartitioning [1470]. repeated [584, 455, 468]. repeated-loess [455, 468]. replacement [1385, 1871]. Reply [254]. Representation [1952, 359, 1148, 1690, 132, 2127]. representations [1680, 2115, 1119]. Representative [1774, 1615]. reproducible [1794]. reproducing [1119]. requires [1670]. Resample [1480, 1288]. Resample-smoothing [1480]. resample-stable [1288]. Resampling [404, 2025, 295, 3, 1581]. Resampling-based [2025]. research [81, 82, 83, 87, 89, 90, 92, 1794]. residue [2156]. Resistant [393]. Resolution [1038, 871, 357, 443, 1173]. resonance [1896]. respect [200]. Response [1004, 1184, 2107, 1129, 1473, 1902, 1052, 1746, 870, 1973, 1429, 1762, 391, 2078, 1889, 1994]. response-type [1994]. responses [282, 370, 2186, 2130, 2163, 1917, 1963]. Restricted [782, 1316, 1136, 1184, 1535, 1371, 2152, 1549]. restrictive [1799]. results [257, 74, 992, 1121, 28, 241, 260]. retailing [580]. retraction [21]. retrieval [1835]. Retrospective [1094, 1500]. return [1763]. returns [1990]. Reverse [912, 1353]. reverse-time [1353]. Reversed [2035]. Reversible [830, 1234, 718, 2108, 717, 1179, 1542, 535]. Review [791, 1192, 690, 342, 790, 229, 2064, 178, 292, 22, 110, 152, 70, 654, 220, 98, 2208, 116, 97, 127]. reviews [61, 108, 197, 128, 60, 221]. revised [154]. revisited [589, 1759, 919, 44, 434]. reweighting [1374]. ridge [1517, 1750, 1637, 2083, 2064]. ridge-penalized [1637]. ridges [1890]. Riemann [1395, 390]. Riemannian [1697, 1972]. right [1164, 1048, 242, 2083]. risk [498, 1631, 1861, 1749, 1528, 1151, 732]. risks [1943, 2025]. rival [1230]. Robert [1061]. Robust [1607, 1532, 1523, 2125, 1130, 1773, 1275, 631, 1266, 1785, 441, 613, 753, 2204, 916, 1445, 1099, 1899, 683, 375, 2169, 2102, 2017, 67, 2045, 891, 969, 2192, 1430, 1679, 2143, 1022, 1583, 1998, 2082, 2158, 240, 463, 1638, 1682, 1374, 1043, 1950, 1634, 1047, 815, 1979, 2074, 1589, 2148, 2020, 1777, 1778, 1289, 2168, 1342, 1235, 1360, 2053, 1963, 1822]. Robustness [219, 854, 1470, 733, 330, 1174]. ROC [2072, 970, 808]. role [205, 1354, 610]. root [1001]. Rotation [542, 1926]. rotations [1611]. Roughness [2089]. rounding [1947]. row [332]. row-column [332]. Roy [780]. Rule [1713, 768, 1198, 179, 1037]. Rule-based [1713]. rules [205, 656, 1828]. run [727]. run-length [727]. RVGA [1853].

S [1192, 1067, 1065]. saddle [1920]. saddlepoint [295, 442, 836]. SAEM [616, 1029, 1977, 2037, 722]. safe [499]. same [1886]. sample [1595, 560, 1615, 1893, 1278, 1933, 1886, 388, 701, 820, 1672, 1200, 836, 1709, 1710, 1541, 1171, 1547, 2111, 2096]. sample-path [836]. sampled [1752, 863]. Sampler [458, 1197, 1624, 149, 1068, 1664, 113, 863, 1559, 1675, 1067, 618, 208, 300, 12, 1188, 1547, 1542, 1699, 1862]. samplers [1608, 1789, 1791, 1230, 675, 1885, 1432, 1907, 280, 1297]. samples [146, 1125, 820, 1241, 365]. Sampling [243, 1535, 228, 1138, 592, 1585, 218, 134, 1442, 1578, 543, 953, 1823, 1841, 1665, 1210, 1457, 935, 587, 819, 2216, 1760, 1714, 678, 361, 1470, 1015, 296, 1272, 1606, 634, 1780, 679, 615, 1886, 556, 826, 1536, 1072, 362, 1868, 1755, 1380, 983, 1252, 723, 1603, 1875, 1488, 1452, 1729, 1589, 1425, 192, 1474, 1493, 2033, 775, 2030, 1657, 1395, 1592, 442, 187, 1026, 2223, 1720, 2050, 1447, 473, 818, 1280, 392, 547, 133, 1753, 1612, 1071, 917, 1812, 2121, 2146, 1805, 366, 1774, 1626, 1560, 1944, 927, 1385, 581, 2210, 732, 1439, 368, 801, 1179, 169, 969, 729]. sampling [1482, 1055, 1180]. SAP [1132]. Sarathy [496]. SARIMA [1594]. SAS [889]. satellite [1752, 1995]. satellites [1096]. Saturated [325]. Saunders [2065, 1842]. SB [2098]. SB-ETAS [2098]. Scalable [1867, 1600, 1078, 1077, 1918, 1121, 2009, 1827, 1723, 2098]. scalar [1509, 2158, 2074, 1716, 1136, 1308, 1901]. scalar-on-function [1509, 2074, 1716, 1136]. Scale [2136, 1607, 961, 1629, 1156, 1995, 1264, 1693, 920, 2014, 1489, 2114, 2160, 2152, 1747, 1984, 1971, 1257, 2120, 1613, 1828, 1441, 1586, 1352, 1834]. scales [1326]. scaling [813, 4, 1981, 1811, 1710]. scan [1206]. scatter [712, 106]. schedules [96]. scheduling [1198]. scheme [1198, 1097]. schemes [1239, 1419, 165, 2195, 547]. Schmidt [2010]. Schön [1069]. Schrödinger [1786]. science [1115]. scientific [130, 349]. score [2202, 1199, 2017]. scores [263, 1631, 2156, 2062]. Scoring [1649, 11, 1621, 1794, 1551]. Scott [984]. Screen [2036]. screening [1275, 1156, 2209, 2014, 2220, 657, 1782, 150, 1790, 2003, 1300]. SDE [1660]. SDEs [1537]. Search [1912, 326, 334, 1216, 1271, 106, 1920, 1784, 1173, 1485, 786]. seasonal [845, 1052, 2223, 741, 1483, 2178]. seasonality [1506]. secant [75]. second [938, 1343]. second-order [1343]. section [1242]. sectional [1001]. Secure [939]. seemingly [1223, 1264]. seen [919]. Segmental [1340]. Segmentation [817, 1090, 653, 1599, 1910]. segmented [48]. segments [785]. select [2036]. selected [560]. Selecting [1453, 1716]. Selection [1437, 1739, 1679, 1997, 1523, 1330, 907, 114, 1253, 1225, 2125, 1022, 111, 736, 1760, 2075, 393, 2153, 2015, 1382, 787, 1922, 1821, 1509, 1925, 349, 1497, 425, 1855, 900, 112, 742, 446, 2058, 1883, 952, 944, 804, 989, 1449, 2182, 2103, 2087, 743, 2135, 1484, 976, 931, 1149, 1203, 2006, 1222, 104, 179, 1810, 1646, 1106, 2118, 1757, 2029, 467, 1594, 2204, 942, 1472, 1306, 2056, 1999, 1977, 2037, 559, 765, 1854, 744, 683, 2079, 1031, 1207, 1286, 1520, 883, 1811, 1723, 657, 1973, 1748, 2026, 667, 1968, 350, 2129, 1724, 1387, 859, 1224, 1907, 1602, 1541, 1648, 214]. selection [1188, 739, 1041, 482, 1168, 1554, 2213]. selection-induced [2056]. selection/averaging [1907]. selectivity [783]. selector [2104]. selectors [860, 297]. Self [1252, 1600, 1515, 1971, 1622, 151, 241]. self-excitatory [1600]. self-exciting [1515, 1622]. Self-healing [1252]. self-updating [1971]. self-validating [151, 241]. SEM [532]. semantic [1995]. semantics [30]. Semi [760, 1260, 454, 1594, 531, 751, 1304, 2161, 1532, 1583, 157, 1053, 1032, 2202, 662, 816, 445, 887, 1734, 828, 1940, 180, 2049, 990, 2001, 1834]. Semi-automated [1594]. semi-explicit [2161]. semi-infinite [1053]. semi-linear [887]. semi-Markov [1032, 2202, 1734, 990]. semi-mechanistic [1304]. semi-nonparametric [828]. Semi-Parametric [454, 760, 1260, 531, 751, 1532, 157, 662, 816, 445, 1940, 180, 2049, 2001, 1834]. semi-supervised [1583]. semidefinite [1372, 1464]. Semiparametric [414, 2222, 1601, 1516, 409, 1454, 898, 1010, 879, 1272, 2041, 1638, 1627, 829, 2217, 716, 1706, 925, 731, 1921, 714]. semiparametricmodel [1682]. sensitive [663, 233]. sensitivity [865, 1527, 361, 484, 848, 875, 1981, 876]. sensor [759, 673, 2039]. sentiment [2208]. sentiment-topic [2208]. separability [748]. separable [419, 2096]. Separating [175]. separation [183, 1132]. sequence [921, 1706]. sequences [1877, 65, 826, 1995, 521, 686, 1996, 583, 722]. Sequential [1623, 872, 1117, 873, 826, 1539, 1251, 2108, 1052, 1753, 918, 1050, 927, 1712, 1194, 1588, 1330, 1181, 1658, 871, 679, 1002, 892, 362, 336, 598, 1593, 551, 435, 1353, 1880, 1639, 1228, 921, 689, 1572, 909, 649, 1698, 1086, 2093, 2034, 1699]. Series [1192, 563, 403, 407, 404, 573, 790, 405, 1460, 419, 2076, 1506, 1938, 1172, 1752, 1512, 1335, 53, 1643, 1094, 123, 848, 931, 881, 1450, 698, 773, 539, 2218, 1051, 1481, 2000, 1734, 427, 929, 1908, 1820, 1340, 1344, 386, 274, 768, 1614, 1945]. Servers [501]. set [2038, 355, 1824, 1337, 1673, 1639, 750, 260]. set-marking [750]. Sets [456, 499, 1950, 2062, 1433]. settings [738, 1562, 1577]. several [2064, 1591, 1231]. sex [972]. shadow [763, 1317]. shallow [1738]. shallow-water [1738]. Shannon [1755, 1910]. Shape [1796, 1104, 1607, 299, 536, 1574, 289, 2152, 1549, 1154]. shape-restricted [2152, 1549]. shaped [1356]. Shapiro [35]. Shapley [2073]. sharing [901, 1482]. sharp [1918].

Sherman [1105]. Shewhart [727]. shortest [1123]. shot [1744]. Shrinkage [2105, 1039, 2142, 740, 224, 346, 742, 1202, 942, 443, 683, 1031, 301, 1249, 1171, 1994]. shrinkage-type [1171]. Shroud [956]. side [2174]. sign [765]. signal [2151, 1022, 1449, 2044]. signals [1978]. signed [2154]. significance [250, 620, 161, 418]. Silhouette [793]. similar [1130]. similarity [1866]. similarity-based [1866]. Simple [297, 76, 277, 1626, 1750, 160, 280]. simpler [1611]. simplex [513, 1544]. simplexes [1201]. simplicial [2173]. simplicial-simplicial [2173]. simplified [1369]. Simplifying [236]. Simpson [1198]. SIMT [1626]. simulate [1576, 2055]. simulated [1150, 48, 634, 298, 244, 154, 46, 580, 18, 248, 96, 1503]. Simulating [1580, 402, 113, 2154, 159]. Simulation [335, 458, 433, 508, 1123, 1143, 242, 2205, 5, 147, 543, 424, 115, 823, 444, 2155, 789, 1629, 265, 1513, 1731, 550, 1686, 1787, 1952, 2201, 848, 79, 338, 628, 605, 1957, 1353, 1362, 464, 573, 480, 493, 1348, 1754, 267, 2098, 121, 209, 450]. Simulation-based [433]. Simulation-efficient [1123]. Simulations [36, 1534, 192, 625, 472, 1656]. Simultaneous [1164, 1211, 2026, 572, 433, 112, 931, 1594, 632, 1520, 2214]. simultaneously [65]. Single [79, 471, 146, 94, 1278, 1389, 2010, 1006, 948, 949, 1106, 1545, 1793, 960, 1927, 1277, 1315]. single-index [1389, 2010, 948, 1106, 1545, 960, 1927, 1277, 1315]. single-index-coefficient [949]. single-loss [1006]. Single-pass [471]. singly [423]. singular [1202, 2102]. SIR [974, 1436]. site [2132]. six [332]. size [1933, 712, 1497, 1171, 1828]. SiZer [666]. sizes [1255, 1584, 1257]. skeletons [632]. sketch [1561]. sketching [1817, 1888]. skew [842, 992, 1193, 613, 753, 1017, 825, 1445, 1834, 2136]. skew- [1017, 825]. skew-normal [1445, 2136]. skewed [1882]. Sklar [1717]. slab [2042]. Slice [543, 775, 1657, 1944, 801]. Sliced [688, 1249, 958]. Slope [853, 1702]. small [1775, 961, 585, 756, 836, 991, 1578, 729]. small- [961]. small-time [836]. smaller [1595]. SMC [1865, 1419, 700, 1907]. Smirnov [1334]. Smooth [851, 1259, 674, 533, 1295, 1854, 828, 394]. smoothed [1245, 1246, 970, 1222]. Smoothing [987, 1997, 371, 2189, 372, 393, 896, 219, 742, 766, 1631, 2226, 58, 1784, 2223, 1480, 1343, 1939, 1132, 1448, 55, 2046, 2112, 2183, 941, 1302, 2127, 538]. smoothness [1831, 448, 443, 858]. Snipping [1043]. snooker [677]. Sobol’ [1391]. social [1074, 1073]. softly [1836]. softly-penalized [1836]. Software [292, 22, 152, 70, 220, 97, 127, 17, 626, 121, 230, 36]. sojourn [94]. solution [1200, 1428, 811]. solutions [1307, 609, 1498]. solver [1806]. solvers [1495, 1497, 1619]. solving [249, 1459, 1145]. Some [95, 668, 1095, 260, 84, 654, 105, 388, 338, 761, 992, 1101, 834, 1591, 930, 142, 36]. sounds [164]. Source [499, 1926]. sources [461, 568]. Space [719, 1580, 1951, 748, 1497, 679, 1641, 123, 1159, 954, 1664, 1729, 962, 1894, 1279, 432, 175, 918, 866, 759, 1800, 1119, 581, 1526, 2008, 1309, 1262, 1360, 1678, 594]. space-filling [1262]. Space-time [719, 1580, 748, 759, 1309]. Spaced [456, 301]. spaces [590, 1829, 1084, 976, 1885, 1812, 927, 1385, 2096]. Sparse [1185, 1702, 745, 2097, 1127, 1928, 2218, 2079, 738, 1196, 1798, 1791, 2010, 2162, 1468, 1846, 1449, 2074, 1484, 1691, 1839, 1926, 1175, 1992, 1437, 2042, 1436, 2169, 837, 1676, 1236, 1361, 2046, 1324, 1346, 1377, 1376, 1983, 2117, 1187, 1742]. sparsely [863]. sparsity [2082, 1395, 1743, 1807]. Spatial [492, 692, 834, 1916, 1282, 302, 543, 791, 1025, 1128, 1096, 1606, 212, 776, 1265, 1489, 2118, 1318, 134, 917, 637, 1317, 1724, 2061, 1887, 520, 1586, 568, 619]. Spatially [1756, 1403, 1599, 1341, 1746, 1282]. spatio [2216, 1995, 475, 1691, 473, 1706, 1367]. spatio-temporal [2216, 1995, 475, 1691, 473, 1706, 1367]. Spatiotemporal [1664, 2086, 1872, 633]. Spearman [1881]. Special [1114, 671, 1242, 1490, 1058]. specific [832, 445]. specification [1287, 1357, 1443, 942, 786]. specifications [1174]. specified [720]. speckled [642]. Spectral [457, 1537, 881, 573, 2141, 2069, 1426, 1444, 1718, 1363, 1708, 1944, 639]. spectrum [1236]. speed [2088, 583]. spelling [11]. sphere [1576, 1513, 1091]. Spherical [1163, 1511, 2028, 2197]. Spike [2042]. Spline [1779, 1210, 896, 2089, 1426, 1796, 2128, 1220, 839, 858, 1277]. splines [2189, 393, 766, 781, 1054, 1784, 1132, 1369, 898, 1259, 1302]. Split [1759, 1003, 1683]. splitting [823, 1158, 245, 341, 78, 1547]. spread [603]. SPSS [41]. spurious [1109]. square [772]. squared [1215, 2156, 1423]. squares [2105, 334, 2073, 740, 155, 1702, 2074, 478, 2114, 2029, 2173, 943, 2001, 1929]. Stab [2129]. Stab-GKnock [2129]. Stability [1233, 1207, 1225, 2006, 385, 1359]. Stable [1046, 398, 1327, 1355, 1025, 659, 774, 1068, 1787, 2201, 1325, 2052, 1288, 1705, 1067, 1632, 1151, 1754]. stage [1898, 1253, 1638, 1682, 435, 1557]. stagenested [326]. Stairs [361]. stalactite [56]. Standard [211, 107, 1448]. standardized [1262]. standards [87]. Stander [497]. STAR [1263]. starting [111]. Stata [342]. State [1823, 1218, 1951, 1865, 1497, 679, 1641, 123, 1159, 954, 1664, 1125, 1148, 1894, 175, 918, 55, 1385, 581, 351, 1678, 594]. State-dependent [1823, 1218, 1148]. state-of-the-art [55]. state-space [1951, 1497, 1641, 954, 1664, 1894, 918, 1678]. states [1804, 1653]. static [1593, 628]. stationary [335, 1335, 475, 1718, 881, 1746, 1001, 884, 1414, 1891, 1827, 870, 1344, 1311, 274]. Statistic [1739, 1893, 780, 43, 809]. Statistical [485, 130, 1307, 452, 500, 1829, 903, 357, 358, 122, 261, 656, 960, 2181, 178, 81, 601, 82, 1120, 283, 1180, 494, 84, 7, 498, 1072, 462, 244, 1030, 176, 235, 1631, 87, 2081, 1478, 2218, 156, 342, 90, 30, 26, 132, 55, 502, 1317, 2046, 135, 1231, 1856, 1983]. statisticians [2007]. Statistics [402, 85, 1115, 89, 470, 790, 92, 139, 560, 1330, 907, 249, 83, 205, 1566, 1715, 1758, 388, 86, 141, 544, 1000, 772, 752, 1674, 2056, 119, 145, 138, 1075, 143, 91, 747, 144, 278, 1533, 1076, 142, 1818, 1171, 790, 791]. status [2132]. STEC [1309]. Steepest [432]. Stein [208]. step [1553, 1919, 1497, 984, 1805, 657, 649, 476, 2003, 1828]. step-size [1497]. steplength [1551]. stepwise [1153]. stereographic [2175]. stick [1609]. stick-breaking [1609]. Sticky [1791]. Stiefel [993]. Stingo [1079]. STM [1130]. Stochastic [787, 1015, 1435, 1848, 796, 1695, 567, 2031, 1827, 1257, 1969, 1795, 1588, 1531, 1456, 1841, 1216, 1878, 1345, 652, 845, 2196, 1461, 1849, 1625, 596, 1731, 777, 654, 1409, 416, 1530, 1462, 1139, 1419, 1915, 1589, 848, 1600, 2180, 1471, 1467, 2002, 2009, 2134, 1299, 1097, 945, 1592, 1635, 2070, 799, 1052, 1418, 1802, 1832, 876, 921, 936, 1769, 1632, 1656, 1570, 836, 2121, 1098, 1918, 31, 1121, 786, 1296, 2110, 1828, 1730, 438, 619, 1834]. Stochastically [1171]. stock [2052]. Stoehr [1075]. Stoffer [1192]. Stone [253]. stopped [978, 1353]. stopping [179, 1288, 124]. storage [471]. stored [1856]. story [381]. Straightforward [941]. strategies [102, 2157, 1074, 1849, 1733, 1073, 607, 435, 1218, 268, 754, 1121, 2053]. strategy [600, 2036, 2112]. stratification [1774, 874]. stratified [615, 391]. stratum [1271]. stream [168]. streaming [1788, 1967, 1413, 632, 2183]. streams [1319, 1130, 1959, 1227, 1486]. strength [235, 686]. string [821]. stripes [763]. Strong [1496]. strongly [1720]. structural [1623, 845, 777, 931, 2046, 2191]. structurally [1028]. Structure [1811, 1877, 1287, 2221, 1440, 1191, 1190, 1844, 562, 1522, 179, 1418, 374, 1708, 1876, 271, 383, 1676, 1187, 2137]. Structure-based [1811]. structure-preserving [179]. Structured [1290, 2067, 1484, 1434, 742, 1211, 995, 1301, 559, 918, 1916, 755, 2193, 1550, 1323, 833, 1803, 1568, 1613, 450, 1726]. structures [1184, 1598, 812, 79, 1137, 356, 1585]. Student [1913, 1598, 1147, 1466, 842]. Student- [1913, 1598, 1147, 1466, 842]. Studentized [544]. studies [634, 985, 2025, 391]. study [2066, 1941, 975, 635, 2203, 154, 1131, 411, 567, 841, 5, 1439, 1541]. studying [120]. stumps [396]. sub [1261, 1447]. sub-sampling [1447]. sub-tree [1261]. subblock [337]. Subgraph [1835]. subgroup [559]. subject [1821, 832, 1338]. subject-specific [832]. subjects [330]. subordinators [2201]. subsamples [1225]. Subsampling [1593, 2001, 2174, 2159, 2085, 2057, 2160, 2172, 1792, 1561]. subsampling-based [2057]. subset [190, 1922, 2103, 2087, 1999]. subsets [1028]. subspace [2151, 1540, 1640, 1747, 744, 1741]. subspaces [1167, 1438]. substitution [169]. success [629]. Sufficient [795, 2100, 752, 2117]. sum [1215, 1644, 278]. summarizing [625]. Summary [1758, 907, 1566, 1715, 1125, 1674, 1075, 1533, 1076]. sums [1823, 1665, 2126, 455, 468, 1006, 478, 390]. supersaturated [1962]. Supervised [1390, 2115, 1583, 1899]. Support [2112, 1253, 840, 1675, 833, 522, 1900]. Suppressing [233]. suppression [462, 506, 759]. SURE [1954]. SURE-tuned [1954]. surface [760, 824, 489, 870]. surfaces [602]. surrogate [1658, 2082, 1758, 1975, 2169, 2212, 1332]. surrogate-based [2212]. surveillance [1130, 2102]. survey [1371, 133]. survival [1666, 2065, 2020, 1203, 559, 738, 1008, 786, 971, 1991, 2001]. swap [1218]. swarm [966, 1134, 2080]. Swendsen [113]. Switched [1920]. switching [1258, 2012, 754]. SYGRAPH [49]. Symbolic [388, 381, 386, 120]. symmetric [932, 829, 290, 1388, 193]. symmetrically [1851]. Symmetry [1494, 1909, 922]. synchronization [541]. synthesis [1896]. Synthetic [357, 1532, 1405, 2090, 1965]. SYSTAT [49]. SYSTAT/SYGRAPH [49]. system [21, 25, 28]. Systematic [1782]. Systemic [2071]. systems [1775, 20, 1548, 654, 1024, 954, 551, 87, 19, 269, 26, 1918, 502, 1968, 1923, 450].

T [1069, 1785, 516]. t-distributed [1785]. table [233]. tables [499, 68, 772, 1012, 741, 290, 203, 161, 150, 209, 729]. tabular [1965]. tabulated [500]. tail [2076, 2165, 2126, 115, 442]. tailed [1936, 982, 1204, 728, 1047, 1006, 1980, 831]. tailweight [1047]. Takamura [901]. tale [72]. Taming [2092]. tangent [1535]. target [1107]. targets [1395, 1675]. task [2168]. taut [821]. Taylor [836]. teaching [249]. technique [2031]. techniques [52, 488, 284, 1088, 29, 217, 19, 1203, 436, 156, 738, 737, 201, 594]. Teh [1067]. Temperature [96]. temperatures [1218]. tempered [827, 1452, 1787, 2201, 218, 1754]. tempering [955, 940, 851, 725, 1218, 1663, 1503, 1667]. Template [2097, 415, 722]. templates [275]. temporal [1457, 2216, 1995, 475, 1691, 473, 1706, 1367]. temporally [1409, 2176]. ten [86, 89, 90, 92]. tenability [142]. tensor [2189, 1536, 1701, 1904, 941, 1302]. tensor-variate [1904]. term [2031]. test [188, 771, 840, 1893, 1180, 2116, 484, 388, 820, 1833, 544, 769, 1000, 661, 236, 884, 1891, 834, 35, 1709, 394, 2023, 2181, 2096, 1315]. Testing [168, 524, 2052, 1909, 194, 565, 405, 2184, 558, 1330, 2143, 964, 250, 776, 1292, 701, 337, 1238, 1124, 1148, 814, 1007, 2167, 1112, 1318, 411, 1200, 1240, 2215, 409, 367, 2111, 1352]. tests [166, 748, 1775, 1773, 712, 1083, 417, 715, 1958, 1045, 414, 1534, 1882, 772, 57, 542, 2099, 37, 2000, 1136, 290, 887, 922, 1591, 161, 649, 418]. text [849, 32]. textual [1345, 1461, 1618]. texture [356, 1311]. textures [1415]. th [170]. their [1320, 770, 132, 839, 1838]. theorem [1740]. theorems [769]. theoretic [1573, 395]. theoretical [1162, 74, 495, 496]. Theory [1192, 446, 1777, 1879, 1973]. there [1670]. Thermodynamic [1313, 1198, 1472]. three [326, 2031, 810]. three-term [2031]. three-way [810]. Threshold [4, 1025, 443]. Threshold-range [4]. Thresholded [457, 234]. thresholding [631, 686, 185, 1491]. thresholds [683, 2220]. tilting [1989]. Time [1192, 403, 407, 404, 573, 405, 1141, 1553, 1580, 1460, 2161, 94, 1506, 748, 1938, 1514, 1775, 2107, 1019, 1272, 2041, 1799, 1752, 1512, 1513, 1335, 53, 1641, 1643, 526, 1728, 123, 848, 1575, 931, 881, 1450, 1353, 698, 1635, 2070, 773, 962, 539, 2139, 2218, 1653, 1051, 1481, 2000, 1557, 1734, 1632, 427, 929, 1567, 1842, 1806, 836, 759, 1908, 1916, 719, 1820, 1340, 1344, 386, 1206, 1923, 1948, 1338, 695, 1309, 274, 1602, 1565, 2183, 768, 1972, 1614, 1991, 2001, 1195, 1945]. time-average [1195]. time-aware [1938]. Time-course [1141]. time-parallel [1806]. time-series [1752]. time-to-event [1557, 1916, 1948]. time-varying [1632, 1923, 1565]. time/space [962]. times [634, 1475, 1803]. Timing [1525]. TMB [2097]. Tobit [1847, 2047, 1147]. Toeplitz [1597]. tomography [601, 358]. tool [361, 2040, 1317]. tools [591, 355, 1478, 1430]. topic [1345, 1461, 1571, 2208]. topic-based [1571]. topics [1878, 1618]. Topology [1958]. Topology-driven [1958]. Toroidal [1890, 1921]. torus [1421]. Total [2011, 758]. totals [233]. tournament [1863]. trace [1781]. Tracking [1375, 1660, 1137, 1107]. Tractable [577, 1245, 1246, 1219]. Tracy [1817]. traditional [19]. traffic [855, 1567, 1515]. train [1536]. training [1799, 2077, 1930, 260]. trait [1020]. trajectories [527, 2139, 1373, 1965]. Trajectory [1409]. trans [636, 1896]. Trans-cGAN [1896]. trans-dimensional [636]. transdimensional [1458]. transfer [427, 2219]. Transform [529, 1120, 681, 720, 274, 1269]. Transformation [1510, 1932, 1382, 541, 1048, 825, 1820, 898]. transformations [149, 1539, 1810, 1762]. Transformed [933, 1097]. transformer [1896]. transformer-Unet-based [1896]. transforming [662]. transforms [2012]. transition [1157, 674, 1772, 836]. transitions [827, 218]. transmission [759]. transport [1656]. transportation [1430]. treatment [2094, 2020, 2025, 899]. treatments [1650]. Tree [1434, 1673, 559, 1613, 518, 1261, 577, 1414, 758, 1323, 1803, 1948, 2008, 2110, 901, 1732]. tree-based [758, 1948, 1732]. Tree-structured [1434, 559, 1613, 1323, 1803]. Trees [1399, 2047, 760, 181, 27, 1245, 1246, 1683, 204, 1384, 245, 1949, 1753, 1616, 341, 78]. trend [1506, 1130, 1024, 2223, 175]. trend-coherence [1130]. trend-renewal [1024]. trends [232]. trial [435]. trials [1765, 689, 1650]. triangulation [225, 968]. triangulations [299]. trimmed [334]. trimming [1266]. trivariate [525]. true [1007]. Truncated [1813, 789, 595, 242, 147, 1815]. truncation [1197, 1183]. truncations [1303]. trust [1697]. trust-region [1697]. try [928, 1606, 581, 1862, 615]. TSP [49]. Tukey [663, 2213]. tune [2199]. tuned [1954]. Tuning [827, 2120, 1453, 1849, 1387, 1822]. Turin [956]. turning [1576]. tutorial [672, 961, 522, 103, 639]. Tweedie [563, 647, 749, 973]. Twin [736, 79]. twin-heaps [79]. twisting [1348]. twisting-based [1348]. Two [982, 166, 685, 984, 1898, 1253, 1893, 2123, 1271, 1773, 29, 706, 1638, 1682, 299, 388, 701, 972, 820, 74, 915, 233, 1557, 290, 921, 1891, 1200, 26, 1650, 1790, 1793, 2111, 2145, 2096, 1204]. Two-dimensional [685, 233, 1891]. two-level [2123, 1790]. two-phase [1271, 1793]. two-sample [1893, 388, 701, 820, 1200, 2111, 2096]. two-sex [972]. two-stage [1898, 1253, 1638, 1682, 1557]. Two-tailed [982, 1204]. two-way [290]. type [1580, 1010, 1773, 766, 842, 2068, 761, 1000, 1706, 237, 1776, 1754, 2049, 990, 1970, 1171, 1994, 1482, 1305, 2083, 478, 2047]. types [1637].

uCPD [2224, 2177]. ultrahigh [1275, 1300]. ultrametric [2032]. ultrasonography [289]. umbrella [771, 1252]. un-normalised [1267]. unadjusted [1625]. Unbalanced [821, 1830, 478]. Unbiased [2122, 1303, 1617, 1725, 593, 696, 2140]. uncertain [136]. Uncertainty [1652, 1655, 2212, 1553, 865, 2157, 426, 1660, 1191, 1190, 1465, 1090, 743, 1458, 1696, 1153]. Unconstrained [210, 1763]. uncovering [1618]. undecimated [274]. under-dispersed [1584]. Understanding [1049, 610, 32]. underwater [164]. undirected [1914, 1079, 2084, 1080]. Unequal [294, 1255]. unequally [301]. Unet [1896]. unevenly [1302]. unfolding [2187]. unidimensional [200]. unified [1155, 120, 1145]. Uniform [1775, 1085, 1201, 65, 257, 1885, 682]. uniformity [840, 1709]. Uniforms [14, 658]. unifying [1495, 742, 1211]. unigrams [1618]. unimodal [980]. Union [1200]. Union-intersection [1200]. unit [1045, 1647, 1001]. units [1404]. Univariate [980, 1893, 882, 1768, 148]. universal [2143, 1021, 1809, 1288]. unknown [1951, 840, 575, 505, 1138, 1653, 614, 618, 1529, 1050]. unlabeled [1690, 2147]. Unlabelled [1883]. unnormalised [1120]. unobserved [1210, 548]. unrelated [1223, 1264]. Unrestricted [1670, 1445]. Unsupervised [1910, 1995]. unsuspected [106]. update [487]. updater [677]. updates [1079, 1080, 1387]. Updating [1698, 1971]. upper [1861]. Use [454, 478, 1884, 1320, 1278, 250, 1072, 226, 363, 387, 365, 1101, 1071, 1838, 1154, 150]. useful [355]. user [1571]. uses [140]. Using [2199, 963, 1223, 1601, 383, 512, 1654, 146, 485, 1069, 709, 760, 1122, 2189, 1025, 1333, 1225, 1035, 527, 115, 1319, 181, 426, 879, 1147, 1420, 1401, 1382, 849, 1470, 634, 1535, 615, 1440, 2080, 1070, 168, 524, 425, 429, 1548, 1426, 2162, 517, 286, 802, 123, 779, 2174, 1831, 416, 1465, 1915, 521, 1379, 848, 1458, 2097, 1399, 838, 2094, 2132, 1198, 2226, 1493, 1326, 1375, 875, 1325, 2020, 544, 1265, 781, 1054, 625, 413, 1809, 1226, 1397, 1905, 1647, 1810, 2060, 1964, 752, 516, 613, 753, 1017, 1749, 2029, 1026, 717, 1909, 1051, 1747, 1705, 1306]. using [473, 831, 1368, 1977, 2037, 218, 624, 461, 1143, 1854, 683, 1939, 1364, 1366, 1529, 1581, 2220, 375, 1806, 658, 1153, 1987, 912, 719, 275, 1722, 700, 1369, 1710, 1091, 1748, 2190, 2026, 1744, 124, 67, 350, 271, 2098, 2129, 1676, 1257, 990, 1179, 1790, 1329, 1328, 12, 1259, 1093, 193, 302, 1089, 1188, 1991, 1921, 1441, 1332].

VAE [1769]. validated [350]. validating [151, 241]. Validation [559, 817, 2119, 2043, 779, 363, 364, 195, 846, 2163, 158, 196, 1368, 1329, 1328, 1900]. validatory [1213]. validity [1582]. Value [506, 1906, 236, 2215, 2102, 1428]. valued [1576, 1021, 1980, 1450, 2004, 1579, 2113, 1850, 1945]. values [251, 417, 1440, 1479, 157, 1202, 472, 37, 366, 267, 837]. vanilla [1907]. Var [1738]. variability [215, 1282]. Variable [1855, 989, 1203, 1810, 1106, 2029, 1306, 1854, 1679, 1523, 371, 114, 1253, 2125, 1986, 2015, 1382, 787, 1821, 1509, 1925, 1479, 425, 446, 1047, 1848, 952, 1804, 159, 1283, 1449, 743, 979, 686, 1670, 931, 645, 2020, 625, 1809, 179, 436, 1213, 2118, 1757, 942, 1442, 2220, 902, 1416, 1520, 1973, 2026, 2129, 1387, 859, 1177, 1602, 1541, 1188, 739, 482, 1554, 1394]. variables [1273, 1021, 1823, 1665, 2126, 1215, 1919, 444, 1998, 1821, 1463, 1129, 706, 1354, 701, 1006, 226, 1906, 1928, 1113, 1442, 1417, 147, 1119, 1748, 278, 259, 43, 1579]. Variance [1789, 589, 1556, 1680, 848, 959, 608, 1514, 693, 593, 934, 2054, 1848, 1241, 781, 441, 1860, 965, 2031, 1342, 210, 180, 1448, 649, 512, 2110, 1195]. variance-corrected [2054]. variance-covariance [210]. variances [166]. variate [1201, 2005, 1021, 149, 75, 996, 1902, 2185, 73, 1646, 5, 1904]. Variates [14, 1456, 1886, 192, 1018, 169]. variation [394]. Variational [1847, 1801, 1743, 1232, 711, 856, 1405, 1236, 1677, 1152, 2214, 1991, 1554, 1937, 1941, 2088, 347, 1809, 1694, 1853, 1381, 1939, 1827, 1257, 1361, 1568, 2093, 2034, 1630, 1770, 835, 1736, 1969, 1552, 2109, 1807, 1698, 1642]. variogram [475]. varying [1237, 1434, 2225, 1599, 1632, 1039, 1923, 1565, 960, 1927]. varying-coefficient [960, 1927]. vast [1995]. Vecchia [1621, 2131]. vector [1576, 335, 833, 522, 1676, 2112, 1900]. vector-valued [1576]. vectorization [217]. vectorization/parallelization [217]. vectors [1484, 658, 1591, 2111]. version [590, 1824, 248, 616]. versus [845, 749]. vertex [1194]. vertices [1345]. Vervaat [1952]. very [450]. VGA [1694]. VGAL [2093, 2034]. via [558, 2140, 2122, 2005, 1355, 1531, 1532, 893, 1197, 817, 611, 1680, 1841, 1863, 1181, 1611, 823, 279, 1986, 2075, 1088, 1134, 1636, 1912, 2010, 263, 1855, 498, 1412, 1507, 2116, 1758, 1883, 952, 1804, 1266, 1890, 1303, 1621, 1370, 1729, 1703, 1669, 1779, 1127, 2088, 1888, 509, 347, 1250, 1825, 910, 1926, 1334, 1353, 772, 472, 576, 2084, 2128, 1928, 1381, 2218, 825, 2012, 1472, 1033, 1999, 1092, 2168, 1031, 1960, 449, 1173, 1153, 2115, 2193, 1740, 1098, 1119, 1118, 581, 1423, 1387, 2142, 1786, 14, 169, 971, 1249, 1745, 1041, 807, 1464, 1561, 1726, 2181, 438, 619, 1725]. video [2102]. view [1495]. Vine [1358, 1801, 1365, 1437, 1369]. vines [1054]. visualizable [1046]. visualization [763]. visualizing [1035]. Vlasov [1402, 2121, 2146]. volatility [760, 2180, 1947, 1632]. Volterra [2002]. Volume [255, 440, 1812]. Voronoi [1480, 637]. voting [2115]. vs [1368, 974].

W [1067]. WAIC [1213, 1368, 1329, 1328]. Wald [1773]. Wald-type [1773]. walk [287, 506, 1710]. Wang [113]. warpings [2211]. warranty [1943]. Wasserstein [1059, 2095, 1060, 1768]. waste [1410]. waste-recycling [1410]. water [1738]. Wavelet [574, 1679, 1161, 1875, 427, 301, 419, 524, 546, 631, 234, 1331, 534, 1610, 443, 1414, 185, 683, 1891, 1311, 274, 1614]. Wavelet-based [1679, 1161, 1614]. wavelet-packet [1610]. wavelets [821, 164, 1491]. way [290, 810, 572]. weak [2072]. weaver [1412]. Wehrly [1240]. Weibull [1852, 2195, 2145]. Weight [1503, 1482]. Weight-preserving [1503]. Weighted [748, 1518, 1034, 1215, 2179, 266, 1731, 703, 2176, 1334, 943]. weighting [1615, 1224]. weights [1281, 1499, 1097, 1568, 943]. well [366]. were [1595]. Which [1463]. while [1898]. white [1891]. Wichmann [66]. widely [1293, 1131]. Widom [1817]. width [1022]. Wilk [35]. Wilkinson [1077]. win [1863]. WinBUGS [374]. wind [1660, 646]. Winding [361]. window [1141]. windows [1022]. wise [1043, 1112, 1724]. Wishart [780, 1910]. with/without [1902]. within [1559]. Without [1385, 1715, 1902, 2132, 1672]. Without-replacement [1385]. workflow [2061]. worlds [585]. wrapped [1746].

XploRe [55].

years [86, 89, 90, 92]. Yor [1714].

zag [1780, 1985]. Zero [965, 2164]. zeroinflated [2164]. zeros [693, 1993, 1889]. zig [1780, 1985]. zig-zag [1780, 1985]. zigzag [1720]. zones [776].

References

Hand:1991:E

[1] David J. Hand. Editorial. Statistics and Computing, 1(1):i–ii, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                              tronic). URL http://link.springer.
                                              com/accesspage/article/10.1007/
                                              BF01890831.

Dellaportas:1991:PEI

[2] Petros Dellaportas and David Wright. Positive embedded integration in Bayesian analysis. Statistics and Computing, 1(1):1–12, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                              tronic). URL http://link.springer.
                                              com/article/10.1007/BF01890832.

Do:1991:QRR

[3] Kim-Anh Do and Peter Hall. Quasirandom resampling for the bootstrap. Statistics and Computing, 1 (1):13–22, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01890833.

Fisch:1991:TRS

[4] Robert Fisch, Janko Gravner, and David Griffeath. Threshold-range scaling of excitable cellular automata. Statistics and Computing, 1(1):23– 39, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01890834.

Ringrose:1991:SSC

[5] T. J. Ringrose and W. J. Krzanowski. Simulation study of confidence regions for canonical variate analysis. Statistics and Computing, 1(1):41–46, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01890835.

Cleveland:1991:CML

[6] William S. Cleveland and Eric Grosse. Computational methods for local regression. Statistics and Computing, 1(1):47–62, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL ftp://cm.bell-labs.com/cm/
cs/doc/91/4-04.ps.gz;   http://
link.springer.com/article/10.1007/
BF01890836.

Daly:1991:SSC

[7] Fergus Daly. SC — statistical calculator. Statistics and Computing, 1 (1):63–70, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01890837.

Everitt:1991:BP

[8] B. S. Everitt. BMDP PC-90. Statistics and Computing, 1(1):71–73, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01890838.

Anonymous:1991:HCa

[9] Anonymous. Help & contacts. Statistics and Computing, 1(1):??, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Michalewicz:1991:GAN

[10] Zbigniew Michalewicz and Cezary Z. Janikow. Genetic algorithms for numerical optimization. Statistics and Computing, 1(2):75–91, December 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01889983.

Church:1991:PSS

[11] Kenneth W. Church and William A. Gale. Probability scoring for spelling correction. Statistics and Computing, 1(2):93–103, December 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01889984.

Verdinelli:1991:BAO

[12] Isabella Verdinelli and Larry Wasserman. Bayesian analysis of outlier problems using the Gibbs sampler. Statistics and Computing, 1(2):105–117, December 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889985.

Carlin:1991:IMC

[13] Bradley P. Carlin and Alan E. Gelfand. An iterative Monte Carlo method for nonconjugate Bayesian analysis. Statistics and Computing, 1(2):119–128, December 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889986.

Wakefield:1991:EGR

[14] J. C. Wakefield, A. E. Gelfand, and A. F. M. Smith. Efficient generation of random variates via the ratioof-uniforms method. Statistics and Computing, 1(2):129–133, December 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889987;
 http://www.springerlink.com/content/
w1vt431r722p17j8/.

Weinberg:1991:EI

[15] Richard Weinberg. Envisioning information. Statistics and Computing, 1(2):135, December 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01889988.

Anonymous:1991:HCb

[16] Anonymous. Help & contacts. Statistics and Computing, 1(2):??, December 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Macleod:1992:PEN

[17] Allan J. Macleod. Performance evaluation of normal distribution software. Statistics and Computing, 2(1):1–5, March 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01890543.

Kjaerulff:1992:ODP

[18] Uffe Kjærulff. Optimal decomposition of probabilistic networks by simulated annealing. Statistics and Computing, 2(1):7–17, March 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01890544.

Hirschberg:1992:CEM

[19] Joseph G. Hirschberg. A computationally efficient method for bootstrapping systems of demand equations: A comparison to traditional techniques. Statistics and Computing, 2(1):19–24, March 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01890545.

Dawid:1992:AGP

[20] A. P. Dawid. Applications of a general propagation algorithm for probabilistic expert systems. Statistics and Computing, 2(1):25–36, March 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01890546.

Cowell:1992:FRE

[21] R. G. Cowell and A. P. Dawid. Fast retraction of evidence in a probabilistic expert system. Statistics and Computing, 2(1):37–40, March 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01890547.

Bedrick:1992:SR

[22] E. J. Bedrick. Software review. Statistics and Computing, 2(1):41–45, March 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01890548.

Anonymous:1992:HCa

[23] Anonymous. Help & contacts. Statistics and Computing, 2(1):??, March 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Anonymous:1992:E

[24] Anonymous.  Editorial.  Statis-

tics and Computing, 2(2):i–ii, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01889581.

Lorenzen:1992:DES

[25] Thomas J. Lorenzen, Lynn T. Truss, W. Scott Spangler, and Andrew B. Parker William T. Corpus. DEXPERT: an expert system for the design of experiments. Statistics and Computing, 2(2):47–54, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01889582.

Raes:1992:ITC

[26] Jan F. M. Raes. Inside two commercially available statistical expert systems. Statistics and Computing, 2(2):55–62, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01889583.

Buntine:1992:LCT

[27] Wray Buntine. Learning classification trees. Statistics and Computing, 2(2):63–73, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01889584.

Tsujino:1992:KAI

[28] Katsuhiko Tsujino and Shogo Nishida. A knowledge acquisition inductive system driven by empirical interpretation of derived results. Statistics and Computing, 2(2):75–81, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889585.

Crawford:1992:ATP

[29] Stuart L. Crawford and Robert M. Fung. An analysis of two probabilistic model induction techniques. Statistics and Computing, 2(2):83–90, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889586.

Pearl:1992:SSC

[30] Judea Pearl and Thomas S. Verma. A statistical semantics for causation. Statistics and Computing, 2(2):91–95, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889587.

Smyth:1992:ASC

[31] Padhraic Smyth. Admissible stochastic complexity models for classification problems. Statistics and Computing, 2(2):97–104, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01889588.

Goldman:1992:PTU

[32] Robert P. Goldman and Eugene Charniak. Probabilistic text understanding. Statistics and Computing, 2 (2):105–114, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01889589.

Anonymous:1992:FP

[33] Anonymous. Forthcoming papers. Statistics and Computing, 2(2):115, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01889590.

Anonymous:1992:HCb

[34] Anonymous. Help & contacts. Statistics and Computing, 2(2):??, June 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Royston:1992:ASW

[35] Patrick Royston. Approximating the Shapiro–Wilk W -test for nonnormality. Statistics and Computing, 2 (3):117–119, September 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01891203.

VanPul:1992:SJM

[36] Mark Van Pul. Simulations on the Jelinski–Moranda model of software reliability; application of some parametric bootstrap methods. Statistics and Computing, 2(3):121–136, September 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01891204.

Lockhart:1992:CAV

[37] R. Lockhart and T. Swartz. Computing asymptotic p-values for EDF tests. Statistics and Computing, 2(3): 137–141, September 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01891205.

Celeux:1992:DRD

[38] Gilles Celeux and Abdallah Mkhadri. Discrete regularized discriminant analysis. Statistics and Computing, 2(3): 143–151, September 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01891206.

Nason:1992:MM

[39] G. P. Nason and Robin Sibson. Measuring multimodality. Statistics and Computing, 2(3):153–160, September 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01891207.

DiCiccio:1992:AAI

[40] Thomas J. DiCiccio, Michael A. Martin, and G. Alastair Young. Analytical approximations for iterated bootstrap confidence intervals. Statistics and Computing, 2(3):161–171, September 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01891208.

Vaughan:1992:SM

[41] Jonathan Vaughan. SPSS for the Macintosh. Statistics and Computing, 2 (3):173–177, September 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01891209.

Anonymous:1992:HCc

[42] Anonymous. Help & contacts. Statistics and Computing, 2(3):??, September 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

VanderGenugten:1992:DQN

[43] B. B. Van der Genugten. Density of the quotient of non-negative quadratic forms in normal variables with application to the F -statistic. Statistics and Computing, 2(4):179–182, December 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889677.

Tibshirani:1992:PCR

[44] Robert Tibshirani. Principal curves revisited. Statistics and Computing, 2 (4):183–190, December 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01889678.

Hwarng:1992:BML

[45] H. Brian Hwarng and Norma Faris Hubele. Boltzmann machines that learn to recognize patterns on control charts. Statistics and Computing, 2 (4):191–202, December 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01889679.

Ingrassia:1992:CBS

[46] Salvatore Ingrassia. A comparison between the simulated annealing and the EM algorithms in normal mixture decompositions. Statistics and Computing, 2(4):203–211, December 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889680.

Dewey:1992:AFD

[47] Michael E. Dewey. Algorithms for frequency distributions: efficiency and generality comparisons. Statistics and Computing, 2(4):213–220, December 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889681.

Atkinson:1992:SAS

[48] A. C. Atkinson. A segmented algorithm for simulated annealing. Statistics and Computing, 2(4):221–230, December 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889682.

Jarrett:1992:SSM

[49] Jeffrey Jarrett. SYSTAT/SYGRAPH and Micro–TSP. Statistics and Computing, 2(4):231–236, December 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01889683.

Anonymous:1992:HCd

[50] Anonymous. Help & contacts. Statistics and Computing, 2(4):??, December 1992. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Anonymous:1993:Ea

[51] Anonymous.  Editorial.  Statis-

tics and Computing, 3(1):i, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00146946.

Bohning:1993:ATF

[52] Dankmar Böhning. Acceleration techniques in fixed-point methods for finding percentage points. Statistics and Computing, 3(1):1–5, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00146947.

Delampady:1993:HBA

[53] M. Delampady, I. M. L. Yee, and J. V. Zidek. Hierarchical Bayesian analysis of a discrete time series of Poisson counts. Statistics and Computing, 3(1):7–15, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00146948.

Aitkin:1993:PBF

[54] Murray Aitkin. Posterior Bayes factor analysis for an exponential regression model. Statistics and Computing, 3(1):17–22, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00146949.

Schimek:1993:SCE

[55] M. G. Schimek and K. G. Schmaranz. The statistical computing environment XploRe and state-of-the-art density and regression smoothing. Statistics and Computing, 3(1):23–26, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00146950.

Atkinson:1993:SPD

[56] A. C. Atkinson and H.-M. Mulira. The stalactite plot for the detection of multivariate outliers. Statistics and Computing, 3(1):27–35, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00146951.

Krzanowski:1993:PTC

[57] W. J. Krzanowski. Permutational tests for correlation matrices. Statistics and Computing, 3(1):37–44, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00146952.

Jones:1993:KSF

[58] M. C. Jones and I. S. Bradbury. Kernel smoothing for finite populations. Statistics and Computing, 3(1):45–50, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00146953.

Thomas:1993:NFC

[59] Alun Thomas. A note on the fourcolourability of pedigrees and its consequences for probability calculations. Statistics and Computing, 3(1):51–54, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00146954.

Seligson:1993:BR

[60] Richard Seligson, Othar Hansson, Andrew Mayer, and Gerhard Holt. Book reviews. Statistics and Computing, 3 (1):55, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00146955.

Anonymous:1993:BRA

[61] Anonymous. Book reviews announcement. Statistics and Computing, 3(1): 58, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00146956.

Anonymous:1993:Eb

[62] Anonymous.  Erratum.  Statis-

tics and Computing, 3(1):59, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF00146957.

Anonymous:1993:HCa

[63] Anonymous. Help & contacts. Statistics and Computing, 3(1):??, March 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Anonymous:1993:Ec

[64] Anonymous. Editorial. Statistics and Computing, 3(2):i, June 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00153063.

Anderson:1993:CCB

[65] N. H. Anderson and D. M. Titterington. Cross-correlation between simultaneously generated sequences of pseudorandom uniform deviates. Statistics and Computing, 3(2):61–65, June 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00153064.

DeMatteis:1993:LRC

[66] A. De Matteis and S. Pagnutti. Long-range correlation analysis of the Wichmann–Hill random number generator. Statistics and Computing, 3(2):67–70, June 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00153065.

Smith:1993:RCH

[67] D. J. Smith, T. C. Bailey, and A. G. Munford. Robust classification of high-dimensional data using artificial neural networks. Statistics and Computing, 3(2):71–81, June 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00153066.

Granville:1993:EAE

[68] V. Granville and E. Schifflers. Efficient algorithms for exact inference in 2 × 2 contingency tables. Statistics and Computing, 3(2):83–87, June 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00153067.

Aitkin:1993:AMD

[69] Murray Aitkin and Camil Fuchs. An analysis of models for the dilution and adulteration of fruit juice. Statistics and Computing, 3(2):89–99, June 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00153068.

Dusoir:1993:SR

[70] Tony Dusoir. Software review. Statistics and Computing, 3(2):101–102, June 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

                                        com/accesspage/article/10.1007/
                                        BF00153069.

Anonymous:1993:HCb

[71] Anonymous. Help & contacts. Statistics and Computing, 3(2):??, June 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Arslan:1993:DCE

[72] Olcay Arslan, Patrick D. L. Constable, and John T. Kent. Domains of convergence for the EM algorithm: a cautionary tale in a location estimation problem. Statistics and Computing, 3(3):103–108, September 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00147772.

Laud:1993:RVG

[73] Purushottam W. Laud, Paul Ramgopal, and Adrian F. M. Smith. Random variate generation from Ddistributions. Statistics and Computing, 3(3):109–112, September 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF00147773.

Hadi:1993:FTR

[74] Ali S. Hadi and Hans Nyquist. Further theoretical results and a comparison between two methods for approximating eigenvalues of perturbed covariance matrices. Statistics and Computing, 3(3):113–123, September 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00147774.

Devroye:1993:RVG

[75] Luc Devroye. On random variate generation for the generalized hyperbolic secant distributions. Statistics and Computing, 3(3):125–134, September 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00147775.

Jones:1993:SBC

[76] M. C. Jones. Simple boundary correction for kernel density estimation. Statistics and Computing, 3(3): 135–146, September 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00147776.

Anonymous:1993:HCc

[77] Anonymous. Help & contacts. Statistics and Computing, 3(3):??, September 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Taylor:1993:BDS

[78] Paul C. Taylor and Bernard W. Silverman. Block diagrams and splitting criteria for classification trees. Statistics and Computing, 3(4):147–161, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00141771.

Hatzinger:1993:STH

[79] R. Hatzinger and W. Panny. Single and twin-heaps as natural data structures for percentile point simulation algorithms. Statistics and Computing, 3(4):163–170, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00141772.

Bowman:1993:DBE

[80] Adrian Bowman and Peter Foster. Density based exploration of bivariate data. Statistics and Computing, 3 (4):171–177, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00141773.

Anonymous:1993:FSR

[81] Anonymous. The future of statistical research. Statistics and Computing, 3(4):179, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF00141774.

Barndorff-Nielsen:1993:IAF

[82] Ole E. Barndorff-Nielsen. Important areas for future statistical research. Statistics and Computing, 3 (4):181, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF00141775.

Chambers:1993:GLS

[83] John M. Chambers. Greater or lesser statistics: a choice for future research. Statistics and Computing, 3 (4):182–184, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00141776.

Copas:1993:SIS

[84] J. B. Copas. On some important statistical problems. Statistics and Computing, 3(4):185–187, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00141777.

Efron:1993:SC

[85] Bradley Efron. Statistics in the 21st century. Statistics and Computing, 3 (4):188–190, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00141778.

Gower:1993:NTY

[86] J. C. Gower. The next ten years in statistics? Statistics and Computing, 3 (4):191–193, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00141779.

Hibbert:1993:NRE

[87] Jack Hibbert. The need for research into the effectiveness of international statistical systems and standards. Statistics and Computing, 3 (4):194–196, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00141780.

Legendre:1993:RDM

[88] Pierre Legendre. Real data are messy. Statistics and Computing, 3 (4):197–199, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00141781.

Moore:1993:SRN

[89] David S. Moore. Statistics research: the next ten years. Statistics and Computing, 3(4):200–201, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF00141782.

Nelder:1993:MIA

[90] J. A. Nelder. The most important areas of statistical research in the next ten years. Statistics and Computing, 3 (4):202–203, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00141783.

Rubin:1993:FS

[91] Donald B. Rubin. The future of statistics. Statistics and Computing, 3(4):204, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00141784.

Smith:1993:SRN

[92] Richard L. Smith. Statistics research for the next ten years. Statistics and Computing, 3(4):205–208, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00141785.

Anonymous:1993:HCd

[93] Anonymous. Help & contacts. Statistics and Computing, 3(4):??, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Ball:1994:NEO

[94] F. G. Ball and G. F. Yeo. Numerical evaluation of observed sojourn time distributions for a single ion channel incorporating time interval omission. Statistics and Computing, 4(1):1–12, March 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00143919.

Eslava:1994:SCP

[95] G. Eslava and F. H. C. Marriott. Some criteria for projection pursuit. Statistics and Computing, 4(1):13–20, March 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00143920.

Stander:1994:TSS

[96] Julian Stander and Bernard W. Silverman. Temperature schedules for simulated annealing. Statistics and Computing, 4(1):21–32, March 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00143921.

Madigan:1994:SR

[97] David Madigan. Software review. Statistics and Computing, 4(1):33–39, March 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00143922.

Jones:1994:BR

[98] M. C. Jones. Book review. Statistics and Computing, 4(1):41–46, March 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00143923.

Anonymous:1994:HCa

[99] Anonymous. Help & contacts. Statistics and Computing, 4(1):??, March 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Hand:1994:E

[100] David J. Hand. Editorial. Statistics and Computing, 4(2):47, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF00175351.

Michalewicz:1994:EC

[101] Zbigniew Michalewicz. Evolutionary computation. Statistics and Computing, 4(2):49–50, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00175352.

Back:1994:BAE

[102] Thomas Bäck and Frank Hoffmeister. Basic aspects of evolution strategies. Statistics and Computing, 4(2):51–63, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00175353.

Whitley:1994:GAT

[103] Darrell Whitley. A genetic algorithm tutorial. Statistics and Computing, 4(2):65–85, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00175354.

Koza:1994:GPM

[104] John R. Koza. Genetic programming as a means for programming computers by natural selection. Statistics and Computing, 4(2):87–112, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00175355.

Fogel:1994:EPI

[105] David B. Fogel. Evolutionary programming: an introduction and some current directions. Statistics and Computing, 4(2):113–129, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00175356.

Glover:1994:GAS

[106] Fred Glover. Genetic algorithms and scatter search: unsuspected potentials. Statistics and Computing, 4 (2):131–140, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00175357.

Michalewicz:1994:NSM

[107] Zbigniew Michalewicz. Non-standard methods in evolutionary computation. Statistics and Computing, 4 (2):141–155, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00175358.

Hand:1994:BR

[108] David J. Hand. Book reviews. Statistics and Computing, 4(2):157–159, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00175359.

Anonymous:1994:HCb

[109] Anonymous. Help & contacts. Statistics and Computing, 4(2):??, June 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Carothers:1994:CAC

[110] Andrew Carothers and Jim Piper. Computer-aided classification of human chromosomes: a review. Statistics and Computing, 4(3):161–171, September 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00142568.

Brooks:1994:ASP

[111] S. P. Brooks and B. J. T. Morgan. Automatic starting point selection for function optimization. Statistics and Computing, 4(3):173–177, September 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00142569.

Dumouchel:1994:ASP

[112] William H. Dumouchel and Thomas P. Lane. Automatic selection of the proper family for simultaneous confidence intervals. Statistics and Computing, 4(3):179–187, September 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00142570.

Gray:1994:SPG

[113] A. J. Gray. Simulating posterior Gibbs distributions: a comparison of the Swendsen–Wang and Gibbs sampler methods. Statistics and Computing, 4(3):189–201, September 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00142571.

Barrett:1994:CFV

[114] Bruce E. Barrett and J. Brian Gray. A computational framework for variable selection in multivariate regression. Statistics and Computing, 4(3): 203–212, September 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00142572.

Blackwell:1994:EET

[115] Paul Blackwell. The efficient estimation of tail probabilities for extremes of moving average processes using conditional simulation. Statistics and Computing, 4(3):213–218, September 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00142573.

Lunneborg:1994:BR

[116] Clifford Lunneborg and David J. Hand. Book review. Statistics and Computing, 4(3):219–220, September 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF00142574.

Anonymous:1994:HCc

[117] Anonymous. Help & contacts. Statistics and Computing, 4(3):??, September 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Anonymous:1994:E

[118] Anonymous.  Editorial.  Statis-

tics and Computing, 4(4):i, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF00156744.

Nelder:1994:SLM

[119] J. A. Nelder. The statistics of linear models: back to basics. Statistics and Computing, 4(4):221–234, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF00156745. See

comments [138, 139, 141, 142, 143, 144] and rejoinder [145].

Stafford:1994:SCU

[120] James E. Stafford, David F. Andrews, and Yong Wang. Symbolic computation: a unified approach to studying likelihood. Statistics and Computing, 4(4):235–245, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00156746.

Talay:1994:PSP

[121] Denis Talay. Presto: a software package for the simulation of diffusion processes. Statistics and Computing, 4 (4):247–251, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00156747.

Lyu:1994:SAP

[122] J. Lyu and A. Gunasekaran. Statistical analysis of a portable parallel nonlinear programming algorithm. Statistics and Computing, 4(4):253–258, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00156748.

Fruhwirth-Schnatter:1994:ASS

[123] Sylvia Frühwirth-Schnatter. Applied state space modelling of non-Gaussian time series using integration-based Kalman filtering. Statistics and Computing, 4(4):259–269, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00156749.

Skouras:1994:ELM

[124] K. Skouras, C. Goutis, and M. J. Bramson. Estimation in linear models using gradient descent with early stopping. Statistics and Computing, 4(4): 271–278, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00156750.

Forster:1994:CMD

[125] Jonathan J. Forster and Allan M. Skene. Calculation of marginal densities for parameters of multinomial distributions. Statistics and Computing, 4 (4):279–286, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00156751.

Diaconis:1994:GCR

[126] Persi Diaconis and Susan Holmes. Gray codes for randomization procedures. Statistics and Computing, 4 (4):287–302, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00156752.

Zeh:1994:SR

[127] Judith E. Zeh. Software review. Statistics and Computing, 4(4):303–309, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00156753.

Kendall:1994:BR

[128] Wilfrid S. Kendall. Book reviews. Statistics and Computing, 4(4):311– 314, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00156754.

Anonymous:1994:HCd

[129] Anonymous. Help & contacts. Statistics and Computing, 4(4):??, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Anonymous:1995:SSD

[130] Anonymous. Statistical and scientific database management. Statistics and Computing, 5(1):1, March 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00140661.

French:1995:I

[131] James C. French. Introduction. Statistics and Computing, 5(1):2, March 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00140662.

Rafanelli:1995:ASD

[132] Maurizio Rafanelli. Aggregate statistical data: models for their representation. Statistics and Computing, 5(1):3–24, March 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00140663.

Olken:1995:RSD

[133] Frank Olken and Doron Rotem. Random sampling from databases: a survey. Statistics and Computing, 5 (1):25–42, March 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00140664.

Olken:1995:SSD

[134] Frank Olken and Doron Rotem. Sampling from spatial databases. Statistics and Computing, 5(1):43–57, March 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00140665.

Tansel:1995:QLS

[135] Abdullah Uz Tansel. Query languages for statistical databases. Statistics and Computing, 5(1):59–72, March 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00140666.

Klensin:1995:WME

[136] John C. Klensin. When the metadata exceed the data: data management with uncertain data. Statistics and Computing, 5(1):73–84, March 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00140667.

Anonymous:1995:HCa

[137] Anonymous. Help & contacts. Statistics and Computing, 5(1):??, March 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Nelder:1995:SLM

[138] J. A. Nelder. The statistics of linear models: back to basics. Statistics and Computing, 5(2):i, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00143933. See

[119].

Aitkin:1995:CJN

[139] Murray Aitkin. Comments on J. A. Nelder: ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):85–86, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF00143934. See

[119].

Lindsey:1995:ULL

[140] J. K. Lindsey. The uses and limits of linear models. Statistics and Computing, 5(2):87–89, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00143935.

Gower:1995:CJN

[141] John Gower. Comments on J. A. Nelder, ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):91–92, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF00143936. See

[119].

VanEeuwijk:1995:TDS

[142] F. A. Van Eeuwijk. On the tenability of distinctions: some comments on ‘The statistics of linear models: back to basics’ by John Nelder. Statistics and Computing, 5(2):93–95, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/BF00143937. See

[119].

Rodriguez:1995:CJN

[143] Robert Rodriguez, Randall Tobias, and Russell Wolfinger. Comments on J. A.

Nelder, ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):97–101, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00143938. See [119].

Searle:1995:CJN

[144] Shayle R. Searle. Comments on J. A. Nelder, ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):103–107, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00143939. See [119].

Nelder:1995:RCS

[145] J. A. Nelder. Rejoinder to comments on ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):109–111, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00143940. See [119].

Aitkin:1995:PMC

[146] Murray Aitkin. Probability model choice in single samples from exponential families using Poisson log-linear modelling, and model comparison using Bayes and posterior Bayes factors. Statistics and Computing, 5 (2):113–120, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00143941.

Robert:1995:STN

[147] Christian P. Robert. Simulation of truncated normal variables. Statistics and Computing, 5(2):121–125, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00143942.

Zoppe:1995:PPU

[148] Alice Zoppè. Principal points of univariate continuous distributions. Statistics and Computing, 5 (2):127–132, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00143943.

Dellaportas:1995:RVT

[149] Petros Dellaportas. Random variate transformations in the Gibbs sampler: issues of efficiency and convergence. Statistics and Computing, 5(2):133–140, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00143944.

VanEeuwijk:1995:UDB

[150] F. A. Van Eeuwijk and L. C. P. Keizer. On the use of diagnostic biplots in model screening for genotype by environment tables. Statistics and Computing, 5(2):141–153, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00143945.

Wang:1995:SVN

[151] Morgan C. Wang and William J. Kennedy. A self-validating numerical method for computation of central and non-central F probabilities and percentiles. Statistics and Computing, 5(2):155–163, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00143946.

Dalgleish:1995:SRB

[152] Lenard I. Dalgleish. Software review: Bootstrapping and jacknifing with BOJA. Statistics and Computing, 5(2):165–174, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00143947.

Anonymous:1995:HCb

[153] Anonymous. Help & contacts. Statistics and Computing, 5(2):??, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Hurn:1995:SSA

[154] Merrilee Hurn and Christopher Jennison. A study of simulated annealing and a revised cascade algorithm for image reconstruction. Statistics and Computing, 5(3):175–190, September 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00142660.

Denham:1995:IPL

[155] M. C. Denham. Implementing partial least squares. Statistics and Computing, 5(3):191–202, September 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00142661.

Luchian:1995:APT

[156] Henri Luchian and Daniel Stamate. Answer-perturbation techniques for the protection of statistical databases. Statistics and Computing, 5(3):203– 213, September 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00142662.

Dixon:1995:SPM

[157] Mark J. Dixon and Jonathan A. Tawn. A semi-parametric model for multivariate extreme values. Statistics and Computing, 5(3):215–225, September 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00142663.

Mertens:1995:ECV

[158] Bart Mertens, Tom Fearn, and Michael Thompson. The efficient crossvalidation of principal components applied to principal component regression. Statistics and Computing, 5(3): 227–235, September 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00142664.

Gauderman:1995:MSF

[159] W. James Gauderman. A method for simulating familial disease data with variable age at onset and genetic and environmental effects. Statistics and Computing, 5(3):237–243, September 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00142665.

Simonoff:1995:SAA

[160] Jeffrey S. Simonoff. A simple, automatic and adaptive bivariate density estimator based on conditional densities. Statistics and Computing, 5(3): 245–252, September 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00142666.

Smith:1995:ECT

[161] Peter W. F. Smith and John W. McDonald. Exact conditional tests for incomplete contingency tables: estimating attained significance levels. Statistics and Computing, 5(3): 253–256, September 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00142667.

Anonymous:1995:HCc

[162] Anonymous. Help & contacts. Statistics and Computing, 5(3):??, September 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Davison:1995:AEC

[163] A. C. Davison, D. V. Hinkley, and B. J. Worton. Accurate and efficient construction of bootstrap likelihoods. Statistics and Computing, 5 (4):257–264, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00162498.

Powell:1995:AWP

[164] K. J. Powell, T. Sapatinas, T. C. Bailey, and W. J. Krzanowski. Application of wavelets to the pre-processing of underwater sounds. Statistics and Computing, 5(4):265–273, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00162499.

Harbron:1995:HAF

[165] Chris Harbron. Heuristic algorithms for finding inexpensive elimination schemes. Statistics and Computing, 5 (4):275–287, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00162500.

Baker:1995:TPT

[166] Rose D. Baker. Two permutation tests of equality of variances. Statistics and Computing, 5(4):289–296, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00162501.

Dey:1995:BPA

[167] Dipak K. Dey, Lynn Kuo, and Sujit K. Sahu. A Bayesian predictive approach to determining the number of components in a mixture distribution. Statistics and Computing, 5(4):297–305, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00162502.

Davies:1995:TRS

[168] Neville Davies, Ed Dawson, Helen Gustafson, and A. N. Pettitt. Testing for randomness in stream ciphers using the binary derivative. Statistics and Computing, 5(4):307–310, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00162503.

Walker:1995:GRV

[169] Stephen Walker. Generating random variates from D-distributions via substitution sampling. Statistics and Computing, 5(4):311–315, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00162504.

Rotondi:1995:CMG

[170] Renata Rotondi and Silvia Drappo. A clustering method for global optimization based on the k-th nearest neighbour. Statistics and Computing, 5(4): 317–326, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00162505.

Goldstein:1995:BLC

[171] Michael Goldstein and David A. Wooff. Bayes linear computation: concepts, implementation and programs. Statistics and Computing, 5(4):327–341, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00162506.

Hazelton:1995:IMC

[172] Martin Hazelton. Improved Monte Carlo inference for models with additive error. Statistics and Computing, 5 (4):343–350, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00162507.

Anonymous:1995:HCd

[173] Anonymous. Help & contacts. Statistics and Computing, 5(4):??, December 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Hand:1996:E

[174] David J. Hand. Editorial. Statistics and Computing, 6(1):1, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00161568.

Ostermark:1996:STC

[175] Ralf Östermark. Separating trend and cyclical dynamics in state space models with exogenous inputs. Statistics and Computing, 6(1):3–10, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00161569.

Froeschl:1996:MAS

[176] K. A. Froeschl. A metadata approach to statistical query processing. Statistics and Computing, 6(1):11–29, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00161570.

Eichenauer-Herrmann:1996:MEI

[177] Jürgen Eichenauer-Herrmann. Modified explicit inversive congruential pseudorandom numbers with power of 2 modulus. Statistics and Computing, 6(1):31–36, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00161571.

Adams:1996:RPP

[178] N. M. Adams, S. P. J. Kirby, P. Harris, and D. B. Clegg. A review of parallel processing for statistical computation. Statistics and Computing, 6(1):37–49, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00161572.

Krzanowski:1996:SRS

[179] W. J. Krzanowski. A stopping rule for structure-preserving variable selection. Statistics and Computing, 6(1):51–56, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00161573.

Rigby:1996:SPA

[180] R. A. Rigby and D. M. Stasinopoulos. A semi-parametric additive model for variance heterogeneity. Statistics and Computing, 6(1):57–65, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00161574.

Bradford:1996:MCM

[181] Russell Bradford and Alun Thomas. Markov chain Monte Carlo methods for family trees using a parallel processor. Statistics and Computing, 6(1):67–75, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00161575.

Laud:1996:MCM

[182] P. W. Laud, A. F. M. Smith, and P. Damien. Monte Carlo methods for approximating a posterior hazard rate process. Statistics and Computing, 6(1):77–83, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00161576.

Lawrence:1996:MSM

[183] C. J. Lawrence and W. J. Krzanowski. Mixture separation for mixed-mode data. Statistics and Computing, 6 (1):85–92, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00161577.

Anonymous:1996:HCa

[184] Anonymous. Help & contacts. Statistics and Computing, 6(1):??, March 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Ogden:1996:CPA

[185] Todd Ogden and Emanuel Parzen. Change-point approach to data analytic wavelet thresholding. Statistics and Computing, 6(2):93–99, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00162519.

Cowles:1996:AMC

[186] Mary Kathryn Cowles. Accelerating Monte Carlo Markov chain convergence for cumulative-link generalized linear models. Statistics and Computing, 6(2):101–111, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00162520.

Liu:1996:MIS

[187] Jun S. Liu. Metropolized independent sampling with comparisons to rejection sampling and importance sampling. Statistics and Computing, 6 (2):113–119, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00162521.

Aitkin:1996:NTP

[188] Murray Aitkin, Steve Finch, Nancy Mendell, and Henry Thode. A new test for the presence of a normal mixture distribution based on the posterior Bayes factor. Statistics and Computing, 6(2):121–125, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00162522.

Aitkin:1996:HEG

[189] Murray Aitkin and Irit Aitkin. A hybrid EM/Gauss–Newton algorithm for maximum likelihood in mixture distributions. Statistics and Computing, 6(2):127–130, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00162523.

Barrett:1996:CDS

[190] Bruce E. Barrett and J. Brian Gray. Computation of determinantal subset influence in regression. Statistics and Computing, 6(2):131–138, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00162524.

Modarres:1996:BPG

[191] Reza Modarres. Bootstrap power of the generalized correlation coefficient. Statistics and Computing, 6 (2):139–145, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00162525.

Hesterberg:1996:CVI

[192] Tim Hesterberg. Control variates and importance sampling for efficient bootstrap simulations. Statistics and Computing, 6(2):147–157, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00162526.

Webb:1996:ANL

[193] Andrew R. Webb. An approach to non-linear principal components analysis using radially symmetric kernel functions. Statistics and Computing, 6(2):159–168, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00162527.

Malvestuto:1996:TIH

[194] F. M. Malvestuto. Testing implication of hierarchical log-linear models for probability distributions. Statistics and Computing, 6(2):169–176, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00162528.

Krzanowski:1996:ECV

[195] W. J. Krzanowski. Efficient crossvalidation of principal components. Statistics and Computing, 6(2):177, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF00162529.

Mertens:1996:ECV

[196] B. J. A. Mertens, T. Fearn, and M. Thompson. Efficient crossvalidation of principal components applied to principal component regression. Statistics and Computing, 6(2): 178, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00162530.

Jones:1996:BR

[197] M. C. Jones, Catherine B. Hurley, and Telba Z. Irony. Book reviews. Statistics and Computing, 6(2):179–184, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00162531.

Anonymous:1996:E

[198] Anonymous.  Erratum.  Statis-

tics and Computing, 6(2):185, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00162532.

Anonymous:1996:HCb

[199] Anonymous. Help & contacts. Statistics and Computing, 6(2):??, June 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Rowe:1996:ACP

[200] Stephen Rowe. An algorithm for computing principal points with respect to a loss function in the unidimensional case. Statistics and Computing, 6(3):187–190, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00140863.

Zelterman:1996:BTP

[201] Daniel Zelterman, Chap T. Le, and Thomas A. Louis. Bootstrap techniques for proportional hazards models with censored observations. Statistics and Computing, 6(3):191–199, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00140864.

Smyth:1996:PAM

[202] Gordon K. Smyth. Partitioned algorithms for maximum likelihood and other non-linear estimation. Statistics and Computing, 6(3):201–216, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00140865.

Mirkin:1996:CCT

[203] Boris Mirkin. Clustering for contingency tables: boxes and partitions. Statistics and Computing, 6(3): 217–229, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00140866.

Keprta:1996:NBC

[204] Stanislav Keprta. Non-binary classification trees. Statistics and Computing, 6(3):231–243, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/BF00140867.

Cools:1996:REI

[205] Ronald Cools and Petros Dellaportas. The role of embedded integration rules in Bayesian statistics. Statistics and Computing, 6(3):245–250, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00140868.

Aitkin:1996:GML

[206] Murray Aitkin. A general maximum likelihood analysis of overdispersion in generalized linear models. Statistics and Computing, 6(3):251–262, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00140869.

deMoraes:1996:IDP

[207] Aipore R. de Moraes and Ian R. Dunsmore. Influential data points in predictive logistic models. Statistics and Computing, 6(3):263–268, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00140870.

Rosenthal:1996:AGS

[208] Jeffrey S. Rosenthal. Analysis of the Gibbs sampler for a model related to James–Stein estimators. Statistics and Computing, 6(3):269–275, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00140871.

Vounatsou:1996:BAC

[209] P. Vounatsou and A. F. M. Smith. Bayesian analysis of contingency tables: a simulation and graphics-based approach. Statistics and Computing, 6 (3):277–287, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00140872.

Pinheiro:1996:UPV

[210] José C. Pinheiro and Douglas M. Bates. Unconstrained parametrizations for variance-covariance matrices. Statistics and Computing, 6(3):289–296, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00140873.

Roverato:1996:SEP

[211] Alerto Roverato and Joe Whittaker. Standard errors for the parameters of graphical Gaussian models. Statistics and Computing, 6(3):297–302, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00140874.

Craig:1996:PLE

[212] Peter S. Craig and Allan H. Seheult. Pseudo-likelihood estimation for a class of spatial Markov chains. Statistics and Computing, 6(3):303–311, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00140875.

Anonymous:1996:HCc

[213] Anonymous. Help & contacts. Statistics and Computing, 6(3):??, September 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Wedelin:1996:EEM

[214] Dag Wedelin. Efficient estimation and model selection in large graphical models. Statistics and Computing, 6(4): 313–323, December 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00143552.

Lin:1996:CPA

[215] Xun-Guo Lin and Alun Pope. Coverage plots for assessing the variability of estimated contours of a density. Statistics and Computing, 6(4):325–336, December 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00143553.

Vines:1996:FBM

[216] S. K. Vines, W. R. Gilks, and P. Wild. Fitting Bayesian multiple random effects models. Statistics and Computing, 6(4):337–346, December 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00143554.

Doallo-Biempica:1996:EVP

[217] R. Doallo-Biempica, B. B. Fraguela-Rodríguez, and A. Quintela-Del-Río. Evaluation of vectorization/ parallelization techniques: application to nonparametric curve estimation. Statistics and Computing, 6(4): 347–351, December 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00143555.

Neal:1996:SMD

[218] Radford M. Neal. Sampling from multimodal distributions using tempered transitions. Statistics and Computing, 6(4):353–366, December 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00143556.

Chan:1996:RAD

[219] Karen Chan and Alison Gray. Robustness of automated data choices of smoothing parameter in image regularization. Statistics and Computing, 6 (4):367–377, December 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00143557.

Jeffers:1996:SR

[220] J. N. R. Jeffers. Software review. Statistics and Computing, 6(4): 379–386, December 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00143558.

Singpurwalla:1996:BR

[221] Nozer D. Singpurwalla, Robert Gentleman, Willem J. Heiser, and Dibyen Majumdar. Book reviews. Statistics and Computing, 6(4):387–391, December 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00143559.

Anonymous:1996:HCd

[222] Anonymous. Help & contacts. Statistics and Computing, 6(4):??, December 1996. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Bensmail:1997:IMB

[223] Halima Bensmail, Gilles Celeux, Adrian E. Raftery, and Christian P. Robert. Inference in model-based cluster analysis. Statistics and Computing, 7(1):1–10, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1018510926151.

Cheng:1997:SLL

[224] Ming-Yen Cheng, Peter Hall, and D. M. Titterington. On the shrinkage of local linear curve estimators. Statistics and Computing, 7(1):11– 17, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1018549127542.

Larranaga:1997:DBN

[225] Pedro Larrañaga, Cindy M. H. Kuijpers, Mikel Poza, and Roberto H. Murga. Decomposing Bayesian networks: triangulation of the moral graph with genetic algorithms. Statistics and Computing, 7(1):19–34, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1018553211613.

Hurn:1997:DUA

[226] Merrilee Hurn. Difficulties in the use of auxiliary variables in Markov chain Monte Carlo methods. Statistics and Computing, 7(1):35–44, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1018505328451.

Brodley:1997:ACA

[227] Carla E. Brodley and Padhraic Smyth. Applying classification algorithms in practice. Statistics and Computing, 7(1):45–56, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1018557312521.

Gamerman:1997:SPD

[228] Dani Gamerman. Sampling from the posterior distribution in generalized linear mixed models. Statistics and Computing, 7(1):57–68, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1018509429360.

Webb:1997:ROO

[229] Andrew Webb, Paul R. Yarnold, and Robert C. Soltysik. Review of ODA 1.0 optimal data analysis for DOSTM. Statistics and Computing, 7(1):69–73, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1018561413430.

Thioulouse:1997:AMA

[230] Jean Thioulouse, Daniel Chessel, Sylvain Dole’dec, and Jean-Michel Olivier. ADE-4: a multivariate analysis and graphical display software. Statistics and Computing, 7 (1):75–83, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1018513530268.

Anonymous:1997:HCa

[231] Anonymous. Help & contacts. Statistics and Computing, 7(1):??, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Krzanowski:1997:RTD

[232] W. J. Krzanowski. Recent trends and developments in computational multivariate analysis. Statistics and Computing, 7(2):87–99, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1018565514339.

Malvestuto:1997:SMT

[233] F. M. Malvestuto and M. Moscarini. Suppressing marginal totals from a two-dimensional table to protect sensitive information. Statistics and Computing, 7(2):101–114, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1018517631177.

Hall:1997:NPB

[234] Peter Hall, Spiridon Penev, Gérard Kerkyacharian, and Dominique Picard. Numerical performance of block thresholded wavelet estimators. Statistics and Computing, 7(2):115– 124, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1018569615247.

Gustafson:1997:ASM

[235] H. M. Gustafson, E. P. Dawson, and J. Dj. Golić. Automated statistical methods for measuring the strength of block ciphers. Statistics and Computing, 7(2):125–135, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1018521732085.

Mato:1997:SCV

[236] A. Silva Mato and A. Martín Andrés. Simplifying the calculation of the P - value for Barnard’s test and its derivatives. Statistics and Computing, 7 (2):137–143, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1018573716156.

Oskrochi:1997:ETA

[237] G. R. Oskrochi and R. B. Davies. An EM-type algorithm for multivariate mixture models. Statistics and Computing, 7(2):145–151, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1018525800226.

Carlin:1997:BEB

[238] Bradley P. Carlin and Thomas A. Louis. Bayes and empirical Bayes methods for data analysis. Statistics and Computing, 7(2):153–154, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1018577817064.

Anonymous:1997:HCb

[239] Anonymous. Help & contacts. Statistics and Computing, 7(2):??, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Cox:1997:RLD

[240] Trevor F. Cox and Kim F. Pearce. A robust logistic discrimination model.

Statistics and Computing, 7(3):155– 161, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1018530001135.

Wang:1997:ANI

[241] Ouhong Wang and William J. Kennedy. Application of numerical interval analysis to obtain self-validating results for multivariate probabilities in a massively parallel environment. Statistics and Computing, 7(3):163–171, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1018582017973.

Philippe:1997:SRL

[242] Anne Philippe. Simulation of right and left truncated gamma distributions by mixtures. Statistics and Computing, 7(3):173–181, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1018534102043.

Chu:1997:SBA

[243] Hui-May Chu and Lynn Kuo. Sampling based approach for one-hit and multi-hit models in quantal bioassay. Statistics and Computing, 7 (3):183–192, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1018586118882.

Franconi:1997:CGA

[244] Luisa Franconi and Christopher Jennison. Comparison of a genetic algorithm and simulated annealing in an application to statistical image reconstruction. Statistics and Computing, 7(3):193–207, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1018538202952.

Mola:1997:FSP

[245] Francesco Mola and Roberta Siciliano. A fast splitting procedure for classification trees. Statistics and Computing, 7(3):209–216, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1018590219790.

Anonymous:1997:HCc

[246] Anonymous. Help & contacts. Statistics and Computing, 7(3):??, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Ostland:1997:EQM

[247] M. Ostland and B. Yu. Exploring quasi Monte Carlo for marginal density approximation. Statistics and Computing, 7(4):217–228, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1018542303861.

Lavielle:1997:SAV

[248] M. Lavielle and E. Moulines. A simulated annealing version of the EM algorithm for non-Gaussian deconvolution. Statistics and Computing, 7 (4):229–236, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1018594320699.

Bowman:1997:BBI

[249] Adrian Bowman, James Currall, and Richard Lyall. The birds and the bees: interactive graphics and problem solving in the teaching of statistics. Statistics and Computing, 7 (4):237–246, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1018546404769.

Dempster:1997:DUL

[250] A. P. Dempster. The direct use of likelihood for significance testing. Statistics and Computing, 7 (4):247–252, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1018598421607.

Aitkin:1997:CVP

[251] Murray Aitkin. The calibration of P - values, posterior Bayes factors and the AIC from the posterior distribution of the likelihood. Statistics and Computing, 7(4):253–261, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1018550505678. See com-

mentary [253, 252].

Stone:1997:DPD

[252] M. Stone. Discussion of papers by Dempster and Aitkin. Statistics and Computing, 7(4):263–264, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1018502622516. See [251, 253].

Dempster:1997:CPM

[253] A. P. Dempster. Commentary on the paper by Murray Aitkin, and on discussion by Mervyn Stone. Statistics and Computing, 7(4):265–269, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1018554606586. See [251, 252].

Aitkin:1997:R

[254] Murray Aitkin. Reply. Statistics and Computing, 7(4):271–272, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1018506723425.

Anonymous:1997:IV

[255] Anonymous. Indexes (volume 7, 1997). Statistics and Computing, 7 (4):273–278, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1018558707495.

Anonymous:1997:HCd

[256] Anonymous. Help & contacts. Statistics and Computing, 7(4):??, ???? 1997. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Anthony:1998:PGF

[257] Martin Anthony. Probabilistic ‘generalization’ of functions and dimensionbased uniform convergence results.

Statistics and Computing, 8(1):5–14, ???? 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008801224333.

Mackay:1998:IMM

[258] David J. C. Mackay and Ryo Takeuchi. Interpolation models with multiple hyperparameters. Statistics and Computing, 8(1):15–23, ???? 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008862908404.

Tibshirani:1998:CVR

[259] Robert Tibshirani and Geoffrey Hinton. Coaching variables for regression and classification. Statistics and Computing, 8(1):25–33, ???? 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008815025242.

Wolpert:1998:SRC

[260] David H. Wolpert, Emanuel Knill, and Tal Grossman. Some results concerning off-training-set and IID error for the Gibbs and the Bayes optimal generalizers. Statistics and Computing, 8(1):35–54, ???? 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008867009312.

Mace:1998:SMA

[261] C. W. H. Mace and A. C. C. Coolen. Statistical mechanical analysis of the dynamics of learning in perceptrons.

Statistics and Computing, 8(1):55– 88, ???? 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008896910704.

Anonymous:1998:HCa

[262] Anonymous. Help & contacts. Statistics and Computing, 8(1):??, ???? 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Cowell:1998:MRP

[263] Robert G. Cowell. Mixture reduction via predictive scores. Statistics and Computing, 8(2):97–103, June 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008921815757.

Coles:1998:ICD

[264] Stuart Coles. Inference for circular distributions and processes. Statistics and Computing, 8(2):105– 113, June 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008930032595.

Cowles:1998:SAC

[265] Mary Kathryn Cowles and Jeffrey S. Rosenthal. A simulation approach to convergence rates for Markov chain Monte Carlo algorithms. Statistics and Computing, 8(2):115–124, June 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008982016666.

Chu:1998:DEA

[266] Moody T. Chu and Nickolay T. Trendafilov. On a differential equation approach to the weighted orthogonal Procrustes problem. Statistics and Computing, 8(2):125–133, June 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008934100736.

Settimi:1998:CSM

[267] R. Settimi and P. G. Blackwell. Conditional simulation for moving average processes with discrete or continuous values. Statistics and Computing, 8(2):135–144, June 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008986117574.

Robert:1998:RSH

[268] Christian P. Robert and D. M. Titterington. Reparameterization strategies for hidden Markov models and Bayesian approaches to maximum likelihood estimation. Statistics and Computing, 8(2):145–158, June 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008938201645.

Nilsson:1998:EAF

[269] D. Nilsson. An efficient algorithm for finding the M most probable configurations in probabilistic expert systems. Statistics and Computing, 8 (2):159–173, June 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1008990218483.

Anonymous:1998:HCb

[270] Anonymous. Help & contacts. Statistics and Computing, 8(2):??, June 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Solka:1998:MSA

[271] Jeffrey L. Solka, Edward J. Wegman, Carey E. Priebe, Wendy L. Poston, and George W. Rogers. Mixture structure analysis using the Akaike information criterion and the bootstrap. Statistics and Computing, 8 (3):177–188, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1008924323509.

Andrews:1998:IFP

[272] David F. Andrews and James E. Stafford. Iterated full partitions. Statistics and Computing, 8(3):189– 192, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008940924418.

Rousseeuw:1998:CLD

[273] Peter J. Rousseeuw and Anja Struyf. Computing location depth and regression depth in higher dimensions. Statistics and Computing, 8(3):193– 203, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008945009397.

Walden:1998:MPU

[274] A. T. Walden and A. Contreras Cristan. Matching pursuit by undecimated discrete wavelet transform for non-stationary time series of arbitrary length. Statistics and Computing, 8 (3):205–219, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008901226235.

Rue:1998:IPD

[275] Håvard Rue and Oddvar K. Husby. Identification of partly destroyed objects using deformable templates. Statistics and Computing, 8(3):221– 228, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008953210305.

Nobile:1998:HMC

[276] Agostino Nobile. A hybrid Markov chain for the Bayesian analysis of the multinomial probit model. Statistics and Computing, 8(3):229–242, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008905311214.

Jensen:1998:SMF

[277] Claus Skaanning Jensen. A simple method for finding a legal configuration in complex Bayesian networks. Statistics and Computing, 8 (3):243–251, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1008909428052.

Silver:1998:OSI

[278] Edward A. Silver, Daniel Costa, and Willard Zangwill. Order statistics of independent identically distributed variables when the sum is known. Statistics and Computing, 8(3):253– 265, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008961428961.

Brooks:1998:QCA

[279] S. P. Brooks. Quantitative convergence assessment for Markov chain Monte Carlo via cusums. Statistics and Computing, 8(3):267–274, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008965613031.

Yu:1998:LMS

[280] Bin Yu and Per Mykland. Looking at Markov samplers through cusum path plots: a simple diagnostic idea. Statistics and Computing, 8(3):275– 286, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008917713940.

Anonymous:1998:HCc

[281] Anonymous. Help & contacts. Statistics and Computing, 8(3):??, August 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Aitkin:1998:RMB

[282] Murray Aitkin and Marco Alfó. Regression models for binary longitudinal responses. Statistics and Computing, 8 (4):289–307, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008847820371.

Benn:1998:MPC

[283] A. Benn and R. Kulperger. Massively parallel computing: a statistical application. Statistics and Computing, 8 (4):309–318, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008868404442.

Brooks:1998:CAT

[284] Stephen P. Brooks and Gareth O. Roberts. Convergence assessment techniques for Markov chain Monte Carlo. Statistics and Computing, 8 (4):319–335, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008820505350.

Denison:1998:BM

[285] D. G. T. Denison, B. K. Mallick, and A. F. M. Smith. Bayesian MARS. Statistics and Computing, 8 (4):337–346, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008824606259.

Foster:1998:EMD

[286] Peter Foster. Exploring multivariate data using directions of high density. Statistics and Computing, 8(4): 347–355, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008828723097.

Gustafson:1998:GWM

[287] Paul Gustafson. A guided walk Metropolis algorithm. Statistics and Computing, 8(4):357–364, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008880707168.

Kennedy:1998:BQN

[288] Marc Kennedy. Bayesian quadrature with non-normal approximating functions. Statistics and Computing, 8(4): 365–375, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008832824006.

Luan:1998:SDN

[289] Jian’an Luan, Julian Stander, and David Wright. On shape detection in noisy images with particular reference to ultrasonography. Statistics and Computing, 8(4):377–389, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008884808076.

Mcdonald:1998:ETT

[290] John W. Mcdonald, David C. Deroure, and Danius T. Michaelides. Exact tests for two-way symmetric contingency tables. Statistics and Computing, 8(4): 391–399, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008841109893.

Anonymous:1998:FDA

[291] Anonymous. Functional data analysis. Statistics and Computing, 8(4): 401–403, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017116429458.

Anonymous:1998:SR

[292] Anonymous. Software review. Statistics and Computing, 8(4):405–406, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1017109630366.

Anonymous:1998:HCd

[293] Anonymous. Help & contacts. Statistics and Computing, 8(4):??, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Busing:1999:DJU

[294] Frank M. T. A. Busing, Erik Meijer, and Rien Van Der Leeden. Delete-m jackknife for unequal m. Statistics and Computing, 9(1):3– 8, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008800423698.

Canty:1999:ISA

[295] Angelo J. Canty and A. C. Davison. Implementation of saddlepoint approximations in resampling problems. Statistics and Computing, 9(1): 9–15, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008801807768.

Chib:1999:MSH

[296] Siddhartha Chib and Bradley P. Carlin. On MCMC sampling in hierarchical longitudinal models. Statistics and Computing, 9(1):17–26, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008853808677.

DeBeer:1999:SEN

[297] C. F. De Beer and J. W. H. Swanepoel. Simple and effective number-of-bins circumference selectors for a histogram. Statistics and Computing, 9(1):27– 35, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008858025515.

Eccleston:1999:DOC

[298] J. Eccleston and D. Whitaker. On the design of optimal change-over experiments through multi-objective simulated annealing. Statistics and Computing, 9(1):37–42, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008810109585.

Faghihi:1999:PSA

[299] Mohammed Reza Faghihi, Charles C. Taylor, and Ian L. Dryden. Procrustes shape analysis of triangulations of a two coloured point pattern. Statistics and Computing, 9(1): 43–53, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008862126424.

Sahu:1999:CEA

[300] Sujit K. Sahu and Gareth O. Roberts. On convergence of the EM algorithm and the Gibbs sampler. Statistics and Computing, 9(1):55– 64, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008814227332.

Sardy:1999:WSU

[301] Sylvain Sardy, Donald B. Percival, Andrew G. Bruce, Hong-Ye Gao, and Werner Stuetzle. Wavelet shrinkage for unequally spaced data. Statistics and Computing, 9(1):65– 75, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008818328241.

Weir:1999:SMB

[302] I. S. Weir and A. N. Pettitt. Spatial modelling for binary data using a hidden conditional autoregressive Gaussian process: a multivariate extension of the probit model. Statistics and Computing, 9(1):77– 86, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008822429149.

Anonymous:1999:HCa

[303] Anonymous. Help & contacts. Statistics and Computing, 9(1):??, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Kuk:1999:PFA

[304] Anthony Y. C. Kuk and Yuk W. Cheng. Pointwise and functional approximations in Monte Carlo maximum likelihood estimation. Statistics and Computing, 9(2):91–99, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008800715000.

Marron:1999:ILB

[305] J. S. Marron and F. Udina. Interactive local bandwidth choice. Statistics and Computing, 9(2):101–110, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008890315909.

Willse:1999:IFM

[306] Alan Willse and Robert J. Boik. Identifiable finite mixtures of location models for clustering mixed-mode data. Statistics and Computing, 9 (2):111–121, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008842432747.

Anonymous:1999:JHF

[307] Anonymous. Jerome H. Friedman and Nicholas I. Fisher. Statistics and Computing, 9(2):123–143, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1017121607248.

Friedman:1999:BHH

[308] Jerome H. Friedman and Nicholas I. Fisher. Bump hunting in highdimensional data. Statistics and Computing, 9(2):123–143, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1023/A%3A1008894516817.

See discussion [309, 310, 311, 312, 313, 314, 315, 316, 317, 318, 319, 320].

Kloesgen:1999:DPF

[309] Willi Kloesgen. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):143–144, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1008846600887. See [308].

Huber:1999:DPF

[310] Peter J. Huber. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):144– 146, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008898617726. See [308].

Piatetsky-Shapiro:1999:DPF

[311] Gregory Piatetsky-Shapiro. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):146, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1008802818634. See [308].

Scott:1999:DPF

[312] David W. Scott. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):146–147, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1008854802705. See [308].

Feelders:1999:DPF

[313] A. J. Feelders. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):147–148, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1008806919543. See [308].

Titterington:1999:DPF

[314] D. M. Titterington. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):148–149, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1008811020452. See [308].

Smyth:1999:DPF

[315] Padhraic Smyth. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):149–150, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1008863004522. See [308].

Ridgeway:1999:DPF

[316] Greg Ridgeway, Thomas Richardson, and David Madigan. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9 (2):150–152, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008815121360. See [308].

Stone:1999:DPF

[317] Glenn Stone. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):152–153, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1008867105431. See [308].

Burges:1999:DPF

[318] Chris Burges. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):153–154, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1008819222269. See [308].

Jurgens:1999:DPF

[319] Marcus Jürgens and Hans-J. Lenz. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):154–155, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1023/A%3A1008871206339.

See [308].

Lunn:1999:DPF

[320] Daniel Lunn. Discussion on the paper by Friedman and Fisher. Statistics and Computing, 9(2):155–156, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1008823323177. See [308].

Anonymous:1999:HCb

[321] Anonymous. Help & contacts. Statistics and Computing, 9(2):??, April 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Hand:1999:E

[322] David J. Hand. Editorial. Statistics and Computing, 9(3):167, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1008985627239.

Eccleston:1999:GEC

[323] John Eccleston. Guest editorial: Computational issues in experimental design. Statistics and Computing, 9(3): 169, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1008957611309.

Burgess:1999:IAF

[324] Leonie Burgess and Deborah J. Street. An interchange algorithm for four factor orthogonal main effect plans. Statistics and Computing, 9(3):171– 177, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008909728147.

Dean:1999:SME

[325] A. M. Dean and N. R. Draper. Saturated main-effect designs for factorial experiments. Statistics and Computing, 9(3):179–185, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008961712217.

Delgado:1999:SOD

[326] Jaime Delgado and Hari Iyer. Search for optimal designs in a three stagenested random model. Statistics and Computing, 9(3):187–193, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008913829056.

Elliott:1999:ADF

[327] L. J. Elliott, J. A. Eccleston, and R. J. Martin. An algorithm for the design of factorial experiments when the data are correlated. Statistics and Computing, 9(3):195–201, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1008965829964.

John:1999:PLD

[328] J. A. John and E. R. Williams. Partially-latinized designs. Statistics and Computing, 9(3):203–207, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008969914035.

Jones:1999:COD

[329] B. Jones and J. Wang. Constructing optimal designs for fitting pharmacokinetic models. Statistics and Computing, 9(3):209–218, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1008922030873.

Low:1999:ARC

[330] J. L. Low, S. M. Lewis, and P. Prescott. Assessing robustness of crossover designs to subjects dropping out. Statistics and Computing, 9(3):219– 227, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008974031782.

Martin:1999:EDC

[331] R. J. Martin, M. C. Bursnall, and E. C. Stillman. Efficient designs for constrained mixture experiments. Statistics and Computing, 9(3):229– 237, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008978116760.

Russell:1999:CSM

[332] K. G. Russell. A comparison of six methods of analysing row-column designs with inter-block information. Statistics and Computing, 9(3):239– 246, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008934318578.

Anonymous:1999:HCc

[333] Anonymous. Help & contacts. Statistics and Computing, 9(3):??, July 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Atkinson:1999:CLT

[334] A. C. Atkinson and T.-C. Cheng. Computing least trimmed squares regression with the forward search. Statistics and Computing, 9(4):251–263, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008942604045.

Chan:1999:SSG

[335] G. Chan and A. T. A. Wood. Simulation of stationary Gaussian vector fields. Statistics and Computing, 9(4): 265–268, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008903804954.

Edgoose:1999:MMC

[336] T. Edgoose and L. Allison. MML Markov classification of sequential data. Statistics and Computing, 9(4): 269–278, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008907921792.

Gustafson:1999:MTS

[337] H. M. Gustafson, E. P. Dawson, J. Dj. Golić, and A. N. Pettitt. Methods for testing subblock patterns. Statistics and Computing, 9(4): 279–286, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008959905863.

Hobert:1999:PSS

[338] James P. Hobert, Christian P. Robert, and D. M. Titterington. On perfect simulation for some mixtures of distributions. Statistics and Computing, 9 (4):287–298, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008912006771.

Jorgensen:1999:DEA

[339] Murray Jorgensen. A dynamic EM algorithm for estimating mixture proportions. Statistics and Computing, 9 (4):299–302, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008916123610.

Karlis:1999:IEA

[340] Dimitris Karlis and Evdokia Xekalaki. Improving the EM algorithm for mixtures. Statistics and Computing, 9(4):

303–307, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008968107680.

Shih:1999:FSC

[341] Y.-S. Shih. Families of splitting criteria for classification trees. Statistics and Computing, 9(4):309–315, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008920224518.

Lumley:1999:RSS

[342] Thomas Lumley. Review of the Stata statistical package. Statistics and Computing, 9(4):317–318, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1008972208589.

Anonymous:1999:HCd

[343] Anonymous. Help & contacts. Statistics and Computing, 9(4):??, November 1999. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Richardson:2000:GE

[344] Thomas S. Richardson. Guest editorial. Statistics and Computing, 10(1):3–4, January 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1008987631331.

Baxter:2000:FOC

[345] Rohan A. Baxter and Jonathan J. Oliver. Finding overlapping components with MML. Statistics and Computing, 10(1):5–16, January 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008928315401.

Chipman:2000:HPB

[346] Hugh Chipman and Robert E. Mc-Culloch. Hierarchical priors for Bayesian CART shrinkage. Statistics and Computing, 10(1):17–24, January 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008980332240.

Jaakkola:2000:BPE

[347] Tommi S. Jaakkola and Michael I. Jordan. Bayesian parameter estimation via variational methods. Statistics and Computing, 10(1):25–37, January 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008932416310.

Kontkanen:2000:PDB

[348] P. Kontkanen, P. Myllymäki, T. Silander, H. Tirri, and P. Grünwald. On predictive distributions and Bayesian networks. Statistics and Computing, 10(1):39–54, January 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008984400380.

Chickering:2000:CSE

[349] David Maxwell Chickering and David Heckerman. A comparison of scientific and engineering criteria for Bayesian model selection. Statistics and Computing, 10(1):55–62, January 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008936501289.

Smyth:2000:MSP

[350] Padhraic Smyth. Model selection for probabilistic clustering using cross-validated likelihood. Statistics and Computing, 10(1):63–72, January 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008940618127.

Wallace:2000:MCM

[351] Chris S. Wallace and David L. Dowe. MML clustering of multistate, Poisson, von Mises circular and Gaussian distributions. Statistics and Computing, 10(1):73–83, January 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008992619036.

Anonymous:2000:HCa

[352] Anonymous. Help & contacts. Statistics and Computing, 10(1):??, January 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Berman:2000:GEI

[353] Mark Berman. Guest editorial: Image analysis. Statistics and Computing, 10(2):91–93, April 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008996708002.

Murtagh:2000:IPT

[354] F. Murtagh and J.-L. Starck. Image processing through multiscale analysis and measurement noise modeling. Statistics and Computing, 10 (2):95–103, April 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008938224840.

Breen:2000:MMU

[355] Edmond J. Breen, Ronald Jones, and Hugues Talbot. Mathematical morphology: a useful set of tools for image analysis. Statistics and Computing, 10(2):105–120, April 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008990208911.

Jeulin:2000:RTM

[356] Dominique Jeulin. Random texture models for material structures. Statistics and Computing, 10(2):121– 132, April 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008942325749.

Kuttikkad:2000:SMA

[357] Shyam Kuttikkad and Rama Chellappa. Statistical modeling and analysis of high-resolution synthetic aperture radar images. Statistics and Computing, 10(2):133–145, April 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008994309819.

Leahy:2000:SAQ

[358] Richard M. Leahy and Jinyi Qi. Statistical approaches in quantitative positron emission tomography. Statistics and Computing, 10(2):147– 165, April 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008946426658.

Jain:2000:ORR

[359] Anil K. Jain and Chitra Dorai. 3D object recognition: Representation and matching. Statistics and Computing, 10(2):167–182, April 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008998410728.

Anonymous:2000:HCb

[360] Anonymous. Help & contacts. Statistics and Computing, 10(2):??, April 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Chan:2000:WSS

[361] Karen Chan, Andrea Saltelli, and Stefano Tarantola. Winding stairs: a sampling tool to compute sensitivity indices. Statistics and Computing, 10(3):187–196, July 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008950625967.

Doucet:2000:SMC

[362] Arnaud Doucet, Simon Godsill, and Christophe Andrieu. On sequential Monte Carlo sampling methods for Bayesian filtering. Statistics and Computing, 10(3):197–208, July 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008935410038.

Jonathan:2000:UCV

[363] P. Jonathan, W. J. Krzanowski, and W. V. McCarthy. On the use of cross-validation to assess performance in multivariate prediction. Statistics and Computing, 10(3):209– 229, July 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008987426876.

Kent:2000:GOG

[364] J. T. Kent and M. Mohammadzadeh. Global optimization of the generalized cross-validation criterion. Statistics and Computing, 10(3):231–236, July 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008939510946.

Murdoch:2000:EUE

[365] Duncan J. Murdoch and Jeffrey S. Rosenthal. Efficient use of exact samples. Statistics and Computing, 10(3):237–243, July 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008991527785.

Read:2000:EMW

[366] Robert Read, Lyn Thomas, and Alan Washburn. Estimating means when sampling gives probabilities as well as values or “Looking a gift horse in the mouth”. Statistics and Computing, 10(3):245–252, July 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008995628693.

Shipley:2000:PPT

[367] Bill Shipley. A permutation procedure for testing the equality of pattern hypotheses across groups involving correlation or covariance matrices. Statistics and Computing, 10 (3):253–257, July 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008943611855.

Thomas:2000:MLA

[368] Alun Thomas, Alexander Gutin, Victor Abkevich, and Aruna Bansal. Multilocus linkage analysis by blocked Gibbs sampling. Statistics and Computing, 10(3):259–269, July 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008947712763.

Anonymous:2000:HCc

[369] Anonymous. Help & contacts. Statistics and Computing, 10(3):??, July 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Alfo:2000:RCM

[370] Marco Alfò and Murray Aitkin. Random coefficient models for binary longitudinal responses with attrition. Statistics and Computing, 10(4):279–287, October 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008999824193.

Asparoukhov:2000:NPS

[371] O. Asparoukhov and W. J. Krzanowski. Non-parametric smoothing of the location model in mixed variable discrimination. Statistics and Computing, 10 (4):289–297, October 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008973308264.

Brewer:2000:BML

[372] Mark J. Brewer. A Bayesian model for local smoothing in kernel density estimation. Statistics and Computing, 10 (4):299–309, October 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008925425102.

Goldstein:2000:BLA

[373] M. Goldstein and D. J. Wilkinson. Bayes linear analysis for graphical models: The geometric approach to local computation and interpretive graphics. Statistics and Computing, 10 (4):311–324, October 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008977409172.

Lunn:2000:WBM

[374] David J. Lunn, Andrew Thomas, Nicky Best, and David Spiegelhalter. Win-BUGS — a Bayesian modelling framework: Concepts, structure, and extensibility. Statistics and Computing, 10 (4):325–337, October 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008929526011.

Peel:2000:RMM

[375] D. Peel and G. J. McLachlan. Robust mixture modelling using the t distribution. Statistics and Computing, 10 (4):339–348, October 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008981510081.

Steele:2000:IBE

[376] Brian M. Steele and David A. Patterson. Ideal bootstrap estimation of expected prediction error for knearest neighbor classifiers: Applications for classification and error assessment. Statistics and Computing, 10 (4):349–355, October 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008933626919.

Anonymous:2000:HCd

[377] Anonymous. Help & contacts. Statistics and Computing, 10(4):??, October 2000. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Stafford:2001:GE

[378] Jamie Stafford. Guest editorial. Statistics and Computing, 11(1):5–6, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1026597513526.

Andrews:2001:AEM

[379] D. F. Andrews. Asymptotic expansions of moments and cumulants. Statistics and Computing, 11(1):7–16, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1026549630364.

Bellio:2001:CAP

[380] Ruggero Bellio and Alessandra R. Brazzale. A computer algebra package for approximate conditional inference. Statistics and Computing, 11 (1):17–24, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1026501714434.

Kendall:2001:SIC

[381] Wilfrid S. Kendall. Symbolic Itô calculus in AXIOM: An ongoing story. Statistics and Computing, 11(1):25– 35, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1026553731272.

Pistone:2001:GBF

[382] G. Pistone, E. Riccomagno, and Henry P. Wynn. Gröbner bases and factorisation in discrete probability and Bayes. Statistics and Computing, 11 (1):37–46, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1026505815343.

Stafford:2001:UIM

[383] James E. Stafford. Using intersection matrices to identify graphical structure. Statistics and Computing, 11 (1):47–55, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1026557832181.

Mcleod:2001:MLE

[384] A. I. Mcleod and B. Quenneville. Mean likelihood estimators. Statistics and Computing, 11(1):57–65, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1026509916251.

Ronchetti:2001:BSH

[385] Elvezio Ronchetti and Laura Ventura. Between stability and higher-order asymptotics. Statistics and Computing, 11(1):67–73, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1026562000322.

Smith:2001:SCC

[386] Bruce Smith and Christopher Field. Symbolic cumulant calculations for frequency domain time series. Statistics and Computing, 11(1):75–82, January 2001. CODEN STACE3.

ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1026514117160.

Kolassa:2001:BCR

[387] John E. Kolassa. Bounding convergence rates for Markov chains: An example of the use of computer algebra. Statistics and Computing, 11 (1):83–87, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1026566101230.

Gatto:2001:SCA

[388] Riccardo Gatto. Symbolic computation for approximating distributions of some families of one and two-sample nonparametric test statistics. Statistics and Computing, 11 (1):89–95, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1026518218069.

Anonymous:2001:HCa

[389] Anonymous. Help & contacts. Statistics and Computing, 11(1):??, January 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Philippe:2001:RSM

[390] Anne Philippe and Christian P. Robert. Riemann sums for MCMC estimation and convergence monitoring. Statistics and Computing, 11(2):103– 115, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1008926514119.

Tang:2001:EPC

[391] Man-Lai Tang. Exact power computation for stratified dose-response studies. Statistics and Computing, 11(2):117–124, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008971130958.

Neal:2001:AIS

[392] Radford M. Neal. Annealed importance sampling. Statistics and Computing, 11(2):125–139, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008923215028.

Cantoni:2001:RSS

[393] Eva Cantoni and Elvezio Ronchetti. Resistant selection of the smoothing parameter for smoothing splines. Statistics and Computing, 11(2):141– 146, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008975231866.

Tusell:2001:PTR

[394] Fernando Tusell. A permutation test for randomness with power against smooth variation. Statistics and Computing, 11(2):147–154, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008927315937.

Malvestuto:2001:HTA

[395] F. M. Malvestuto. A hypergraphtheoretic analysis of collapsibility and decomposability for extended log-linear models. Statistics and Computing, 11(2):155–169, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008979300007.

Denison:2001:BBS

[396] David G. T. Denison. Boosting with Bayesian stumps. Statistics and Computing, 11(2):171–178, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008931416845.

Brooks:2001:BAF

[397] S. P. Brooks. On Bayesian analyses and finite mixtures for proportions. Statistics and Computing, 11 (2):179–190, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1008983500916.

Lauritzen:2001:SLC

[398] Steffen L. Lauritzen and Frank Jensen. Stable local computation with conditional Gaussian distributions. Statistics and Computing, 11(2):191–203, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1008935617754.

Anonymous:2001:HCb

[399] Anonymous. Help & contacts. Statistics and Computing, 11(2):??, April 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Wolff:2001:GEI

[400] Rodney Wolff. Guest editorial: Introduction. Statistics and Computing, 11(3):211–212, July 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1023/A%3A1016645019314.

Lawrance:2001:CBD

[401] Anthony J. Lawrance. Chaos: But not in both directions! Statistics and Computing, 11(3):213–216, July 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1016696103384.

Diamond:2001:SSC

[402] Phil Diamond and Alexei Pokrovskii. The statistics of simulating chaos. Statistics and Computing, 11(3):217– 228, July 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1016648220222.

Finkenstadt:2001:CDA

[403] Bärbel F. Finkenstädt, Qiwei Yao, and Howell Tong. A conditional density approach to the order determination of time series. Statistics and Computing, 11(3):229–240, July 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016600304293.

Golia:2001:RAC

[404] Silvia Golia and Marco Sandri. A resampling algorithm for chaotic time series. Statistics and Computing, 11 (3):241–255, July 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016652321131.

Small:2001:TTS

[405] Michael Small, Kevin Judd, and Alistair Mees. Testing time series for nonlinearity. Statistics and Computing, 11(3):257–268, July 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016604405201.

Berliner:2001:MCB

[406] L. Mark Berliner. Monte Carlo based ensemble forecasting. Statistics and Computing, 11(3):269–275, July 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1016656422040.

Geegan:2001:PCT

[407] Dominique Geégan and Rolf Tschernig. Prediction of chaotic time series in the presence of measurement error: the importance of initial conditions. Statistics and Computing, 11(3):277–284, July 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016608506110.

Anonymous:2001:HCc

[408] Anonymous. Help & contacts. Statistics and Computing, 11(3):??, July 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Schimek:2001:GES

[409] Michael G. Schimek. Guest editorial: Semiparametric function estimation and testing. Statistics and Computing, 11(4):291–292, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1011973117955.

Staniswalis:2001:EGP

[410] Joan G. Staniswalis and Peter F. Thall. An explanation of generalized profile likelihoods. Statistics and Computing, 11(4):293–298, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1011912802026.

Muller:2001:ETG

[411] Marlene Müller. Estimation and testing in generalized partial linear models — a comparative study. Statistics and Computing, 11(4):299–309, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1011981314532.  See discussion

[412].

Staniswalis:2001:DPD

[412] Joan G. Staniswalis. Discussion of the paper by Dr. M. Müller. Statistics and Computing, 11(4):311–312, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1011964818864. See [411].

Kohn:2001:NRU

[413] Robert Kohn, Michael Smith, and David Chan. Nonparametric regression using linear combinations of basis functions. Statistics and Computing, 11(4):313–322, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1011916902934.

Grund:2001:SLF

[414] Birgit Grund and Jörg Polzehl. Semiparametric lack-of-fit tests in an additive hazard regression model. Statistics and Computing, 11(4):323–335, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1011968919773.

Hurn:2001:PED

[415] Merrilee Hurn, Ingelin Steinsland, and Håvard Rue. Parameter estimation for a deformable template model. Statistics and Computing, 11 (4):337–346, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1011921103843.

Gibson:2001:LES

[416] Gavin J. Gibson and Eric Renshaw. Likelihood estimation for stochastic compartmental models using Markov chain methods. Statistics and Computing, 11(4):347–358, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1011973120681.

Crowder:2001:CVT

[417] Martin Crowder. Corrected p-values for tests based on estimated nuisance parameters. Statistics and Computing, 11 (4):359–365, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1011925204752.

Yao:2001:BEA

[418] Qiwei Yao, Wenyang Zhang, and Howell Tong. Bootstrap estimation of actual significance levels for tests based on estimated nuisance parameters. Statistics and Computing, 11 (4):367–371, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1011977221590.

Amato:2001:AWS

[419] Umberto Amato and Anestis Antoniadis. Adaptive wavelet series estimation in separable nonparametric regression models. Statistics and Computing, 11(4):373–394, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1011929305660.

Anonymous:2001:HCd

[420] Anonymous. Help & contacts. Statistics and Computing, 11(4):??, October 2001. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Hand:2002:BGF

[421] David J. Hand. It’s been great-and the future looks even better. Statistics and Computing, 12(1):5, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1013121618984.

Oldford:2002:E

[422] R. Wayne Oldford. Editorial. Statistics and Computing, 12(1):7, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1013108003055.

Butler:2002:CDD

[423] Ronald W. Butler and Marc S. Paolella. Calculating the density and distribution function for the singly and doubly noncentral F. Statistics and Computing, 12(1):9–16, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1013160019893.

Barone:2002:RMC

[424] Piero Barone, Giovanni Sebastiani, and Julian Stander. Over-relaxation methods and coupled Markov chains for Monte Carlo simulation. Statistics and Computing, 12(1):17–26, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1013112103963.

Dellaportas:2002:BMV

[425] Petros Dellaportas, Jonathan J. Forster, and Ioannis Ntzoufras. On Bayesian model and variable selection using MCMC. Statistics and Computing, 12(1):27–36, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1013164120801.

Bunnin:2002:OPU

[426] F. O. Bunnin, Y. Guo, and Y. Ren. Option pricing under model and parameter uncertainty using predictive densities. Statistics and Computing, 12(1):37–44, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1013116204872.

Nason:2002:WPT

[427] Guy P. Nason and Theofanis Sapatinas. Wavelet packet transfer function modelling of nonstationary time series. Statistics and Computing, 12 (1):45–56, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1013168221710.

Rayner:2002:NML

[428] G. D. Rayner and H. L. MacGillivray. Numerical maximum likelihood estimation for the g-and-k and generalized g-and-h distributions. Statistics and Computing, 12(1):57–75, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1013120305780.

Doucet:2002:MMP

[429] Arnaud Doucet, Simon J. Godsill, and Christian P. Robert. Marginal maximum a posteriori estimation using Markov chain Monte Carlo. Statistics and Computing, 12(1):77–84, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1013172322619.

Anonymous:2002:HCa

[430] Anonymous. Help & contacts. Statistics and Computing, 12(1):??, January 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Alba:2002:IFE

[431] Enrique Alba and José M. Troya. Improving flexibility and efficiency by adding parallelism to genetic algorithms. Statistics and Computing, 12(2):91–114, April 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014803900897.

Molchanov:2002:SDA

[432] Ilya Molchanov and Sergei Zuyev. Steepest descent algorithms in a space of measures. Statistics and Computing, 12(2):115–123, April 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014878317736.

DeLuna:2002:SBI

[433] Xavier De Luna and Marc G. Genton. Simulation-based inference for simultaneous processes on regular lattices. Statistics and Computing, 12 (2):125–134, April 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014830401806.

Trendafilov:2002:GRP

[434] Nickolay T. Trendafilov. GIPSCAL revisited. A projected gradient approach. Statistics and Computing, 12(2):135–145, April 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014882518644.

Kadane:2002:HMC

[435] Joseph B. Kadane and Pantelis K. Vlachos. Hybrid methods for calculating optimal few-stage sequential strategies: Data monitoring for a clinical trial. Statistics and Computing, 12(2):147–152, April 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014834602714.

Lancaster:2002:LVT

[436] Gillian Lancaster and Mick Green. Latent variable techniques for categorical data. Statistics and Computing, 12(2):153–161, April 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014886619553.

Aitkin:2002:GML

[437] Murray Aitkin and Roberto Rocci. A general maximum likelihood analysis of measurement error in generalized linear models. Statistics and Computing, 12(2):163–174, April 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014838703623.

Zhu:2002:AGL

[438] Hong-Tu Zhu and Sik-Yum Lee. Analysis of generalized linear mixed models via a stochastic approximation algorithm with Markov chain Monte–Carlo method. Statistics and Computing, 12(2):175–183, April 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014890720461.

Anonymous:2002:HCb

[439] Anonymous. Help & contacts. Statistics and Computing, 12(2):??, April 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Croux:2002:LAM

[440] Christophe Croux, Gentiane Haesbroeck, and Peter J. Rousseeuw.

Location adjustment for the minimum volume ellipsoid estimator. Statistics and Computing, 12(3):191– 200, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020713207683.

Lee:2002:RVE

[441] Youngjo Lee. Robust variance estimators for fixed-effect estimates with hierarchical-likelihood. Statistics and Computing, 12(3):201–207, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020790524521.

Lee:2002:HAB

[442] Stephen M. S. Lee and Irene O. L. Wong. A hybrid approach based on saddlepoint and importance sampling methods for bootstrap tail probability estimation. Statistics and Computing, 12(3):209–217, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1020742625430.

Nason:2002:CWS

[443] Guy P. Nason. Choice of wavelet smoothness, primary resolution and threshold in wavelet shrinkage. Statistics and Computing, 12(3):219–227, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020746709500.

Cai:2002:PSC

[444] Yuzhi Cai and Wilfrid S. Kendall. Perfect simulation for correlated Poisson random variables conditioned to be positive. Statistics and Computing, 12(3):229–243, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1020798726338.

Mackenzie:2002:MHR

[445] Todd Mackenzie and Michal Abrahamowicz. Marginal and hazard ratio specific random data generation: Applications to semi-parametric bootstrapping. Statistics and Computing, 12(3):245–252, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1020750810409.

Fearn:2002:BDT

[446] T. Fearn, P. J. Brown, and P. Besbeas. A Bayesian decision theory approach to variable selection for discrimination. Statistics and Computing, 12(3):253–260, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1020702927247.

Ventura:2002:NPB

[447] Valérie Ventura. Non-parametric bootstrap recycling. Statistics and Computing, 12(3):261–273, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1020754911317.

Lim:2002:CMS

[448] Kian Guan Lim and Qin Xiao. Computing maximum smoothness forward rate curves. Statistics and Computing, 12(3):275–279, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1020707028156.

Phillips:2002:LAD

[449] Robert F. Phillips. Least absolute deviations estimation via the EM algorithm. Statistics and Computing, 12(3):281–285, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1020759012226.

Wilkinson:2002:CSH

[450] Darren J. Wilkinson and Stephen K. H. Yeung. Conditional simulation from highly structured Gaussian systems, with application to blocking-MCMC for the Bayesian analysis of very large linear models. Statistics and Computing, 12(3):287–300, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1020711129064.

Anonymous:2002:HCc

[451] Anonymous. Help & contacts. Statistics and Computing, 12(3):??, July 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Dinwoodie:2002:AMP

[452] I. H. Dinwoodie. Algebraic methods for polynomial statistical models. Statistics and Computing, 12 (4):307–314, October 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1020775726595.

Viele:2002:MML

[453] Kert Viele and Barbara Tong. Modeling with mixtures of linear regressions. Statistics and Computing, 12 (4):315–330, October 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1020779827503.

Hutson:2002:SPQ

[454] Alan D. Hutson. A semi-parametric quantile function estimator for use in bootstrap estimation procedures. Statistics and Computing, 12(4):331– 338, October 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020783911574.

Foster:2002:ARL

[455] David H. Foster. Automatic repeatedloess decomposition of data consisting of sums of oscillatory curves. Statistics and Computing, 12(4):339–351, October 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020736012482.

Pettitt:2002:CAG

[456] A. N. Pettitt, I. S. Weir, and A. G. Hart. A conditional autoregressive Gaussian process for irregularly spaced multivariate data with application to modelling large sets of binary data. Statistics and Computing, 12(4):353–367, October 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1020792130229.

Allcroft:2002:SEA

[457] David J. Allcroft and Chris A. Glasbey. A spectral estimator of Arma parameters from thresholded data. Statistics and Computing, 12(4):369–376, October 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020796314300.

Cowles:2002:MSC

[458] Mary Kathryn Cowles. MCMC sampler convergence rates for hierarchical normal linear models: A simulation approach. Statistics and Computing, 12 (4):377–389, October 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1020700515208.

Aykroyd:2002:AGD

[459] R. G. Aykroyd. Approximations for Gibbs distribution normalising constants. Statistics and Computing, 12 (4):391–397, October 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1020752516117.

Anonymous:2002:HCd

[460] Anonymous. Help & contacts. Statistics and Computing, 12(4):??, October 2002. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Neuenschwander:2003:GRN

[461] Daniel Neuenschwander and Hansmartin Zeuner. Generating random numbers of prescribed distribution using physical sources. Statistics and Computing, 13(1):5–11, February 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1021999708104.

Fischetti:2003:PCS

[462] Matteo Fischetti and Juan-José Salazar-González. Partial cell suppression: a new methodology for statistical disclosure control. Statistics and Computing, 13(1):13–21, February 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1021927424942.

Croux:2003:FMM

[463] C. Croux, P. Filzmoser, G. Pison, and P. J. Rousseeuw. Fitting multiplicative models by robust alternating regressions. Statistics and Computing, 13(1):23–36, February 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1021979409012.

ONeill:2003:PSR

[464] Philip D. O’Neill. Perfect simulation for Reed–Frost epidemic models. Statistics and Computing, 13 (1):37–44, February 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1021931526759.

Ng:2003:CNB

[465] S. K. Ng and G. J. McLachlan. On the choice of the number of blocks with the incremental EM algorithm for the fitting of normal mixtures. Statistics and Computing, 13(1):45– 55, February 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1021987710829.

Hojbjerre:2003:PLD

[466] Malene Højbjerre. Profile likelihood in directed graphical models from BUGS output. Statistics and Computing, 13 (1):57–66, February 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1021939828576.

Longford:2003:AMS

[467] Nicholas T. Longford. An alternative to model selection in ordinary regression. Statistics and Computing, 13 (1):67–80, February 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1021995912647.

Foster:2003:ARL

[468] David H. Foster. Automatic repeatedloess decomposition of data consisting of sums of oscillatory curves. Statistics and Computing, 13(1):81, February 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

                                        com/accesspage/article/10.1023/
                                        A%3A1021990220424.

Anonymous:2003:HCa

[469] Anonymous. Help & contacts. Statistics and Computing, 13(1):??, February 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Oldford:2003:ESC

[470] R. W. Oldford. Editorial: Statistics and computing: Having an impact. Statistics and Computing, 13(2): 87–89, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1023291706390.

Liechty:2003:SPL

[471] John C. Liechty, Dennis K. J. Lin, and James P. McDermott. Singlepass low-storage arbitrary quantile estimation for massive datasets. Statistics and Computing, 13(2):91–100, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1023296123228.

Kuk:2003:ACD

[472] Anthony Y. C. Kuk. Automatic choice of driving values in Monte Carlo likelihood approximation via posterior simulations. Statistics and Computing, 13(2):101–109, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1023248207299.

Marion:2003:ELS

[473] Glenn Marion, Gavin Gibson, and Eric Renshaw. Estimating likelihoods for spatio-temporal models using importance sampling. Statistics and Computing, 13(2):111–119, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1023200324137.

Kolassa:2003:AAC

[474] John E. Kolassa. Algorithms for approximate conditional inference. Statistics and Computing, 13(2):121– 126, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1023252308207.

Fernandez-Casal:2003:FST

[475] Rubén Fernández-Casal, Wenceslao González-Manteiga, and Manuel Febrero-Bande. Flexible spatiotemporal stationary variogram models. Statistics and Computing, 13 (2):127–136, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1023204525046.

VanDyk:2003:OSL

[476] David A. Van Dyk and Ruoxi Tang. The one-step-late PXEM algorithm. Statistics and Computing, 13(2):137– 152, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1023256509116.

Miller:2003:ECL

[477] Kim Miller, Suneeta Ramaswami, Peter Rousseeuw, J. Antoni Sellarès, Diane Souvaine, Ileana Streinu, and Anja Struyf. Efficient computation of location depth contours by methods of computational geometry. Statistics and Computing, 13(2):153– 162, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1023208625954.

Langsrud:2003:AUD

[478] Øyvind Langsrud. ANOVA for unbalanced data: Use Type II instead of Type III sums of squares. Statistics and Computing, 13(2):163– 167, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1023260610025.

Forster:2003:MCM

[479] Jonathan J. Forster, John W. Mc-Donald, and Peter W. F. Smith. Markov chain Monte Carlo exact inference for binomial and multinomial logistic regression models. Statistics and Computing, 13(2):169–177, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1023212726863.

Philippe:2003:PSP

[480] Anne Philippe and Christian P. Robert. Perfect simulation of positive Gaussian distributions. Statistics and Computing, 13(2):179–186, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1023264710933.

Anonymous:2003:HCb

[481] Anonymous. Help & contacts. Statistics and Computing, 13(2):??, April 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Yau:2003:EVS

[482] Paul Yau and Robert Kohn. Estimation and variable selection in nonparametric heteroscedastic regression. Statistics and Computing, 13 (3):191–208, August 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1024293931757.

Brewer:2003:DIN

[483] Mark J. Brewer. Discretisation for inference on normal mixture models. Statistics and Computing, 13 (3):209–219, August 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1024214615828.

Fang:2003:IGF

[484] Shoufan Fang, George Z. Gertner, Svetlana Shinkareva, Guangxing Wang, and Alan Anderson. Improved generalized Fourier amplitude sensitivity test (FAST) for model assessment. Statistics and Computing, 13 (3):221–226, August 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1024266632666.

Aitkin:2003:SMA

[485] Murray Aitkin and Rob Foxall. Statistical modelling of artificial neural networks using the multi-layer perceptron. Statistics and Computing, 13 (3):227–239, August 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1024218716736.

Jones:2003:DML

[486] M. C. Jones and I. Koch. Dependence maps: Local dependence in practice. Statistics and Computing, 13 (3):241–255, August 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1024270700807.

Bohning:2003:EAG

[487] Dankmar Böhning. The EM algorithm with gradient function update for discrete mixtures with known (fixed) number of components. Statistics and Computing, 13(3):257–265, August 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1024222817645.

Bolton:2003:PTN

[488] Richard J. Bolton, David J. Hand, and Andrew R. Webb. Projection techniques for nonlinear principal component analysis. Statistics and Computing, 13(3):267–276, August 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1024274801715.

Joe:2003:NOS

[489] Harry Joe and John C. Nash. Numerical optimization and surface estimation with imprecise function evaluations. Statistics and Computing, 13 (3):277–286, August 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1024226918553.

Anonymous:2003:HCc

[490] Anonymous. Help & contacts. Statistics and Computing, 13(3):??, August 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Willenborg:2003:GEI

[491] Leon Willenborg. Guest editorial: Introduction. Statistics and Computing, 13(4):291–293, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1025655703469.

Franconi:2003:SNS

[492] Luisa Franconi and Julian Stander. Spatial and non-spatial model-based protection procedures for the release of business microdata. Statistics and Computing, 13(4):295–305, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025654520307.

Polettini:2003:MES

[493] Silvia Polettini. Maximum entropy simulation for microdata protection. Statistics and Computing, 13(4):307– 320, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025606604377.

Burridge:2003:IPS

[494] Jim Burridge. Information preserving statistical obfuscation. Statistics and Computing, 13(4):321–327, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025658621216.

Muralidhar:2003:TBP

[495] Krishnamurty Muralidhar and Rathindra Sarathy. A theoretical basis for perturbation methods. Statistics and Computing, 13(4):329–335, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%

3A1025610705286. See comment [496] and rejoinder [497].

Polettini:2003:CTB

[496] Silvia Polettini and Julian Stander. A comment on “A theoretical basis for perturbation methods” by Krishnamurty Muralidhar and Rathindra Sarathy. Statistics and Computing, 13 (4):337–338, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/

article/10.1023/A%3A1025662722124.

See [495, 497].

Muralidhar:2003:RCP

[497] Krishnamurty Muralidhar and Rathindra Sarathy. A rejoinder to the comments by Polettini and Stander. Statistics and Computing, 13(4):339–342, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025614806194. See [496].

Domingo-Ferrer:2003:DRA

[498] Josep Domingo-Ferrer and Vicenç Torra. Disclosure risk assessment in statistical microdata protection via advanced record linkage. Statistics and Computing, 13(4):343–354, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025666923033.

Cuppen:2003:SDP

[499] Menno Cuppen and Leon Willenborg. Source data perturbation and consistent sets of safe tables. Statistics and Computing, 13(4):355–362, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025619007103.

Dobra:2003:PCH

[500] Adrian Dobra, Alan F. Karr, and Ashish P. Sanil. Preserving confidentiality of high-dimensional tabulated data: Statistical and computational issues. Statistics and Computing, 13 (4):363–370, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1025671023941.

Reiter:2003:MDR

[501] Jerome P. Reiter. Model diagnostics for remote access regression servers. Statistics and Computing, 13(4):371– 380, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1025623108012.

Schouten:2003:RAS

[502] Barry Schouten and Marc Cigrang. Remote access systems for statistical analysis of microdata. Statistics and Computing, 13(4):381–389, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1025675124850.

Anonymous:2003:HCd

[503] Anonymous. Help & contacts. Statistics and Computing, 13(4):??, October 2003. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

LEcuyer:2004:DLR

[504] Pierre L’Ecuyer and Renée Touzin. On the Deng–Lin random number generators and related methods. Statistics and Computing, 14(1):5–9, January 2004. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL

                                        http://dl.acm.org/citation.cfm?
                                        id=961292.961302;   http://link.

springer.com/article/10.1023/B%
3ASTCO.0000009417.88960.81.

Fearnhead:2004:PFM

[505] Paul Fearnhead. Particle filters for mixture models with an unknown number of components. Statistics and Computing, 14(1):11–21, January 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000009418.04621.cd.

Gustafson:2004:VDE

[506] Paul Gustafson, Ying C. MacNab, and Sijin Wen. On the value of derivative evaluations and random walk suppression in Markov chain Monte Carlo algorithms. Statistics and Computing, 14(1):23–38, January 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3ASTCO.0000009413.87656.
ef.

Trendafilov:2004:OPP

[507] Nickolay T. Trendafilov and G. A. Watson. The l_{1} oblique Procrustes problem. Statistics and Computing, 14(1):39–51, January 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3ASTCO.0000009415.14785.
2a.

Kent:2004:SCB

[508] John T. Kent, Patrick D. L. Constable, and Fikret Er. Simulation for the complex Bingham distribution. Statistics and Computing, 14 (1):53–57, January 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/B%3ASTCO.0000009414.14099.
                                        03.

Hsiao:2004:BMI

[509] Chuhsing Kate Hsiao, Su-Yun Huang, and Ching-Wei Chang. Bayesian marginal inference via candidate’s formula. Statistics and Computing, 14 (1):59–66, January 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/B%3ASTCO.0000009416.78796.
                                        78.

Protassov:2004:EBM

[510] Rostislav S. Protassov. EM-based maximum likelihood parameter estimation for multivariate generalized hyperbolic distributions with fixed λ. Statistics and Computing, 14(1):67– 77, January 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/B%
                                        3ASTCO.0000009419.12588.da.

Anonymous:2004:HCa

[511] Anonymous. Help & contacts. Statistics and Computing, 14(1):??, January 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Waddington:2004:UCP

[512] D. Waddington and R. Thompson. Using a correlated probit model approximation to estimate the variance for binary matched pairs. Statistics and Computing, 14(2):83–90, April 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000021406.25797.98.

Huang:2004:CSM

[513] Xin Huang and Christopher G. Small. Calculating the simplex median. Statistics and Computing, 14(2):91–98, April 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000021407.41456.1e.

Gorsich:2004:DNI

[514] David J. Gorsich and Marc G. Genton. On the discretization of nonparametric isotropic covariogram estimators. Statistics and Computing, 14(2):99–108, April 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3ASTCO.0000021408.63640.
d8.

Bognar:2004:BIP

[515] Matthew A. Bognar and Mary Kathryn Cowles. Bayesian inference for pairwise interacting point processes. Statistics and Computing, 14(2):109–117, April 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000021409.73461.b9.

Lin:2004:BAM

[516] Tsung I. Lin, Jack C. Lee, and Huey F. Ni. Bayesian analysis of mixture modelling using the multivariate t distribution. Statistics and Computing, 14(2):119–130, April 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/B%3ASTCO.0000021410.33077.
                                        10.

Feuerverger:2004:ALD

[517] Andrey Feuerverger and Jeffrey S. Rosenthal. Achieving limiting distributions for Markov chains using back buttons. Statistics and Computing, 14(2):131–141, April 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/B%3ASTCO.0000021411.12404.
                                        e4.

Wainwright:2004:TCB

[518] Martin Wainwright, Tommi Jaakkola, and Alan Willsky. Tree consistency and bounds on the performance of the max-product algorithm and its generalizations. Statistics and Computing, 14(2):143–166, April 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/B%3ASTCO.0000021412.33763.
                                        d5.

Anonymous:2004:HCb

[519] Anonymous. Help & contacts. Statistics and Computing, 14(2):??, April 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Whiley:2004:PAM

[520] Matt Whiley and Simon P. Wilson. Parallel algorithms for Markov chain Monte Carlo methods in latent spatial Gaussian models. Statistics and Computing, 14(3):171–179, August 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000035299.51541.5e.

Gomez:2004:NDS

[521] M. Gómez and C. Bielza. Node deletion sequences in influence diagrams using genetic algorithms. Statistics and Computing, 14(3):181–198, August 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000035300.39728.03.

Smola:2004:TSV

[522] Alex J. Smola and Bernhard Schölkopf. A tutorial on support vector regression. Statistics and Computing, 14 (3):199–222, August 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3ASTCO.0000035301.49549.
88.

Tsionas:2004:BIM

[523] Efthymios G. Tsionas. Bayesian inference for multivariate gamma distributions. Statistics and Computing, 14 (3):223–233, August 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3ASTCO.0000035302.87186.
be.

DeCanditiis:2004:TAJ

[524] Daniela De Canditiis and Theofanis Sapatinas. Testing for additivity and joint effects in multivariate nonparametric regression using Fourier and wavelet methods. Statistics and Computing, 14(3):235–249, August 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/B%
                                        3ASTCO.0000035303.24825.b3.

Genz:2004:NCR

[525] Alan Genz. Numerical computation of rectangular bivariate and trivariate normal and t probabilities. Statistics and Computing, 14(3):251–260, August 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/B%
                                        3ASTCO.0000035304.20635.31.

Fearnhead:2004:FRC

[526] Paul Fearnhead. Filtering recursions for calculating likelihoods for queues based on inter-departure time data. Statistics and Computing, 14 (3):261–266, August 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/B%3ASTCO.0000035305.92337.
                                        80.

Biernacki:2004:IEU

[527] Christophe Biernacki. Initializing EM using the properties of its trajectories in Gaussian mixtures. Statistics and Computing, 14(3):267–279, August 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1023/B%
                                        3ASTCO.0000035306.77434.31.

Anonymous:2004:HCc

[528] Anonymous. Help & contacts. Statistics and Computing, 14(3):??, August 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Dinwoodie:2004:TMH

[529] Ian H. Dinwoodie, Laura Felicia Matusevich, and Ed Mosteig. Transform methods for the hypergeometric distribution. Statistics and Computing, 14 (4):287–297, October 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3ASTCO.0000039478.96565.
a5.

Lin:2004:FCA

[530] Chin-Tsai Lin, Chie-Bein Chen, and Wen-Hsiang Wu. Fuzzy clustering algorithm for latent class model. Statistics and Computing, 14(4):299–310, October 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000039479.56180.d5.

Podlich:2004:SPE

[531] Heather M. Podlich, Malcolm J. Faddy, and Gordon K. Smyth. Semiparametric extended Poisson process models for count data. Statistics and Computing, 14(4):311–321, October 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000039480.66002.5a.

Dias:2004:ECE

[532] José G. Dias and Michel Wedel. An empirical comparison of EM, SEM and MCMC performance for problematic Gaussian mixture likelihoods.

Statistics and Computing, 14(4):323– 332, October 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000039481.32211.5a.

Maire:2004:PAM

[533] Sylvain Maire. Polynomial approximations of multivariate smooth functions from quasi-random data. Statistics and Computing, 14(4):333–336, October 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/article/10.1023/B%
3ASTCO.0000039482.91826.ce.

Lee:2004:APW

[534] Thomas C. M. Lee and Hee-Seok Oh. Automatic polynomial wavelet regression. Statistics and Computing, 14 (4):337–341, October 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3ASTCO.0000039483.71153.
8d.

Zhang:2004:LMG

[535] Zhihua Zhang, Kap Luk Chan, Yiming Wu, and Chibiao Chen. Learning a multivariate Gaussian mixture model with the reversible jump MCMC algorithm. Statistics and Computing, 14 (4):343–355, October 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3ASTCO.0000039484.36470.
41.

Faraway:2004:MCS

[536] Julian J. Faraway. Modeling continuous shape change for facial animation. Statistics and Computing, 14 (4):357–363, October 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3ASTCO.0000039485.52129.
ed.

Anonymous:2004:HCd

[537] Anonymous. Help & contacts. Statistics and Computing, 14(4):??, October 2004. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Zhao:2005:ECQ

[538] G. H. Zhao, K. L. Teo, and K. S. Chan. Estimation of conditional quantiles by a new smoothing approximation of asymmetric loss functions. Statistics and Computing, 15(1):5–11, January 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
4785-9.

Liang:2005:BNN

[539] Faming Liang. Bayesian neural networks for nonlinear time series forecasting. Statistics and Computing, 15(1):13–29, January 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-4786-8.

Shi:2005:HGP

[540] J. Q. Shi, R. Murray-Smith, and D. M. Titterington. Hierarchical Gaussian process mixtures for regression. Statistics and Computing, 15(1):31–41, January 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
4787-7.

Hilliam:2005:CCS

[541] Rechel M. Hilliam and Anthony J. Lawrance. Chaos communication synchronization: Combatting noise by distribution transformation. Statistics and Computing, 15(1):43–52, January 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
4788-6.

Langsrud:2005:RT

[542] Øyvind Langsrud. Rotation tests. Statistics and Computing, 15(1):53–60, January 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
4789-5.

Agarwal:2005:SSS

[543] Deepak K. Agarwal and Alan E. Gelfand. Slice sampling for simulation based fitting of spatial data models. Statistics and Computing, 15(1):61–69, January 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
4790-z.

Kabaila:2005:CEC

[544] Paul Kabaila. Computation of exact confidence intervals from discrete data using Studentized test statistics. Statistics and Computing, 15(1):71–78, January 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
4791-y.

Anonymous:2005:HCa

[545] Anonymous. Help & contacts. Statistics and Computing, 15(1):??, January 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Eckley:2005:ECD

[546] Idris A. Eckley and Guy P. Nason. Efficient computation of the discrete autocorrelation wavelet inner product matrix. Statistics and Computing, 15(2):83–92, April 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-6200-y.

Nott:2005:ESS

[547] David J. Nott, Anthony Y. C. Kuk, and Hiep Duc. Efficient sampling schemes for Bayesian MARS models with many predictors. Statistics and Computing, 15(2):93–101, April 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-6201-x.

Giakoumatos:2005:BAU

[548] Stefanos G. Giakoumatos, Petros Dellaportas, and Dimitris N. Politis. Bayesian analysis of the unobserved ARCH model. Statistics and Computing, 15(2):103–111, April 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-005-
                                        6202-9.

DiMarzio:2005:KDC

[549] M. Di Marzio and C. C. Taylor. Kernel density classification and boosting: an L_{2} analysis. Statistics and Computing, 15(2):113–123, April 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-005-6203-8.

Fearnhead:2005:DSD

[550] Paul Fearnhead. Direct simulation for discrete mixture distributions. Statistics and Computing, 15(2):125–133, April 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-005-
                                        6204-7.

Guo:2005:NSM

[551] Dong Guo, Xiaodong Wang, and Rong Chen. New sequential Monte Carlo methods for nonlinear dynamic systems. Statistics and Computing, 15 (2):135–147, April 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-005-6846-5.

Anonymous:2005:HCb

[552] Anonymous. Help & contacts. Statistics and Computing, 15(2):??, April 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Tutz:2005:LC

[553] G. Tutz and H. Binder. Localized classification. Statistics and Computing, 15(3):155–166, July 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-1305-x.

Sweeting:2005:APD

[554] Trevor Sweeting and Samer Kharroubi. Application of a predictive distribution formula to Bayesian computation for incomplete data models. Statistics and Computing, 15(3):167–178, July 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
1306-9.

Schlattmann:2005:BNC

[555] Peter Schlattmann. On bootstrapping the number of components in finite mixtures of Poisson distributions. Statistics and Computing, 15 (3):179–188, July 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-1307-8.

Desgagne:2005:ISG

[556] Alain Desgagné and Jean-François Angers. Importance sampling with the generalized exponential power density. Statistics and Computing, 15 (3):189–196, July 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-1308-7.

Frolich:2005:MEO

[557] Markus Frölich. Matching estimators and optimal bandwidth choice. Statistics and Computing, 15(3):197–215, July 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
1309-6.

Aitkin:2005:BPN

[558] Murray Aitkin, Richard J. Boys, and Tom Chadwick. Bayesian point null hypothesis testing via the posterior likelihood ratio. Statistics and Computing, 15(3):217–230, July 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-1310-0.

Negassa:2005:TSS

[559] Abdissa Negassa, Antonio Ciampi, Michal Abrahamowicz, Stanley Shapiro, and Jean-François Boivin. Treestructured subgroup analysis for censored survival data: Validation of computationally inexpensive model selection criteria. Statistics and Computing, 15(3):231–239, July 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-1311-z.

Arnold:2005:ESM

[560] Dirk V. Arnold and Hans-Georg Beyer. Expected sample moments of concomitants of selected order statistics. Statistics and Computing, 15(3):241–250, July 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11222-005-
1312-y.

Anonymous:2005:HCc

[561] Anonymous. Help & contacts. Statistics and Computing, 15(3):??, July 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Karlis:2005:MPR

[562] Dimitris Karlis and Loukia Meligkotsidou. Multivariate Poisson regression with covariance structure. Statistics and Computing, 15(4):255–265, October 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
4069-4.

Dunn:2005:SET

[563] Peter K. Dunn and Gordon K. Smyth. Series evaluation of Tweedie exponential dispersion model densities. Statistics and Computing, 15(4):267–280, October 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
4070-y.

Butler:2005:APM

[564] Ronald W. Butler and Andrew T. A. Wood. Approximation of power in multivariate analysis. Statistics and Computing, 15(4):281–287, October 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-005-
4071-x.

Roca-Pardinas:2005:TIG

[565] J. Roca-Pardiñas, C. Cadarso-Suárez, and W. González-Manteiga. Testing for interactions in generalized additive models: Application to SO_{2} pollution data. Statistics and Computing, 15 (4):289–299, October 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-4072-9.

Einbeck:2005:LPC

[566] Jochen Einbeck, Gerhard Tutz, and Ludger Evers. Local principal curves. Statistics and Computing, 15 (4):301–313, October 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-4073-8.

Neal:2005:CSN

[567] Peter Neal and Gareth Roberts. A case study in non-centering for data augmentation: Stochastic epidemics. Statistics and Computing, 15 (4):315–327, October 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-4074-7.

Zhang:2005:BMI

[568] Feng Zhang, Bani Mallick, and Zhujun Weng. A Bayesian method for identifying independent sources of nonrandom spatial patterns. Statistics and Computing, 15(4):329–339, October 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11222-005-
4075-6.

Knoth:2005:AAC

[569] Sven Knoth. Accurate ARL computation for EWMA-S^{2} control charts. Statistics and Computing, 15(4): 341–352, October 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-005-3393-z.

Anonymous:2005:HCd

[570] Anonymous. Help & contacts. Statistics and Computing, 15(4):??, October 2005. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

McAuliffe:2006:NEB

[571] Jon D. McAuliffe, David M. Blei, and Michael I. Jordan. Nonparametric empirical Bayes for the Dirichlet process mixture model. Statistics and Computing, 16(1):5–14, January 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-5196-2.

Wolfsegger:2006:SCI

[572] Martin J. Wolfsegger and Thomas Jaki. Simultaneous confidence intervals by iteratively adjusted alpha for relative effects in the one-way layout. Statistics and Computing, 16(1):15–23, January 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
5197-1.

Percival:2006:ESG

[573] Donald B. Percival and William L. B. Constantine. Exact simulation of Gaussian time series from nonparametric spectral estimates with application to bootstrapping. Statistics and Computing, 16(1):25–35, January 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
5198-0.

Amato:2006:WKP

[574] Umberto Amato, Anestis Antoniadis, and Marianna Pensky. Wavelet kernel penalized estimation for nonequispaced design regression. Statistics and Computing, 16(1):37–55, January 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
5283-4.

Dellaportas:2006:MMN

[575] Petros Dellaportas and Ioulia Papageorgiou. Multivariate mixtures of normals with unknown number of components. Statistics and Computing, 16(1):57–68, January 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-5338-6.

Lee:2006:FAR

[576] Youngjo Lee and John A. Nelder. Fitting via alternative random-effect models. Statistics and Computing, 16 (1):69–75, January 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-5534-4.

Meila:2006:TBL

[577] Marina Meila and Tommi Jaakkola. Tractable Bayesian learning of tree belief networks. Statistics and Computing, 16(1):77–92, January 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-006-5535-
3.

Eidsvik:2006:DMH

[578] Jo Eidsvik and HåKon Tjelmeland. On directional Metropolis–Hastings algorithms. Statistics and Computing, 16 (1):93–106, January 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-5536-2.

Anonymous:2006:HCa

[579] Anonymous. Help & contacts. Statistics and Computing, 16(1):??, January 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Jank:2006:ESM

[580] Wolfgang Jank. Efficient simulated maximum likelihood with an application to online retailing. Statistics and Computing, 16(2):111–124, June 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
6890-9.

So:2006:BAN

[581] Mike K. P. So. Bayesian analysis of nonlinear and non-Gaussian state space models via multiple-try sampling methods. Statistics and Computing, 16(2):125–141, June 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-6891-8.

Nunes:2006:ALN

[582] Matthew A. Nunes, Marina I. Knight, and Guy P. Nason. Adaptive lifting for nonparametric regression. Statistics and Computing, 16(2):143–159, June 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
6560-y.

Poland:2006:MCS

[583] Jan Poland and Marcus Hutter. MDL convergence speed for Bernoulli sequences. Statistics and Computing, 16(2):161–175, June 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-6746-3.

Bernholt:2006:MRM

[584] T. Bernholt, R. Fried, U. Gather, and I. Wegener. Modified repeated median filters. Statistics and Computing, 16(2):177–192, June 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-8449-1.

Guan:2006:MCM

[585] Yongtao Guan, Roland Fleißner, Paul Joyce, and Stephen M. Krone. Markov chain Monte Carlo in small worlds. Statistics and Computing, 16(2):193–202, June 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-6966-6.

Fearnhead:2006:EEB

[586] Paul Fearnhead. Exact and efficient Bayesian inference for multiple changepoint problems. Statistics and Computing, 16(2):203–213, June 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
8450-8.

Bouguila:2006:PBE

[587] Nizar Bouguila, Djemel Ziou, and Ernest Monga. Practical Bayesian estimation of a finite beta mixture through Gibbs sampling and its applications. Statistics and Computing, 16(2):215–225, June 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-8451-7.

Anonymous:2006:HCb

[588] Anonymous. Help & contacts. Statistics and Computing, 16(2):??, June 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Alfo:2006:VCM

[589] Marco Alfò and Murray Aitkin. Variance component models for longitudinal count data with baseline information: epilepsy data revisited. Statistics and Computing, 16(3):231– 238, September 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-7072-5.

Braak:2006:MCM

[590] Cajo J. F. Ter Braak. A Markov chain Monte Carlo version of the genetic algorithm differential evolution: easy Bayesian computing for real parameter spaces. Statistics and Computing, 16 (3):239–249, September 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-8769-1.

Alba:2006:CAM

[591] Enrique Alba and Enrique Domínguez. Comparative analysis of modern optimization tools for the p-median problem. Statistics and Computing, 16(3): 251–260, September 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-8079-7.

Kume:2006:SCD

[592] Alfred Kume and Stephen G. Walker. Sampling from compositional and directional distributions. Statistics and Computing, 16(3):261–265, September 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
8077-9.

Chan:2006:MVU

[593] Kenny Y. F. Chan, Stephen M. S. Lee, and Kai W. Ng. Minimum variance unbiased estimation based on bootstrap iterations. Statistics and Computing, 16(3):267–277, September 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
8078-8.

Zoeter:2006:DAI

[594] Onno Zoeter and Tom Heskes. Deterministic approximate inference techniques for conditionally Gaussian state space models. Statistics and Computing, 16(3):279–292, September 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
8770-8.

Cobb:2006:APD

[595] Barry R. Cobb, Prakash P. Shenoy, and Rafael Rumí. Approximating probability density functions in hybrid Bayesian networks with mixtures of truncated exponentials. Statistics and Computing, 16(3):293–308, September 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
8175-8.

Demiris:2006:CFO

[596] Nikolaos Demiris and Philip D. O’Neill. Computation of final outcome probabilities for the generalised stochastic epidemic. Statistics and Computing, 16 (3):309–317, September 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-8320-4.

Anonymous:2006:HCc

[597] Anonymous. Help & contacts. Statistics and Computing, 16(3):??, September 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Golightly:2006:BSI

[598] Andrew Golightly and Darren J. Wilkinson. Bayesian sequential inference for nonlinear multivariate diffusions. Statistics and Computing, 16 (4):323–338, December 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-006-9392-x.

Haario:2006:DEA

[599] Heikki Haario, Marko Laine, Antonietta Mira, and Eero Saksman. DRAM: Efficient adaptive MCMC. Statistics and Computing, 16(4):339–354, December 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-006-
                                        9438-0.

Corander:2006:BML

[600] Jukka Corander, Mats Gyllenberg, and Timo Koski. Bayesian model learning based on a parallel MCMC strategy. Statistics and Computing, 16(4): 355–362, December 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-006-9391-y.

Aykroyd:2006:FSE

[601] Robert G. Aykroyd and Brain A. Cattle. A flexible statistical and efficient computational approach to object location applied to electrical tomography. Statistics and Computing, 16(4): 363–375, December 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-9619-x.

Bowman:2006:DDN

[602] A. W. Bowman, A. Pope, and B. Ismail. Detecting discontinuities in nonparametric regression curves and surfaces. Statistics and Computing, 16 (4):377–390, December 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-9618-y.

Gibson:2006:BEP

[603] G. J. Gibson, W. Otten, J. A. N. Filipe, A. Cook, G. Marion, and C. A. Gilligan. Bayesian estimation for percolation models of disease spread in plant populations. Statistics and Computing, 16(4):391–402, December 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
0019-z.

Anonymous:2006:HCd

[604] Anonymous. Help & contacts. Statistics and Computing, 16(4):??, December 2006. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Kendall:2007:CBB

[605] W. S. Kendall, J.-M. Marin, and C. P. Robert. Confidence bands for Brownian motion and applications to Monte Carlo simulation. Statistics and Computing, 17(1):1–10, March 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
9001-z.

Hand:2007:OPP

[606] David J. Hand, Wojtek J. Krzanowski, and Martin J. Crowder. Optimal predictive partitioning. Statistics and Computing, 17(1):11–21, March 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
9003-x.

Goswami:2007:LSE

[607] Gopi Goswami and Jun S. Liu. On learning strategies for evolutionary Monte Carlo. Statistics and Computing, 17(1):23–38, March 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-9002-y.

Bock:2007:EIE

[608] M. Bock, A. W. Bowman, and B. Ismail. Estimation and inference for error variance in bivariate nonparametric regression. Statistics and Computing, 17(1):39–47, March 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-9000-0.

Lee:2007:HLP

[609] Youngjo Lee, John A. Nelder, and Maengseok Noh. H-likelihood: problems and solutions. Statistics and Computing, 17(1):49–55, March 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-006-9006-
7.

Mitra:2007:URF

[610] Sinjini Mitra, Nicole A. Lazar, and Yanxi Liu. Understanding the role of facial asymmetry in human face identification. Statistics and Computing, 17(1):57–70, March 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-9004-9.

Azzalini:2007:CND

[611] Adelchi Azzalini and Nicola Torelli. Clustering via nonparametric density estimation. Statistics and Computing, 17(1):71–80, March 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-9010-y.

Anonymous:2007:HCa

[612] Anonymous. Help & contacts. Statistics and Computing, 17(1):??, March 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Lin:2007:RMM

[613] Tsung I. Lin, Jack C. Lee, and Wan J. Hsieh. Robust mixture modeling using the skew t distribution. Statistics and Computing, 17(2):81–92, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11222-006-
9005-8.

Meligkotsidou:2007:BMP

[614] Loukia Meligkotsidou. Bayesian multivariate Poisson mixtures with an unknown number of components. Statistics and Computing, 17(2):93–107, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
9008-5.

Craiu:2007:AMT

[615] Radu V. Craiu and Christiane Lemieux. Acceleration of the Multiple-Try Metropolis algorithm using antithetic and stratified sampling. Statistics and Computing, 17(2):109–120, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
9009-4.

Lavielle:2007:PEV

[616] Marc Lavielle and Cristian Meza. A parameter expansion version of the SAEM algorithm. Statistics and Computing, 17(2):121–130, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
9007-6.

Leslie:2007:GAH

[617] David S. Leslie, Robert Kohn, and David J. Nott. A general approach to heteroscedastic linear regression. Statistics and Computing, 17(2):131–146, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-9013-8.

Nobile:2007:BFM

[618] Agostino Nobile and Alastair T. Fearnside. Bayesian finite mixtures with an unknown number of components: The allocation sampler. Statistics and Computing, 17(2):147–162, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
9014-7.

Zhu:2007:MLS

[619] Hongtu Zhu, Minggao Gu, and Bradley Peterson. Maximum likelihood from spatial random effects models via the stochastic approximation expectation maximization algorithm. Statistics and Computing, 17(2):163–177, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-006-
9012-9.

Ganguli:2007:FSG

[620] B. Ganguli and M. P. Wand. Feature significance in generalized additive models. Statistics and Computing, 17(2):179–192, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-006-9011-x.

Bhattacharya:2007:MFI

[621] Sourabh Bhattacharya, Alan E. Gelfand, and Kent E. Holsinger. Model fitting and inference under latent equilibrium processes. Statistics and Computing, 17(2):193–208, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-006-
                                        9015-6.

Anonymous:2007:HCb

[622] Anonymous. Help & contacts. Statistics and Computing, 17(2):??, June 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Same:2007:OCE

[623] Allou Samé, Christophe Ambroise, and Gérard Govaert. An online classification EM algorithm based on the mixture model. Statistics and Computing, 17(3):209–218, September 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-007-
                                        9017-z.

Neil:2007:IHB

[624] Martin Neil, Manesh Tailor, and David Marquez. Inference in hybrid Bayesian networks using dynamic discretization. Statistics and Computing, 17(3): 219–233, September 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-007-9018-y.

Kerman:2007:MSP

[625] Jouni Kerman and Andrew Gelman. Manipulating and summarizing posterior simulations using random variable objects. Statistics and Computing, 17 (3):235–244, September 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9020-4.

Stevens:2007:EAM

[626] Richard J. Stevens and Trevor J. Sweeting. Estimation across multiple models with application to Bayesian computing and software development. Statistics and Computing, 17(3):245– 252, September 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9026-y.

Jang:2007:CAM

[627] Woncheol Jang and Martin Hendry. Cluster analysis of massive datasets in astronomy. Statistics and Computing, 17(3):253–262, September 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-007-9027-
x.

Jasra:2007:PBS

[628] Ajay Jasra, David A. Stephens, and Christopher C. Holmes. On population-based simulation for static inference. Statistics and Computing, 17 (3):263–279, September 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9028-9.

Jokiel-Rokita:2007:MEP

[629] Alicja Jokiel-Rokita and Ryszard Magiera. Minimax estimation of a probability of success under LINEX loss. Statistics and Computing, 17(3):

281–291, September 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-007-9032-0.

Anonymous:2007:HCc

[630] Anonymous. Help & contacts. Statistics and Computing, 17(3):??, September 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Gannaz:2007:REW

[631] Irène Gannaz. Robust estimation and wavelet thresholding in partially linear models. Statistics and Computing, 17 (4):293–310, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-007-9019-x.

McDermott:2007:DSS

[632] James P. McDermott, G. Jogesh Babu, John C. Liechty, and Dennis K. J. Lin. Data skeletons: simultaneous estimation of multiple quantiles for massive streaming datasets with applications to density estimation. Statistics and Computing, 17(4):311–321, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-007-
                                        9021-3.

Yan:2007:PMB

[633] Jun Yan, Mary Kathryn Cowles, Shaowen Wang, and Marc P. Armstrong. Parallelizing MCMC for Bayesian spatiotemporal geostatistical models. Statistics and Computing, 17 (4):323–335, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9022-2.

Choi:2007:OST

[634] Leena Choi, Brian Caffo, and Charles Rohde. Optimal sampling times in bioequivalence studies using a simulated annealing algorithm. Statistics and Computing, 17(4):337–347, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-007-
9023-1.

Delaigle:2007:FPC

[635] A. Delaigle and I. Gijbels. Frequent problems in calculating integrals and optimizing objective functions: a case study in density deconvolution. Statistics and Computing, 17 (4):349–355, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9024-0.

Sisson:2007:DBD

[636] S. A. Sisson and Y. Fan. A distance-based diagnostic for transdimensional Markov chains. Statistics and Computing, 17(4):357–367, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-007-
9025-z.

Skare:2007:BAS

[637] Øivind Skare, Jesper Møller, and Eva B. Vedel Jensen. Bayesian analysis of spatial point processes in the neighbourhood of Voronoi networks. Statistics and Computing, 17 (4):369–379, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-007-9029-8.

Preisser:2007:DDM

[638] John S. Preisser and Jamie Perin. Deletion diagnostics for marginal mean and correlation model parameters in estimating equations. Statistics and Computing, 17(4):381–393, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-007-
                                        9031-1.

vonLuxburg:2007:TSC

[639] Ulrike von Luxburg. A tutorial on spectral clustering. Statistics and Computing, 17(4):395–416, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-007-
                                        9033-z.

Anonymous:2007:HCd

[640] Anonymous. Help & contacts. Statistics and Computing, 17(4):??, December 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Fruhwirth-Schnatter:2008:BPC

[641] Sylvia Frühwirth-Schnatter and Regina Tüchler. Bayesian parsimonious covariance estimation for hierarchical linear mixed models. Statistics and Computing, 18(1):1–13, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9030-2.

Gambini:2008:AED

[642] Juliana Gambini, Marta E. Mejail, Julio Jacobo-Berlles, and Alejandro C. Frery. Accuracy of edge detection methods with local information in speckled imagery. Statistics and Computing, 18(1):15–26, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-007-
9034-y.

Karlis:2008:EBM

[643] Dimitris Karlis and Panagiotis Tsiamyrtzis. Exact Bayesian modeling for bivariate Poisson data and extensions. Statistics and Computing, 18(1):27–40, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-007-
9035-x.

Hansen:2008:CCA

[644] Pierre Hansen and Nenad Mladenović. Complement to a comparative analysis of heuristics for the p-median problem. Statistics and Computing, 18(1):41–46, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-007-
9036-9.

Johansen:2008:PMM

[645] Adam M. Johansen, Arnaud Doucet, and Manuel Davy. Particle methods for maximum likelihood estimation in latent variable models. Statistics and Computing, 18(1):47–57, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-007-
                                        9037-8.

Pinson:2008:LLR

[646] Pierre Pinson, Henrik Aa. Nielsen, Henrik Madsen, and Torben S. Nielsen. Local linear regression with adaptive orthogonal fitting for the wind power application. Statistics and Computing, 18(1):59–71, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-007-9038-7.

Dunn:2008:ETE

[647] Peter K. Dunn and Gordon K. Smyth. Evaluation of Tweedie exponential dispersion model densities by Fourier inversion. Statistics and Computing, 18(1):73–86, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-007-9039-6.

Binder:2008:CMF

[648] Harald Binder and Gerhard Tutz. A comparison of methods for the fitting of generalized additive models. Statistics and Computing, 18(1):87–99, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-007-
                                        9040-0.

Su:2008:MSV

[649] Zheng Su, Jiaqiao Hu, and Wei Zhu. Multi-step variance minimization in sequential tests. Statistics and Computing, 18(1):101–108, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-007-
9041-z.

Anonymous:2008:HCa

[650] Anonymous. Help & contacts. Statistics and Computing, 18(1):??, March 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Zhao:2008:MEF

[651] J.-H. Zhao, Philip L. H. Yu, and Qibao Jiang. ML estimation for factor analysis: EM or non-EM? Statistics and Computing, 18(2):109–123, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-007-
9042-y.

Boys:2008:BID

[652] R. J. Boys, D. J. Wilkinson, and T. B. L. Kirkwood. Bayesian inference for a discretely observed stochastic kinetic model. Statistics and Computing, 18(2):125–135, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9043-x.

Alfo:2008:FMM

[653] Marco Alfò, Luciano Nieddu, and Donatella Vicari. A finite mixture model for image segmentation. Statistics and Computing, 18(2):137–150, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-007-
9044-9.

Fearnhead:2008:CMC

[654] Paul Fearnhead. Computational methods for complex stochastic systems: a review of some alternatives to MCMC. Statistics and Computing, 18(2):151–171, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9045-8.

Daudin:2008:MMR

[655] J.-J. Daudin, F. Picard, and S. Robin. A mixture model for random graphs. Statistics and Computing, 18(2): 173–183, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9046-7.

Weiss:2008:SMI

[656] Christian H. Weiß. Statistical mining of interesting association rules. Statistics and Computing, 18(2):185–194, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-007-
9047-6.

Sauerbrei:2008:IAS

[657] Willi Sauerbrei, Norbert Holländer, and Anika Buchholz. Investigation about a screening step in model selection. Statistics and Computing, 18(2):195–208, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9048-5.

Perez:2008:EGR

[658] C. J. Pérez, J. Martín, C. Rojano, and F. J. Girón. Efficient generation of random vectors by using the ratio-of-uniforms method with ellipsoidal envelopes. Statistics and Computing, 18(2):209–217, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-007-9049-4.

Besbeas:2008:IES

[659] P. Besbeas and B. J. T. Morgan. Improved estimation of the stable laws. Statistics and Computing, 18 (2):219–231, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9050-6.

Anonymous:2008:HCb

[660] Anonymous. Help & contacts. Statistics and Computing, 18(2):??, June 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Lo:2008:LRT

[661] Yungtai Lo. A likelihood ratio test of a homoscedastic normal mixture against a heteroscedastic normal mixture. Statistics and Computing, 18(3): 233–240, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9052-4.

Koekemoer:2008:SPM

[662] Gerhard Koekemoer and Jan W. H. Swanepoel. A semi-parametric method for transforming data to normality. Statistics and Computing, 18(3): 241–257, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9053-3.

Bremner:2008:OSA

[663] David Bremner, Dan Chen, John Iacono, Stefan Langerman, and Pat Morin. Output-sensitive algorithms for Tukey depth and related problems. Statistics and Computing, 18(3): 259–266, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9054-2.

Shi:2008:CPC

[664] J. Q. Shi and B. Wang. Curve prediction and clustering with mixtures of Gaussian process functional regression models. Statistics and Computing, 18 (3):267–283, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9055-1.

McNicholas:2008:PGM

[665] Paul David McNicholas and Thomas Brendan Murphy. Parsimonious Gaussian mixture models. Statistics and Computing, 18(3):285–296, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s11222-008-
9056-0.

Gonzalez-Manteiga:2008:SMI

[666] W. González-Manteiga, M. D. Martínez-Miranda, and R. Raya-Miranda. SiZer map for inference with additive models. Statistics and Computing, 18(3): 297–312, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9057-z.

Silva:2008:CMD

[667] Ralph dos Santos Silva and Hedibert Freitas Lopes. Copula, marginal distributions and model selection: a Bayesian note. Statistics and Computing, 18(3):313–320, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9058-y.

Gatto:2008:SCA

[668] Riccardo Gatto. Some computational aspects of the generalized von Mises distribution. Statistics and Computing, 18(3):321–331, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9060-4.

Baharev:2008:CNN

[669] Ali Baharev and Sándor Kemény. On the computation of the noncentral F and noncentral beta distribution. Statistics and Computing, 18(3): 333–340, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9061-3.

Anonymous:2008:HCc

[670] Anonymous. Help & contacts. Statistics and Computing, 18(3):??, September 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Fearnhead:2008:ESI

[671] Paul Fearnhead. Editorial: Special issue on adaptive Monte Carlo methods. Statistics and Computing, 18(4): 341–342, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-008-9107-
                                        6.

Andrieu:2008:TAM

[672] Christophe Andrieu and Johannes Thoms. A tutorial on adaptive MCMC. Statistics and Computing, 18 (4):343–373, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9110-y.

Ren:2008:AEM

[673] Yuan Ren, Yu Ding, and Faming Liang. Adaptive evolutionary Monte Carlo algorithm for optimization with applications to sensor placement problems. Statistics and Computing, 18 (4):375–390, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9079-6.

Gerlach:2008:BIM

[674] Richard Gerlach and Cathy W. S. Chen. Bayesian inference and model comparison for asymmetric smooth transition heteroskedastic models. Statistics and Computing, 18(4):391– 408, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9063-1.

Keith:2008:AIS

[675] Jonathan M. Keith, Dirk P. Kroese, and George Y. Sofronov. Adaptive independence samplers. Statistics and Computing, 18(4):409–420, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9070-2.

Cai:2008:MHA

[676] Bo Cai, Renate Meyer, and François Perron. Metropolis–Hastings algorithms with adaptive proposals. Statistics and Computing, 18(4):421– 433, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9051-5.

terBraak:2008:DEM

[677] Cajo J. F. ter Braak and Jasper A. Vrugt. Differential evolution Markov chain with snooker updater and fewer chains. Statistics and Computing, 18 (4):435–446, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-008-9104-9.pdf.

Cappe:2008:AIS

[678] Olivier Cappé, Randal Douc, Arnaud Guillin, Jean-Michel Marin, and Christian P. Robert. Adaptive importance sampling in general mixture classes. Statistics and Computing, 18 (4):447–459, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9059-x.

Cornebise:2008:AMS

[679] Julien Cornebise, Éric Moulines, and Jimmy Olsson. Adaptive methods for sequential importance sampling with application to state space models. Statistics and Computing, 18(4): 461–480, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9089-4.

Anonymous:2008:HCd

[680] Anonymous. Help & contacts. Statistics and Computing, 18(4):??, December 2008. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Knight:2009:NLT

[681] Marina I. Knight and Guy P. Nason. A ‘nondecimated’ lifting transform. Statistics and Computing, 19(1):1–16, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-008-
                                        9062-2.

Roy:2009:ICQ

[682] Debasish Roy, Geoff Nicholls, and Colin Fox. Imaging convex quadrilateral inclusions in uniform conductors from electrical boundary measurements. Statistics and Computing, 19(1):17–26, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9064-0.

Oh:2009:RWS

[683] Hee-Seok Oh, Donghoh Kim, and Yongdai Kim. Robust wavelet shrinkage using robust selection of thresholds. Statistics and Computing, 19(1):27–34, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9066-y.

Berlinet:2009:PAE

[684] A. Berlinet and C. Roland. Parabolic acceleration of the EM algorithm. Statistics and Computing, 19(1):35–47, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9067-x.

Glasbey:2009:TDG

[685] C. A. Glasbey. Two-dimensional generalisations of dynamic programming for image analysis. Statistics and Computing, 19(1):49–56, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9068-9.

Heaton:2009:ATS

[686] T. J. Heaton. Adaptive thresholding of sequences with locally variable strength. Statistics and Computing, 19(1):57–71, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9071-1.

Karlis:2009:MBC

[687] Dimitris Karlis and Anais Santourian. Model-based clustering with nonelliptically contoured distributions. Statistics and Computing, 19(1):73–83, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-008-
                                        9072-0.

Bernard-Michel:2009:GRS

[688] Caroline Bernard-Michel, Laurent Gardes, and Stéphane Girard. Gaussian regularized sliced inverse regression. Statistics and Computing, 19 (1):85–98, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9073-z.

Orawo:2009:BSA

[689] Luke Akong’o Orawo and J. Andrés Christen. Bayesian sequential analysis for multiple-arm clinical trials. Statistics and Computing, 19(1):99–109, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-008-
                                        9074-y.

Chopin:2009:BRJ

[690] Nicolas Chopin. Book review: Jim Albert: Bayesian computation with R. Statistics and Computing, 19 (1):111–112, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-008-9069-
8.

Anonymous:2009:HCa

[691] Anonymous. Help & contacts. Statistics and Computing, 19(1):??, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Furrer:2009:SMF

[692] Reinhard Furrer and Stephan R. Sain. Spatial model fitting for large datasets with applications to climate and microarray problems. Statistics and Computing, 19(2):113–128, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9075-x.

Chafai:2009:NME

[693] Djalil Chafaï and Didier Concordet. A new method for the estimation of variance matrix with prescribed zeros in nonlinear mixed effects models. Statistics and Computing, 19(2):129–138, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9076-9.

Wang:2009:EAC

[694] Hsiuying Wang. Exact average coverage probabilities and confidence coefficients of confidence intervals for discrete distributions. Statistics and Computing, 19(2):139–148, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9077-8.

Tsai:2009:NNN

[695] Henghsiu Tsai and Kung-Sik Chan. A note on the non-negativity of continuous-time ARMA and GARCH processes. Statistics and Computing, 19(2):149–153, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9078-7.

DiNardo:2009:NMF

[696] E. Di Nardo, G. Guarino, and D. Senato. A new method for fast computing unbiased estimators of cumulants. Statistics and Computing, 19 (2):155–165, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9080-0.

Kume:2009:FBD

[697] A. Kume and S. G. Walker. On the Fisher–Bingham distribution. Statistics and Computing, 19(2):167–172, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9081-z.

Lansangan:2009:PCA

[698] Joseph Ryan G. Lansangan and Erniel B. Barrios. Principal components analysis of nonstationary time series data. Statistics and Computing, 19(2):173–187, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9082-y.

Allingham:2009:BEQ

[699] D. Allingham, R. A. R. King, and K. L. Mengersen. Bayesian estimation of quantile distributions. Statistics and Computing, 19(2):189–201, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9083-x.

Saha:2009:GPD

[700] S. Saha, P. K. Mandal, Y. Boers, H. Driessen, and A. Bagchi. Gaussian proposal density using moment matching in SMC methods. Statistics and Computing, 19(2):203–208, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
008-9084-9.pdf.

Giancristofaro:2009:PAT

[701] Rosa Arboretti Giancristofaro, Stefano Bonnini, and Fortunato Pesarin. A permutation approach for testing heterogeneity in two-sample categorical variables. Statistics and Computing, 19(2):209–216, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9085-8.

Anonymous:2009:HCb

[702] Anonymous. Help & contacts. Statistics and Computing, 19(2):??, June 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Hazelton:2009:NDD

[703] Martin L. Hazelton and Berwin A. Turlach. Nonparametric density deconvolution by weighted kernel estimators. Statistics and Computing, 19(3): 217–228, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9086-7.

Dryden:2009:FPC

[704] Ian L. Dryden, Li Bai, Christopher J. Brignell, and Linlin Shen. Factored principal components analysis, with applications to face recognition. Statistics and Computing, 19(3): 229–238, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9087-6.

Tutz:2009:PRC

[705] Gerhard Tutz and Jan Ulbricht. Penalized regression with correlation-based penalty. Statistics and Computing, 19 (3):239–253, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9088-5.

Delicado:2009:MNL

[706] Pedro Delicado and Marcelo Smrekar. Measuring non-linear dependence for two random variables distributed along a curve. Statistics and Computing, 19 (3):255–269, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9090-y.

Chen:2009:ABQ

[707] Colin Chen and Keming Yu. Automatic Bayesian quantile regression curve fitting. Statistics and Computing, 19(3): 271–281, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9091-x.

Chatelain:2009:PLE

[708] Florent Chatelain, Sophie Lambert-Lacroix, and Jean-Yves Tourneret. Pairwise likelihood estimation for multivariate mixed Poisson models generated by Gamma intensities. Statistics and Computing, 19(3): 283–301, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-008-9092-9.

Armstrong:2009:BCM

[709] Helen Armstrong, Christopher K. Carter, Kin Foon Kevin Wong, and Robert Kohn. Bayesian covariance matrix estimation using a mixture of decomposable graphical models. Statistics and Computing, 19(3): 303–316, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9093-8.

Zhang:2009:NMC

[710] Chun-Xia Zhang and Jiang-She Zhang. A novel method for constructing ensemble classifiers. Statistics and Computing, 19(3):317–327, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-008-
                                        9094-7.

McGrory:2009:VBE

[711] C. A. McGrory, D. M. Titterington, R. Reeves, and A. N. Pettitt. Variational Bayes for estimating the parameters of a hidden Potts model. Statistics and Computing, 19 (3):329–340, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9095-6.

Cerioli:2009:CSM

[712] Andrea Cerioli, Marco Riani, and Anthony C. Atkinson. Controlling the size of multivariate outlier tests with the MCD estimator of scatter. Statistics and Computing, 19(3): 341–353, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9096-5.

Anonymous:2009:HCc

[713] Anonymous. Help & contacts. Statistics and Computing, 19(3):??, September 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Yang:2009:PSD

[714] Ying Yang. Penalized semiparametric density estimation. Statistics and Computing, 19(4):355–366, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9097-4.

Cuesta-Albertos:2009:PBT

[715] Juan A. Cuesta-Albertos, Antonio Cuevas, and Ricardo Fraiman. On projection-based tests for directional and compositional data. Statistics and Computing, 19(4):367–380, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9098-3.

Maruotti:2009:SAH

[716] Antonello Maruotti and Tobias Rydén. A semiparametric approach to hidden Markov models under longitudinal observations. Statistics and Computing, 19(4):381–393, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9099-2.

Lunn:2009:GRJ

[717] David J. Lunn, Nicky Best, and John C. Whittaker. Generic reversible jump MCMC using graphical models. Statistics and Computing, 19(4): 395–408, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9100-0.

Fan:2009:AER

[718] Y. Fan, G. W. Peters, and S. A. Sisson. Automating and evaluating reversible jump MCMC proposal distributions. Statistics and Computing, 19 (4):409–421, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9101-z.

Renshaw:2009:STG

[719] Eric Renshaw and Carlos Comas. Space-time generation of high intensity patterns using growth-interaction processes. Statistics and Computing, 19 (4):423–437, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-008-9102-y.

Ridout:2009:GRN

[720] M. S. Ridout. Generating random numbers from a distribution specified by its Laplace transform. Statistics and Computing, 19(4):439–450, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-008-
                                        9103-x.

Viroli:2009:BIN

[721] Cinzia Viroli. Bayesian inference in non-Gaussian factor analysis. Statistics and Computing, 19(4):451–463, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9105-8.

Richard:2009:SAE

[722] Frédéric J. P. Richard, Adeline M. M. Samson, and Charles A. Cuénod. A SAEM algorithm for the estimation of template and deformation parameters in medical image sequences. Statistics and Computing, 19(4):465–478, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9106-7.

Fruhwirth-Schnatter:2009:IAM

[723] Sylvia Frühwirth-Schnatter, Rudolf Frühwirth, Leonhard Held, and Håvard Rue. Improved auxiliary mixture sampling for hierarchical models of non-Gaussian data. Statistics and Computing, 19(4):479–492, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-008-
9109-4.

Anonymous:2009:HCd

[724] Anonymous. Help & contacts. Statistics and Computing, 19(4):??, December 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Gramacy:2010:IT

[725] Robert Gramacy, Richard Samworth, and Ruth King. Importance tempering. Statistics and Computing, 20(1):1–7, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11222-008-
                                        9108-5.

Greselin:2010:CME

[726] F. Greselin and S. Ingrassia. Constrained monotone EM algorithms for mixtures of multivariate t distributions. Statistics and Computing, 20(1):9–22, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-008-
                                        9112-9.

Capizzi:2010:ERL

[727] Giovanna Capizzi and Guido Masarotto. Evaluation of the run-length distribution for a combined Shewhart– EWMA control chart. Statistics and Computing, 20(1):23–33, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-008-
                                        9113-8.

Charpentier:2010:BKQ

[728] Arthur Charpentier and Abder Oulidi. Beta kernel quantile estimators of heavy-tailed loss distributions. Statistics and Computing, 20(1):35–55, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-009-
                                        9114-2.

Wicker:2010:PSA

[729] Nicolas Wicker. Perfect sampling algorithm for small m × n contingency tables. Statistics and Computing, 20 (1):57–61, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9115-1.

Blum:2010:NLR

[730] Michael G. B. Blum and Olivier François. Non-linear regression models for approximate Bayesian computation. Statistics and Computing, 20 (1):63–73, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9116-0.

Wang:2010:MLC

[731] Yong Wang. Maximum likelihood computation for fitting semiparametric mixture models. Statistics and Computing, 20(1):75–86, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9117-z.

Tang:2010:ACI

[732] Man-Lai Tang and Maozai Tian. Approximate confidence interval construction for risk difference under inverse sampling. Statistics and Computing, 20(1):87–98, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9118-y.

Critchley:2010:RAC

[733] Frank Critchley, Michaël Schyns, Gentiane Haesbroeck, Cécile Fauconnier, Guobing Lu, Richard A. Atkinson, and Dong Qian Wang. A relaxed approach to combinatorial problems in robustness and diagnostics. Statistics and Computing, 20(1):99–115, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-009-
                                        9119-x.

Anonymous:2010:HCa

[734] Anonymous. Help & contacts. Statistics and Computing, 20(1):??, January 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Tutz:2010:GER

[735] Gerhard Tutz. Guest editorial: Regularisation methods in regression and classification. Statistics and Computing, 20(2):117–118, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        010-9171-6.pdf.

Buhlmann:2010:TBI

[736] Peter Bühlmann and Torsten Hothorn. Twin boosting: improved feature selection and prediction. Statistics and Computing, 20(2):119–138, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-009-
                                        9148-5.

Schmid:2010:ERT

[737] Matthias Schmid, Sergej Potapov, Annette Pfahlberg, and Torsten Hothorn. Estimation and regularization techniques for regression models with multidimensional prediction functions. Statistics and Computing, 20 (2):139–150, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9162-7.

Porzelius:2010:SRT

[738] Christine Porzelius, Martin Schumacher, and Harald Binder. Sparse regression techniques in low-dimensional survival data settings. Statistics and Computing, 20(2):151–163, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9155-6.

Xu:2010:RBV

[739] Jinfeng Xu, Chenlei Leng, and Zhiliang Ying. Rank-based variable selection with censored data. Statistics and Computing, 20(2):165–176, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9126-y.

Beran:2010:LSS

[740] Rudolf Beran and Lutz Dümbgen. Least squares and shrinkage estimation under bimonotonicity constraints. Statistics and Computing, 20(2):177–189, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9124-0.

Marx:2010:BMM

[741] Brian D. Marx, Paul H. C. Eilers, Jutta Gampe, and Roland Rau. Bilinear modulation models for seasonal tables of counts. Statistics and Computing, 20(2):191–202, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-009-9144-9.pdf.

Fahrmeir:2010:BRS

[742] Ludwig Fahrmeir, Thomas Kneib, and Susanne Konrath. Bayesian regularisation in structured additive regression: a unifying perspective on shrinkage, smoothing and predictor selection. Statistics and Computing, 20 (2):203–219, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9158-3.

Hans:2010:MUV

[743] Chris Hans. Model uncertainty and variable selection in Bayesian lasso regression. Statistics and Computing, 20(2):221–229, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9160-9.

Obozinski:2010:JCS

[744] Guillaume Obozinski, Ben Taskar, and Michael I. Jordan. Joint covariate selection and joint subspace selection for multiple classification problems. Statistics and Computing, 20 (2):231–252, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-008-9111-x.pdf.

Hebiri:2010:SCP

[745] Mohamed Hebiri. Sparse conformal predictors. Statistics and Computing, 20(2):253–266, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9167-2.

Anonymous:2010:HCb

[746] Anonymous. Help & contacts. Statistics and Computing, 20(2):??, April 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Rubinshtein:2010:OLP

[747] Evgenia Rubinshtein and Anuj Srivastava. Optimal linear projections for enhancing desired data statistics. Statistics and Computing, 20(3):267–282, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9120-4.

Bevilacqua:2010:WCL

[748] M. Bevilacqua, J. Mateu, E. Porcu, H. Zhang, and A. Zini. Weighted composite likelihood-based tests for space-time separability of covariance functions. Statistics and Computing, 20(3):283–293, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9121-3.

Lee:2010:OBV

[749] Youngjo Lee and Il Do Ha. Orthodox BLUP versus h-likelihood methods for inferences about random effects in Tweedie mixed models. Statistics and Computing, 20(3):295–303, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9122-2.

Myllymaki:2010:BIG

[750] Mari Myllymäki and Antti Penttinen. Bayesian inference for Gaussian excursion set generated Cox processes with set-marking. Statistics and Computing, 20(3):305–315, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9123-1.

Rogers:2010:SPA

[751] Simon Rogers, Mark Girolami, and Tamara Polajnar. Semi-parametric analysis of multi-rater data. Statistics and Computing, 20(3):317–334, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9125-z.

Liechty:2010:MSS

[752] Merrill W. Liechty and Matthew Tibbits. Multivariate sufficient statistics using Kronecker products. Statistics and Computing, 20(3):335–341, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9127-x.

Lin:2010:RMM

[753] Tsung-I Lin. Robust mixture modeling using multivariate skew t distributions. Statistics and Computing, 20(3):343–356, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9128-9.

Sperrin:2010:PRS

[754] M. Sperrin, T. Jaki, and E. Wit. Probabilistic relabelling strategies for the label switching problem in Bayesian mixture models. Statistics and Computing, 20(3):357–366, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9129-8.

Roca-Pardinas:2010:FEG

[755] Javier Roca-Pardiñas and Stefan Sperlich. Feasible estimation in generalized structured models. Statistics and Computing, 20(3):367–379, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
009-9130-2.pdf.

Lee:2010:PPI

[756] Eun-Kyung Lee and Dianne Cook. A projection pursuit index for large p small n data. Statistics and Computing, 20(3):381–392, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9131-1.

Anonymous:2010:HCc

[757] Anonymous. Help & contacts. Statistics and Computing, 20(3):??, July 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Sandri:2010:ACB

[758] Marco Sandri and Paola Zuccolotto. Analysis and correction of bias in total decrease in node impurity measures for tree-based algorithms. Statistics and Computing, 20(4):393–407, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9132-0.

Puggioni:2010:AST

[759] Gavino Puggioni and Alan E. Gelfand. Analyzing space-time sensor network data under suppression and failure in transmission. Statistics and Computing, 20(4):409–419, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9133-z.

Audrino:2010:SPF

[760] Francesco Audrino and Dominik Colangelo. Semi-parametric forecasts of the implied volatility surface using regression trees. Statistics and Computing, 20(4):421–434, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9134-y.

Iqbal:2010:MES

[761] Farhat Iqbal and Kanchan Mukherjee. M-estimators of some GARCHtype models; computation and application. Statistics and Computing, 20 (4):435–445, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9135-x.

Ahn:2010:EMC

[762] Kwang Woo Ahn and Kung-Sik Chan. Efficient Markov chain Monte Carlo with incomplete multinomial data. Statistics and Computing, 20 (4):447–456, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9136-9.

Leisch:2010:NGS

[763] Friedrich Leisch. Neighborhood graphs, stripes and shadow plots for cluster visualization. Statistics and Computing, 20(4):457–469, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9137-8.

Scrucca:2010:DRM

[764] Luca Scrucca. Dimension reduction for model-based clustering. Statistics and Computing, 20(4):471–484, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9138-7.

Nott:2010:SBL

[765] David J. Nott and Li Jialiang. A sign based loss approach to model selection in nonparametric regression. Statistics and Computing, 20(4):485–498, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11222-009-
                                        9139-6.

Gijbels:2010:PSR

[766] I. Gijbels and A. Verhasselt. Psplines regression smoothing and difference type of penalty. Statistics and Computing, 20(4):499–511, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-009-
                                        9140-0.

Anonymous:2010:HCd

[767] Anonymous. Help & contacts. Statistics and Computing, 20(4):??, October 2010. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Weiss:2011:RGC

[768] Christian H. Weiß. Rule generation for categorical time series with Markov assumptions. Statistics and Computing, 21(1):1–16, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-009-9141-z.

Kojadinovic:2011:GFT

[769] Ivan Kojadinovic and Jun Yan. A goodness-of-fit test for multivariate multiparameter copulas based on multiplier central limit theorems. Statistics and Computing, 21(1):17–30, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-009-
                                        9142-y.

Poleto:2011:MDM

[770] Frederico Z. Poleto, Julio M. Singer, and Carlos Daniel Paulino. Missing data mechanisms and their implications on the analysis of categorical data. Statistics and Computing, 21(1):31–43, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9143-x.

Basso:2011:PTU

[771] Dario Basso and Luigi Salmaso. A permutation test for umbrella alternatives. Statistics and Computing, 21(1):45–54, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9145-8.

Krampe:2011:AMS

[772] Anne Krampe, Maria Kateri, and Sonja Kuhnt. Asymmetry models for square contingency tables: exact tests via algebraic statistics. Statistics and Computing, 21(1):55–67, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9146-7.

Lau:2011:MCM

[773] John W. Lau and Mike K. P. So. A Monte Carlo Markov chain algorithm for a class of mixture time series models. Statistics and Computing, 21(1):69–81, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-009-9147-6.

Duan:2011:SEI

[774] Jin-Chuan Duan and Andras Fulop. A stable estimator of the information matrix under EM for dependent data. Statistics and Computing, 21(1):83–91, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-009-
                                        9149-4.

Kalli:2011:SSM

[775] Maria Kalli, Jim E. Griffin, and Stephen G. Walker. Slice sampling mixture models. Statistics and Computing, 21(1):93–105, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-009-9150-y.

Gabriel:2011:ETZ

[776] Edith Gabriel, Denis Allard, and Jean-Noël Bacro. Estimating and testing zones of abrupt change for spatial data. Statistics and Computing, 21 (1):107–120, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-009-9151-x.

Dupuy:2011:NPE

[777] J.-F. Dupuy, J.-M. Loubes, and E. Maza. Non parametric estimation of the structural expectation of a stochastic increasing function. Statistics and Computing, 21(1):121–136, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9152-9.

Anonymous:2011:HCa

[778] Anonymous. Help & contacts. Statistics and Computing, 21(1):??, January 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Fushiki:2011:EPE

[779] Tadayoshi Fushiki. Estimation of prediction error by using K-fold crossvalidation. Statistics and Computing, 21(2):137–146, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9153-8.

Butler:2011:EDC

[780] Ronald W. Butler and Robert L. Paige. Exact distributional computations for Roy’s statistic and the largest eigenvalue of a Wishart distribution. Statistics and Computing, 21(2):147–157, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9154-7.

Kauermann:2011:FVE

[781] Göran Kauermann and Michael Wegener. Functional variance estimation using penalized splines with principal component analysis. Statistics and Computing, 21(2):159–171, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-009-
9156-5.

Bellio:2011:RLI

[782] Ruggero Bellio and Alessandra R. Brazzale. Restricted likelihood inference for generalized linear mixed models. Statistics and Computing, 21(2):173–183, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-009-9157-4.

Alfo:2011:FMM

[783] Marco Alfò, Antonello Maruotti, and Giovanni Trovato. A finite mixture model for multivariate counts under endogenous selectivity. Statistics and Computing, 21(2):185–202, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-009-
                                        9159-2.

Domijan:2011:BKP

[784] Katarina Domijan and Simon P. Wilson. Bayesian kernel projections for classification of high dimensional data. Statistics and Computing, 21 (2):203–216, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-009-9161-8.

Fearnhead:2011:EBA

[785] Paul Fearnhead and Zhen Liu. Efficient Bayesian analysis of multiple changepoint models with dependence across segments. Statistics and Computing, 21(2):217–229, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9163-6.

Wagner:2011:BES

[786] Helga Wagner. Bayesian estimation and stochastic model specification search for dynamic survival models. Statistics and Computing, 21 (2):231–246, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-009-9164-5.

Chen:2011:SMP

[787] Ray-Bing Chen, Chi-Hsiang Chu, Te-You Lai, and Ying Nian Wu. Stochastic matching pursuit for Bayesian variable selection. Statistics and Computing, 21(2):247–259, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-009-9165-4.pdf.

Zhou:2011:QNA

[788] Hua Zhou, David Alexander, and Kenneth Lange. A quasi-Newton acceleration for high-dimensional optimization algorithms. Statistics and Computing, 21(2):261–273, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-009-9166-3.pdf.

Chopin:2011:FST

[789] Nicolas Chopin. Fast simulation of truncated Gaussian distributions. Statistics and Computing, 21 (2):275–288, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-009-9168-1.

Robert:2011:BRJ

[790] Christian P. Robert. Book review: James E. Gentle: Computational statistics (Statistics and Computing Series). Statistics and Computing, 21(2):289–291, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-010-9189-9.

Allard:2011:BRG

[791] Denis Allard. Book review: A. E. Gelfand, P. J. Diggle, M. Fuentes, P. Guttorp (eds.): Handbook of spatial statistics. Statistics and Computing, 21(2):293–294, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-010-9211-
                                        2.

Anonymous:2011:HCb

[792] Anonymous. Help & contacts. Statistics and Computing, 21(2):??, April 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Menardi:2011:DBS

[793] Giovanna Menardi. Density-based silhouette diagnostics for clustering methods. Statistics and Computing, 21(3):295–308, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-010-9169-0.

Bove:2011:BFP

[794] Daniel Sabanés Bové and Leonhard Held. Bayesian fractional polynomials. Statistics and Computing, 21 (3):309–324, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-010-9170-7.

Zhang:2011:SDR

[795] Li-Mei Zhang, Li-Ping Zhu, and Li-Xing Zhu. Sufficient dimension reduction in regressions through cumulative Hessian directions. Statistics and Computing, 21(3):325–334, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9172-5.

Jasra:2011:SBA

[796] Ajay Jasra and Christopher C. Holmes. Stochastic boosting algorithms. Statistics and Computing, 21(3):335–347, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9173-4.

Corrado:2011:EDM

[797] Charles J. Corrado. The exact distribution of the maximum, minimum and the range of multinomial/Dirichlet and multivariate hypergeometric frequencies. Statistics and Computing, 21(3):349–359, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-010-9174-3.

Andrews:2011:EMM

[798] Jeffrey L. Andrews and Paul D. Mc-Nicholas. Extending mixtures of multivariate t-factor analyzers. Statistics and Computing, 21(3):361–373, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9175-2.

Liang:2011:AES

[799] Faming Liang. Annealing evolutionary stochastic approximation Monte Carlo for global optimization. Statistics and Computing, 21(3):375–393, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9176-1.

Chen:2011:CER

[800] Cathy W. S. Chen, Jennifer S. K. Chan, Richard Gerlach, and William Y. L. Hsieh. A comparison of estimators for regression models with change points. Statistics and Computing, 21(3):395–414, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-010-9177-0.

Tibbits:2011:PMS

[801] Matthew M. Tibbits, Murali Haran, and John C. Liechty. Parallel multivariate slice sampling. Statistics and Computing, 21(3):415–430, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s11222-010-
9178-z.

Friel:2011:CUD

[802] N. Friel and A. N. Pettitt. Classification using distance nearest neighbours. Statistics and Computing, 21 (3):431–437, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-010-9179-y.

Meligkotsidou:2011:FNH

[803] Loukia Meligkotsidou and Petros Dellaportas. Forecasting with nonhomogeneous hidden Markov models. Statistics and Computing, 21 (3):439–449, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-010-9180-5.

Ghosh:2011:GSM

[804] Samiran Ghosh. On the grouped selection and model complexity of the adaptive elastic net. Statistics and Computing, 21(3):451–462, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-010-9181-4.

Anonymous:2011:HCc

[805] Anonymous. Help & contacts. Statistics and Computing, 21(3):??, July 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Atchade:2011:CFE

[806] Yves F. Atchadé. A computational framework for empirical Bayes inference. Statistics and Computing, 21 (4):463–473, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-010-9182-3.

Yu:2011:ODC

[807] Yaming Yu. D-optimal designs via a cocktail algorithm. Statistics and Computing, 21(4):475–481, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
010-9183-2.pdf.

Rodriguez-Alvarez:2011:RCC

[808] María Xosé Rodríguez-Álvarez, Javier Roca-Pardiñas, and Carmen Cadarso-Suárez. ROC curve and covariates: extending induced methodology to the non-parametric framework. Statistics and Computing, 21(4):483–499, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9184-1.

Wang:2011:NFS

[809] Chun-Chao Wang and Yi-Ting Hwang. A new functional statistic for multivariate normality. Statistics and Computing, 21(4):501–509, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9186-z.

Viroli:2011:FMM

[810] Cinzia Viroli. Finite mixtures of matrix normal distributions for classifying three-way data. Statistics and Computing, 21(4):511–522, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9188-x.

Soffritti:2011:MLR

[811] Gabriele Soffritti and Giuliano Galimberti. Multivariate linear regression with non-normal errors: a solution based on mixture models. Statistics and Computing, 21(4):523–536, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9190-3.

Chiquet:2011:IMG

[812] Julien Chiquet, Yves Grandvalet, and Christophe Ambroise. Inferring multiple graphical structures. Statistics and Computing, 21(4):537–553, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9191-2.

Atchade:2011:TOS

[813] Yves F. Atchadé, Gareth O. Roberts, and Jeffrey S. Rosenthal. Towards optimal scaling of Metropolis-coupled Markov chain Monte Carlo. Statistics and Computing, 21(4):555–568, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9192-1.

Hwang:2011:EFC

[814] Yi-Ting Hwang, Shih-Kai Chu, and Shyh-Tyan Ou. Evaluations of FDRcontrolling procedures in multiple hypothesis testing. Statistics and Computing, 21(4):569–583, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9193-0.

Garcia-Escudero:2011:ENG

[815] L. A. García-Escudero, A. Gordaliza, C. Matrán, and A. Mayo-Iscar. Exploring the number of groups in robust model-based clustering. Statistics and Computing, 21(4):585–599, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9194-z.

Ma:2011:GEE

[816] Jun Ma, Sigurbjorg Gudlaugsdottir, and Graham Wood. Generalized EM estimation for semi-parametric mixture distributions with discretized nonparametric component. Statistics and Computing, 21(4):601–612, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9195-y.

Arlot:2011:SMH

[817] Sylvain Arlot and Alain Celisse. Segmentation of the mean of heteroscedastic data via cross-validation. Statistics and Computing, 21(4):613–632, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9196-x.

Martino:2011:GAR

[818] Luca Martino and Joaquín Míguez. A generalization of the adaptive rejection sampling algorithm. Statistics and Computing, 21(4):633–647, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9197-9.

Brewer:2011:DNS

[819] Brendon J. Brewer, Livia B. Pártay, and Gábor Csányi. Diffusive nested sampling. Statistics and Computing, 21(4):649–656, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-010-9198-8.pdf.

Gurevich:2011:TSE

[820] Gregory Gurevich and Albert Vexler. A two-sample empirical likelihood ratio test based on samples entropy. Statistics and Computing, 21(4):657–670, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9199-7.

Cho:2011:MIT

[821] Haeran Cho and Piotr Fryzlewicz. Multiscale interpretation of taut string estimation and its connection to unbalanced Haar wavelets. Statistics and Computing, 21(4):671–681, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9200-5.

Anonymous:2011:HCd

[822] Anonymous. Help & contacts. Statistics and Computing, 21(4):??, October 2011. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Botev:2012:EMC

[823] Zdravko I. Botev and Dirk P. Kroese. Efficient Monte Carlo simulation via the generalized splitting method. Statistics and Computing, 22(1):1–16, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9201-4.

Chu:2012:EJL

[824] Chih-Kang Chu, Jhao-Siang Siao, Lih-Chung Wang, and Wen-Shuenn Deng. Estimation of 2D jump location curve and 3D jump location surface in nonparametric regression. Statistics and Computing, 22(1):17–31, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9203-2.

Lo:2012:FMM

[825] Kenneth Lo and Raphael Gottardo. Flexible mixture modeling via the multivariate t distribution with the Box– Cox transformation: an alternative to the skew-t distribution. Statistics and Computing, 22(1):33–52, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9204-1.

Dinwoodie:2012:SIS

[826] Ian H. Dinwoodie. Sequential importance sampling of binary sequences. Statistics and Computing, 22(1):53–63, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9205-0.

Behrens:2012:TTT

[827] Gundula Behrens, Nial Friel, and Merrilee Hurn. Tuning tempered transitions. Statistics and Computing, 22 (1):65–78, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-010-9206-z.

Papageorgiou:2012:MGL

[828] Georgios Papageorgiou and John Hinde. Multivariate generalized linear mixed models with semi-nonparametric and smooth nonparametric random effects densities. Statistics and Computing, 22(1):79–92, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-010-9207-
y.

Kottas:2012:BSM

[829] Athanasios Kottas and Gilbert W. Fellingham. Bayesian semiparametric modeling and inference with mixtures of symmetric distributions. Statistics and Computing, 22(1):93–106, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        010-9208-x.pdf.

Forster:2012:RJM

[830] Jonathan J. Forster, Roger C. Gill, and Antony M. Overstall. Reversible jump methods for generalised linear models and generalised linear mixed models. Statistics and Computing, 22 (1):107–120, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-010-9210-3.

Meza:2012:ENM

[831] Cristian Meza, Felipe Osorio, and Rolando De la Cruz. Estimation in nonlinear mixed-effects models using heavy-tailed distributions. Statistics and Computing, 22(1):121–139, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9212-1.

Farcomeni:2012:QRL

[832] Alessio Farcomeni. Quantile regression for longitudinal data based on latent Markov subject-specific parameters. Statistics and Computing, 22 (1):141–152, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-010-9213-0.

Slawski:2012:SEN

[833] Martin Slawski. The structured elastic net for quantile regression and support vector classification. Statistics and Computing, 22(1):153–168, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9214-z.

Piras:2012:STS

[834] Gianfranco Piras and Nancy Lozano-Gracia. Spatial J-test: some Monte Carlo evidence. Statistics and Computing, 22(1):169–183, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-010-9215-
y.

Wu:2012:NVB

[835] Burton Wu, Clare A. McGrory, and Anthony N. Pettitt. A new variational Bayesian algorithm with application to human mobility pattern modeling. Statistics and Computing, 22 (1):185–203, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-010-9217-9.

Preston:2012:ATD

[836] S. P. Preston and Andrew T. A. Wood. Approximation of transition densities of stochastic differential equations by saddlepoint methods applied to smalltime Ito–Taylor sample-path expansions. Statistics and Computing, 22 (1):205–217, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-010-9218-8.

Stadler:2012:MVS

[837] Nicolas Städler and Peter Bühlmann. Missing values: sparse inverse covariance estimation and an extension to sparse regression. Statistics and Computing, 22(1):219–235, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9219-7.

Hooker:2012:PBR

[838] Giles Hooker and Saharon Rosset. Prediction-based regularization using data augmented regression. Statistics and Computing, 22(1):237–249, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9220-1.

Redd:2012:COS

[839] Andrew Redd. A comment on the orthogonalization of B-spline basis functions and their derivatives. Statistics and Computing, 22(1):251–257, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9221-0.

Berrendero:2012:MUT

[840] José R. Berrendero, Antonio Cuevas, and Beatriz Pateiro-López. A multivariate uniformity test for the case of unknown support. Statistics and Computing, 22(1):259–271, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s11222-010-
9222-z.

Ning:2012:CSN

[841] Jianhui Ning and Philip E. Cheng. A comparison study of nonparametric imputation methods. Statistics and Computing, 22(1):273–285, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9223-y.

Ho:2012:MLI

[842] Hsiu J. Ho, Saumyadipta Pyne, and Tsung I. Lin. Maximum likelihood inference for mixtures of skew Studentt-normal distributions through practical EM-type algorithms. Statistics and Computing, 22(1):287–299, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9225-9.

Anonymous:2012:HCa

[843] Anonymous. Help & contacts. Statistics and Computing, 22(1):??, January 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Yao:2012:MBL

[844] Weixin Yao. Model based labeling for mixture models. Statistics and Computing, 22(2):337–347, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9226-8.

Caporale:2012:DVS

[845] Guglielmo Maria Caporale, Juncal Cunado, and Luis A. Gil-Alana. Deterministic versus stochastic seasonal fractional integration and structural breaks. Statistics and Computing, 22 (2):349–358, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9227-2.

Lamnisos:2012:CVP

[846] D. Lamnisos, J. E. Griffin, and M. F. J. Steel. Cross-validation prior choice in Bayesian probit regression with many covariates. Statistics and Computing, 22(2):359–373, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9228-1.

Kaban:2012:NPD

[847] Ata Kabán. Non-parametric detection of meaningless distances in high dimensional data. Statistics and Computing, 22(2):375–385, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9229-0.

Harris:2012:VDN

[848] T. J. Harris and W. Yu. Variance decompositions of nonlinear time series using stochastic simulation and sensitivity analysis. Statistics and Computing, 22(2):387–396, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11222-011-
9230-7.

Chen:2012:DCC

[849] Jian Chen and Jeffrey S. Rosenthal. Decrypting classical cipher text using Markov chain Monte Carlo. Statistics and Computing, 22(2):397–413, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9232-5.

Wyse:2012:BCC

[850] Jason Wyse and Nial Friel. Block clustering with collapsed latent block models. Statistics and Computing, 22 (2):415–428, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9233-4.

Campbell:2012:SFT

[851] David Campbell and Russell J. Steele. Smooth functional tempering for nonlinear differential equation models. Statistics and Computing, 22 (2):429–443, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9234-3.

Trendafilov:2012:DDI

[852] Nickolay T. Trendafilov. DIND-SCAL: direct INDSCAL. Statistics and Computing, 22(2):445–454, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9235-2.

Baudry:2012:SHO

[853] Jean-Patrick Baudry, Cathy Maugis, and Bertrand Michel. Slope heuristics: overview and implementation. Statistics and Computing, 22(2):455–470, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9236-1.

Boudt:2012:GRC

[854] Kris Boudt, Jonathan Cornelissen, and Christophe Croux. The Gaussian rank correlation estimator: robustness properties. Statistics and Computing, 22 (2):471–483, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9237-0.

Lung-Yut-Fong:2012:DDL

[855] A. Lung-Yut-Fong, C. Lévy-Leduc, and O. Cappé. Distributed detection/localization of change-points in high-dimensional network traffic data. Statistics and Computing, 22 (2):485–496, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9240-5.

Nott:2012:VAH

[856] David J. Nott, Minh-Ngoc Tran, and Chenlei Leng. Variational approximation for heteroscedastic linear models and matching pursuit algorithms. Statistics and Computing, 22(2):497–512, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9243-2.

Amrein:2012:REP

[857] Michael Amrein and Hans R. Künsch. Rate estimation in partially observed Markov jump processes with measurement errors. Statistics and Computing, 22(2):513–526, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9244-1.

Yanagihara:2012:NIO

[858] Hirokazu Yanagihara. A non-iterative optimization method for smoothness in penalized spline regression. Statistics and Computing, 22(2):527–544, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9245-0.

Tutz:2012:NEL

[859] Gerhard Tutz and Sebastian Petry. Nonparametric estimation of the link function including variable selection. Statistics and Computing, 22 (2):545–561, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9246-z.

Achilleos:2012:LBS

[860] Achilleas Achilleos and Aurore Delaigle. Local bandwidth selectors for deconvolution kernel density estimation. Statistics and Computing, 22 (2):563–577, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9247-y.

Caron:2012:LCD

[861] François Caron, Arnaud Doucet, and Raphael Gottardo. On-line changepoint detection and parameter estimation with application to genomic data. Statistics and Computing, 22 (2):579–595, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9248-x.

Baghishani:2012:HDC

[862] Hossein Baghishani, Håvard Rue, and Mohsen Mohammadzadeh. On a hybrid data cloning method and its application in generalized linear mixed models. Statistics and Computing, 22 (2):597–613, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9254-z.

Lindstrom:2012:RBS

[863] Erik Lindström. A regularized bridge sampler for sparsely sampled diffusions. Statistics and Computing, 22 (2):615–623, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9255-y.

Anonymous:2012:HCb

[864] Anonymous. Help & contacts. Statistics and Computing, 22(2):??, March 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Antoniadis:2012:MCE

[865] Anestis Antoniadis and Alberto Pasanisi. Modeling of computer experiments for uncertainty propagation and sensitivity analysis. Statistics and Computing, 22(3):677–679, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
011-9282-8.pdf.

Pronzato:2012:DCE

[866] Luc Pronzato and Werner G. Müller. Design of computer experiments: space filling and beyond. Statistics and Computing, 22(3):681–701, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9242-3.

Auffray:2012:MDN

[867] Yves Auffray, Pierre Barbillon, and Jean-Michel Marin. Maximin design on non hypercube domains and kernel interpolation. Statistics and Computing, 22(3):703–712, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9273-9.

Gramacy:2012:CNM

[868] Robert B. Gramacy and Herbert K. H. Lee. Cases for the nugget in modeling computer experiments. Statistics and Computing, 22(3):713–722, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11222-010-
9224-x.

Muehlenstaedt:2012:DDK

[869] Thomas Muehlenstaedt, Olivier Roustant, Laurent Carraro, and Sonja Kuhnt. Data-driven kriging models based on FANOVA-decomposition. Statistics and Computing, 22(3): 723–738, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9259-7.

Sambucini:2012:CRS

[870] Valeria Sambucini. Confidence regions for the stationary point of a quadratic response surface based on the asymptotic distribution of its MLE. Statistics and Computing, 22(3):739–751, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9202-3.

Cohen:2012:ASD

[871] Serge Cohen, Sébastien Déjean, and Sébastien Gadat. Adaptive sequential design for regression on multiresolution bases. Statistics and Computing, 22(3):753–772, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-011-9267-
7.

Bect:2012:SDC

[872] Julien Bect, David Ginsbourger, Ling Li, Victor Picheny, and Emmanuel Vazquez. Sequential design of computer experiments for the estimation of a probability of failure. Statistics and Computing, 22(3):773–793, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9241-4.

Cerou:2012:SMC

[873] F. Cérou, P. Del Moral, T. Furon, and A. Guyader. Sequential Monte Carlo for rare event estimation. Statistics and Computing, 22(3):795–808, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9231-6.

Zuniga:2012:AAD

[874] Miguel Munoz Zuniga, Josselin Garnier, Emmanuel Remy, and Etienne de Rocquigny. Analysis of adaptive directional stratification for the controlled estimation of rare event probabilities. Statistics and Computing, 22(3):809–821, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9277-5.

Jourdan:2012:GSA

[875] Astrid Jourdan. Global sensitivity analysis using complex linear models. Statistics and Computing, 22 (3):823–831, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9239-y.

Marrel:2012:GSA

[876] Amandine Marrel, Bertrand Iooss, Sébastien Da Veiga, and Mathieu Ribatet. Global sensitivity analysis of stochastic computer models with joint metamodels. Statistics and Computing, 22(3):833–847, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9274-8.

Anonymous:2012:HCc

[877] Anonymous. Help & contacts. Statistics and Computing, 22(3):??, May 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Dellaportas:2012:CGM

[878] Petros Dellaportas and Mohsen Pourahmadi. Cholesky-GARCH models with applications to finance. Statistics and Computing, 22(4):849–855, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9251-2.

Cao:2012:EGS

[879] Jiguo Cao. Estimating generalized semiparametric additive models using parameter cascading. Statistics and Computing, 22(4):857–865, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9252-1.

Sexton:2012:BCM

[880] Joseph Sexton and Petter Laake. Boosted coefficient models. Statistics and Computing, 22(4):867–876, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
011-9253-0.pdf.

Knight:2012:SEL

[881] Marina I. Knight, Matthew A. Nunes, and Guy P. Nason. Spectral estimation for locally stationary time series with missing observations. Statistics and Computing, 22(4):877–895, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9256-x.

Chopin:2012:FEM

[882] Nicolas Chopin, Tony Lelièvre, and Gabriel Stoltz. Free energy methods for Bayesian inference: efficient exploration of univariate Gaussian mixture posteriors. Statistics and Computing, 22(4):897–916, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9257-9.

Rigaill:2012:EPD

[883] G. Rigaill, E. Lebarbier, and S. Robin. Exact posterior distributions and model selection criteria for multiple change-point detection problems. Statistics and Computing, 22 (4):917–929, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9258-8.

Mrkvicka:2012:GFT

[884] Tomás Mrkvicka, Samuel Soubeyrand, and Joël Chadoeuf. Goodness-of-fit test of the mark distribution in a point process with non-stationary marks. Statistics and Computing, 22(4):931–943, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9263-y.

Paquet:2012:HMO

[885] Ulrich Paquet, Blaise Thomson, and Ole Winther. A hierarchical model for ordinal matrix factorization. Statistics and Computing, 22(4):945–957, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9264-x.

Lee:2012:MRA

[886] Woojoo Lee and Youngjo Lee. Modifications of REML algorithm for HGLMs. Statistics and Computing, 22(4):959–966, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9265-9.

Meintanis:2012:GFT

[887] Simos G. Meintanis and Jochen Einbeck. Goodness-of-fit tests in semilinear models. Statistics and Computing, 22(4):967–979, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9266-8.

Aston:2012:IDM

[888] J. A. D. Aston, J. Y. Peng, and D. E. K. Martin. Implied distributions in multiple change point problems. Statistics and Computing, 22(4):981–993, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9268-6.

Gilmour:2012:JLD

[889] Steven G. Gilmour. John Lawson: Design and analysis of experiments with SAS. Statistics and Computing, 22(4):995–996, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-011-9238-
z.

Anonymous:2012:HCd

[890] Anonymous. Help & contacts. Statistics and Computing, 22(4):??, July 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Vihola:2012:RAM

[891] Matti Vihola. Robust adaptive Metropolis algorithm with coerced acceptance rate. Statistics and Computing, 22(5):997–1008, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9269-5.

DelMoral:2012:ASM

[892] Pierre Del Moral, Arnaud Doucet, and Ajay Jasra. An adaptive sequential Monte Carlo method for approximate Bayesian computation. Statistics and Computing, 22(5):1009–1020, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9271-y.

Andrews:2012:MBC

[893] Jeffrey L. Andrews and Paul D. Mc-Nicholas. Model-based clustering, classification, and discriminant analysis via mixtures of multivariate tdistributions. Statistics and Computing, 22(5):1021–1029, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9272-x.

Chan:2012:ICE

[894] Joshua C. C. Chan and Dirk P. Kroese. Improved cross-entropy method for estimation. Statistics and Computing, 22(5):1031–1040, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9275-7.

Marshall:2012:AAL

[895] Tristan Marshall and Gareth Roberts. An adaptive approach to Langevin MCMC. Statistics and Computing, 22(5):1041–1057, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-011-9276-

Cao:2012:ECD

[896] Jiguo Cao, Jing Cai, and Liangliang Wang. Estimating curves and derivatives with parametric penalized spline smoothing. Statistics and Computing, 22(5):1059–1067, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9278-4.

Tran:2012:PL

[897] Minh-Ngoc Tran, David J. Nott, and Chenlei Leng. The predictive lasso. Statistics and Computing, 22(5):1069– 1084, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9279-3.

Song:2012:STM

[898] Xin-Yuan Song and Zhao-Hua Lu. Semiparametric transformation models with Bayesian P -splines. Statistics and Computing, 22(5):1085–1098, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9280-x.

vanValkenhoef:2012:APM

[899] Gert van Valkenhoef, Tommi Tervonen, Bert de Brock, and Hans Hillege. Algorithmic parameterization of mixed treatment comparisons. Statistics and Computing, 22(5):1099–1111, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

com/content/pdf/10.1007/s11222-
011-9281-9.pdf.

Donnet:2012:EBP

[900] Sophie Donnet and Jean-Michel Marin. An empirical Bayes procedure for the selection of Gaussian graphical models. Statistics and Computing, 22(5): 1113–1123, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9285-5.

Xu:2012:IHT

[901] Ping-Feng Xu, Jianhua Guo, and Man-Lai Tang. An improved Hara– Takamura procedure by sharing computations on junction tree in Gaussian graphical models. Statistics and Computing, 22(5):1125–1133, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9286-4.

Poon:2012:LVM

[902] Wai-Yin Poon and Hai-Bin Wang. Latent variable models with ordinal categorical covariates. Statistics and Computing, 22(5):1135–1154, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9290-8.

Kontorovich:2012:SEB

[903] Leonid (Aryeh) Kontorovich. Statistical estimation with bounded memory. Statistics and Computing, 22(5): 1155–1164, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9293-5.

Anonymous:2012:HCe

[904] Anonymous. Help & contacts. Statistics and Computing, 22(5):??, September 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Celeux:2012:ABC

[905] Gilles Celeux. Approximate Bayesian computation methods. Statistics and Computing, 22(6):1165–1166, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
012-9350-8.pdf.

Marin:2012:ABC

[906] Jean-Michel Marin, Pierre Pudlo, Christian P. Robert, and Robin J. Ryder. Approximate Bayesian computational methods. Statistics and Computing, 22(6):1167–1180, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
011-9288-2.pdf.

Barnes:2012:CAC

[907] Chris P. Barnes, Sarah Filippi, Michael P. H. Stumpf, and Thomas Thorne. Considerate approaches to constructing summary statistics for ABC model selection. Statistics and Computing, 22(6):1181–1197, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11222-012-
                                        9335-7.

McVinish:2012:IAQ

[908] R. McVinish. Improving ABC for quantile distributions. Statistics and Computing, 22(6):1199–1207, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9209-9.

Peters:2012:SMC

[909] G. W. Peters, Y. Fan, and S. A. Sisson. On sequential Monte Carlo, partial rejection control and approximate Bayesian computation. Statistics and Computing, 22(6):1209–1222, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-012-
                                        9315-y.

Jasra:2012:FAB

[910] Ajay Jasra, Sumeetpal S. Singh, James S. Martin, and Emma McCoy. Filtering via approximate Bayesian computation. Statistics and Computing, 22(6):1223–1237, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-010-
                                        9185-0.

Neal:2012:ELF

[911] Peter Neal. Efficient likelihoodfree Bayesian computation for household epidemics. Statistics and Computing, 22(6):1239–1256, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-010-
9216-x.

Rau:2012:REG

[912] Andrea Rau, Florence Jaffrézic, Jean-Louis Foulley, and R. W. Doerge. Reverse engineering gene regulatory networks using approximate Bayesian computation. Statistics and Computing, 22(6):1257–1271, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9309-1.

Nott:2012:EKF

[913] David J. Nott, Lucy Marshall, and Tran Minh Ngoc. The ensemble Kalman filter is an ABC algorithm. Statistics and Computing, 22(6):1273– 1276, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9300-x.

Anonymous:2012:HCf

[914] Anonymous. Help & contacts. Statistics and Computing, 22(6):??, November 2012. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Kolossiatis:2013:BNM

[915] M. Kolossiatis, J. E. Griffin, and M. F. J. Steel. On Bayesian nonparametric modelling of two correlated distributions. Statistics and Computing, 23(1):1–15, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9283-7.

Maadooliat:2013:REC

[916] Mehdi Maadooliat, Mohsen Pourahmadi, and Jianhua Z. Huang. Robust estimation of the correlation matrix of longitudinal data. Statistics and Computing, 23(1):17–28, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9284-6.

Peyrard:2013:MBA

[917] Nathalie Peyrard, Régis Sabbadin, Daniel Spring, Barry Brook, and Ralph MacNally. Model-based adaptive spatial sampling for occurrence map construction. Statistics and Computing, 23(1):29–42, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9287-3.

Prado:2013:SPL

[918] Raquel Prado and Hedibert F. Lopes. Sequential parameter learning and filtering in structured autoregressive state-space models. Statistics and Computing, 23(1):43–57, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9289-1.

Corander:2013:SYB

[919] Jukka Corander, Yaqiong Cui, Timo Koski, and Jukka Sirén. Have I seen you before? Principles of Bayesian predictive classification revisited. Statistics and Computing, 23(1):59–73, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9291-7.

Lambert:2013:NAL

[920] Philippe Lambert. Nonparametric additive location-scale models for interval censored data. Statistics and Computing, 23(1):75–90, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9292-6.

Miguez:2013:CTS

[921] Joaquín Míguez, Dan Crisan, and Petar M. Djurić. On the convergence of two sequential Monte Carlo methods for maximum a posteriori sequence estimation and stochastic global optimization. Statistics and Computing, 23(1):91–107, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9294-4.

Mulder:2013:BTC

[922] Joris Mulder and Jean-Paul Fox. Bayesian tests on components of the compound symmetry covariance matrix. Statistics and Computing, 23 (1):109–122, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-011-9295-3.pdf.

Griffin:2013:AMH

[923] Jim E. Griffin and Stephen G. Walker. On adaptive Metropolis– Hastings methods. Statistics and Computing, 23(1):123–134, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9296-2.

Anonymous:2013:HCa

[924] Anonymous. Help & contacts. Statistics and Computing, 23(1):??, January 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Sobotka:2013:CIS

[925] Fabian Sobotka, Göran Kauermann, Linda Schulze Waltrup, and Thomas Kneib. On confidence intervals for semiparametric expectile regression. Statistics and Computing, 23 (2):135–148, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-011-9297-1.pdf.

Hsieh:2013:NAM

[926] Ping-Hung Hsieh. A nonparametric assessment of model adequacy based on Kullback–Leibler divergence. Statistics and Computing, 23(2):149–162, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9298-0.

Schafer:2013:SMC

[927] Christian Schäfer and Nicolas Chopin. Sequential Monte Carlo on large binary sampling spaces. Statistics and Computing, 23(2):163–184, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9299-z.

Casarin:2013:IMT

[928] Roberto Casarin, Radu Craiu, and Fabrizio Leisen. Interacting multiple try algorithms with different proposal distributions. Statistics and Computing, 23(2):185–200, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9301-9.

Park:2013:MIA

[929] Jin-Hong Park. Multiple-index approach to multiple autoregressive time series model. Statistics and Computing, 23(2):201–208, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9302-8.

Trendafilov:2013:EFP

[930] Nickolay T. Trendafilov, Steffen Unkel, and Wojtek Krzanowski. Exploratory factor and principal component analyses: some new aspects. Statistics and Computing, 23(2):209–220, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9303-7.

Jin:2013:NFM

[931] Baisuo Jin, Xiaoping Shi, and Yuehua Wu. A novel and fast methodology for simultaneous multiple structural break estimation and variable selection for nonstationary time series models. Statistics and Computing, 23 (2):221–231, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9304-6.

Chee:2013:EFM

[932] Chew-Seng Chee and Yong Wang. Estimation of finite mixtures with symmetric components. Statistics and Computing, 23(2):233–249, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-011-
                                        9305-5.

Botts:2013:TDR

[933] Carsten Botts, Wolfgang Hörmann, and Josef Leydold. Transformed density rejection with inflection points. Statistics and Computing, 23(2): 251–260, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-011-9306-4.

Chib:2013:CVE

[934] Siddhartha Chib and Edward Greenberg. On conditional variance estimation in nonparametric regression. Statistics and Computing, 23 (2):261–270, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9307-3.

Botev:2013:MCI

[935] Zdravko I. Botev, Pierre L’Ecuyer, and Bruno Tuffin. Markov chain importance sampling with applications to rare event probability estimation. Statistics and Computing, 23 (2):271–285, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-011-9308-2.

Milner:2013:MCB

[936] Peter Milner, Colin S. Gillespie, and Darren J. Wilkinson. Moment closure based parameter inference of stochastic kinetic models. Statistics and Computing, 23(2):287–295, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9310-8.

Anonymous:2013:HCb

[937] Anonymous. Help & contacts. Statistics and Computing, 23(2):??, March 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Bar-Hen:2013:ESO

[938] A. Bar-Hen, J. Chadoeuf, H. Dessard, and P. Monestiez. Estimating second order characteristics of point processes with known independent noise. Statistics and Computing, 23(3):297–309, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11222-011-
9311-7.

Ghosh:2013:SBM

[939] Joyee Ghosh and Jerome P. Reiter. Secure Bayesian model averaging for horizontally partitioned data. Statistics and Computing, 23(3):311–322, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-011-
9312-6.

Baragatti:2013:PTE

[940] Meïli Baragatti, Agnès Grimaud, and Denys Pommeret. Parallel tempering with equi-energy moves. Statistics and Computing, 23(3):323–339, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9313-0.

Wood:2013:SIR

[941] Simon N. Wood, Fabian Scheipl, and Julian J. Faraway. Straightforward intermediate rank tensor product smoothing in mixed models. Statistics and Computing, 23(3):341–360, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9314-z.

Lykou:2013:BLV

[942] Anastasia Lykou and Ioannis Ntzoufras. On Bayesian lasso variable selection and the specification of the shrinkage parameter. Statistics and Computing, 23(3):361–390, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9316-x.

Wiens:2013:DWL

[943] Douglas P. Wiens. Designs for weighted least squares regression, with estimated weights. Statistics and Computing, 23(3):391–401, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9317-9.

Gatu:2013:FAN

[944] Cristian Gatu and Erricos John Kontoghiorghes. A fast algorithm for nonnegativity model selection. Statistics and Computing, 23(3):403–411, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9318-8.

Kuhn:2013:CSE

[945] Estelle Kuhn and Charles El-Nouty. On a convergent stochastic estimation algorithm for frailty models. Statistics and Computing, 23(3):413–423, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9319-7.

Montoya:2013:REM

[946] Eduardo L. Montoya and Wendy Meiring. On the relative efficiency of a monotone parameter curve estimator in a functional nonlinear model. Statistics and Computing, 23(3):425–436, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-012-
                                        9320-1.

Anonymous:2013:HCc

[947] Anonymous. Help & contacts. Statistics and Computing, 23(3):??, May 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Hu:2013:BQR

[948] Yuao Hu, Robert B. Gramacy, and Heng Lian. Bayesian quantile regression for single-index models. Statistics and Computing, 23(4):437–454, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-012-
                                        9321-0.

Huang:2013:PEL

[949] Zhensheng Huang and Riquan Zhang. Profile empirical-likelihood inferences for the single-index-coefficient regression model. Statistics and Computing, 23(4):455–465, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-012-9322-z.

Fitzpatrick:2013:EBE

[950] Matthew Fitzpatrick and Dobrin Marchev. Efficient Bayesian estimation of the multivariate double chain Markov model. Statistics and Computing, 23(4):467–480, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-012-9323-y.

Oh:2013:BMC

[951] Man-Suk Oh. Bayesian multiple comparison of models for binary data with inequality constraints. Statistics and Computing, 23(4):481–490, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9324-x.

Fouskakis:2013:CIV

[952] D. Fouskakis and I. Ntzoufras. Computation for intrinsic variable selection in normal regression models via expectedposterior prior. Statistics and Computing, 23(4):491–499, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9325-9.

Aune:2013:INM

[953] Erlend Aune, Jo Eidsvik, and Yvo Pokern. Iterative numerical methods for sampling from high dimensional Gaussian distributions. Statistics and Computing, 23(4):501–521, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9326-8.

Gauchi:2013:NPF

[954] Jean-Pierre Gauchi and Jean-Pierre Vila. Nonparametric particle filtering approaches for identification and inference in nonlinear state-space dynamic systems. Statistics and Computing, 23(4):523–533, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-012-9327-7.

Baragatti:2013:LFP

[955] Meïli Baragatti, Agnès Grimaud, and Denys Pommeret. Likelihood-free parallel tempering. Statistics and Computing, 23(4):535–549, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-012-9328-6.

Riani:2013:RAP

[956] Marco Riani, Anthony C. Atkinson, Giulio Fanti, and Fabio Crosilla. Regression analysis with partially labelled regressors: carbon dating of the shroud of Turin. Statistics and Computing, 23(4):551–561, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-012-9329-5.

Anonymous:2013:HCd

[957] Anonymous. Help & contacts. Statistics and Computing, 23(4):??, July 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Yao:2013:ISI

[958] Wei-Ting Yao and Han-Ming Wu. Isometric sliced inverse regression for nonlinear manifold learning. Statistics and Computing, 23(5):563–576, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-012-
                                        9330-z.

Iliopoulos:2013:VRE

[959] George Iliopoulos and Sonia Malefaki. Variance reduction of estimators arising from Metropolis–Hastings algorithms. Statistics and Computing, 23 (5):577–587, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9331-y.

Xue:2013:SIS

[960] Liugen Xue and Zhen Pang. Statistical inference for a single-index varyingcoefficient model. Statistics and Computing, 23(5):589–599, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9332-x.

Chung:2013:TRB

[961] Matthias Chung, Qi Long, and Brent A. Johnson. A tutorial on rankbased coefficient estimation for censored data in small- and large-scale problems. Statistics and Computing, 23 (5):601–614, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9333-9.

Lebrun:2013:ETS

[962] R. Lebrun. Efficient time/space algorithm to compute rectangular probabilities of multinomial, multivariate hypergeometric and multivariate Pólya distributions. Statistics and Computing, 23(5):615–623, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9334-8.

Galimberti:2013:UCI

[963] Giuliano Galimberti and Gabriele Soffritti. Using conditional independence for parsimonious model-based Gaussian clustering. Statistics and Computing, 23(5):625–638, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9336-6.

Bazyari:2013:PPT

[964] Abouzar Bazyari and Fortunato Pesarin. Parametric and permutation testing for multivariate monotonic alternatives. Statistics and Computing, 23(5):639–652, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9338-4.

Mira:2013:ZVM

[965] Antonietta Mira, Reza Solgi, and Daniele Imparato. Zero variance Markov chain Monte Carlo for Bayesian estimators. Statistics and Computing, 23(5):653–662, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9344-6.

Chen:2013:OLH

[966] Ray-Bing Chen, Dai-Ni Hsieh, Ying Hung, and Weichung Wang. Optimizing Latin hypercube designs by particle swarm. Statistics and Computing, 23 (5):663–676, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9363-3.

Anonymous:2013:HCe

[967] Anonymous. Help & contacts. Statistics and Computing, 23(5):??, September 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Liebscher:2013:RDT

[968] Steffen Liebscher, Thomas Kirschstein, and Claudia Becker. RDELA — a Delaunay-triangulation-based, location and covariance estimator with high breakdown point. Statistics and Computing, 23(6):677–688, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9337-5.

Welsh:2013:RMB

[969] A. H. Welsh and Douglas P. Wiens. Robust model-based sampling designs. Statistics and Computing, 23 (6):689–701, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9339-3.

Jokiel-Rokita:2013:NER

[970] Alicja Jokiel-Rokita and Michal Pulit. Nonparametric estimation of the ROC curve based on smoothed empirical distribution functions. Statistics and Computing, 23(6):703–712, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
012-9340-x.pdf.

Wang:2013:ECN

[971] Yong Wang and Stephen M. Taylor. Efficient computation of nonparametric survival functions via a hierarchical mixture formulation. Statistics and Computing, 23(6):713–725, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9341-9.

Gonzalez:2013:PBI

[972] Miguel González, Cristina Gutiérrez, and Rodrigo Martínez. Parametric Bayesian inference for Y -linked twosex branching models. Statistics and Computing, 23(6):727–741, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9342-8.

Zhang:2013:LBB

[973] Yanwei Zhang. Likelihood-based and Bayesian methods for Tweedie compound Poisson linear mixed models. Statistics and Computing, 23(6): 743–757, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9343-7.

Petetin:2013:OSA

[974] Yohan Petetin and François Desbouvries. Optimal SIR algorithm vs. fully adapted auxiliary particle filter: a non asymptotic analysis. Statistics and Computing, 23(6):759–775, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9345-5.

Cornillon:2013:IBR

[975] P.-A. Cornillon, N. Hengartner, N. Jegou, and E. Matzner-Løber. Iterative bias reduction: a comparative study. Statistics and Computing, 23 (6):777–791, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9346-4.

Hirose:2013:ESB

[976] Yoshihiro Hirose and Fumiyasu Komaki. Edge selection based on the geometry of dually flat spaces for Gaussian graphical models. Statistics and Computing, 23(6):793–800, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9347-3.

Anonymous:2013:HCf

[977] Anonymous. Help & contacts. Statistics and Computing, 23(6):??, November 2013. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Jasra:2014:BPI

[978] Ajay Jasra, Nikolas Kantas, and Adam Persing. Bayesian parameter inference for partially observed stopped processes. Statistics and Computing, 24(1):1–20, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9348-2.

Hapfelmeier:2014:NVI

[979] Alexander Hapfelmeier, Torsten Hothorn, Kurt Ulm, and Carolin Strobl. A new variable importance measure for random forests with missing data. Statistics and Computing, 24(1):21–34, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9349-1.

Rodriguez:2014:UBN

[980] Carlos E. Rodríguez and Stephen G. Walker. Univariate Bayesian nonparametric mixture modeling with unimodal kernels. Statistics and Computing, 24(1):35–49, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-012-9351-
7.

Liu:2014:CPD

[981] Xiaohui Liu and Yijun Zuo. Computing projection depth and its associated estimators. Statistics and Computing, 24(1):51–63, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9352-6.

Andres:2014:TTA

[982] A. Martín Andrés and M. Álvarez Hernández. Two-tailed approximate confidence intervals for the ratio of proportions. Statistics and Computing, 24(1):65–75, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9353-5. See erra-

tum [1204].

Fiorentini:2014:EMS

[983] Gabriele Fiorentini, Christophe Planas, and Alessandro Rossi. Efficient MCMC sampling in dynamic mixture models. Statistics and Computing, 24(1):77–89, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9354-4.

Mrkvicka:2014:TSE

[984] T. Mrkvicka, M. Muska, and J. Kubecka. Two step estimation for Neyman–Scott point process with inhomogeneous cluster centers. Statistics and Computing, 24(1):91–100, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9355-3.

Haines:2014:COD

[985] Linda M. Haines and Allan E. Clark. The construction of optimal designs for dose-escalation studies. Statistics and Computing, 24(1):101–109, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9356-2.

Wang:2014:EME

[986] L. Wang, J. Cao, J. O. Ramsay, D. M. Burger, C. J. L. Laporte, and J. K. Rockstroh. Estimating mixed-effects differential equation models. Statistics and Computing, 24(1):111–121, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9357-1.

Leng:2014:SCE

[987] Chenlei Leng and Weiping Zhang. Smoothing combined estimating equations in quantile regression for longitudinal data. Statistics and Computing, 24(1):123–136, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9358-0.

Anonymous:2014:HCa

[988] Anonymous. Help & contacts. Statistics and Computing, 24(1):??, January 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Groll:2014:VSG

[989] Andreas Groll and Gerhard Tutz. Variable selection for generalized linear mixed models by L_{1}-penalized estimation. Statistics and Computing, 24 (2):137–154, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9359-z.

Titman:2014:EPS

[990] Andrew C. Titman. Estimating parametric semi-Markov models from panel data using phase-type approximations. Statistics and Computing, 24 (2):155–164, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9360-6.

Scealy:2014:FKM

[991] J. L. Scealy and A. H. Welsh. Fitting Kent models to compositional data with small concentration. Statistics and Computing, 24(2):165–179, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9361-5.

Lee:2014:FMM

[992] Sharon Lee and Geoffrey J. McLachlan. Finite mixtures of multivariate skew t-distributions: some recent and new results. Statistics and Computing, 24(2):181–202, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9362-4.

Browne:2014:OSM

[993] Ryan P. Browne and Paul D. McNicholas. Orthogonal Stiefel manifold optimization for eigen-decomposed covariance parameter estimation in mixture models. Statistics and Computing, 24(2):203–210, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9364-2.

Cahoy:2014:PEF

[994] Dexter O. Cahoy and Federico Polito. Parameter estimation for fractional birth and fractional death processes. Statistics and Computing, 24(2): 211–222, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9365-1.

Lang:2014:MSA

[995] Stefan Lang, Nikolaus Umlauf, Peter Wechselberger, Kenneth Harttgen, and Thomas Kneib. Multilevel structured additive regression. Statistics and Computing, 24(2):223–238, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9366-0.

Devroye:2014:RVG

[996] Luc Devroye. Random variate generation for the generalized inverse Gaussian distribution. Statistics and Computing, 24(2):239–246, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9367-z.

Aune:2014:PEH

[997] Erlend Aune, Daniel P. Simpson, and Jo Eidsvik. Parameter estimation in high dimensional Gaussian distributions. Statistics and Computing, 24 (2):247–263, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9368-y.

Young:2014:MRC

[998] Derek S. Young. Mixtures of regressions with changepoints. Statistics and Computing, 24(2):265–281, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-012-
9369-x.

Anonymous:2014:HCb

[999] Anonymous. Help & contacts. Statistics and Computing, 24(2):??, March 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Konietschke:2014:BPP

[1000] Frank Konietschke and Markus Pauly.

Bootstrapping and permuting paired t-test type statistics. Statistics and Computing, 24(3):283–296, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
012-9370-4.pdf.

Meligkotsidou:2014:BMD

[1001] Loukia Meligkotsidou, Elias Tzavalis, and Ioannis D. Vrontos. A Bayesian method of distinguishing unit root from stationary processes based on panel data models with cross-sectional dependence. Statistics and Computing, 24(3):297–315, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-012-9371-3.

Cornebise:2014:ASM

[1002] Julien Cornebise, Eric Moulines, and Jimmy Olsson. Adaptive sequential Monte Carlo by means of mixture of experts. Statistics and Computing, 24(3):317–337, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-012-9372-2.

Shahbaba:2014:SHM

[1003] Babak Shahbaba, Shiwei Lan, Wesley O. Johnson, and Radford M. Neal. Split Hamiltonian Monte Carlo. Statistics and Computing, 24(3):339–349, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-012-
                                        9373-1.

Ferrari:2014:RIC

[1004] Davide Ferrari, Matteo Borrotti, and Davide De March. Response improvement in complex experiments by coinformation composite likelihood optimization. Statistics and Computing, 24(3):351–363, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9374-8.

Akbilgic:2014:NHR

[1005] Oguz Akbilgic, Hamparsum Bozdogan, and M. Erdal Balaban. A novel hybrid RBF neural networks model as a forecaster. Statistics and Computing, 24(3):365–375, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9375-7.

Hernandez:2014:PSH

[1006] Lorenzo Hernández, Jorge Tejero, Alberto Su, and Santiago Carrillo-Menéndez. Percentiles of sums of heavy-tailed random variables: beyond the single-loss approximation. Statistics and Computing, 24(3):377–397, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9376-6.

Hwang:2014:ENT

[1007] Yi-Ting Hwang, Hsun-Chih Kuo, Chun-Chao Wang, and Meng Feng Lee. Estimating the number of true null hypotheses in multiple hypothesis testing. Statistics and Computing, 24(3):399–416, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9377-5.

Viviani:2014:GLM

[1008] Sara Viviani, Marco Alfó, and Dimitris Rizopoulos. Generalized linear mixed joint model for longitudinal and survival outcomes. Statistics and Computing, 24(3):417–427, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9378-4.

Lee:2014:BAB

[1009] Seokho Lee and Jianhua Z. Huang. A biclustering algorithm for binary matrices based on penalized Bernoulli likelihood. Statistics and Computing, 24(3):429–441, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9379-3.

Arribas-Gil:2014:LTE

[1010] Ana Arribas-Gil, Karine Bertin, Cristian Meza, and Vincent Rivoirard. LASSO-type estimators for semiparametric nonlinear mixed-effects models estimation. Statistics and Computing, 24(3):443–460, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9380-x.

Geraci:2014:LQM

[1011] Marco Geraci and Matteo Bottai. Linear quantile mixed models. Statistics and Computing, 24(3):461–479, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9381-9.

Kuhnt:2014:ODC

[1012] Sonja Kuhnt, Fabio Rapallo, and André Rehage. Outlier detection in contingency tables based on minimal patterns. Statistics and Computing, 24(3):481–491, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9382-8.

Anonymous:2014:HCc

[1013] Anonymous. Help & contacts. Statistics and Computing, 24(3):??, May 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Volant:2014:HMM

[1014] Stevenn Volant, Caroline Bérard, Marie-Laure Martin-Magniette, and Stéphane Robin. Hidden Markov models with mixtures as emission distributions. Statistics and Computing, 24(4):493–504, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9383-7.

Cheon:2014:SAM

[1015] Sooyoung Cheon, Faming Liang, Yuguo Chen, and Kai Yu. Stochastic approximation Monte Carlo importance sampling for approximating exact conditional probabilities. Statistics and Computing, 24(4):505–520, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9384-6.

Wang:2014:CDA

[1016] Hao Wang. Coordinate descent algorithm for covariance graphical lasso. Statistics and Computing, 24 (4):521–529, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9385-5.

Lin:2014:FMM

[1017] Tsung-I Lin, Hsiu J. Ho, and Chia-Rong Lee. Flexible mixture modelling using the multivariate skew-t-normal distribution. Statistics and Computing, 24(4):531–546, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9386-4.

Hormann:2014:GGI

[1018] Wolfgang Hörmann and Josef Leydold. Generating generalized inverse Gaussian random variates. Statistics and Computing, 24(4):547–557, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9387-3.

Chiou:2014:FAF

[1019] Sy Han Chiou, Sangwook Kang, and Jun Yan. Fast accelerated failure time modeling for case-cohort data. Statistics and Computing, 24(4):559–568, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9388-2.

Gollini:2014:MLT

[1020] Isabella Gollini and Thomas Brendan Murphy. Mixture of latent trait analyzers for model-based clustering of categorical data. Statistics and Computing, 24(4):569–588, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9389-1.

Barabesi:2014:NUR

[1021] Lucio Barabesi and Luca Pratelli. A note on a universal random variate generator for integer-valued random variables. Statistics and Computing, 24(4):589–596, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9390-8.

Borowski:2014:OSE

[1022] Matthias Borowski and Roland Fried.

Online signal extraction by robust regression in moving windows with data-adaptive width selection. Statistics and Computing, 24(4):597–613, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9391-7.

Fassino:2014:AIC

[1023] Claudia Fassino, Giovanni Pistone, and Eva Riccomagno. The algebra of interpolatory cubature formulæ for generic nodes. Statistics and Computing, 24(4):615–632, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9392-6.

Franz:2014:PTR

[1024] Jürgen Franz, Alicja Jokiel-Rokita, and Ryszard Magiera. Prediction in trend-renewal processes for repairable systems. Statistics and Computing, 24(4):633–649, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-013-9393-5.pdf.

Bacro:2014:ESM

[1025] Jean-Noel Bacro and Carlo Gaetan. Estimation of spatial max-stable models using threshold exceedances. Statistics and Computing, 24(4):651–662, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9394-4.

Lloyd:2014:CHA

[1026] Chris J. Lloyd and Degui Li. Computing highly accurate confidence limits from discrete data using importance sampling. Statistics and Computing, 24(4):663–673, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9409-1.

Anonymous:2014:HCd

[1027] Anonymous. Help & contacts. Statistics and Computing, 24(4):??, July 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Leger:2014:DSH

[1028] Jean-Benoist Leger, Corinne Vacher, and Jean-Jacques Daudin. Detection of structurally homogeneous subsets in graphs. Statistics and Computing, 24 (5):675–692, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9395-3.

Lavielle:2014:ISA

[1029] Marc Lavielle and Cyprien Mbogning. An improved SAEM algorithm for maximum likelihood estimation in mixtures of non linear mixed effects models. Statistics and Computing, 24(5): 693–707, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9396-2.

Friel:2014:IPP

[1030] Nial Friel, Merrilee Hurn, and Jason Wyse. Improving power posterior estimation of statistical evidence. Statistics and Computing, 24 (5):709–723, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9397-1.

Pan:2014:RES

[1031] Jianxin Pan and Chao Huang. Random effects selection in generalized linear mixed models via shrinkage penalty function. Statistics and Computing, 24 (5):725–738, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9398-0.

Economou:2014:MIB

[1032] Theodoros Economou, Trevor C. Bailey, and Zoran Kapelan. MCMC implementation for Bayesian hidden semi-Markov models with illustrative applications. Statistics and Computing, 24 (5):739–752, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9399-z.

Menardi:2014:ACN

[1033] Giovanna Menardi and Adelchi Azzalini. An advancement in clustering via nonparametric density estimation. Statistics and Computing, 24(5): 753–767, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9400-x.

daSilva:2014:GWN

[1034] Alan Ricardo da Silva and Thais Carvalho Valadares Rodrigues. Geographically weighted negative binomial regression-incorporating overdispersion. Statistics and Computing, 24 (5):769–783, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9401-9.

Berentsen:2014:RVD

[1035] Geir Drage Berentsen and Dag Tjøstheim. Recognizing and visualizing departures from independence in bivariate data using local Gaussian correlation. Statistics and Computing, 24 (5):785–801, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9402-8.

Yu:2014:AAP

[1036] Yi Yu and Yang Feng. APPLE: approximate path for penalized likelihood estimators. Statistics and Computing, 24 (5):803–819, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9403-7.

Shaharanee:2014:EOF

[1037] I. N. M. Shaharanee and F. Hadzic. Evaluation and optimization of frequent, closed and maximal association rule based classification. Statistics and Computing, 24(5):821–843, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9404-6.

Lee:2014:RDM

[1038] Shih-Chang Lee, Ning-Ning Pang, and Wen-Jer Tzeng. Resolution dependence of the maximal information coefficient for noiseless relationship. Statistics and Computing, 24(5):845–852, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9405-5.

Peng:2014:SEV

[1039] Limin Peng, Jinfeng Xu, and Nancy Kutner. Shrinkage estimation of varying covariate effects based on quantile regression. Statistics and Computing, 24(5):853–869, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9406-4.

Jiang:2014:MMC

[1040] Dingfeng Jiang and Jian Huang. Majorization minimization by coordinate descent for concave penalized generalized linear models. Statistics and Computing, 24(5):871–883, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9407-3.

Yao:2014:ALL

[1041] Yonggang Yao and Yoonkyung Lee. Another look at linear programming for feature selection via methods of regularization. Statistics and Computing, 24(5):885–905, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9408-2.

Anonymous:2014:HCe

[1042] Anonymous. Help & contacts. Statistics and Computing, 24(5):??, September 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Farcomeni:2014:SRM

[1043] Alessio Farcomeni. Snipping for robust k-means clustering under componentwise contamination. Statistics and Computing, 24(6):907–919, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9410-8.

Marschner:2014:CEA

[1044] Ian C. Marschner. Combinatorial EM algorithms. Statistics and Computing, 24(6):921–940, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9411-7.

Finos:2014:PTB

[1045] Livio Finos and Dario Basso. Permutation tests for between-unit fixed effects in multivariate generalized linear mixed models. Statistics and Computing, 24 (6):941–952, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9412-6.

Biernacki:2014:SVG

[1046] Christophe Biernacki and Alexandre Lourme. Stable and visualizable Gaussian parsimonious clustering models. Statistics and Computing, 24(6): 953–969, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9413-5.

Forbes:2014:NFM

[1047] Florence Forbes and Darren Wraith. A new family of multivariate heavy-tailed distributions with variable marginal amounts of tailweight: application to robust clustering. Statistics and Computing, 24(6):971–984, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9414-4.

Li:2014:ELI

[1048] Jianbo Li, Zhensheng Huang, and Heng Lian. Empirical likelihood inference for general transformation models with right censored data. Statistics and Computing, 24(6):985–995, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9415-3.

Gelman:2014:UPI

[1049] Andrew Gelman, Jessica Hwang, and Aki Vehtari. Understanding predictive information criteria for Bayesian models. Statistics and Computing, 24(6): 997–1016, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9416-2.

Ross:2014:SCD

[1050] Gordon J. Ross. Sequential change detection in the presence of unknown parameters. Statistics and Computing, 24(6):1017–1030, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9417-1.

Maharaj:2014:CCT

[1051] Elizabeth Ann Maharaj. Classification of cyclical time series using complex demodulation. Statistics and Computing, 24(6):1031–1046, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9418-0.

Lin:2014:SBI

[1052] Junjing Lin and Michael Ludkovski. Sequential Bayesian inference in hidden Markov stochastic kinetic models with application to detection and response to seasonal epidemics. Statistics and Computing, 24(6):1047–1062, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s11222-013-
9419-z.

Duarte:2014:SIP

[1053] Belmiro P. M. Duarte and Weng Kee Wong. A semi-infinite programming based algorithm for finding minimax optimal designs for nonlinear models. Statistics and Computing, 24(6): 1063–1080, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9420-6.

Kauermann:2014:FPC

[1054] Göran Kauermann and Christian Schellhase. Flexible pair-copula estimation in D-vines using bivariate penalized splines. Statistics and Computing, 24(6):1081–1100, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9421-5.

Yang:2014:ASB

[1055] Hongxia Yang, Fei Liu, Chunlin Ji, and David Dunson. Adaptive sampling for Bayesian geospatial models. Statistics and Computing, 24(6): 1101–1110, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9422-4.

Prates:2014:GLM

[1056] Marcos O. Prates, Denise R. Costa, and Victor H. Lachos. Generalized linear mixed models for correlated binary data with t-link. Statistics and Computing, 24(6):1111–1123, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9423-3.

Anonymous:2014:HCf

[1057] Anonymous. Help & contacts. Statistics and Computing, 24(6):??, November 2014. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Mira:2015:ISI

[1058] Antonietta Mira and Christian P. Robert. An introduction to the special issue “Joint IMS–ISBA meeting — MCMSki 4”. Statistics and Computing, 25(1):1, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-014-9536-3.pdf.

Haario:2015:IQB

[1059] Heikki Haario. Introduction to “Quantitative bounds of convergence for geometrically ergodic Markov Chain in the Wasserstein distance with application to the Metropolis adjusted Langevin algorithm” by A. Durmus, É. Moulines. Statistics and Computing, 25(1):3, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        014-9529-2.pdf. See [1060].

Durmus:2015:QBC

[1060] Alain Durmus and Éric Moulines. Quantitative bounds of convergence for geometrically ergodic Markov chain in the Wasserstein distance with application to the Metropolis Adjusted Langevin Algorithm. Statistics and Computing, 25(1):5–19, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-

9511-z. See [1059].

Friel:2015:IPP

[1061] Nial Friel. Introduction to “Preprocessing for approximate Bayesian computation in image analysis” by M. Moores, C. Drovandi, K. Mengersen, C. Robert. Statistics and Computing, 25(1):21, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
014-9533-6.pdf. See [1062].

Moores:2015:PPA

[1062] Matthew T. Moores, Christopher C.

Drovandi, Kerrie Mengersen, and Christian P. Robert. Pre-processing for approximate Bayesian computation in image analysis. Statistics and Computing, 25(1):23–33, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-

9525-6. See [1061].

Carlin:2015:ICB

[1063] Bradley Carlin. Introduction to “Cuts in Bayesian graphical models” by M. Plummer. Statistics and Computing, 25(1):35, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        014-9538-1.pdf. See [1064].

Plummer:2015:CBG

[1064] Martyn Plummer. Cuts in Bayesian graphical models. Statistics and Computing, 25(1):37–43, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s11222-014-9503-

z. See [1063].

Rue:2015:IFM

[1065] Håvard Rue. Introduction to “Fast matrix computations for functional additive models” by S. Barthelmé. Statistics and Computing, 25(1):45, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        014-9527-4.pdf. See [1066].

Barthelme:2015:FMC

[1066] Simon Barthelmé. Fast matrix computations for functional additive models. Statistics and Computing, 25(1):47–63, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-

9490-0. See [1065].

Mueller:2015:ICS

[1067] Peter Mueller. Introduction to “On a class of σ-stable Poisson–Kingman models and an effective marginalized sampler” by S. Favaro, M. Lomeli, Y. W. Teh. Statistics and Computing, 25(1):65–66, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-014-9537-

2. See [1068].

Favaro:2015:CSP

[1068] S. Favaro, M. Lomeli, and Y. W.

Teh. On a class of σ-stable Poisson– Kingman models and an effective marginalized sampler. Statistics and Computing, 25(1):67–78, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-

9499-4. See [1067].

Andrieu:2015:IPM

[1069] Christophe Andrieu. Introduction to “Particle Metropolis–Hastings using gradient and Hessian information” by J. Dahlin, F. Lindsten, T. Schön. Statistics and Computing, 25(1):79, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
014-9528-3.pdf. See [1070].

Dahlin:2015:PMH

[1070] Johan Dahlin, Fredrik Lindsten, and Thomas B. Schön. Particle Metropolis– Hastings using gradient and Hessian information. Statistics and Computing, 25(1):81–92, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9510-0. See [1069].

Peluso:2015:IUM

[1071] Stefano Peluso. Introduction to “On the use of Markov chain Monte Carlo methods for the sampling of mixture models” by R. Douc, F. Maire, J. Olsson. Statistics and Computing, 25 (1):93–94, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-014-9532-

7. See [1072].

Douc:2015:UMC

[1072] Randal Douc, Florian Maire, and Jimmy Olsson. On the use of Markov chain Monte Carlo methods for the sampling of mixture models: a statistical perspective. Statistics and Computing, 25(1):95–110, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-

9526-5. See [1071].

Friel:2015:IEC

[1073] Nial Friel. Introduction to “Efficient computational strategies for doubly intractable problems with applications to Bayesian social networks” by A. Caimo, A. Mira. Statistics and Computing, 25(1):111, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
014-9535-4.pdf. See [1074].

Caimo:2015:ECS

[1074] Alberto Caimo and Antonietta Mira. Efficient computational strategies for doubly intractable problems with applications to Bayesian social networks. Statistics and Computing, 25 (1):113–125, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9516-7. See [1073].

Robert:2015:IAA

[1075] Christian P. Robert. Introduction to “Adaptive ABC model choice and geometric summary statistics for hidden Gibbs random fields” by J. Stoehr, P. Pudlo, L. Cucala. Statistics and Computing, 25(1):127, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
014-9530-9.pdf. See [1076].

Stoehr:2015:AAM

[1076] Julien Stoehr, Pierre Pudlo, and Lionel Cucala. Adaptive ABC model choice and geometric summary statistics for hidden Gibbs random fields. Statistics and Computing, 25(1):129–141, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-

9514-9. See [1075].

Ryder:2015:ISI

[1077] Robin J. Ryder. Introduction to “Scalable inference for Markov processes with intractable likelihoods” by J. Owen, D. Wilkinson, C. Gillespie. Statistics and Computing, 25(1):143, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
014-9534-5.pdf. See [1078].

Owen:2015:SIM

[1078] Jamie Owen, Darren J. Wilkinson, and Colin S. Gillespie. Scalable inference for Markov processes with intractable likelihoods. Statistics and Computing, 25(1):145–156, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9524-7;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        014-9524-7.pdf. See [1077].

Fox:2015:IEL

[1079] Colin Fox. Introduction to “Efficient local updates for undirected graphical models” by F. Stingo, G. Marchetti. Statistics and Computing, 25(1):157, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9531-8;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        014-9531-8.pdf. See [1080].

Stingo:2015:ELU

[1080] Francesco Stingo and Giovanni M. Marchetti. Efficient local updates for undirected graphical models. Statistics and Computing, 25(1):159–171, January 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-

9541-6. See [1079].

Breheny:2015:GDA

[1081] Patrick Breheny and Jian Huang. Group descent algorithms for nonconvex penalized linear and logistic regression models with grouped predictors. Statistics and Computing, 25 (2):173–187, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9424-2.

Guo:2015:FDA

[1082] Mengmeng Guo, Lan Zhou, Jianhua Z. Huang, and Wolfgang Karl Härdle. Functional data analysis of generalized regression quantiles. Statistics and Computing, 25(2):189–202, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9425-1.

Corain:2015:IPM

[1083] Livio Corain and Luigi Salmaso.

Improving power of multivariate combination-based permutation tests. Statistics and Computing, 25(2): 203–214, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9426-0.

Greselin:2015:MLE

[1084] Francesca Greselin and Salvatore Ingrassia. Maximum likelihood estimation in constrained parameter spaces for mixtures of factor analyzers. Statistics and Computing, 25(2):215–226, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9427-z.

Kuipers:2015:URG

[1085] Jack Kuipers and Giusi Moffa. Uniform random generation of large acyclic digraphs. Statistics and Computing, 25(2):227–242, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9428-y.

Verge:2015:PIS

[1086] Christelle Vergé, Cyrille Dubarry, Pierre Del Moral, and Eric Moulines. On parallel implementation of sequential Monte Carlo methods: the island particle model. Statistics and Computing, 25(2):243–260, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9429-x.

Martella:2015:PPM

[1087] Francesca Martella, Donatella Vicari, and Maurizio Vichi. Partitioning predictors in multivariate regression models. Statistics and Computing, 25 (2):261–272, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9430-4.

Chen:2015:CIC

[1088] Ray-Bing Chen, Meihui Guo, Wolfgang K. Härdle, and Shih-Feng Huang. COPICA-independent component analysis via copula techniques. Statistics and Computing, 25 (2):273–288, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9431-3.

White:2015:PAB

[1089] S. R. White, T. Kypraios, and S. P.

Preston. Piecewise Approximate Bayesian Computation: fast inference for discretely observed Markov models using a factorised posterior distribution. Statistics and Computing, 25 (2):289–301, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-013-9432-2.pdf.

Guedon:2015:SUM

[1090] Yann Guédon. Segmentation uncertainty in multiple change-point models. Statistics and Computing, 25 (2):303–320, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9433-1.

Sei:2015:CNC

[1091] Tomonari Sei and Alfred Kume. Calculating the normalising constant of the Bingham distribution on the sphere using the holonomic gradient method. Statistics and Computing, 25(2):321–332, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-013-9434-0.

Neal:2015:EBI

[1092] Peter Neal and Theodore Kypraios. Exact Bayesian inference via data augmentation. Statistics and Computing, 25(2):333–347, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-013-9435-z.pdf.

Wang:2015:NMD

[1093] Xuxu Wang and Yong Wang. Nonparametric multivariate density estimation using mixtures. Statistics and Computing, 25(2):349–364, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9436-y.

Fried:2015:RBO

[1094] Roland Fried, Inoncent Agueusop, Björn Bornkamp, Konstantinos Fokianos, Jana Fruth, and Katja Ickstadt. Retrospective Bayesian outlier detection in INGARCH series. Statistics and Computing, 25(2):365–374, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9437-x.

Tomasson:2015:SCA

[1095] Helgi Tómasson. Some computational aspects of Gaussian CARMA modelling. Statistics and Computing, 25 (2):375–387, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9438-9.

Chakraborty:2015:ASM

[1096] Avishek Chakraborty, Swarup De, Kenneth P. Bowman, Huiyan Sang, Marc G. Genton, and Bani K. Mallick. An adaptive spatial model for precipitation data from multiple satellites over large regions. Statistics and Computing, 25(2):389–405, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9439-8.

Koblents:2015:PMC

[1097] Eugenia Koblents and Joaquín Míguez.

A population Monte Carlo scheme with transformed weights and its application to stochastic kinetic models. Statistics and Computing, 25(2):407–425, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9440-2.

Sarkka:2015:PIP

[1098] Simo Särkkä, Jouni Hartikainen, Isambi Sailon Mbalawata, and Heikki Haario. Posterior inference on parameters of stochastic differential equations via non-linear Gaussian filtering and adaptive MCMC. Statistics and Computing, 25(2):427–437, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9441-1.

Maronna:2015:RNP

[1099] Ricardo A. Maronna, Fernanda Méndez, and Víctor J. Yohai. Robust nonlinear principal components. Statistics and Computing, 25(2):439–448, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-013-
                                        9442-0.

Sansonnet:2015:MPI

[1100] Laure Sansonnet and Christine Tuleau-Malot. A model of Poissonian interactions and detection of dependence. Statistics and Computing, 25 (2):449–470, March 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-013-9443-z.

Naranjo:2015:BAS

[1101] L. Naranjo, C. J. Pérez, and J. Martín. Bayesian analysis of some models that use the asymmetric exponential power distribution. Statistics and Computing, 25(3):497–514, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9449-1.

Bartolucci:2015:IMH

[1102] Francesco Bartolucci and Alessio Farcomeni. Information matrix for hidden Markov models with covariates. Statistics and Computing, 25(3):515–526, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9450-8.

Graversen:2015:CAD

[1103] Therese Graversen and Steffen Lauritzen. Computational aspects of DNA mixture analysis. Statistics and Computing, 25(3):527–541, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9451-7.

Pya:2015:SCA

[1104] Natalya Pya and Simon N. Wood.

Shape constrained additive models. Statistics and Computing, 25 (3):543–559, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-013-9448-7.pdf.

Ruiz:2015:EIE

[1105] Elias D. Nino Ruiz, Adrian Sandu, and Jeffrey Anderson. An efficient implementation of the ensemble Kalman filter based on an iterative Sherman–Morrison formula. Statistics and Computing, 25(3):561–577, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9454-4.

Li:2015:VSE

[1106] Gaorong Li, Peng Lai, and Heng Lian. Variable selection and estimation for partially linear single-index models with longitudinal data. Statistics and Computing, 25(3):579–593, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9447-8.

Yildirim:2015:CGM

[1107] Sinan Yildirim, Lan Jiang, Sumeetpal S. Singh, and Thomas A. Dean. Calibrating the Gaussian multi-target tracking model. Statistics and Computing, 25(3):595–608, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9456-2.

Hansen:2015:NLB

[1108] Niels Richard Hansen. Nonparametric likelihood based estimation of linear filters for point processes. Statistics and Computing, 25(3):609–618, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9452-6.

Garcia-Escudero:2015:ASL

[1109] L. A. García-Escudero, A. Gordaliza, C. Matrán, and A. Mayo-Iscar. Avoiding spurious local maximizers in mixture modeling. Statistics and Computing, 25(3):619–633, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9455-3.

Su:2015:FPQ

[1110] Steve Su. Flexible parametric quantile regression model. Statistics and Computing, 25(3):635–650, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9457-1.

Hu:2015:BQR

[1111] Yuao Hu, Kaifeng Zhao, and Heng Lian. Bayesian quantile regression for partially linear additive models. Statistics and Computing, 25(3):651–668, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-013-
9446-9.

Machado:2015:DAC

[1112] Alexei Manso Correa Machado. Dependence aliasing and the control of family-wise error rate in multiple hypothesis testing. Statistics and Computing, 25(3):669–681, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9459-z.

Marques:2015:DLC

[1113] Filipe J. Marques, Carlos A. Coelho, and Miguel de Carvalho. On the distribution of linear combinations of independent Gumbel random variables. Statistics and Computing, 25 (3):683–701, May 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9453-5.

Anonymous:2015:ENA

[1114] Anonymous. Editor’s note: 25th anniversary special issue. Statistics and Computing, 25(4):703, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-015-
9584-3;   http://link.springer.
com/content/pdf/10.1007/s11222-
015-9584-3.pdf.

Hand:2015:SCG

[1115] David J. Hand. Statistics and computing: the genesis of data science. Statistics and Computing, 25(4):705–711, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-015-
                                        9565-6;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        015-9565-6.pdf.

Baudry:2015:EM

[1116] Jean-Patrick Baudry and Gilles Celeux. EM for mixtures. Statistics and Computing, 25(4):713–726, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s11222-015-9561-
                                        x;   http://link.springer.com/content/
                                        pdf/10.1007/s11222-015-9561-x.pdf.

Beskos:2015:SMC

[1117] Alexandros Beskos, Ajay Jasra, Ege A. Muzaffer, and Andrew M. Stuart. Sequential Monte Carlo methods for Bayesian elliptic inverse problems. Statistics and Computing, 25 (4):727–737, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-015-9556-7;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-015-9556-7.pdf.

Silva:2015:BIP

[1118] Ricardo Silva and Alfredo Kalaitzis. Bayesian inference via projections. Statistics and Computing, 25(4): 739–753, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-015-9557-6;   http:
//link.springer.com/content/pdf/
10.1007/s11222-015-9557-6.pdf.

Scholkopf:2015:CFR

[1119] Bernhard Schölkopf, Krikamol Muandet, Kenji Fukumizu, Stefan Harmeling, and Jonas Peters. Computing functions of random variables via reproducing kernel Hilbert space representations. Statistics and Computing, 25(4):755–766, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-015-9558-5;   http:
//link.springer.com/content/pdf/
10.1007/s11222-015-9558-5.pdf.

Barthelme:2015:PTU

[1120] Simon Barthelmé and Nicolas Chopin.

The Poisson transform for unnormalised statistical models. Statistics and Computing, 25(4):767–780, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-015-
9559-4;   http://link.springer.
com/content/pdf/10.1007/s11222-
015-9559-4.pdf.

Toulis:2015:SES

[1121] Panos Toulis and Edoardo M. Airoldi.

Scalable estimation strategies based on stochastic approximations: classical results and new insights. Statistics and Computing, 25(4):781–795, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-015-
                                        9560-y;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        015-9560-y.pdf.

Bacallado:2015:FPU

[1122] Sergio Bacallado, Persi Diaconis, and Susan Holmes. De Finetti priors using Markov chain Monte Carlo computations. Statistics and Computing, 25(4):797–808, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-015-9562-9;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-015-9562-9.pdf.

Liu:2015:SES

[1123] Ying Liu, Andrew Gelman, and Tian Zheng. Simulation-efficient shortest probability intervals. Statistics and Computing, 25(4):809–819, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-015-
                                        9563-8;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        015-9563-8.pdf.

Hennig:2015:FPB

[1124] Christian Hennig and Chien-Ju Lin. Flexible parametric bootstrap for testing homogeneity against clustering and assessing the number of clusters. Statistics and Computing, 25(4):821–833, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-015-
                                        9566-5;   http://link.springer.

com/content/pdf/10.1007/s11222-
015-9566-5.pdf.

Green:2015:BCS

[1125] Peter J. Green, Krzysztof Latuszyński, Marcelo Pereyra, and Christian P. Robert. Bayesian computation: a summary of the current state, and samples backwards and forwards. Statistics and Computing, 25(4):835–862, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-015-
9574-5;   http://link.springer.
com/content/pdf/10.1007/s11222-
015-9574-5.pdf.

Anonymous:2015:HCa

[1126] Anonymous. Help & contacts. Statistics and Computing, 25(4):??, July 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Hirose:2015:SEN

[1127] Kei Hirose and Michio Yamamoto.

Sparse estimation via nonconcave penalized likelihood in factor analysis model. Statistics and Computing, 25 (5):863–875, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9458-0.

Bevilacqua:2015:CCL

[1128] Moreno Bevilacqua and Carlo Gaetan. Comparing composite likelihood methods based on pairs for spatial Gaussian random fields. Statistics and Computing, 25(5):877–892, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9460-6.

Deleforge:2015:HDR

[1129] Antoine Deleforge, Florence Forbes, and Radu Horaud. High-dimensional regression with Gaussian mixtures and partially-latent response variables. Statistics and Computing, 25 (5):893–911, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9461-5.

Borowski:2015:ROS

[1130] Matthias Borowski, Dennis Busse, and Roland Fried. Robust onlinesurveillance of trend-coherence in multivariate data streams: the similar trend monitoring (STM) procedure. Statistics and Computing, 25(5): 913–928, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9462-4.

Mononen:2015:CSW

[1131] Tommi Mononen. A case study of the widely applicable Bayesian information criterion and its optimality. Statistics and Computing, 25(5): 929–940, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9463-3.

Rodriguez-Alvarez:2015:FSP

[1132] María Xosé Rodríguez-Álvarez, Dae-Jin Lee, Thomas Kneib, María Durbán, and Paul Eilers. Fast smoothing parameter separation in multidimensional generalized P -splines: the SAP algorithm. Statistics and Computing, 25 (5):941–957, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9464-2.

Chacon:2015:ERA

[1133] José E. Chacón and Tarn Duong. Efficient recursive algorithms for functionals based on higher order derivatives of the multivariate Gaussian density. Statistics and Computing, 25(5): 959–974, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9465-1.

Chen:2015:MOD

[1134] Ray-Bing Chen, Shin-Perng Chang, Weichung Wang, Heng-Chih Tung, and Weng Kee Wong. Minimax optimal designs via particle swarm optimization methods. Statistics and Computing, 25 (5):975–988, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9466-0.

Trinh:2015:BCA

[1135] Giang Trinh and Alan Genz. Bivariate conditioning approximations for multivariate normal probabilities. Statistics and Computing, 25(5): 989–996, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9468-y.

McLean:2015:RLR

[1136] Mathew W. McLean, Giles Hooker, and David Ruppert. Restricted likelihood ratio tests for linearity in scalaron-function regression. Statistics and Computing, 25(5):997–1008, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9473-1.

IodiceDEnza:2015:LDT

[1137] Alfonso Iodice D’Enza and Angelos Markos. Low-dimensional tracking of association structures in categorical data. Statistics and Computing, 25(5): 1009–1022, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9470-4.

Hastie:2015:SDP

[1138] David I. Hastie, Silvia Liverani, and Sylvia Richardson. Sampling from Dirichlet process mixture models with unknown concentration parameter: mixing issues in large data implementations. Statistics and Computing, 25(5): 1023–1037, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9471-3;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-014-9471-3.pdf.

Golightly:2015:DAP

[1139] Andrew Golightly, Daniel A. Henderson, and Chris Sherlock. Delayed acceptance particle MCMC for exact inference in stochastic kinetic models. Statistics and Computing, 25(5): 1039–1055, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9469-x.

Anonymous:2015:HCb

[1140] Anonymous. Help & contacts. Statistics and Computing, 25(5):??, September 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Vujacic:2015:TCW

[1141] Ivan Vujacić, Itai Dattner, Javier González, and Ernst Wit. Time-course window estimator for ordinary differential equations linear in the parameters. Statistics and Computing, 25(6): 1057–1070, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9486-9.

Pircalabelu:2015:FIC

[1142] Eugen Pircalabelu, Gerda Claeskens, and Lourens Waldorp. A focused information criterion for graphical models. Statistics and Computing, 25(6): 1071–1092, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9504-y.

Nguyen:2015:SBC

[1143] Minh Khoa Nguyen, Steve Phelps, and Wing Lon Ng. Simulation based calibration using extended balanced augmented empirical likelihood.

Statistics and Computing, 25(6):1093– 1112, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9506-9.

Ahipasaoglu:2015:FOA

[1144] Selin Damla Ahipasaoglu. A firstorder algorithm for the A-optimal experimental design problem: a mathematical programming approach. Statistics and Computing, 25(6):1113– 1127, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9476-y.

Yang:2015:FUA

[1145] Yi Yang and Hui Zou. A fast unified algorithm for solving group-lasso penalize learning problems. Statistics and Computing, 25(6):1129–1141, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9498-5.

Bouveyron:2015:KDA

[1146] C. Bouveyron, M. Fauvel, and S. Girard. Kernel discriminant analysis and clustering with parsimonious Gaussian process models. Statistics and Computing, 25(6):1143–1162, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9505-x.

Castro:2015:LBI

[1147] Luis Mauricio Castro, Denise Reis Costa, Marcos Oliveira Prates, and Victor Hugo Lachos. Likelihood-based inference for Tobit confirmatory factor analysis using the multivariate Student-t distribution. Statistics and Computing, 25(6):1163–1183, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9502-0.

Holzmann:2015:HMM

[1148] Hajo Holzmann and Florian Schwaiger.

Hidden Markov models with statedependent mixtures: minimal representation, model testing and applications to clustering. Statistics and Computing, 25(6):1185–1200, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9481-1.

Keribin:2015:ESL

[1149] Christine Keribin, Vincent Brault, Gilles Celeux, and Gérard Govaert. Estimation and selection for the latent block model on categorical data. Statistics and Computing, 25(6):1201– 1216, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9472-2.

Albert:2015:SAA

[1150] Carlo Albert, Hans R. Künsch, and Andreas Scheidegger. A simulated annealing approach to approximate Bayes computations. Statistics and Computing, 25(6):1217–1232, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9507-8.

Robinson:2015:PCF

[1151] G. K. Robinson. Practical computing for finite moment log-stable distributions to model financial risk. Statistics and Computing, 25(6):1233– 1246, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9478-9.

Waldmann:2015:VAG

[1152] Elisabeth Waldmann and Thomas Kneib. Variational approximations in geoadditive latent Gaussian regression: mean and quantile regression. Statistics and Computing, 25(6): 1247–1263, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9480-2.

Picheny:2015:MOU

[1153] Victor Picheny. Multiobjective optimization using Gaussian process emulators via stepwise uncertainty reduction. Statistics and Computing, 25(6): 1265–1280, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9477-x.

Rubio:2015:LEU

[1154] F. J. Rubio. Letter to the Editor:

On the use of improper priors for the shape parameters of asymmetric exponential power models. Statistics and Computing, 25(6):1281–1287, November 2015. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9479-8.

Hofner:2016:UFC

[1155] Benjamin Hofner, Thomas Kneib, and Torsten Hothorn. A unified framework of constrained regression. Statistics and Computing, 26(1–2):1–14, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9520-y.

Cuervo:2016:ODL

[1156] Daniel Palhazi Cuervo, Peter Goos, and Kenneth Sörensen. Optimal design of large-scale screening experiments: a critical look at the coordinate-exchange algorithm. Statistics and Computing, 26(1–2):15–28, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9467-z.

Alquier:2016:NMC

[1157] P. Alquier, N. Friel, R. Everitt, and A. Boland. Noisy Monte Carlo: convergence of Markov chains with approximate transition kernels. Statistics and Computing, 26(1–2):29–47, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9521-x.

Faraway:2016:DDS

[1158] Julian J. Faraway. Does data splitting improve prediction? Statistics and Computing, 26(1–2):49–60, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9522-9.

Gassiat:2016:IFS

[1159] E. Gassiat, A. Cleynen, and S. Robin. Inference in finite state space non parametric hidden Markov models and applications. Statistics and Computing, 26(1–2):61–71, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9523-8.

Heinzl:2016:AMM

[1160] Felix Heinzl and Gerhard Tutz. Additive mixed models with approximate Dirichlet process mixtures: the EM approach. Statistics and Computing, 26 (1–2):73–92, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9475-z.

Dubossarsky:2016:WBG

[1161] E. Dubossarsky, J. H. Friedman, J. T. Ormerod, and M. P. Wand. Waveletbased gradient boosting. Statistics and Computing, 26(1–2):93–105, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9474-0.

Aletti:2016:KOD

[1162] Giacomo Aletti, Caterina May, and Chiara Tommasi. KL-optimum designs: theoretical properties and practical computation. Statistics and Computing, 26(1–2):107–117, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9515-8.

Gagnon:2016:SRA

[1163] Jacob Gagnon, Hua Liang, and Anna Liu. Spherical radial approximation for nested mixed effects models. Statistics and Computing, 26(1–2):119–130, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9483-z.

Bang:2016:SEN

[1164] Sungwan Bang, HyungJun Cho, and Myoungshic Jhun. Simultaneous estimation for non-crossing multiple quantile regression with right censored data. Statistics and Computing, 26(1– 2):131–147, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9482-0.

Chowdhary:2016:DFI

[1165] K. Chowdhary and H. N. Najm. Data free inference with processed data products. Statistics and Computing, 26 (1–2):149–169, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9484-y.

Prangle:2016:LA

[1166] Dennis Prangle. Lazy ABC. Statistics and Computing, 26(1–2):171–185, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9544-3.

Nomura:2016:EGO

[1167] Noboru Nomura. Evaluation of Gaussian orthant probabilities based on orthogonal projections to subspaces. Statistics and Computing, 26(1–2): 187–197, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-014-9487-8.

You:2016:GDF

[1168] Chong You, Samuel Müller, and John T. Ormerod. On generalized degrees of freedom with application in linear mixed models selection. Statistics and Computing, 26(1–2):199–210, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9488-7.

Simonnet:2016:CAA

[1169] Eric Simonnet. Combinatorial analysis of the adaptive last particle method. Statistics and Computing, 26(1–2): 211–230, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9489-6.

Wang:2016:MCA

[1170] Junshan Wang and Ajay Jasra. Monte Carlo algorithms for computing αpermanents. Statistics and Computing, 26(1–2):231–248, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-014-9491-
z.

Wei:2016:SOB

[1171] Bei Wei, Stephen M. S. Lee, and Xiyuan Wu. Stochastically optimal bootstrap sample size for shrinkagetype statistics. Statistics and Computing, 26(1–2):249–262, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-014-9493-
x.

Cleynen:2016:CCP

[1172] A. Cleynen and S. Robin. Comparing change-point location in independent series. Statistics and Computing, 26 (1–2):263–276, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9492-y.

Phoa:2016:SMG

[1173] Frederick Kin Hing Phoa, Tai-Chi Wang, and Shu-Ching Lin. A search of maximum generalized resolution quaternary-code designs via integer linear programming. Statistics and Computing, 26(1–2):277–283, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9496-7.

Viallon:2016:RGF

[1174] Vivian Viallon, Sophie Lambert-Lacroix, Hölger Hoefling, and Franck Picard. On the robustness of the generalized fused lasso to prior specifications. Statistics and Computing, 26 (1–2):285–301, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9497-6.

Malsiner-Walli:2016:MBC

[1175] Gertraud Malsiner-Walli, Sylvia Frühwirth-Schnatter, and Bettina Grün. Modelbased clustering based on sparse finite Gaussian mixtures. Statistics and Computing, 26(1–2):303–324, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9500-2;   http://link.springer.
com/content/pdf/10.1007/s11222-
014-9500-2.pdf.

Critchley:2016:CFG

[1176] F. Critchley and P. Marriott. Computing with Fisher geodesics and extended exponential families. Statistics and Computing, 26(1–2):325–332, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9501-1.

Vitoratou:2016:EBJ

[1177] Silia Vitoratou, Ioannis Ntzoufras, and Irini Moustaki. Explaining the behavior of joint and marginal Monte Carlo estimators in latent variable models with independence assumptions. Statistics and Computing, 26 (1–2):333–348, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9495-8.

Fallaize:2016:EBI

[1178] Christopher J. Fallaize and Theodore Kypraios. Exact Bayesian inference for the Bingham distribution. Statistics and Computing, 26(1–2):349–360, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9508-7.

Tran:2016:AMH

[1179] Minh-Ngoc Tran, Michael K. Pitt, and Robert Kohn. Adaptive Metropolis– Hastings sampling using reversible dependent mixture proposals. Statistics and Computing, 26(1–2):361–381, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9509-6.

Buchner:2016:STN

[1180] Johannes Buchner. A statistical test for Nested Sampling algorithms. Statistics and Computing, 26(1–2):383–392, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11222-014-
                                        9512-y.

Bhadra:2016:APA

[1181] Anindya Bhadra and Edward L. Ionides. Adaptive particle allocation in iterated sequential Monte Carlo via approximating meta-models. Statistics and Computing, 26(1–2):393–407, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9513-x.

Xiao:2016:FCE

[1182] Luo Xiao, Vadim Zipunnikov, David Ruppert, and Ciprian Crainiceanu. Fast covariance estimation for highdimensional functional data. Statistics and Computing, 26(1–2):409–421, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9485-x.

Griffin:2016:ATM

[1183] J. E. Griffin. An adaptive truncation method for inference in Bayesian nonparametric models. Statistics and Computing, 26(1–2):423–441, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-014-
                                        9519-4.

Azevedo:2016:BLI

[1184] Caio L. N. Azevedo, Jean-Paul Fox, and Dalton F. Andrade. Bayesian longitudinal item response modeling with restricted covariance pattern structures. Statistics and Computing, 26 (1–2):443–460, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9518-5.

Chen:2016:SRR

[1185] Lisha Chen and Jianhua Z. Huang.

Sparse reduced-rank regression with covariance estimation. Statistics and Computing, 26(1–2):461–470, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9517-6.

Pulkkinen:2016:NKD

[1186] Seppo Pulkkinen. Nonlinear kernel density principal component analysis with application to climate data. Statistics and Computing, 26(1–2): 471–492, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9539-0.

Zhang:2016:BSC

[1187] Lin Zhang, Abhra Sarkar, and Bani K.

Mallick. Bayesian sparse covariance decomposition with a graphical structure. Statistics and Computing, 26(1– 2):493–510, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9540-7.

White:2016:BVS

[1188] Arthur White, Jason Wyse, and Thomas Brendan Murphy. Bayesian variable selection for latent class analysis using a collapsed Gibbs sampler. Statistics and Computing, 26(1– 2):511–527, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-014-9542-5.

Ranalli:2016:MMO

[1189] Monia Ranalli and Roberto Rocci. Mixture models for ordinal data: a pairwise likelihood approach. Statistics and Computing, 26(1–2):529–547, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-
9543-4.

Durand:2016:LLS

[1190] Jean-Baptiste Durand and Yann Guédon. Localizing the latent structure canonical uncertainty: entropy profiles for hidden Markov models. Statistics and Computing, 26(1–2):549–567, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-014-

9494-9. See erratum [1191].

Durand:2016:ELL

[1191] Jean-Baptiste Durand and Yann Guédon. Erratum to: Localizing the latent structure canonical uncertainty: entropy profiles for hidden Markov models. Statistics and Computing, 26(1–2):569–570, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-015-
9586-1;   http://link.springer.
com/content/pdf/10.1007/s11222-
015-9586-1.pdf. See [1190].

Caron:2016:BRN

[1192] François Caron. Book review: Nonlinear Time Series: Theory, methods, and applications with R examples, by Randal Douc, Eric Moulines, David S. Stoffer. Statistics and Computing, 26 (1–2):571–572, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9575-
4;   http://link.springer.com/article/
10.1007/s11222-015-9575-4.

Lee:2016:FMC

[1193] Sharon X. Lee and Geoffrey J. McLachlan. Finite mixtures of canonical fundamental skew t-distributions. Statistics and Computing, 26(3):573–589, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-015-
9545-x.

Vaisman:2016:SMC

[1194] Radislav Vaisman, Zdravko I. Botev, and Ad Ridder. Sequential Monte Carlo for counting vertex covers in general graphs. Statistics and Computing, 26(3):591–607, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-015-9546-9.

Yau:2016:NRE

[1195] Chun Yip Yau and Kin Wai Chan.

New recursive estimators of the timeaverage variance constant. Statistics and Computing, 26(3):609–627, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-015-
9548-7.

Wang:2016:SOD

[1196] Yanhong Wang, Yixin Fang, and Junhui Wang. Sparse optimal discriminant clustering. Statistics and Computing, 26(3):629–639, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-015-9547-8.

Argiento:2016:BGS

[1197] Raffaele Argiento, Ilaria Bianchini, and Alessandra Guglielmi. A blocked Gibbs sampler for NGG-mixture models via a priori truncation. Statistics and Computing, 26(3):641–661, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-015-
9549-6.

Hug:2016:ASS

[1198] Sabine Hug, Michael Schwarzfischer, Jan Hasenauer, Carsten Marr, and Fabian J. Theis. An adaptive scheduling scheme for calculating Bayes factors with thermodynamic integration using Simpson’s rule. Statistics and Computing, 26(3):663–677, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-015-9550-0.

Ruli:2016:ABC

[1199] Erlis Ruli, Nicola Sartori, and Laura Ventura. Approximate Bayesian computation with composite score functions. Statistics and Computing, 26 (3):679–692, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-015-9551-z.

Pesarin:2016:UIP

[1200] Fortunato Pesarin, Luigi Salmaso, Eleonora Carrozzo, and Rosa Arboretti. Union-intersection permutation solution for two-sample equivalence testing. Statistics and Computing, 26(3):693–701, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-015-9552-y.

Ahmadi-Javid:2016:EAR

[1201] Amir Ahmadi-Javid and Asghar Moeini. An economical acceptancerejection algorithm for uniform random variate generation over constrained simplexes. Statistics and Computing, 26(3):703–713, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-015-9553-x.

Josse:2016:ASS

[1202] Julie Josse and Sylvain Sardy. Adaptive shrinkage of singular values. Statistics and Computing, 26(3):715–724, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-015-
9554-9.

Khan:2016:VSS

[1203] Md Hasinur Rahaman Khan and J. Ewart H. Shaw. Variable selection for survival data with a class of adaptive elastic net techniques. Statistics and Computing, 26(3):725–741, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-015-
9555-8.

Andres:2016:ETT

[1204] A. Martín Andrés and M. Álvarez Hernández. Erratum to: Two-tailed approximate confidence intervals for the ratio of proportions. Statistics and Computing, 26(3):743–744, May 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
015-9619-9.pdf. See [982].

Pereyra:2016:PMC

[1205] Marcelo Pereyra. Proximal Markov chain Monte Carlo algorithms. Statistics and Computing, 26(4):745–760, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
015-9567-4.pdf.

Spade:2016:CPE

[1206] David A. Spade. A computational procedure for estimation of the mixing time of the random-scan Metropolis algorithm. Statistics and Computing, 26(4):761–781, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9568-

Perthame:2016:SFS

[1207] Émeline Perthame, Chloé Friguet, and David Causeur. Stability of feature selection in classification issues for highdimensional correlated data. Statistics and Computing, 26(4):783–796, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
015-9569-2.pdf.

Oates:2016:EEM

[1208] Chris J. Oates, Jim Q. Smith, Sach Mukherjee, and James Cussens. Exact estimation of multiple directed acyclic graphs. Statistics and Computing, 26(4):797–811, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9570-
9.

Anderson-Bergman:2016:CLC

[1209] Clifford Anderson-Bergman and Yaming Yu. Computing the log concave NPMLE for interval censored data. Statistics and Computing, 26 (4):813–826, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9571-
8.

Bhadra:2016:ESU

[1210] Anindya Bhadra and Raymond J. Carroll. Exact sampling of the unobserved covariates in Bayesian spline models for measurement error problems. Statistics and Computing, 26(4):827–840, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9572-7.

Klein:2016:SIS

[1211] Nadja Klein and Thomas Kneib. Simultaneous inference in structured additive conditional copula regression models: a unifying Bayesian approach. Statistics and Computing, 26(4):841–860, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9573-6.

Hasnat:2016:MBH

[1212] Md. Abul Hasnat, Olivier Alata, and Alain Trémeau. Model-based hierarchical clustering with Bregman divergences and fishers mixture model: application to depth image analysis. Statistics and Computing, 26(4):861– 880, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s11222-015-9576-3.

Li:2016:ACV

[1213] Longhai Li, Shi Qiu, Bei Zhang, and Cindy X. Feng. Approximating cross-validatory predictive evaluation in Bayesian latent variable models with integrated IS and WAIC. Statistics and Computing, 26(4):881–897, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9577-2.

Ridgway:2016:CGO

[1214] James Ridgway. Computation of Gaussian orthant probabilities in high dimension. Statistics and Computing, 26(4):899–916, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9578-
1.

Bodenham:2016:CEA

[1215] Dean A. Bodenham and Niall M.

Adams. A comparison of efficient approximations for a weighted sum of chi-squared random variables. Statistics and Computing, 26(4):917–928, July 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
015-9583-4.pdf.

Biernacki:2016:MBC

[1216] Christophe Biernacki and Julien Jacques. Model-based clustering of multivariate ordinal data relying on a stochastic binary search algorithm. Statistics and Computing, 26 (5):929–943, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9585-
2.

Hooker:2016:MAF

[1217] Giles Hooker and Steven Roberts. Maximal autocorrelation functions in functional data analysis. Statistics and Computing, 26(5):945–950, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9582-5.

Lacki:2016:SDS

[1218] Mateusz Krzysztof Lacki and Blazej Miasojedow. State-dependent swap strategies and automatic reduction of number of temperatures in adaptive parallel tempering algorithm. Statistics and Computing, 26(5):951– 964, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-015-9579-0.pdf.

Mazo:2016:FTC

[1219] Gildas Mazo, Stéphane Girard, and Florence Forbes. A flexible and tractable class of one-factor copulas. Statistics and Computing, 26(5): 965–979, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9580-
7.

Radice:2016:CRS

[1220] Rosalba Radice, Giampiero Marra, and Małgorzata Wojtyś. Copula regression spline models for binary outcomes. Statistics and Computing, 26 (5):981–995, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9581-
6.

Tang:2016:MQB

[1221] Xingyu Tang and Heng Lian. Mean and quantile boosting for partially linear additive models. Statistics and Computing, 26(5):997–1008, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9592-3.

Kleppe:2016:BSP

[1222] Tore Selland Kleppe and Hans J.

Skaug. Bandwidth selection in presmoothed particle filters. Statistics and Computing, 26(5):1009–1024, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9591-4.

Galimberti:2016:UMS

[1223] Giuliano Galimberti, Elena Scardovi, and Gabriele Soffritti. Using mixtures in seemingly unrelated linear regression models with non-normal errors. Statistics and Computing, 26(5): 1025–1038, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9587-
0.

Tutz:2016:INN

[1224] Gerhard Tutz and Dominik Koch.

Improved nearest neighbor classifiers by weighting and selection of predictors. Statistics and Computing, 26(5): 1039–1057, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-015-9588-
                                        z.

Beinrucker:2016:ESS

[1225] Andre Beinrucker, Ürün Dogan, and Gilles Blanchard. Extensions of stability selection using subsamples of observations and covariates. Statistics and Computing, 26(5):1059–1077, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9589-y.

Kosmidis:2016:MBC

[1226] Ioannis Kosmidis and Dimitris Karlis. Model-based clustering using copulas with applications. Statistics and Computing, 26(5):1079–1099, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9590-5.

Hofmeyr:2016:DCH

[1227] David P. Hofmeyr, Nicos G. Pavlidis, and Idris A. Eckley. Divisive clustering of high dimensional data streams. Statistics and Computing, 26(5):1101– 1120, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-015-9597-
                                        y.

McGree:2016:PMS

[1228] J. M. McGree, C. C. Drovandi, G. White, and A. N. Pettitt. A pseudomarginal sequential Monte Carlo algorithm for random effects models in Bayesian sequential design. Statistics and Computing, 26(5):1121–1136, September 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s11222-015-9596-z.

Spence:2016:CRI

[1229] Michael A. Spence and Paul G. Blackwell. Coupling random inputs for parameter estimation in complex models. Statistics and Computing, 26(6): 1137–1146, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9593-
2.

Heard:2016:CMC

[1230] Nicholas A. Heard and Melissa J. M.

Turcotte. Convergence of Monte Carlo distribution estimates from rival samplers. Statistics and Computing, 26(6): 1147–1161, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9595-
0.

Tommasi:2016:MMO

[1231] C. Tommasi, R. Martín-Martín, and J. López-Fidalgo. Max-min optimal discriminating designs for several statistical models. Statistics and Computing, 26(6):1163–1172, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9606-1.

Latouche:2016:VBM

[1232] Pierre Latouche and Stéphane Robin. Variational Bayes model averaging for graphon functions and motif frequencies inference in W -graph models. Statistics and Computing, 26(6): 1173–1185, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-015-9607-
                                        0.

Medina-Aguayo:2016:SNM

[1233] F. J. Medina-Aguayo, A. Lee, and G. O. Roberts. Stability of noisy Metropolis–Hastings. Statistics and Computing, 26(6):1187–1211, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11222-

015-9604-3.pdf. See erratum [1359].

Bierkens:2016:NRM

[1234] Joris Bierkens. Non-reversible Metropolis– Hastings. Statistics and Computing, 26(6):1213–1228, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL http://link.springer.com/
                                        content/pdf/10.1007/s11222-015-
                                        9598-x.pdf.

Schmitt:2016:FAR

[1235] Eric Schmitt and Kaveh Vakili. The FastHCS algorithm for robust PCA. Statistics and Computing, 26(6):1229– 1242, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/

article/10.1007/s11222-015-9602-
5.

Tan:2016:VIS

[1236] Linda S. L. Tan, Victor M. H. Ong, David J. Nott, and Ajay Jasra. Variational inference for sparse spectrum Gaussian process regression. Statistics and Computing, 26(6):1243–1261, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s11222-015-9600-7.

Barreto-Souza:2016:GMP

[1237] Wagner Barreto-Souza and Alexandre B. Simas. General mixed Poisson regression models with varying dispersion. Statistics and Computing, 26(6): 1263–1280, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-015-9601-
6.

Gyori:2016:HTM

[1238] Benjamin M. Gyori and Daniel Paulin.

Hypothesis testing for Markov chain Monte Carlo. Statistics and Computing, 26(6):1281–1292, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9594-1.

Fearnhead:2016:ASP

[1239] Paul Fearnhead and Loukia Meligkotsidou. Augmentation schemes for particle MCMC. Statistics and Computing, 26(6):1293–1306, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9603-4.

Pewsey:2016:PBG

[1240] Arthur Pewsey and Shogo Kato. Parametric bootstrap goodness-of-fit testing for Wehrly–Johnson bivariate circular distributions. Statistics and Computing, 26(6):1307–1317, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9605-2.

Helwig:2016:EEV

[1241] Nathaniel E. Helwig. Efficient estimation of variance components in nonparametric mixed-effects models with large samples. Statistics and Computing, 26(6):1319–1336, November 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9610-5.

Anonymous:2017:ENS

[1242] Anonymous. Editor’s note: special section on Bayesian nonparametrics. Statistics and Computing, 27(1):1, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        016-9718-2.pdf.

Arbel:2017:MMF

[1243] Julyan Arbel and Igor Prünster. A moment-matching Ferguson & Klass algorithm. Statistics and Computing, 27(1):3–17, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9676-
8.

Ghosh:2017:FAB

[1244] Sanmitra Ghosh, Srinandan Dasmahapatra, and Koushik Maharatna. Fast approximate Bayesian computation for estimating parameters in differential equations. Statistics and Computing, 27(1):19–38, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9643-
4.

Cipolli:2017:CTA

[1245] William Cipolli III and Timothy Hanson. Computationally tractable approximate and smoothed Pólya trees. Statistics and Computing, 27(1):39–51, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-016-9652-3. See erratum

[1246].

Cipolli:2017:ECT

[1246] William Cipolli III and Timothy Hanson. Erratum to: Computationally tractable approximate and smoothed Pólya trees. Statistics and Computing, 27(1):53, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
016-9686-6.pdf. See [1245].

Pircalabelu:2017:CDA

[1247] Eugen Pircalabelu, Gerda Claeskens, and Irène Gijbels. Copula directed acyclic graphs. Statistics and Computing, 27(1):55–78, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9599-9.

Geppert:2017:RPB

[1248] Leo N. Geppert, Katja Ickstadt, Alexander Munteanu, Jens Quedenfeld, and Christian Sohler. Random projections for Bayesian regression. Statistics and Computing, 27 (1):79–101, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-015-9608-z.pdf.

Wang:2017:MPS

[1249] Tao Wang, Xuerong Meggie Wen, and Lixing Zhu. Multiple-population shrinkage estimation via sliced inverse regression. Statistics and Computing, 27(1):103–114, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-015-9609-
                                        y.

Janhunen:2017:LDD

[1250] Tomi Janhunen, Martin Gebser, Jussi Rintanen, Henrik Nyman, Johan Pensar, and Jukka Corander. Learning discrete decomposable graphical models via constraint optimization. Statistics and Computing, 27(1):115–130, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9611-4.

Griffin:2017:SMC

[1251] J. E. Griffin. Sequential Monte Carlo methods for mixtures with normalized random measures with independent increments priors. Statistics and Computing, 27(1):131–145, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9612-3.

Fort:2017:SHU

[1252] Gersende Fort, Benjamin Jourdain, Tony Lelièvre, and Gabriel Stoltz. Self-healing umbrella sampling: convergence and efficiency. Statistics and Computing, 27(1):147–168, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9613-2.

Becu:2017:BST

[1253] Jean-Michel Bécu, Yves Grandvalet, Christophe Ambroise, and Cyril Dalmasso. Beyond support in twostage variable selection. Statistics and Computing, 27(1):169–179, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9614-1.

Yuan:2017:RRM

[1254] Ting Yuan and Junhui Wang. Reducedrank multi-label classification. Statistics and Computing, 27(1):181–191, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9615-0.

Alonso-Revenga:2017:NIE

[1255] J. M. Alonso-Revenga, N. Martín, and L. Pardo. New improved estimators for overdispersion in models with clustered multinomial data and unequal cluster sizes. Statistics and Computing, 27 (1):193–217, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-015-9616-
                                        z.

Walter:2017:PPB

[1256] Clément Walter. Point process-based Monte Carlo estimation. Statistics and Computing, 27(1):219–236, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9617-y.

Tan:2017:SVI

[1257] Linda S. L. Tan. Stochastic variational inference for large-scale discrete choice models using adaptive batch sizes. Statistics and Computing, 27 (1):237–257, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-015-9618-

Langrock:2017:MSG

[1258] Roland Langrock, Thomas Kneib, Richard Glennie, and Théo Michelot. Markov-switching generalized additive models. Statistics and Computing, 27 (1):259–270, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-015-9620-3.pdf.

Waltrup:2017:SEP

[1259] Linda Schulze Waltrup and Göran Kauermann. Smooth expectiles for panel data using penalized splines. Statistics and Computing, 27(1):271– 282, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s11222-015-9621-2.

Donat:2017:SPB

[1260] Francesco Donat and Giampiero Marra. Semi-parametric bivariate polychotomous ordinal regression. Statistics and Computing, 27(1):283–299, January 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9622-1.

Lilleborge:2017:EDB

[1261] Marie Lilleborge and Jo Eidsvik. Efficient designs for Bayesian networks with sub-tree bounds. Statistics and Computing, 27(2):301–318, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-015-9623-0.

Wahl:2017:SDB

[1262] François Wahl, Cécile Mercadier, and Céline Helbert. A standardized distance-based index to assess the quality of space-filling designs. Statistics and Computing, 27(2):319–329, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9624-z.

Stramer:2017:BIH

[1263] Osnat Stramer, Xiaoyu Shen, and Matthew Bognar. Bayesian inference for Heston–STAR models. Statistics and Computing, 27(2):331–348, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-015-9625-y.

Hadjiantoni:2017:ELS

[1264] Stella Hadjiantoni and Erricos John Kontoghiorghes. Estimating large-scale general linear and seemingly unrelated regressions models after deleting observations. Statistics and Computing, 27(2):349–361, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-016-9626-
                                        5.

Katzfuss:2017:PIM

[1265] Matthias Katzfuss and Dorit Hammerling. Parallel inference for massive distributed spatial data using low-rank models. Statistics and Computing, 27 (2):363–375, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9627-
4.

Garcia-Escudero:2017:REM

[1266] L. A. García-Escudero, A. Gordaliza, F. Greselin, S. Ingrassia, and A. Mayo-Iscar. Robust estimation of mixtures of regressions with random covariates, via trimming and constraints. Statistics and Computing, 27(2):377–402, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-016-9628-3.

Everitt:2017:BMC

[1267] Richard G. Everitt, Adam M. Johansen, Ellen Rowing, and Melina Evdemon-Hogan. Bayesian model comparison with un-normalised likelihoods. Statistics and Computing, 27 (2):403–422, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9629-
2.

Cao:2017:HEP

[1268] Yongtao Cao, Byran J. Smucker, and Timothy J. Robinson. A hybrid elitist pareto-based coordinate exchange algorithm for constructing multi-criteria optimal experimental designs. Statistics and Computing, 27(2):423–437, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-016-9630-9.

Walker:2017:LTI

[1269] Stephen G. Walker. A Laplace transform inversion method for probability distribution functions. Statistics and Computing, 27(2):439–448, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-016-9631-8.

Elfadaly:2017:EDG

[1270] Fadlalla G. Elfadaly and Paul H. Garthwaite. Eliciting Dirichlet and Gaussian copula prior distributions for multinomial models. Statistics and Computing, 27(2):449–467, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        016-9632-7.pdf.

Borrotti:2017:MOC

[1271] Matteo Borrotti, Francesco Sambo, Kalliopi Mylona, and Steven Gilmour. A multi-objective coordinate-exchange two-phase local search algorithm for multi-stratum experiments. Statistics and Computing, 27(2):469–481, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-016-9633-6.

Chiou:2017:RBE

[1272] Sy Han Chiou and Gongjun Xu. Rankbased estimation for semiparametric accelerated failure time model under length-biased sampling. Statistics and Computing, 27(2):483–500, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-016-9634-5.

Audigier:2017:MMI

[1273] Vincent Audigier, François Husson, and Julie Josse. MIMCA: multiple imputation for categorical variables with multiple correspondence analysis. Statistics and Computing, 27 (2):501–518, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9635-
4.

Maidstone:2017:OMC

[1274] Robert Maidstone, Toby Hocking, Guillem Rigaill, and Paul Fearnhead. On optimal multiple changepoint algorithms for large data. Statistics and Computing, 27(2):519–533, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
016-9636-3.pdf.

Cheng:2017:RRS

[1275] Guosheng Cheng, Xingxiang Li, Peng Lai, Fengli Song, and Jun Yu. Robust rank screening for ultrahigh dimensional discriminant analysis. Statistics and Computing, 27(2):535–545, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-016-9637-2.

Alfo:2017:FMQ

[1276] Marco Alfò, Nicola Salvati, and M. Giovanna Ranallli. Finite mixtures of quantile and M-quantile regression models. Statistics and Computing, 27 (2):547–570, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-016-9638-

1. See correction [1942].

Yu:2017:PSE

[1277] Yan Yu, Chaojiang Wu, and Yuankun Zhang. Penalised spline estimation for generalised partially linear single-index models. Statistics and Computing, 27 (2):571–582, March 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-016-9639-
                                        0.

Bornn:2017:USP

[1278] Luke Bornn, Natesh S. Pillai, Aaron Smith, and Dawn Woodard. The use of a single pseudo-sample in approximate Bayesian computation. Statistics and Computing, 27(3):583–590, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-016-9640-7.

Maroufy:2017:MMB

[1279] Vahed Maroufy and Paul Marriott. Mixture models: building a parameter space. Statistics and Computing, 27(3):591–597, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9641-
6.

Martino:2017:LAI

[1280] L. Martino, V. Elvira, D. Luengo, and J. Corander. Layered adaptive importance sampling. Statistics and Computing, 27(3):599–623, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9642-
5.

Amaral:2017:OLN

[1281] Sergio Amaral, Douglas Allaire, and Karen Willcox. Optimal L_{2}-norm empirical importance weights for the change of probability measure. Statistics and Computing, 27(3):625–643, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-016-9644-3.

Romano:2017:SVC

[1282] Elvira Romano, Antonio Balzanella, and Rosanna Verde. Spatial variability clustering for spatially dependent functional data. Statistics and Computing, 27(3):645–658, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9645-
2.

Gregorutti:2017:CVI

[1283] Baptiste Gregorutti, Bertrand Michel, and Philippe Saint-Pierre. Correlation and variable importance in random forests. Statistics and Computing, 27(3):659–678, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s11222-016-9646-
                                        1.

Dass:2017:LBA

[1284] Sarat C. Dass, Jaeyong Lee, Kyoungjae Lee, and Jonghun Park. Laplace based approximate posterior inference for differential equation models. Statistics and Computing, 27(3):679–698, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-016-9647-0.

Devroye:2017:EBC

[1285] Luc Devroye and Claude Gravel. The expected bit complexity of the von Neumann rejection algorithm. Statistics and Computing, 27(3):699–710, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s11222-016-9648-z.

Piironen:2017:CBP

[1286] Juho Piironen and Aki Vehtari. Comparison of Bayesian predictive methods for model selection. Statistics and Computing, 27(3):711–735, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        016-9649-y.pdf.

Arnesen:2017:PSN

[1287] Petter Arnesen and Håkon Tjelmeland. Prior specification of neighbourhood and interaction structure in binary Markov random fields. Statistics and Computing, 27(3):737–756, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-016-9650-5.

Magnanensi:2017:NUR

[1288] Jérémy Magnanensi, Frédéric Bertrand, Myriam Maumy-Bertrand, and Nicolas Meyer. A new universal resamplestable bootstrap-based stopping criterion for PLS component construction. Statistics and Computing, 27 (3):757–774, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-016-9651-4.pdf.

OKeefe:2017:PRL

[1289] Christine M. O’Keefe, Tim Ayre, Sebastien Lucie, Atikur R. Khan, Soomin Song, and Soonmin Kwon. Perturbed robust linear estimating equations for confidentiality protection in remote analysis. Statistics and Computing, 27(3):775–787, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9653-
2.

Chiquet:2017:SRC

[1290] Julien Chiquet, Tristan Mary-Huard, and Stéphane Robin. Structured regularization for conditional Gaussian graphical models. Statistics and Computing, 27(3):789–804, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s11222-016-9654-1.

Lemyre:2017:MBM

[1291] Félix Camirand Lemyre and Jean-François Quessy. Multiplier bootstrap methods for conditional distributions. Statistics and Computing, 27(3):805–821, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9655-
0.

Gandy:2017:QQM

[1292] Axel Gandy and Georg Hahn. Quick-MMCTest: quick multiple Monte Carlo testing. Statistics and Computing, 27(3):823–832, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9656-
z.

Friel:2017:IWA

[1293] N. Friel, J. P. McKeone, C. J. Oates, and A. N. Pettitt. Investigation of the widely applicable Bayesian information criterion. Statistics and Computing, 27(3):833–844, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9657-
y.

Esteban-Bravo:2017:EOE

[1294] Mercedes Esteban-Bravo, Agata Leszkiewicz, and Jose M. Vidal-Sanz. Exact optimal experimental designs with constraints. Statistics and Computing, 27(3):845–863, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s11222-016-9658-
x.

Mkhadri:2017:CDA

[1295] Abdallah Mkhadri, Mohamed Ouhourane, and Karim Oualkacha. A coordinate descent algorithm for computing penalized smooth quantile regression. Statistics and Computing, 27(4):865–883, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Whitaker:2017:IBC

[1296] Gavin A. Whitaker, Andrew Golightly, Richard J. Boys, and Chris Sherlock. Improved bridge constructs for stochastic differential equations. Statistics and Computing, 27(4):885–900, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
016-9660-3.pdf.

Zhu:2017:DBD

[1297] Houying Zhu and Josef Dick. A discrepancy bound for deterministic acceptance-rejection samplers beyond N^{−1/2} in dimension 1. Statistics and Computing, 27(4):901–911, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Brockhaus:2017:BFF

[1298] Sarah Brockhaus, Michael Melcher, Friedrich Leisch, and Sonja Greven. Boosting flexible functional regression models with a high number of functional historical effects. Statistics and Computing, 27(4):913–926, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Karagiannis:2017:PIS

[1299] Georgios Karagiannis, Bledar A. Konomi, Guang Lin, and Faming Liang. Parallel and interacting stochastic approximation annealing algorithms for global optimisation. Statistics and Computing, 27(4):927–945, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Zhou:2017:MFF

[1300] Tingyou Zhou and Liping Zhu. Modelfree feature screening for ultrahigh dimensional censored regression. Statistics and Computing, 27(4):947–961, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Migliorati:2017:SDM

[1301] Sonia Migliorati, Andrea Ongaro, and Gianna S. Monti. A structured Dirichlet mixture model for compositional data: inferential and applicative issues. Statistics and Computing, 27 (4):963–983, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Wood:2017:PSD

[1302] Simon N. Wood. P-splines with derivative based penalties and tensor product smoothing of unevenly distributed data. Statistics and Computing, 27 (4):985–989, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-016-9666-x.pdf.

Georgoulas:2017:UBI

[1303] Anastasis Georgoulas, Jane Hillston, and Guido Sanguinetti. Unbiased Bayesian inference for population Markov jump processes via random truncations. Statistics and Computing, 27(4):991–1002, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-016-9667-9.pdf.

Aderhold:2017:ABI

[1304] Andrej Aderhold, Dirk Husmeier, and Marco Grzegorczyk. Approximate Bayesian inference in semi-mechanistic models. Statistics and Computing, 27 (4):1003–1040, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-016-9668-8.pdf.

Yau:2017:LTA

[1305] Chun Yip Yau and Tsz Shing Hui.

LARS-type algorithm for group lasso. Statistics and Computing, 27(4):1041– 1048, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Marbac:2017:VSM

[1306] Matthieu Marbac and Mohammed Sedki. Variable selection for modelbased clustering using the integrated complete-data likelihood. Statistics and Computing, 27(4):1049–1063, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Conrad:2017:SAD

[1307] Patrick R. Conrad, Mark Girolami, Simo Särkkä, Andrew Stuart, and Konstantinos Zygalakis. Statistical analysis of differential equations: introducing probability measures on numerical solutions. Statistics and Computing, 27(4):1065–1082, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-016-9671-0.pdf.

Muehlenstaedt:2017:CEF

[1308] Thomas Muehlenstaedt, Jana Fruth, and Olivier Roustant. Computer experiments with functional inputs and scalar outputs by a norm-based approach. Statistics and Computing, 27 (4):1083–1097, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Veloso:2017:MSM

[1309] Bráulio M. Veloso, Thais R. Correa, Marcos O. Prates, Gabriel F. Oliveira, and Andréa I. Tavares. MAD–STEC: a method for multiple automatic detection of space-time emerging clusters. Statistics and Computing, 27 (4):1099–1110, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Barbillon:2017:PEC

[1310] Pierre Barbillon, Célia Barthélémy, and Adeline Samson. Parameter estimation of complex mixed models based on meta-model approach. Statistics and Computing, 27(4):1111–1128, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Taylor:2017:MLS

[1311] Sarah L. Taylor, Idris A. Eckley, and Matthew A. Nunes. Multivariate locally stationary 2D wavelet processes with application to colour texture analysis. Statistics and Computing, 27 (4):1129–1143, July 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11222-016-9675-9.pdf.

Laclau:2017:DLB

[1312] Charlotte Laclau and Mohamed Nadif.

Diagonal latent block model for binary data. Statistics and Computing, 27(5):1145–1163, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Vitoratou:2017:TBM

[1313] Silia Vitoratou and Ioannis Ntzoufras.

Thermodynamic Bayesian model comparison. Statistics and Computing, 27 (5):1165–1180, September 2017. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Chen:2017:MDF

[1314] Lu-Hung Chen and Ci-Ren Jiang.

Multi-dimensional functional principal component analysis. Statistics and Computing, 27(5):1181–1192, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Zhu:2017:AMT

[1315] Xuehu Zhu, Xu Guo, and Lixing Zhu. An adaptive-to-model test for partially parametric single-index models. Statistics and Computing, 27(5): 1193–1204, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Doshi-Velez:2017:RIB

[1316] Finale Doshi-Velez and Sinead A. Williamson. Restricted Indian buffet processes. Statistics and Computing, 27(5):1205–1223, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Stoica:2017:ASA

[1317] Radu S. Stoica, Anne Philippe, Pablo Gregori, and Jorge Mateu. ABC Shadow algorithm: a tool for statistical analysis of spatial patterns. Statistics and Computing, 27(5):1225–1238, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Mrkvicka:2017:MMC

[1318] Tomás Mrkvicka, Mari Myllymäki, and Ute Hahn. Multiple Monte Carlo testing, with applications in spatial point processes. Statistics and Computing, 27(5):1239–1255, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Bodenham:2017:CMC

[1319] Dean A. Bodenham and Niall M. Adams. Continuous monitoring for changepoints in data streams using adaptive estimation. Statistics and Computing, 27(5):1257–1270, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Austad:2017:ACB

[1320] Haakon Michael Austad and Håkon Tjelmeland. Approximate computations for binary Markov random fields and their use in Bayesian models. Statistics and Computing, 27(5): 1271–1292, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Haynes:2017:CEN

[1321] Kaylea Haynes, Paul Fearnhead, and Idris A. Eckley. A computationally efficient nonparametric approach for changepoint detection. Statistics and Computing, 27(5):1293–1305, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11222-
016-9687-5.pdf.

Cambou:2017:QRN

[1322] Mathieu Cambou, Marius Hofert, and Christiane Lemieux. Quasi-random numbers for copula models. Statistics and Computing, 27(5):1307–1329, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Schwaller:2017:EBI

[1323] L. Schwaller and S. Robin. Exact Bayesian inference for off-line changepoint detection in tree-structured graphical models. Statistics and Computing, 27(5):1331–1345, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Wang:2017:PEL

[1324] Shanshan Wang and Liming Xiang. Penalized empirical likelihood inference for sparse additive hazards regression with a diverging number of covariates. Statistics and Computing, 27(5): 1347–1364, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Julian-Moreno:2017:FPA

[1325] Guillermo Julián-Moreno, Jorge E. López de Vergara, Iván González, Luis de Pedro, Javier Royuela del Val, and Federico Simmross-Wattenberg. Fast parallel α-stable distribution function evaluation and parameter estimation using OpenCL in GPGPUs. Statistics and Computing, 27(5):1365–1382, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Jansen:2017:MLP

[1326] Maarten Jansen and Mohamed Amghar. Multiscale local polynomial decompositions using bandwidths as scales. Statistics and Computing, 27(5):1383– 1399, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Lin:2017:SPH

[1327] Bingqing Lin, Qihua Wang, Jun Zhang, and Zhen Pang. Stable prediction in high-dimensional linear models. Statistics and Computing, 27(5):1401–1412, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Vehtari:2017:PBM

[1328] Aki Vehtari, Andrew Gelman, and Jonah Gabry. Practical Bayesian model evaluation using leave-one-out crossvalidation and WAIC. Statistics and Computing, 27(5):1413–1432, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). See erratum [1329].

Vehtari:2017:EPB

[1329] Aki Vehtari, Andrew Gelman, and Jonah Gabry. Erratum to: Practical Bayesian model evaluation using leave-one-out cross-validation and WAIC. Statistics and Computing, 27(5):1433, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
content/pdf/10.1007/s11222-016-

9709-3.pdf. See [1328].

Bader:2017:ASL

[1330] Brian Bader, Jun Yan, and Xuebin Zhang. Automated selection of r for the r largest order statistics approach with adjustment for sequential testing. Statistics and Computing, 27(6): 1435–1451, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9697-3;   http:
//link.springer.com/content/pdf/
10.1007/s11222-016-9697-3.pdf.

Knight:2017:WLA

[1331] Marina I. Knight, Guy P. Nason, and Matthew A. Nunes. A wavelet lifting approach to long-memory estimation. Statistics and Computing, 27(6):

1453–1471, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9698-2;   http:
//link.springer.com/content/pdf/
10.1007/s11222-016-9698-2.pdf.

Zhang:2017:HMC

[1332] Cheng Zhang, Babak Shahbaba, and Hongkai Zhao. Hamiltonian Monte Carlo acceleration using surrogate functions with random bases. Statistics and Computing, 27(6):1473– 1490, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9699-1.

Bass:2017:CCP

[1333] Mark R. Bass and Sujit K. Sahu. A comparison of centring parameterisations of Gaussian process-based models for Bayesian computation using MCMC. Statistics and Computing, 27(6):1491–1512, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9700-z;   http://link.springer.
com/content/pdf/10.1007/s11222-
016-9700-z.pdf.

Kordzakhia:2017:AWK

[1334] Nino Kordzakhia, Alexander Novikov, and Bernard Ycart. Approximations for weighted Kolmogorov–Smirnov distributions via boundary crossing probabilities. Statistics and Computing, 27(6):1513–1523, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9701-y;   http://link.springer.
com/content/pdf/10.1007/s11222-
016-9701-y.pdf.

DeYoreo:2017:BNM

[1335] Maria DeYoreo and Athanasios Kottas. A Bayesian nonparametric Markovian model for non-stationary time series. Statistics and Computing, 27(6): 1525–1538, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9702-x.

Waldmann:2017:BRG

[1336] Elisabeth Waldmann, Fabian Sobotka, and Thomas Kneib. Bayesian regularisation in geoadditive expectile regression. Statistics and Computing, 27(6): 1539–1553, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9703-9.

Dunlop:2017:HBL

[1337] Matthew M. Dunlop, Marco A. Iglesias, and Andrew M. Stuart. Hierarchical Bayesian level set inversion. Statistics and Computing, 27(6): 1555–1584, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9704-8.

Titman:2017:NPM

[1338] Andrew C. Titman. Non-parametric maximum likelihood estimation of interval-censored failure time data subject to misclassification. Statistics and Computing, 27(6):1585–1593, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9705-7.

Otneim:2017:LGD

[1339] Håkon Otneim and Dag Tjøstheim. The locally Gaussian density estimator for multivariate data. Statistics and Computing, 27(6):1595–1616, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9706-6.

Same:2017:SDF

[1340] Allou Samé and Gérard Govaert. Segmental dynamic factor analysis for time series of curves. Statistics and Computing, 27(6):1617–1637, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9707-5.

Liu:2017:FPC

[1341] Chong Liu, Surajit Ray, and Giles Hooker. Functional principal component analysis of spatially correlated data. Statistics and Computing, 27(6): 1639–1654, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9708-4;   http:
//link.springer.com/content/pdf/
10.1007/s11222-016-9708-4.pdf.

Perez:2017:FRE

[1342] B. Pérez, I. Molina, A. Thieler, R. Fried, and D. Peña. Fast and robust estimators of variance components in the nested error model. Statistics and Computing, 27(6):1655–1675, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9710-x.

Nguyen:2017:SOI

[1343] Dao Nguyen and Edward L. Ionides. A second-order iterated smoothing algorithm. Statistics and Computing, 27(6): 1677–1692, November 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9711-9.

Shang:2018:BMS

[1344] Han Lin Shang. Bootstrap methods for stationary functional time series. Statistics and Computing, 28(1):1–10, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9712-8.

Bouveyron:2018:STB

[1345] C. Bouveyron, P. Latouche, and R. Zreik. The stochastic topic block model for the clustering of vertices in networks with textual edges. Statistics and Computing, 28(1):11–31, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9713-7.

Wang:2018:ESR

[1346] Tao Wang, Mengjie Chen, Hongyu Zhao, and Lixing Zhu. Estimating a sparse reduction for general regression in high dimensions. Statistics and Computing, 28(1):33–46, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-016-
                                        9714-6.

Jasra:2018:MPF

[1347] Ajay Jasra, Kengo Kamatani, Prince Peprah Osei, and Yan Zhou. Multilevel particle filters: normalizing constant estimation. Statistics and Computing, 28(1):47–60, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-016-9715-5.

Rached:2018:GHR

[1348] Nadhir Ben Rached, Fatma Benkhelifa, Abla Kammoun, Mohamed-Slim Alouini, and Raul Tempone. On the generalization of the hazard rate twisting-based simulation approach. Statistics and Computing, 28(1):61–75, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-016-
                                        9716-4.

Hooker:2018:BBC

[1349] Giles Hooker and Lucas Mentch. Bootstrap bias corrections for ensemble methods. Statistics and Computing, 28(1):77–86, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9717-3.

Richardson:2018:BNP

[1350] Robert Richardson, Athanasios Kottas, and Bruno Sansó. Bayesian non-parametric modeling for integrodifference equations. Statistics and Computing, 28(1):87–101, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9719-1.

Cakiroglu:2018:ORG

[1351] Sera Aylin Cakiroglu. Optimal regular graph designs. Statistics and Computing, 28(1):103–112, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-016-
9720-8.

Zhang:2018:LSK

[1352] Qinyi Zhang, Sarah Filippi, Arthur Gretton, and Dino Sejdinovic. Largescale kernel methods for independence testing. Statistics and Computing, 28 (1):113–130, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-016-9721-7;   http:
//link.springer.com/content/pdf/
10.1007/s11222-016-9721-7.pdf.

Koskela:2018:IRE

[1353] Jere Koskela, Dario Spanò, and Paul A.

Jenkins. Inference and rare event simulation for stopped Markov processes via reverse-time sequential Monte Carlo. Statistics and Computing, 28 (1):131–144, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9722-1.

Galimberti:2018:MRV

[1354] Giuliano Galimberti, Annamaria Manisi, and Gabriele Soffritti. Modelling the role of variables in modelbased cluster analysis. Statistics and Computing, 28(1):145–169, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9723-0.

Ament:2018:AEN

[1355] Sebastian Ament and Michael O’Neil. Accurate and efficient numerical calculation of stable densities via optimized quadrature and asymptotics. Statistics and Computing, 28(1):171–185, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9725-y.

Wang:2018:NML

[1356] Yong Wang and Shabnam Fani. Nonparametric maximum likelihood computation of a U-shaped hazard function. Statistics and Computing, 28 (1):187–200, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9724-z.

Gascon:2018:ASP

[1357] Alberto Gascón and Eugenio F. Sánchez-Úbeda. Automatic specification of piecewise linear additive models: application to forecasting natural gas demand. Statistics and Computing, 28(1):201–217, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9726-x.

Zhang:2018:VCA

[1358] Mimi Zhang and Tim Bedford. Vine copula approximation: a generic method for coping with conditional dependence. Statistics and Computing, 28(1):219–237, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9727-9;   http:
//link.springer.com/content/pdf/
10.1007/s11222-017-9727-9.pdf.

Medina-Aguayo:2018:ESN

[1359] F. J. Medina-Aguayo, A. Lee, and G. O. Roberts. Erratum to: Stability of noisy Metropolis–Hastings. Statistics and Computing, 28(1):239, January 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9755-5;   http://link.springer.
com/content/pdf/10.1007/s11222-
017-9755-5.pdf. See [1233].

Wiens:2018:RDR

[1360] Douglas P. Wiens. I-robust and D-robust designs on a finite design space. Statistics and Computing, 28 (2):241–258, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9728-8.

Tan:2018:GVA

[1361] Linda S. L. Tan and David J. Nott. Gaussian variational approximation with sparse precision matrices. Statistics and Computing, 28(2):259–275, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9729-7.

Leimkuhler:2018:EPM

[1362] Benedict Leimkuhler, Charles Matthews, and Jonathan Weare. Ensemble preconditioning for Markov chain Monte Carlo simulation. Statistics and Computing, 28(2):277–290, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9730-1;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        017-9730-1.pdf.

Norwood:2018:LMC

[1363] Ben Norwood and Rebecca Killick. Long memory and changepoint models: a spectral classification procedure. Statistics and Computing, 28 (2):291–302, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9731-0;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-017-9731-0.pdf.

Otneim:2018:CDE

[1364] Håkon Otneim and Dag Tjøstheim. Conditional density estimation using the local Gaussian correlation. Statistics and Computing, 28(2):303–321, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9732-z.

Killiches:2018:MDV

[1365] Matthias Killiches, Daniel Kraus, and Claudia Czado. Model distances for vine copulas in high dimensions. Statistics and Computing, 28(2):323–341, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9733-y.

Overstall:2018:AFF

[1366] Antony M. Overstall, James M. Mc-Gree, and Christopher C. Drovandi. An approach for finding fully Bayesian optimal designs using normal-based approximations to loss functions. Statistics and Computing, 28(2):343–358, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9734-x;   http://link.springer.
com/content/pdf/10.1007/s11222-
017-9734-x.pdf.

Paci:2018:DMB

[1367] Lucia Paci and Francesco Finazzi.

Dynamic model-based clustering for spatio-temporal data. Statistics and Computing, 28(2):359–374, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9735-9.

Millar:2018:CVM

[1368] Russell B. Millar. Conditional vs marginal estimation of the predictive loss of hierarchical models using WAIC and cross-validation. Statistics and Computing, 28(2):375–385, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9736-8.

Schellhase:2018:ENS

[1369] Christian Schellhase and Fabian Spanhel. Estimating non-simplified vine copulas using penalized splines. Statistics and Computing, 28(2):387–409, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9737-7.

Gutmann:2018:LFI

[1370] Michael U. Gutmann, Ritabrata Dutta, Samuel Kaski, and Jukka Corander. Likelihood-free inference via classification. Statistics and Computing, 28 (2):411–425, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9738-6;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-017-9738-6.pdf.

Espuny-Pujol:2018:GOA

[1371] Ferran Espuny-Pujol, Karyn Morrissey, and Paul Williamson. A global optimisation approach to rangerestricted survey calibration. Statistics and Computing, 28(2):427–439, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9739-5;   http://link.springer.
com/content/pdf/10.1007/s11222-
017-9739-5.pdf.

Duarte:2018:AGS

[1372] Belmiro P. M. Duarte, Weng Kee Wong, and Holger Dette. Adaptive grid semidefinite programming for finding optimal designs. Statistics and Computing, 28(2):441–460, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9741-y.

Sen:2018:CPF

[1373] Deborshee Sen, Alexandre H. Thiery, and Ajay Jasra. On coupling particle filter trajectories. Statistics and Computing, 28(2):461–475, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9740-z.

Dotto:2018:RAR

[1374] Francesco Dotto, Alessio Farcomeni, Luis Angel García-Escudero, and Agustín Mayo-Iscar. A reweighting approach to robust clustering. Statistics and Computing, 28(2):477–493, March 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9742-x.

Jiang:2018:TMM

[1375] Lan Jiang and Sumeetpal S. Singh. Tracking multiple moving objects in images using Markov Chain Monte Carlo. Statistics and Computing, 28 (3):495–510, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9743-9.

Xiao:2018:FCE

[1376] Luo Xiao, Cai Li, William Checkley, and Ciprian Crainiceanu. Fast covariance estimation for sparse functional data. Statistics and Computing, 28(3):511–522, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9744-8;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-017-9744-8.pdf. See

erratum [1377].

Xiao:2018:EFC

[1377] Luo Xiao, Cai Li, William Checkley, and Ciprian Crainiceanu. Erratum to: Fast covariance estimation for sparse functional data. Statistics and Computing, 28(3):523, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9768-0;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        017-9768-0.pdf. See [1376].

Vettori:2018:CDF

[1378] Sabrina Vettori, Raphaël Huser, and Marc G. Genton. A comparison of dependence function estimators in multivariate extremes. Statistics and Computing, 28(3):525–538, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9745-7.

Hand:2018:NUM

[1379] David Hand and Peter Christen. A note on using the F -measure for evaluating record linkage algorithms. Statistics and Computing, 28(3):539–547, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9746-6.

Fasiolo:2018:LIM

[1380] Matteo Fasiolo, Flávio Eler de Melo, and Simon Maskell. Langevin incremental mixture importance sampling. Statistics and Computing, 28 (3):549–561, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9747-5.

Lindo:2018:NEC

[1381] Alexey Lindo, Sergei Zuyev, and Serik Sagitov. Nonparametric estimation for compound Poisson process via variational analysis on measures. Statistics and Computing, 28(3):563–577, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9748-4;   http://link.springer.
com/content/pdf/10.1007/s11222-
017-9748-4.pdf.

Charitidou:2018:OBT

[1382] E. Charitidou, D. Fouskakis, and I. Ntzoufras. Objective Bayesian transformation and variable selection using default Bayes factors. Statistics and Computing, 28(3):579–594, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9749-3.

Chanialidis:2018:EBI

[1383] Charalampos Chanialidis, Ludger Evers, Tereza Neocleous, and Agostino Nobile. Efficient Bayesian inference for COM–Poisson regression models. Statistics and Computing, 28 (3):595–608, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9750-x;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-017-9750-x.pdf.

Meyer:2018:DED

[1384] Daniel W. Meyer. Density estimation with distribution element trees. Statistics and Computing, 28(3):609–632, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9751-9.

Shah:2018:RSP

[1385] Rohan Shah and Dirk P. Kroese. Without-replacement sampling for particle methods on finite state spaces. Statistics and Computing, 28(3):633–652, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9752-8.

Kamper:2018:RGB

[1386] François Kamper, Johan A. du Preez, Sarel J. Steel, and Stephan Wagner. Regularized Gaussian belief propagation. Statistics and Computing, 28(3):653–672, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9753-7.

Thomas:2018:GBD

[1387] Janek Thomas, Andreas Mayr, Bernd Bischl, Matthias Schmid, Adam Smith, and Benjamin Hofner. Gradient boosting for distributional regression: faster tuning and improved variable selection via noncyclical updates. Statistics and Computing, 28(3):673–687, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9754-6.

Paine:2018:ESA

[1388] P. J. Paine, S. P. Preston, M. Tsagris, and Andrew T. A. Wood. An elliptically symmetric angular Gaussian distribution. Statistics and Computing, 28(3):689–697, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9756-4;   http:
//link.springer.com/content/pdf/
10.1007/s11222-017-9756-4.pdf.

Chen:2018:LSI

[1389] Zhi-Yong Chen and Hai-Bin Wang. Latent single-index models for ordinal data. Statistics and Computing, 28(3):699–711, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9757-3.

Nie:2018:SFP

[1390] Yunlong Nie, Liangliang Wang, Baisen Liu, and Jiguo Cao. Supervised functional principal component analysis. Statistics and Computing, 28 (3):713–723, May 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9758-2.

Rugama:2018:REE

[1391] Lluís Antoni Jiménez Rugama and Laurent Gilquin. Reliable error estimation for Sobol’ indices. Statistics and Computing, 28(4):725–738, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9759-1.

Leatherman:2018:CED

[1392] Erin R. Leatherman, Thomas J. Santner, and Angela M. Dean. Computer experiment designs for accurate prediction. Statistics and Computing, 28(4):739–751, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9760-8.

Pazira:2018:EDG

[1393] Hassan Pazira, Luigi Augugliaro, and Ernst Wit. Extended differential geometric LARS for high-dimensional GLMs with general dispersion parameter. Statistics and Computing, 28(4):753–774, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9761-7.

Zhao:2018:PVD

[1394] Junlong Zhao, Hongyu Zhao, and Lixing Zhu. Pivotal variable detection of the covariance matrix and its application to high-dimensional factor models. Statistics and Computing, 28(4):775–793, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9762-6.

Kleppe:2018:MCR

[1395] Tore Selland Kleppe. Modified Cholesky Riemann Manifold Hamiltonian Monte Carlo: exploiting sparsity for fast sampling of high-dimensional targets. Statistics and Computing, 28(4):795–817, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9763-5.

Prangle:2018:REA

[1396] Dennis Prangle, Richard G. Everitt, and Theodore Kypraios. A rare event approach to high-dimensional approximate Bayesian computation. Statistics and Computing, 28(4):819–834, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9764-4;   http://link.springer.
com/content/pdf/10.1007/s11222-
017-9764-4.pdf.

Kume:2018:EML

[1397] A. Kume and T. Sei. On the exact maximum likelihood inference of Fisher–Bingham distributions using an adjusted holonomic gradient method. Statistics and Computing, 28(4):835–847, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9765-3;   http:
//link.springer.com/content/pdf/
10.1007/s11222-017-9765-3.pdf.

Rulliere:2018:NKP

[1398] Didier Rullière, Nicolas Durrande, François Bachoc, and Clément Chevalier. Nested kriging predictions for datasets with a large number of observations. Statistics and Computing, 28(4):849–867, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9766-2.

Hernandez:2018:BAR

[1399] Belinda Hernández, Adrian E. Raftery, Stephen R. Pennington, and Andrew C. Parnell. Bayesian additive regression trees using Bayesian model averaging. Statistics and Computing, 28(4):869–890, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9767-1.

Lee:2018:NEM

[1400] Clement Lee, Andrew Garbett, and Darren J. Wilkinson. A network epidemic model for online community commissioning data. Statistics and Computing, 28(4):891–904, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9770-6;   http://link.springer.
com/content/pdf/10.1007/s11222-
017-9770-6.pdf.

Champion:2018:ILG

[1401] Magali Champion, Victor Picheny, and Matthieu Vignes. Inferring large graphs using ℓ_{1}-penalized likelihood. Statistics and Computing, 28(4):905–921, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9769-z;   http:
//link.springer.com/content/pdf/
10.1007/s11222-017-9769-z.pdf. See

correction [1420].

Haji-Ali:2018:MMI

[1402] Abdul-Lateef Haji-Ali and Raúl Tempone. Multilevel and multiindex Monte Carlo methods for the McKean–Vlasov equation. Statistics and Computing, 28(4):923–935, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9771-5;   http://link.springer.

                                        com/content/pdf/10.1007/s11222-
                                        017-9771-5.pdf.

Davies:2018:FCS

[1403] Tilman M. Davies and Adrian Baddeley. Fast computation of spatially adaptive kernel estimates. Statistics and Computing, 28(4):937–956, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9772-4.

Egidi:2018:RBM

[1404] Leonardo Egidi, Roberta Pappadà, Francesco Pauli, and Nicola Torelli. Relabelling in Bayesian mixture models by pivotal units. Statistics and Computing, 28(4):957–969, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9774-2.

Ong:2018:VBS

[1405] Victor M. H. Ong, David J. Nott, Minh-Ngoc Tran, Scott A. Sisson, and Christopher C. Drovandi. Variational Bayes with synthetic likelihood. Statistics and Computing, 28(4):971–988, July 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9773-3.

Corneli:2018:MCP

[1406] Marco Corneli, Pierre Latouche, and Fabrice Rossi. Multiple change points detection and clustering in dynamic networks. Statistics and Computing, 28(5):989–1007, September 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9775-1.

Page:2018:CCI

[1407] Garritt L. Page and Fernando A.

Quintana. Calibrating covariate informed product partition models. Statistics and Computing, 28(5):1009– 1031, September 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9777-z.

Gomez-Rubio:2018:MCM

[1408] Virgilio Gómez-Rubio and Håvard Rue.

Markov chain Monte Carlo with the Integrated Nested Laplace Approximation. Statistics and Computing, 28(5): 1033–1051, September 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9778-y.

Folia:2018:TIP

[1409] Maria Myrto Folia and Magnus Rattray. Trajectory inference and parameter estimation in stochastic models with temporally aggregated data. Statistics and Computing, 28(5):1053– 1072, September 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9779-x;   http:
//link.springer.com/content/pdf/
10.1007/s11222-017-9779-x.pdf.

Yang:2018:PMC

[1410] Shihao Yang, Yang Chen, Espen Bernton, and Jun S. Liu. On parallelizable Markov chain Monte Carlo algorithms with waste-recycling. Statistics and Computing, 28(5):1073–1081, September 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9780-4.

Wang:2018:NAE

[1411] Xin Wang, Vivekananda Roy, and Zhengyuan Zhu. A new algorithm to estimate monotone nonparametric link functions and a comparison with parametric approach. Statistics and Computing, 28(5):1083–1094, September 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9781-3.

Dong:2018:MLE

[1412] Fanghu Dong and Guosheng Yin. Maximum likelihood estimation for incomplete multinomial data via the weaver algorithm. Statistics and Computing, 28(5):1095–1117, September 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9782-2.

Law:2018:CMO

[1413] Jonathan Law and Darren J. Wilkinson. Composable models for online Bayesian analysis of streaming data. Statistics and Computing, 28(6): 1119–1137, November 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9783-1;   http:
//link.springer.com/content/pdf/
10.1007/s11222-017-9783-1.pdf.

Nelson:2018:LSD

[1414] J. D. B. Nelson, A. J. Gibberd, C. Nafornita, and N. Kingsbury. The locally stationary dual-tree complex wavelet model. Statistics and Computing, 28(6):1139–1154, November 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9784-0;   http://link.springer.
com/content/pdf/10.1007/s11222-
017-9784-0.pdf.

Richard:2018:AHG

[1415] Frédéric J. P. Richard. Anisotropy of Hölder Gaussian random fields: characterization, estimation, and application to image textures. Statistics and Computing, 28(6):1155–1168, November 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9785-z.

Rastelli:2018:OBE

[1416] Riccardo Rastelli and Nial Friel. Optimal Bayesian estimators for latent variable cluster models. Statistics and Computing, 28(6):1169–1186, November 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9786-y;   http://link.springer.

                                        com/content/pdf/10.1007/s11222-
                                        017-9786-y.pdf.

Perez:2018:AVB

[1417] Iker Perez, David Hodge, and Theodore Kypraios. Auxiliary variables for Bayesian inference in multi-class queueing networks. Statistics and Computing, 28(6):1187–1200, November 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9787-x.

Ludkin:2018:DSB

[1418] Matthew Ludkin, Idris Eckley, and Peter Neal. Dynamic stochastic block models: parameter estimation and detection of changes in community structure. Statistics and Computing, 28(6): 1201–1213, November 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9788-9.

Golightly:2018:ESS

[1419] Andrew Golightly and Theodore Kypraios. Efficient SMC^{2} schemes for stochastic kinetic models. Statistics and Computing, 28(6):1215– 1230, November 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9789-8;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-017-9789-8.pdf.

Champion:2018:CIL

[1420] Magali Champion, Victor Picheny, and Matthieu Vignes. Correction to: Inferring large graphs using ℓ_{1}- penalized likelihood. Statistics and Computing, 28(6):1231, November 2018. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9795-x;   http://link.springer.
com/content/pdf/10.1007/s11222-
017-9795-x.pdf. See [1401].

Garcia-Portugues:2019:LDT

[1421] Eduardo García-Portugués, Michael Sørensen, Kanti V. Mardia, and Thomas Hamelryck. Langevin diffusions on the torus: estimation and applications. Statistics and Computing, 29(1):1–22, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9790-2.

Xue:2019:DPM

[1422] Jingnan Xue and Faming Liang.

Double-Parallel Monte Carlo for Bayesian analysis of big data. Statistics and Computing, 29(1):23–32, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9791-1.

Su:2019:BLE

[1423] Lin Su and Howard D. Bondell. Best linear estimation via minimization of relative mean squared error. Statistics and Computing, 29(1):33–42, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9792-0.

Viroli:2019:DGM

[1424] Cinzia Viroli and Geoffrey J. McLachlan. Deep Gaussian mixture models. Statistics and Computing, 29(1):43–51, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9793-z.

He:2019:ANE

[1425] Zhijian He and Lingjiong Zhu. Asymptotic normality of extensible grid sampling. Statistics and Computing, 29 (1):53–65, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-017-9794-y.

Edwards:2019:BNS

[1426] Matthew C. Edwards, Renate Meyer, and Nelson Christensen. Bayesian nonparametric spectral density estimation using B-spline priors. Statistics and Computing, 29(1):67–78, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9796-9.

Bon:2019:FMP

[1427] Joshua J. Bon, Kevin Murray, and Berwin A. Turlach. Fitting monotone polynomials in mixed effects models. Statistics and Computing, 29(1):79–98, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-017-
                                        9797-8.

Schober:2019:PMN

[1428] Michael Schober, Simo Särkkä, and Philipp Hennig. A probabilistic model for the numerical solution of initial value problems. Statistics and Computing, 29(1):99–122, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-017-
9798-7;   http://link.springer.
com/content/pdf/10.1007/s11222-
017-9798-7.pdf.

Spiegel:2019:GAM

[1429] Elmar Spiegel, Thomas Kneib, and Fabian Otto-Sobotka. Generalized additive models with flexible response functions. Statistics and Computing, 29(1):123–138, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-017-9799-6.

delBarrio:2019:RCT

[1430] E. del Barrio, J. A. Cuesta-Albertos, C. Matrán, and A. Mayo-Íscar. Robust clustering tools based on optimal transportation. Statistics and Computing, 29(1):139–160, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9800-z.

Gu:2019:PED

[1431] Jiaying Gu, Fei Fu, and Qing Zhou. Penalized estimation of directed acyclic graphs from discrete data. Statistics and Computing, 29(1):161–176, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-018-
                                        9801-y.

Ma:2019:ISJ

[1432] Yi-An Ma, Emily B. Fox, Tianqi Chen, and Lei Wu. Irreversible samplers from jump and continuous Markov processes. Statistics and Computing, 29 (1):177–202, January 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-018-9802-x.

Zuanetti:2019:BNC

[1433] Daiane Aparecida Zuanetti, Peter Müller, Yitan Zhu, Shengjie Yang, and Yuan Ji. Bayesian nonparametric clustering for large data sets. Statistics and Computing, 29(2):203–215, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-018-
                                        9803-9.

Berger:2019:TSM

[1434] Moritz Berger, Gerhard Tutz, and Matthias Schmid. Tree-structured modelling of varying coefficients. Statistics and Computing, 29(2): 217–229, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-018-9804-8.

Fort:2019:SPG

[1435] Gersende Fort, Edouard Ollier, and Adeline Samson. Stochastic proximalgradient algorithms for penalized mixed models. Statistics and Computing, 29(2):231–253, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.com/
article/10.1007/s11222-018-9805-
7.

Picheny:2019:ISS

[1436] Victor Picheny, Rémi Servien, and Nathalie Villa-Vialaneix. Interpretable sparse SIR for functional data. Statistics and Computing, 29(2):255–267, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9806-6.

Muller:2019:SSV

[1437] Dominik Müller and Claudia Czado.

Selection of sparse vine copulas in high dimensions with the Lasso. Statistics and Computing, 29(2):269–287, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9807-5.

Nomura:2019:OPE

[1438] Noboru Nomura. Orthant probabilities of elliptical distributions from orthogonal projections to subspaces. Statistics and Computing, 29(2):289–300, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9808-4.

Terenin:2019:GAG

[1439] Alexander Terenin, Shawfeng Dong, and David Draper. GPU-accelerated Gibbs sampling: a case study of the Horseshoe Probit model. Statistics and Computing, 29(2):301–310, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9809-3;   http://link.springer.
com/content/pdf/10.1007/s11222-
018-9809-3.pdf.

Cui:2019:LCS

[1440] Ruifei Cui, Perry Groot, and Tom Heskes. Learning causal structure from mixed data with missing values using Gaussian copula models. Statistics and Computing, 29(2):311–333, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9810-x;   http://link.springer.
com/content/pdf/10.1007/s11222-
018-9810-x.pdf.

Yu:2019:RAU

[1441] Philip L. H. Yu and Hang Xu. Rank aggregation using latent-scale distance-based models. Statistics and Computing, 29(2):335–349, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9811-9.

Papageorgiou:2019:SPP

[1442] Ioulia Papageorgiou and Irini Moustaki. Sampling of pairs in pairwise likelihood estimation for latent variable models with categorical observed variables. Statistics and Computing, 29(2):351–365, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9812-8.

Luo:2019:PSB

[1443] Xin Luo and Håkon Tjelmeland. Prior specification for binary Markov mesh models. Statistics and Computing, 29 (2):367–389, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9813-7.

Hofmeyr:2019:MSC

[1444] David P. Hofmeyr, Nicos G. Pavlidis, and Idris A. Eckley. Minimum spectral connectivity projection pursuit. Statistics and Computing, 29(2):391–414, March 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9814-6.

Maleki:2019:RFM

[1445] Mohsen Maleki, Darren Wraith, and Reinaldo B. Arellano-Valle. Robust finite mixture modeling of multivariate unrestricted skew-normal generalized hyperbolic distributions. Statistics and Computing, 29(3):415–428, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9815-5.

Glaws:2019:GCQ

[1446] Andrew Glaws and Paul G. Constantine. Gauss–Christoffel quadrature for inverse regression: applications to computer experiments. Statistics and Computing, 29(3):429–447, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9816-4.

Maire:2019:ISS

[1447] Florian Maire, Nial Friel, and Pierre Alquier. Informed sub-sampling MCMC: approximate Bayesian inference for large datasets. Statistics and Computing, 29(3):449–482, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9817-3.

Rodriguez-Alvarez:2019:EVP

[1448] María Xosé Rodríguez-Álvarez, Maria Durban, Dae-Jin Lee, and Paul H. C. Eilers. On the estimation of variance parameters in non-standard generalised linear mixed models: application to penalised smoothing. Statistics and Computing, 29(3):483–500, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9818-2.

Guastavino:2019:CNE

[1449] Sabrina Guastavino and Federico Benvenuto. A consistent and numerically efficient variable selection method for sparse Poisson regression with applications to learning and signal recovery. Statistics and Computing, 29(3):501–516, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9819-1.

Knight:2019:LME

[1450] Marina I. Knight and Matthew A.

Nunes. Long memory estimation for complex-valued time series. Statistics and Computing, 29(3):517–536, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9820-8;   http://link.springer.
com/content/pdf/10.1007/s11222-
018-9820-8.pdf.

Sugasawa:2019:LMM

[1451] Shonosuke Sugasawa, Genya Kobayashi, and Yuki Kawakubo. Latent mixture modeling for clustered data. Statistics and Computing, 29(3):537–548, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9821-7.

Grabchak:2019:RST

[1452] Michael Grabchak. Rejection sampling for tempered Lévy processes. Statistics and Computing, 29(3):549–558, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9822-6.

Abbruzzo:2019:STP

[1453] Antonino Abbruzzo, Ivan Vujacić, Angelo M. Mineo, and Ernst C. Wit. Selecting the tuning parameter in penalized Gaussian graphical models. Statistics and Computing, 29(3):559–569, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9823-5.

Shen:2019:SBA

[1454] Junshan Shen, Hanjun Yu, Jin Yang, and Chunling Liu. Semiparametric Bayesian analysis for longitudinal mixed effects models with non-normal AR(1) errors. Statistics and Computing, 29(3):571–583, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9824-4.

Cao:2019:HBC

[1455] Jian Cao, Marc G. Genton, David E. Keyes, and George M. Turkiyyah. Hierarchical-block conditioning approximations for high-dimensional multivariate normal probabilities. Statistics and Computing, 29(3):585–598, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9825-3.

Baker:2019:CVS

[1456] Jack Baker, Paul Fearnhead, Emily B. Fox, and Christopher Nemeth. Control variates for stochastic gradient MCMC. Statistics and Computing, 29(3):599–615, May 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9826-2;   http:
//link.springer.com/content/pdf/
10.1007/s11222-018-9826-2.pdf.

Black:2019:ISP

[1457] Andrew J. Black. Importance sampling for partially observed temporal epidemic models. Statistics and Computing, 29(4):617–630, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9827-1.

Heck:2019:QUT

[1458] Daniel W. Heck, Antony M. Overstall, Quentin F. Gronau, and Eric-Jan Wagenmakers. Quantifying uncertainty in transdimensional Markov chain Monte Carlo using discrete Markov models. Statistics and Computing, 29 (4):631–643, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9828-0;   http:
//link.springer.com/content/pdf/
10.1007/s11222-018-9828-0.pdf.

Chae:2019:ASF

[1459] Minwoo Chae, Ryan Martin, and Stephen G. Walker. On an algorithm for solving Fredholm integrals of the first kind. Statistics and Computing, 29(4):645–654, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9829-z.

Alonso:2019:CTS

[1460] Andrés M. Alonso and Daniel Peña.

Clustering time series by linear dependency. Statistics and Computing, 29(4):655–676, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9830-6.

Corneli:2019:DST

[1461] Marco Corneli, Charles Bouveyron, Pierre Latouche, and Fabrice Rossi. The dynamic stochastic topic block model for dynamic networks with textual edges. Statistics and Computing, 29(4):677–695, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9832-4.

Gillespie:2019:ECB

[1462] C. S. Gillespie and R. J. Boys. Efficient construction of Bayes optimal designs for stochastic process models. Statistics and Computing, 29(4):697–706, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9833-3.

Debruyne:2019:OWV

[1463] Michiel Debruyne, Sebastiaan Höppner, Sven Serneels, and Tim Verdonck. Outlyingness: Which variables contribute most? Statistics and Computing, 29(4):707–723, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9831-5.

Yue:2019:ODM

[1464] Yuguang Yue, Lieven Vandenberghe, and Weng Kee Wong. T -optimal designs for multi-factor polynomial regression models via a semidefinite relaxation method. Statistics and Computing, 29(4):725–738, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9834-2.

Giles:2019:DMU

[1465] Michael B. Giles and Takashi Goda.

Decision-making under uncertainty: using MLMC for efficient estimation of EVPPI. Statistics and Computing, 29(4):739–751, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9835-1.

Hartmann:2019:LAN

[1466] Marcelo Hartmann and Jarno Vanhatalo. Laplace approximation and natural gradient for Gaussian process regression with heteroscedastic Studentt model. Statistics and Computing, 29(4):753–773, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9836-0;   http:
//link.springer.com/content/pdf/
10.1007/s11222-018-9836-0.pdf.

Jasra:2019:MPF

[1467] Ajay Jasra, Kody J. H. Law, and Prince Peprah Osei. Multilevel particle filters for Lévy-driven stochastic differential equations. Statistics and Computing, 29(4):775–789, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9837-z;   http://link.springer.
com/content/pdf/10.1007/s11222-

                                        018-9837-z.pdf.  See correction

[1471].

Fop:2019:MBC

[1468] Michael Fop, Thomas Brendan Murphy, and Luca Scrucca. Model-based clustering with sparse covariance matrices. Statistics and Computing, 29 (4):791–819, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-018-9838-y.

Sorbye:2019:AFG

[1469] Sigrunn H. Sørbye, Eirik Myrvoll-Nilsen, and Håvard Rue. An approximate fractional Gaussian noise model with O(n) computational cost. Statistics and Computing, 29(4):821–833, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-018-
                                        9843-1.

Chen:2019:IER

[1470] Xi Chen, Michael Hobson, Saptarshi Das, and Paul Gelderblom. Improving the efficiency and robustness of nested sampling using posterior repartitioning. Statistics and Computing, 29(4):835–850, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-018-9841-3;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-018-9841-3.pdf.

Jasra:2019:CMP

[1471] Ajay Jasra, Kody J. H. Law, and Prince Peprah Osei. Correction to:

Multilevel particle filters for Lévydriven stochastic differential equations. Statistics and Computing, 29(4):851, July 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9839-x;   http://link.springer.
com/content/pdf/10.1007/s11222-
018-9839-x.pdf. See [1467].

Macdonald:2019:MSM

[1472] Benn Macdonald and Dirk Husmeier.

Model selection via marginal likelihood estimation by combining thermodynamic integration and gradient matching. Statistics and Computing, 29(5): 853–867, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9840-4;   http:
//link.springer.com/content/pdf/
10.1007/s11222-018-9840-4.pdf.

Espinosa:2019:CRM

[1473] Javier Espinosa and Christian Hennig.

A constrained regression model for an ordinal response with ordinal predictors. Statistics and Computing, 29(5): 869–890, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9842-2;   http:
//link.springer.com/content/pdf/
10.1007/s11222-018-9842-2.pdf.

Higson:2019:DNS

[1474] Edward Higson, Will Handley, Michael Hobson, and Anthony Lasenby. Dynamic nested sampling: an improved algorithm for parameter estimation and evidence calculation. Statistics and Computing, 29(5):891–913, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-018-
                                        9844-0;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        018-9844-0.pdf.

Kolev:2019:IEM

[1475] Aleksandar A. Kolev and Gordon J. Ross. Inference for ETAS models with non-Poissonian mainshock arrival times. Statistics and Computing, 29 (5):915–931, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-018-9845-z;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-018-9845-z.pdf.

Gunawan:2019:CEB

[1476] D. Gunawan, M.-N. Tran, K. Suzuki, J. Dick, and R. Kohn. Computationally efficient Bayesian estimation of high-dimensional Archimedean copulas with discrete and mixed margins. Statistics and Computing, 29(5): 933–946, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-018-9846-y.

Yildirim:2019:EMD

[1477] Sinan Yildirim and Beyza Ermis. Exact MCMC with differentially private moves. Statistics and Computing, 29 (5):947–963, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-9847-x.

Langsrud:2019:IPR

[1478] Øyvind Langsrud. Information preserving regression-based tools for statistical disclosure control. Statistics and Computing, 29(5):965–976, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
9848-9.

Cui:2019:NBA

[1479] Ruifei Cui, Ioan Gabriel Bucur, Perry Groot, and Tom Heskes. A novel Bayesian approach for latent variable modeling from mixed data with missing values. Statistics and Computing, 29 (5):977–993, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-018-09849-7;   http:
//link.springer.com/content/pdf/
10.1007/s11222-018-09849-7.pdf.

Moradi:2019:RSV

[1480] M. Mehdi Moradi, Ottmar Cronie, Ege Rubak, Raphael Lachieze-Rey, Jorge Mateu, and Adrian Baddeley. Resample-smoothing of Voronoi intensity estimators. Statistics and Computing, 29(5):995–1010, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-018-
09850-0;   http://link.springer.
com/content/pdf/10.1007/s11222-
018-09850-0.pdf.

Maharaj:2019:CIT

[1481] Elizabeth Ann Maharaj, Paulo Teles, and Paula Brito. Clustering of interval time series. Statistics and Computing, 29(5):1011–1034, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-018-
                                        09851-z.

Yamashita:2019:CAH

[1482] Hiroshi Yamashita and Hideyuki Suzuki. Convergence analysis of herded-Gibbs-type sampling algorithms: effects of weight sharing. Statistics and Computing, 29(5):1035– 1053, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09852-6;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-019-09852-6.pdf.

Touron:2019:CML

[1483] Augustin Touron. Consistency of the maximum likelihood estimator in seasonal hidden Markov models. Statistics and Computing, 29(5):1055– 1075, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09854-4.

Heiner:2019:SPS

[1484] Matthew Heiner, Athanasios Kottas, and Stephan Munch. Structured priors for sparse probability vectors with application to model selection in Markov chains. Statistics and Computing, 29(5):1077–1093, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09856-2.

Scutari:2019:LBN

[1485] Marco Scutari, Claudia Vitolo, and Allan Tucker. Learning Bayesian networks from big data with greedy search: computational complexity and efficient implementation. Statistics and Computing, 29(5):1095–1108, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09857-1;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09857-1.pdf.

Plasse:2019:MCD

[1486] Joshua Plasse and Niall M. Adams.

Multiple changepoint detection in categorical data streams. Statistics and Computing, 29(5):1109–1125, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09858-0;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09858-0.pdf.

Lomeli:2019:AMC

[1487] M. Lomelí, M. Rowland, A. Gretton, and Z. Ghahramani. Antithetic and Monte Carlo kernel estimators for partial rankings. Statistics and Computing, 29(5):1127–1147, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09859-z;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        019-09859-z.pdf.

Golightly:2019:ESC

[1488] Andrew Golightly and Chris Sherlock. Efficient sampling of conditioned Markov jump processes. Statistics and Computing, 29(5):1149–1163, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09861-5;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        019-09861-5.pdf.

Li:2019:RAM

[1489] Miaoqi Li and Emily L. Kang. Randomized algorithms of maximum likelihood estimation with spatial autoregressive models for large-scale networks. Statistics and Computing, 29(5): 1165–1179, September 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09862-4.

Girolami:2019:ESE

[1490] M. Girolami, I. C. F. Ipsen, C. J. Oates, A. B. Owen, and T. J. Sullivan. Editorial: special edition on probabilistic numerics. Statistics and Computing, 29(6):1181–1183, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09892-y;   http://link.springer.

com/content/pdf/10.1007/s11222-
019-09892-y.pdf.

Yoo:2019:NTO

[1491] Gene Ryan Yoo and Houman Owhadi.

De-noising by thresholding operator adapted wavelets. Statistics and Computing, 29(6):1185–1201, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09893-x.

Ehler:2019:OMC

[1492] Martin Ehler, Manuel Gräf, and Chris. J. Oates. Optimal Monte Carlo integration on closed manifolds. Statistics and Computing, 29(6): 1203–1214, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09894-w;   http:
//link.springer.com/content/pdf/
10.1007/s11222-019-09894-w.pdf.

Jagadeeswaran:2019:FAB

[1493] R. Jagadeeswaran and Fred J. Hickernell. Fast automatic Bayesian cubature using lattice sampling. Statistics and Computing, 29(6):1215–1229, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09895-9.

Karvonen:2019:SEB

[1494] Toni Karvonen, Simo Särkkä, and Chris. J. Oates. Symmetry exploits for Bayesian cubature methods. Statistics and Computing, 29(6):

1231–1248, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09896-8;   http:
//link.springer.com/content/pdf/
10.1007/s11222-019-09896-8.pdf.

Bartels:2019:PLS

[1495] Simon Bartels, Jon Cockayne, Ilse C. F. Ipsen, and Philipp Hennig. Probabilistic linear solvers: a unifying view. Statistics and Computing, 29(6): 1249–1263, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09897-7;   http:
//link.springer.com/content/pdf/
10.1007/s11222-019-09897-7.pdf.

Lie:2019:SCR

[1496] Han Cheng Lie, A. M. Stuart, and T. J. Sullivan. Strong convergence rates of probabilistic integrators for ordinary differential equations. Statistics and Computing, 29(6):1265–1283, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09898-6.

Chkrebtii:2019:ASS

[1497] Oksana A. Chkrebtii and David A. Campbell. Adaptive step-size selection for state-space probabilistic differential equation solvers. Statistics and Computing, 29(6):1285–1295, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09899-5.

Tronarp:2019:PSO

[1498] Filip Tronarp, Hans Kersting, Simo Särkkä, and Philipp Hennig. Probabilistic solutions to ordinary differential equations as nonlinear Bayesian filtering: a new perspective. Statistics and Computing, 29(6):1297–1315, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09900-1;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09900-1.pdf.

Karvonen:2019:PMB

[1499] Toni Karvonen, Motonobu Kanagawa, and Simo Särkkä. On the positivity and magnitudes of Bayesian quadrature weights. Statistics and Computing, 29(6):1317–1333, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09901-0;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09901-0.pdf.

Oates:2019:MRP

[1500] C. J. Oates and T. J. Sullivan. A modern retrospective on probabilistic numerics. Statistics and Computing, 29(6):1335–1351, November 2019. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09902-z;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09902-z.pdf.

Kobayashi:2020:EGG

[1501] Kei Kobayashi and Henry P. Wynn. Empirical geodesic graphs and CAT(k) metrics for data analysis. Statistics and Computing, 30(1):1–18, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09855-3.

Li:2020:FMM

[1502] Zheyuan Li and Simon N. Wood. Faster model matrix crossproducts for large generalized linear models with discretized covariates. Statistics and Computing, 30(1):19–25, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09864-2;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        019-09864-2.pdf.

Tawn:2020:WPS

[1503] Nicholas G. Tawn, Gareth O. Roberts, and Jeffrey S. Rosenthal. Weightpreserving simulated tempering. Statistics and Computing, 30(1):27–41, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

                                        springer.com/article/10.1007/s11222-
                                        019-09863-3;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        019-09863-3.pdf.

Kosmidis:2020:MMB

[1504] Ioannis Kosmidis, Euloge Clovis Kenne Pagui, and Nicola Sartori. Mean and median bias reduction in generalized linear models. Statistics and Computing, 30(1):43–59, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09860-6;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09860-6.pdf.

Barbillon:2020:ENI

[1505] Pierre Barbillon, Loïc Schwaller, Stéphane Robin, Andrew Flachs, and Glenn Davis Stone. Epidemiologic network inference. Statistics and Computing, 30(1):61–75, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09865-1.

Battaglia:2020:PPA

[1506] Francesco Battaglia, Domenico Cucina, and Manuel Rizzo. Parsimonious periodic autoregressive models for time series with evolving trend and seasonality. Statistics and Computing, 30 (1):77–91, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09866-0.

Duarte:2020:OED

[1507] Belmiro P. M. Duarte, José F. O.

Granjo, and Weng Kee Wong. Optimal exact designs of experiments via Mixed Integer Nonlinear Programming. Statistics and Computing, 30 (1):93–112, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09867-z.

Boudt:2020:MRC

[1508] Kris Boudt, Peter J. Rousseeuw, Steven Vanduffel, and Tim Verdonck. The minimum regularized covariance determinant estimator. Statistics and Computing, 30(1):113–128, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09869-x.

Cheng:2020:NME

[1509] Yafeng Cheng, Jian Qing Shi, and Janet Eyre. Nonlinear mixed-effects scalar-on-function models and variable selection. Statistics and Computing, 30 (1):129–140, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09871-3;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-019-09871-3.pdf.

Hothorn:2020:TBM

[1510] Torsten Hothorn. Transformation boosting machines. Statistics and Computing, 30(1):141–152, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09870-4;   http://link.springer.
                                        com/content/pdf/10.1007/s11222-
                                        019-09870-4.pdf.

Paine:2020:SRM

[1511] P. J. Paine, S. P. Preston, M. Tsagris, and Andrew T. A. Wood. Spherical regression models with general covariates and anisotropic errors. Statistics and Computing, 30(1):153–165, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09872-2;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09872-2.pdf.

Craigmile:2020:ELR

[1512] Peter F. Craigmile and Debashis Mondal. Estimation of long-range dependence in gappy Gaussian time series. Statistics and Computing, 30 (1):167–185, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09874-0.

Cuevas:2020:FES

[1513] Francisco Cuevas, Denis Allard, and Emilio Porcu. Fast and exact simulation of Gaussian random fields defined on the sphere cross time. Statistics and Computing, 30(1):187–194, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09873-1.

Bock:2020:BDM

[1514] Olivier Bock, Xavier Collilieux, François Guillamon, Emilie Lebarbier, and Claire Pascal. A breakpoint detection in the mean model with heterogeneous variance on fixed time intervals. Statistics and Computing, 30 (1):195–207, February 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09853-5.

Price-Williams:2020:NSE

[1515] Matthew Price-Williams and Nicholas A. Heard. Nonparametric self-exciting models for computer network traffic. Statistics and Computing, 30 (2):209–220, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09875-z;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-019-09875-z.pdf.

Leonelli:2020:SBM

[1516] Manuele Leonelli and Dani Gamerman. Semiparametric bivariate modelling with flexible extremal dependence. Statistics and Computing, 30 (2):221–236, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09878-w;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-019-09878-w.pdf.

Glaws:2020:IRR

[1517] Andrew Glaws, Paul G. Constantine, and R. Dennis Cook. Inverse regression for ridge recovery: a datadriven approach for parameter reduction in computer experiments. Statistics and Computing, 30(2):237–253, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09876-y.

Greco:2020:WLM

[1518] Luca Greco and Claudio Agostinelli. Weighted likelihood mixture modeling and model-based clustering. Statistics and Computing, 30(2):255–277, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09881-1.

Rugamer:2020:ILB

[1519] David Rügamer and Sonja Greven.

Inference for L_{2}-boosting. Statistics and Computing, 30(2):279–289, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09882-0.

Ren:2020:MMS

[1520] Sheng Ren, Emily L. Kang, and Jason L. Lu. MCEN: a method of simultaneous variable selection and clustering for high-dimensional multinomial regression. Statistics and Computing, 30(2):291–304, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09880-2.

Akyildiz:2020:NPF

[1521] Ömer Deniz Akyildiz and Joaquín Míguez. Nudging the particle filter. Statistics and Computing, 30 (2):305–330, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09884-y;   http:
//link.springer.com/content/pdf/
10.1007/s11222-019-09884-y.pdf.

Kim:2020:MIS

[1522] Gang-Hoo Kim and Sung-Ho Kim. Marginal information for structure learning. Statistics and Computing, 30(2):331–349, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09877-x.

Andrade:2020:RBM

[1523] Daniel Andrade, Akiko Takeda, and Kenji Fukumizu. Robust Bayesian model selection for variable clustering with the Gaussian graphical model. Statistics and Computing, 30 (2):351–376, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09879-9;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11222-019-09879-9.pdf.

Radivojevic:2020:MHM

[1524] Tijana Radivojević and Elena Akhmatskaya. Modified Hamiltonian Monte Carlo for Bayesian inference. Statistics and Computing, 30(2):377–404, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09885-x.

Javeed:2020:TOD

[1525] Aurya Javeed and Giles Hooker. Timing observations of diffusions. Statistics and Computing, 30(2):405–417, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s11222-019-
09883-z.

Solin:2020:HSM

[1526] Arno Solin and Simo Särkkä. Hilbert space methods for reduced-rank Gaussian process regression. Statistics and Computing, 30(2):419–446, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09886-w;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09886-w.pdf.

Antoniano-Villalobos:2020:NEP

[1527] Isadora Antoniano-Villalobos, Emanuele Borgonovo, and Xuefei Lu. Nonparametric estimation of probabilistic sensitivity measures. Statistics and Computing, 30(2):447–467, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09887-9.

McSwiggan:2020:ERR

[1528] Greg McSwiggan, Adrian Baddeley, and Gopalan Nair. Estimation of relative risk for events on a linear network. Statistics and Computing, 30 (2):469–484, March 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09889-7.

Papastamoulis:2020:CMD

[1529] Panagiotis Papastamoulis. Clustering multivariate data using factor analytic Bayesian mixtures with an unknown number of components. Statistics and Computing, 30(3):485–506, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09891-z.

Giles:2020:MLM

[1530] Michael B. Giles, Mateusz B. Majka, Lukasz Szpruch, Sebastian J. Vollmer, and Konstantinos C. Zygalakis. Multilevel Monte Carlo methods for the approximation of invariant measures of stochastic differential equations. Statistics and Computing, 30(3):507–524, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09890-0;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09890-0.pdf.

Amri:2020:DDS

[1531] Mohamed Reda El Amri, Céline Helbert, Olivier Lepreux, Miguel Muñoz Zuniga, Clémentine Prieur, and Delphine Sinoquet. Data-driven stochastic inversion via functional quantization. Statistics and Computing, 30(3):525–541, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09888-8.

An:2020:RBS

[1532] Ziwen An, David J. Nott, and Christopher Drovandi. Robust Bayesian synthetic likelihood via a semi-parametric approach. Statistics and Computing, 30(3):543–557, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09904-x.

Siren:2020:LDR

[1533] Jukka Sirén and Samuel Kaski. Local dimension reduction of summary statistics for likelihood-free inference. Statistics and Computing, 30(3):559–570, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09905-w;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09905-w.pdf.

Hahn:2020:OAM

[1534] Georg Hahn. Optimal allocation of Monte Carlo simulations to multiple hypothesis tests. Statistics and Computing, 30(3):571–586, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09906-9.

Chua:2020:SMR

[1535] Alvin J. K. Chua. Sampling from manifold-restricted distributions using tangent bundle projections. Statistics and Computing, 30(3):587–602, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09907-8.

Dolgov:2020:ASM

[1536] Sergey Dolgov, Karim Anaya-Izquierdo, Colin Fox, and Robert Scheichl. Approximation and sampling of multivariate probability distributions in the tensor train decomposition. Statistics and Computing, 30(3):603–625, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09910-z;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09910-z.pdf.

Buckwar:2020:SDB

[1537] Evelyn Buckwar, Massimiliano Tamborrino, and Irene Tubikanec. Spectral density-based and measure-preserving ABC for partially observed diffusion processes. an illustration on Hamiltonian SDEs. Statistics and Computing, 30(3):627–648, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09909-6;   http:
//link.springer.com/content/pdf/
10.1007/s11222-019-09909-6.pdf.

Choiruddin:2020:REH

[1538] Achmad Choiruddin, Francisco Cuevas-Pacheco, Jean-François Coeurjolly, and Rasmus Waagepetersen. Regularized estimation for highly multivariate log Gaussian Cox processes. Statistics and Computing, 30(3):649–662, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09911-y.

Everitt:2020:SMC

[1539] Richard G. Everitt, Richard Culliford, Felipe Medina-Aguayo, and Daniel J. Wilson. Sequential Monte Carlo with transformations. Statistics and Computing, 30(3):663–676, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09903-y;   http://link.springer.
com/content/pdf/10.1007/s11222-
019-09903-y.pdf.

Christou:2020:CQS

[1540] Eliana Christou. Central quantile subspace. Statistics and Computing, 30(3):677–695, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09915-8.

Wang:2020:HDR

[1541] Fan Wang, Sach Mukherjee, Sylvia Richardson, and Steven M. Hill. Highdimensional regression in practice: an empirical study of finite-sample prediction, variable selection and ranking. Statistics and Computing, 30 (3):697–719, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-019-09914-9;   http:
//link.springer.com/content/pdf/
10.1007/s11222-019-09914-9.pdf.

Wu:2020:CSN

[1542] Changye Wu and Christian P. Robert.

Coordinate sampler: a non-reversible Gibbs-like MCMC sampler. Statistics and Computing, 30(3):721–730, May 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09913-w.

Nguyen:2020:MBL

[1543] Hien D. Nguyen, Florence Forbes, and Geoffrey J. McLachlan. Minibatch learning of exponential family finite mixture models. Statistics and Computing, 30(4):731–748, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09919-4.

Ongaro:2020:NMM

[1544] Andrea Ongaro, Sonia Migliorati, and Roberto Ascari. A new mixture model on the simplex. Statistics and Computing, 30(4):749–770, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09920-x.

Sang:2020:FSI

[1545] Peijun Sang and Jiguo Cao. Functional single-index quantile regression models. Statistics and Computing, 30(4):771–781, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11222-019-09917-6.

Mendes:2020:FPM

[1546] Eduardo F. Mendes, Christopher K. Carter, David Gunawan, and Robert Kohn. A flexible particle Markov chain Monte Carlo method. Statistics and Computing, 30(4):783–798, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11222-019-
                                        09916-7.

Wong:2020:PBS

[1547] Jackie S. T. Wong, Jonathan J. Forster, and Peter W. F. Smith. Properties of the bridge sampler with a focus on splitting the MCMC sample. Statistics and Computing, 30(4):799–816, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-019-
09918-5.

Dresvyanskiy:2020:DAF

[1548] Denis Dresvyanskiy, Tatiana Karaseva, Vitalii Makogin, Sergei Mitrofanov, Claudia Redenbach, and Evgeny Spodarev. Detecting anomalies in fibre systems using 3-dimensional image data. Statistics and Computing, 30(4):817–837, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-020-09921-1;   http:
//link.springer.com/content/pdf/
10.1007/s11222-020-09921-1.pdf.

Ray:2020:EBS

[1549] Pallavi Ray, Debdeep Pati, and Anirban Bhattacharya. Efficient Bayesian shape-restricted function estimation with constrained Gaussian process priors. Statistics and Computing, 30 (4):839–853, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-020-09922-0.

Santos:2020:NSA

[1550] Bruno Santos and Thomas Kneib. Noncrossing structured additive multipleoutput Bayesian quantile regression models. Statistics and Computing, 30(4):855–869, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-020-09925-x.

Takai:2020:IDF

[1551] Keiji Takai. Incomplete-data Fisher scoring method with steplength adjustment. Statistics and Computing, 30(4):871–886, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-020-09923-z;   http:
//link.springer.com/content/pdf/
10.1007/s11222-020-09923-z.pdf.

Ye:2020:MCC

[1552] Lifeng Ye, Alexandros Beskos, Maria De Iorio, and Jie Hao. Monte Carlo co-ordinate ascent variational inference. Statistics and Computing, 30 (4):887–905, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-020-09924-y;   http:
//link.springer.com/content/pdf/
10.1007/s11222-020-09924-y.pdf.

Abdulle:2020:RTS

[1553] Assyr Abdulle and Giacomo Garegnani. Random time step probabilistic methods for uncertainty quantification in chaotic and geometric numerical integration. Statistics and Computing, 30(4):907–932, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-020-09926-w.

Yu:2020:VDA

[1554] Weichang Yu, John T. Ormerod, and Michael Stewart. Variational discriminant analysis with variable selection. Statistics and Computing, 30(4):933–951, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-020-09928-8.

Muller:2020:MBP

[1555] Raoul Müller, Dominic Schuhmacher, and Jorge Mateu. Metrics and barycenters for point pattern data. Statistics and Computing, 30(4):953–972, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-020-
09932-y;   http://link.springer.
com/content/pdf/10.1007/s11222-
020-09932-y.pdf.

Belomestny:2020:VRM

[1556] D. Belomestny, L. Iosipoi, E. Moulines, A. Naumov, and S. Samsonov. Variance reduction for Markov chains with application to MCMC. Statistics and Computing, 30(4):973–997, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-020-
09931-z.

Mauff:2020:JMM

[1557] Katya Mauff, Ewout Steyerberg, Isabella Kardys, Eric Boersma, and Dimitris Rizopoulos. Joint models with multiple longitudinal outcomes and a time-to-event outcome: a corrected two-stage approach. Statistics and Computing, 30(4):999–1014, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-020-
09927-9;   http://link.springer.
com/content/pdf/10.1007/s11222-
020-09927-9.pdf.

Lu:2020:BNP

[1558] Hongliang Lü, Julyan Arbel, and Florence Forbes. Bayesian nonparametric priors for hidden Markov random fields. Statistics and Computing, 30 (4):1015–1035, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-020-09935-9.

Liu:2020:AMW

[1559] Qiang Liu and Xin T. Tong. Accelerating Metropolis-within-Gibbs sampler with localized computations of differential equations. Statistics and Computing, 30(4):1037–1056, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-020-
09934-w.

Rodrigues:2020:LFA

[1560] G. S. Rodrigues, David J. Nott, and S. A. Sisson. Likelihood-free approximate Gibbs sampling. Statistics and Computing, 30(4):1057–1073, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-020-
09933-x.

Zhang:2020:AIH

[1561] Aijun Zhang, Hengtao Zhang, and Guosheng Yin. Adaptive iterative Hessian sketch via A-optimal subsampling. Statistics and Computing, 30 (4):1075–1090, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:

//link.springer.com/article/10.
1007/s11222-020-09936-8.

Torrecilla:2020:OCG

[1562] José L. Torrecilla, Carlos Ramos-Carreño, Manuel Sánchez-Montañés, and Alberto Suárez. Optimal classification of Gaussian processes in homoand heteroscedastic settings. Statistics and Computing, 30(4):1091–1111, July 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11222-020-
09937-7.

Wickramasuriya:2020:ONN

[1563] Shanika L. Wickramasuriya, Berwin A.

Turlach, and Rob J. Hyndman. Optimal non-negative forecast reconciliation. Statistics and Computing, 30(5): 1167–1182, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09930-0.

Senarathne:2020:LBA

[1564] S. G. J. Senarathne, C. C. Drovandi, and J. M. McGree. A Laplace-based algorithm for Bayesian adaptive design. Statistics and Computing, 30(5): 1183–1208, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09938-6.

Wang:2020:ETV

[1565] Haixu Wang and Jiguo Cao. Estimating time-varying directed neural networks. Statistics and Computing, 30(5): 1209–1220, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09941-x.

Cronie:2020:IHO

[1566] Ottmar Cronie, Mehdi Moradi, and Jorge Mateu. Inhomogeneous higherorder summary statistics for point processes on linear networks. Statistics and Computing, 30(5):1221– 1239, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09942-w.

Passino:2020:CPA

[1567] Francesco Sanna Passino and Nicholas A. Heard. Classification of periodic arrivals in event time data for filtering computer network traffic. Statistics and Computing, 30(5):1241– 1254, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09943-9.

Tan:2020:CSV

[1568] Linda S. L. Tan, Aishwarya Bhaskaran, and David J. Nott. Conditionally structured variational Gaussian approximation with importance weights.

Statistics and Computing, 30(5):1255– 1272, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09944-8.

Augugliaro:2020:CCG

[1569] Luigi Augugliaro, Gianluca Sottile, and Veronica Vinciotti. The conditional censored graphical lasso estimator. Statistics and Computing, 30(5): 1273–1289, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09945-7.

Passino:2020:BEL

[1570] Francesco Sanna Passino and Nicholas A.

Heard. Bayesian estimation of the latent dimension and communities in stochastic blockmodels. Statistics and Computing, 30(5):1291– 1307, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09946-6.

Igarashi:2020:CTB

[1571] Mirai Igarashi and Nobuhiko Terui.

Characterization of topic-based online communities by combining network data and user generated content. Statistics and Computing, 30(5): 1309–1324, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09947-5.

Park:2020:MCM

[1572] Joonha Park and Yves Atchadé. Markov chain Monte Carlo algorithms with sequential proposals. Statistics and Computing, 30(5):1325– 1345, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09948-4.

Ayed:2020:ITA

[1573] Fadhel Ayed, Marco Battiston, and Federico Camerlenghi. An information theoretic approach to post randomization methods under differential privacy. Statistics and Computing, 30(5): 1347–1361, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09949-3.

Garcia-Escudero:2020:MBC

[1574] Luis Angel García-Escudero, Agustín Mayo-Iscar, and Marco Riani. Modelbased clustering with determinant-and-shape constraint. Statistics and Computing, 30(5):1363–1380, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-020-09950-
                                        w.

Jasra:2020:MPF

[1575] Ajay Jasra, Fangyuan Yu, and Jeremy Heng. Multilevel particle filters for the non-linear filtering problem in continuous time. Statistics and Computing, 30(5):1381– 1402, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09951-9.

Alegria:2020:TAC

[1576] Alfredo Alegría, Xavier Emery, and Christian Lantuéjoul. The turning arcs: a computationally efficient algorithm to simulate isotropic vectorvalued Gaussian random fields on the d-sphere. Statistics and Computing, 30(5):1403–1418, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-020-09952-
8.

vanderWurp:2020:GJR

[1577] Hendrik van der Wurp, Andreas Groll, and Rosalba Radice. Generalised joint regression for count data: a penalty extension for competitive settings. Statistics and Computing, 30(5): 1419–1432, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09953-7.

Soize:2020:SBP

[1578] Christian Soize, Roger G. Ghanem, and Christophe Desceliers. Sampling of Bayesian posteriors with a non-Gaussian probabilistic learning on manifolds from a small dataset. Statistics and Computing, 30(5):1433– 1457, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09954-6.

Whitaker:2020:CLM

[1579] T. Whitaker, B. Beranger, and S. A. Sisson. Composite likelihood methods for histogram-valued random variables. Statistics and Computing, 30(5): 1459–1477, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09955-5.

Allard:2020:SST

[1580] Denis Allard, Xavier Emery, and Christian Lantuéjoul. Simulating spacetime random fields with nonseparable Gneiting-type covariance functions. Statistics and Computing, 30(5): 1479–1495, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09956-4.

Park:2020:IHD

[1581] Joonha Park and Edward L. Ionides. Inference on high-dimensional implicit dynamic models using a guided intermediate resampling filter. Statistics and Computing, 30(5): 1497–1522, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09957-3.

Akhanli:2020:CCN

[1582] Serhat Emre Akhanli and Christian Hennig. Comparing clusterings and numbers of clusters by aggregation of calibrated clustering validity indexes. Statistics and Computing, 30(5): 1523–1544, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09958-2.

Cappozzo:2020:AND

[1583] Andrea Cappozzo, Francesca Greselin, and Thomas Brendan Murphy. Anomaly and novelty detection for robust semi-supervised learning. Statistics and Computing, 30(5): 1545–1571, September 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09959-1.

Andersson:2020:ICP

[1584] Claes Andersson and Tomás Mrkvicka.

Inference for cluster point processes with over- or under-dispersed cluster sizes. Statistics and Computing, 30(6): 1573–1590, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09960-8.

Lijoi:2020:SHD

[1585] Antonio Lijoi, Igor Prünster, and Tommaso Rigon. Sampling hierarchies of discrete random structures. Statistics and Computing, 30(6):1591– 1607, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09961-7.

Zammit-Mangion:2020:MSP

[1586] Andrew Zammit-Mangion and Jonathan Rougier. Multi-scale process modelling and distributed computation for spatial data. Statistics and Computing, 30(6):

1609–1627, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09962-6.

Sportisse:2020:ILR

[1587] Aude Sportisse, Claire Boyer, and Julie Josse. Imputation and low-rank estimation with Missing Not At Random data. Statistics and Computing, 30(6): 1629–1643, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09963-5.

Akyildiz:2020:PSM

[1588] Ömer Deniz Akyildiz, Dan Crisan, and Joaquín Míguez. Parallel sequential Monte Carlo for stochastic gradient-free nonconvex optimization. Statistics and Computing, 30(6): 1645–1663, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09964-4.

Hammouda:2020:ISR

[1589] Chiheb Ben Hammouda, Nadhir Ben Rached, and Raúl Tempone. Importance sampling for a robust and efficient multilevel Monte Carlo estimator for stochastic reaction networks. Statistics and Computing, 30(6): 1665–1689, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09965-3.

Hahn:2020:BFC

[1590] Georg Hahn, Paul Fearnhead, and Idris A. Eckley. BayesProject: Fast computation of a projection direction for multivariate changepoint detection. Statistics and Computing, 30(6): 1691–1705, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09966-2.

Roy:2020:SCT

[1591] Angshuman Roy, Soham Sarkar, and Alok Goswami. On some consistent tests of mutual independence among several random vectors of arbitrary dimensions. Statistics and Computing, 30(6):1707– 1723, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09967-1.

Kuhn:2020:PSA

[1592] Estelle Kuhn, Catherine Matias, and Tabea Rebafka. Properties of the stochastic approximation EM algorithm with mini-batch sampling. Statistics and Computing, 30(6):1725– 1739, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09968-0.

Gunawan:2020:SSM

[1593] David Gunawan, Khue-Dung Dang, and Minh-Ngoc Tran. Subsampling sequential Monte Carlo for static Bayesian models. Statistics and Computing, 30(6):1741– 1758, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09969-z.

Lowther:2020:SAS

[1594] Aaron P. Lowther, Paul Fearnhead, and Kjeld Jensen. Semiautomated simultaneous predictor selection for regression-SARIMA models. Statistics and Computing, 30(6): 1759–1778, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09970-6.

Ahfock:2020:APC

[1595] Daniel Ahfock and Geoffrey J. McLachlan. An apparent paradox: a classifier based on a partially classified sample may have smaller expected error rate than that if the sample were completely classified. Statistics and Computing, 30(6):1779– 1790, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09971-5.

Kersting:2020:CRG

[1596] Hans Kersting, T. J. Sullivan, and Philipp Hennig. Convergence rates of Gaussian ODE filters. Statistics and Computing, 30(6):1791– 1816, November 2020. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09972-4.

Dick:2021:TMC

[1597] Josef Dick, Takashi Goda, and Hiroya Murata. Toeplitz Monte Carlo. Statistics and Computing, 31(1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09987-x.

Cao:2021:ELR

[1598] Jian Cao, Marc G. Genton, and George M. Turkiyyah. Exploiting lowrank covariance structures for computing high-dimensional normal and Student-t probabilities. Statistics and Computing, 31(1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-020-09978-
y.

Juan-Albarracin:2021:NLS

[1599] Javier Juan-Albarracín, Elies Fuster-Garcia, and Juan M. García-Gómez. Non-local spatially varying finite mixture models for image segmentation. Statistics and Computing, 31(1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09988-w.

Holbrook:2021:SBI

[1600] Andrew J. Holbrook, Charles E. Loeffler, and Marc A. Suchard. Scalable Bayesian inference for self-excitatory stochastic processes applied to big American gunfire data. Statistics and Computing, 31(1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-020-09980-
4.

Hu:2021:MCU

[1601] Shengwei Hu and Yong Wang. Modal clustering using semiparametric mixtures and mode flattening. Statistics and Computing, 31 (1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09985-z.

Wan:2021:AMM

[1602] Kitty Yuen Yi Wan and Jim E. Griffin. An adaptive MCMC method for Bayesian variable selection in logistic and accelerated failure time regression models. Statistics and Computing, 31(1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09974-2.

Gautier:2021:FSE

[1603] Guillaume Gautier, Rémi Bardenet, and Michal Valko. Fast sampling from β-ensembles. Statistics and Computing, 31(1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09984-0.

Aizenbud:2021:ROM

[1604] Yariv Aizenbud, Boris Landa, and Yoel Shkolnisky. Rank-one multi-reference factor analysis. Statistics and Computing, 31(1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09990-2.

Ding:2021:EKI

[1605] Zhiyan Ding and Qin Li. Ensemble Kalman inversion: meanfield limit and convergence analysis. Statistics and Computing, 31 (1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09976-0.

Cho:2021:PEM

[1606] Wendy K. Tam Cho and Yan Y.

Liu. A parallel evolutionary multipletry Metropolis Markov chain Monte Carlo algorithm for sampling spatial partitions. Statistics and Computing, 31(1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09977-z.

Aeberhard:2021:RFG

[1607] William H. Aeberhard, Eva Cantoni, and Rosalba Radice. Robust fitting for generalized additive models for location, scale and shape. Statistics and Computing, 31(1):??, January 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09979-x.

Akyildiz:2021:CRO

[1608] Ömer Deniz Akyildiz and Joaquín Míguez. Convergence rates for optimised adaptive importance samplers. Statistics and Computing, 31(2):??, March 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09983-1.

Stefanucci:2021:MSB

[1609] Marco Stefanucci and Antonio Canale. Multiscale stick-breaking mixture models. Statistics and Computing, 31(2):??, March 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09991-1.

Meyer:2021:BWP

[1610] Mark J. Meyer, Elizabeth J. Malloy, and Brent A. Coull. Bayesian wavelet-packet historical functional linear models. Statistics and Computing, 31(2):??, March 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09981-3.

Blaser:2021:RAA

[1611] Rico Blaser and Piotr Fryzlewicz. Regularizing axis-aligned ensembles via data rotations that favor simpler learners. Statistics and Computing, 31(2):??, March 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09973-3.

Paananen:2021:IAI

[1612] Topi Paananen, Juho Piironen, and Aki Vehtari. Implicitly adaptive importance sampling. Statistics and Computing, 31(2):??, March 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09982-2.

Tutz:2021:TSS

[1613] Gerhard Tutz and Moritz Berger. Treestructured scale effects in binary and ordinal regression. Statistics and Computing, 31(2):??, March 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09992-0.

Wilson:2021:WBA

[1614] Rebecca E. Wilson, Idris A. Eckley, and Timothy Park. A waveletbased approach for imputation in nonstationary multivariate time series. Statistics and Computing, 31(2):??, March 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-09998-2.

Barratt:2021:ORS

[1615] Shane Barratt, Guillermo Angeris, and Stephen Boyd. Optimal representative sample weighting. Statistics and Computing, 31(2):??, March 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10001-1.

Prado:2021:BAR

[1616] Estevão B. Prado, Rafael A. Moral, and Andrew C. Parnell. Bayesian additive regression trees with model trees. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-09997-3.

Jasra:2021:UEG

[1617] Ajay Jasra, Kody J. H. Law, and Deng Lu. Unbiased estimation of the gradient of the log-likelihood in inverse problems. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-09994-6.

Viroli:2021:DMU

[1618] Cinzia Viroli and Laura Anderlucci. Deep mixtures of unigrams for uncovering topics in textual data. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-020-09989-9.

Tronarp:2021:BOS

[1619] Filip Tronarp, Simo Särkkä, and Philipp Hennig. Bayesian ODE solvers: the maximum a posteriori estimate. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-09993-7.

Karppinen:2021:CPF

[1620] Santeri Karppinen and Matti Vihola. Conditional particle filters with diffuse initial distributions. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09975-1.

Guinness:2021:GPL

[1621] Joseph Guinness. Gaussian process learning via Fisher scoring of Vecchia’s approximation. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-09999-1.

Stindl:2021:AER

[1622] Tom Stindl and Feng Chen. Accelerating the estimation of renewal Hawkes self-exciting point processes. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10002-0.

Agrell:2021:SBO

[1623] Christian Agrell and Kristina Rognlien Dahl. Sequential Bayesian optimal experimental design for structural reliability analysis. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10000-2.

Coullon:2021:ESI

[1624] Jeremie Coullon and Robert J.

Webber. Ensemble sampler for infinite-dimensional inverse problems. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10004-y.

DeBortoli:2021:ESO

[1625] Valentin De Bortoli, Alain Durmus, and Ana F. Vidal. Efficient stochastic optimisation by unadjusted Langevin Monte Carlo. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-020-09986-y.

Ridley:2021:SMR

[1626] Gavin Ridley and Benoit Forget. A simple method for rejection sampling efficiency improvement on SIMT architectures. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10003-z.

Hattab:2021:MMA

[1627] Mohammad W. Hattab and David Ruppert. A mixed model approach to measurement error in semiparametric regression. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10005-x.

Affeldt:2021:RBD

[1628] Séverine Affeldt, Lazhar Labiod, and Mohamed Nadif. Regularized bidirectional co-clustering. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-021-10006-
w.

Cole:2021:LIG

[1629] D. Austin Cole, Ryan B. Christianson, and Robert B. Gramacy. Locally induced Gaussian processes for large-scale simulation experiments. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10007-9.

Wang:2021:PBE

[1630] Yiwei Wang, Jiuhai Chen, and Lulu Kang. Particle-based energetic variational inference. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10009-7.

Hahn:2021:FES

[1631] Georg Hahn, Sharon M. Lutz, and Christoph Lange. A fast and efficient smoothing approach to lasso regression and an application in statistical genetics: polygenic risk scores for chronic obstructive pulmonary disease (COPD). Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10010-0.

Muller:2021:ETV

[1632] Gernot Müller and Sebastian Uhl. Estimation of time-varying autoregressive stochastic volatility models with stable innovations. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-09995-5.

Bierkens:2021:PDM

[1633] Joris Bierkens, Sebastiano Grazzi, and Moritz Schauer. A piecewise deterministic Monte Carlo method for diffusion bridges. Statistics and Computing, 31(3):??, May 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10008-8.

Fischer:2021:REA

[1634] Samuel M. Fischer and Mark A. Lewis. A robust and efficient algorithm to find profile likelihood confidence intervals. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10012-y.

Latz:2021:ASG

[1635] Jonas Latz. Analysis of stochastic gradient descent in continuous time. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10016-8. See cor-

rection [2070].

Chen:2021:MLE

[1636] Yici Chen and Ken’ichiro Tanaka.

Maximum likelihood estimation of the Fisher–Bingham distribution via efficient calculation of its normalizing constant. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10015-9.

Trip:2021:PAR

[1637] Diederik S. Laman Trip and Wessel N.

van Wieringen. A parallel algorithm for ridge-penalized estimation of the multivariate exponential family from data of mixed types. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10013-x.

Denti:2021:TSB

[1638] Francesco Denti, Andrea Cappozzo, and Francesca Greselin. A two-stage Bayesian semiparametric model for novelty detection with robust prior information. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10017-7.

Lyu:2021:EGP

[1639] Xiong Lyu, Mickaël Binois, and Michael Ludkovski. Evaluating Gaussian process metamodels and sequential designs for noisy level set estimation. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10014-w.

Jouvin:2021:BFE

[1640] Nicolas Jouvin, Charles Bouveyron, and Pierre Latouche. A Bayesian Fisher–EM algorithm for discriminative Gaussian subspace clustering. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10018-6.

Duarte:2021:OED

[1641] Belmiro P. M. Duarte, Anthony C. Atkinson, and Nuno M. C. Oliveira. Optimal experimental design for linear time invariant state-space models. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10020-y.

Zhou:2021:EEV

[1642] Feng Zhou, Simon Luo, Zhidong Li, Xuhui Fan, Yang Wang, Arcot Sowmya, and Fang Chen. Efficient EMvariational inference for nonparametric Hawkes process. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10021-x.

Fasano:2021:CFF

[1643] Augusto Fasano, Giovanni Rebaudo, Daniele Durante, and Sonia Petrone.

A closed-form filter for binary time series. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10022-w.

Fort:2021:FIE

[1644] G. Fort, P. Gach, and E. Moulines. Fast incremental expectation maximization for finite-sum optimization: nonasymptotic convergence. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10023-9.

Kejzlar:2021:FCC

[1645] Vojtech Kejzlar, Mookyong Son, Shrijita Bhattacharya, and Tapabrata Maiti. A fast and calibrated computer model emulator: an empirical Bayes approach. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10024-8.

Leluc:2021:CVS

[1646] Rémi Leluc, François Portier, and Johan Segers. Control variate selection for Monte Carlo integration. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10011-z.

Lee:2021:IIA

[1647] Duncan Lee, Kitty Meeks, and William Pettersson. Improved inference for areal unit count data using graph-based optimisation. Statistics and Computing, 31(4):??, July 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10025-7.

Wang:2021:CGG

[1648] Linna Wang, Yichen Qin, and Yang Li. Confidence graphs for graphical model selection. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10027-5.

Maullin-Sapey:2021:FSC

[1649] Thomas Maullin-Sapey and Thomas E. Nichols. Fisher scoring for crossed factor linear mixed models. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-021-10026-
                                        6.

Singh:2021:MMC

[1650] Satya Prakash Singh, Siuli Mukhopadhyay, and Harsh Raj. Min-max crossover designs for two treatments binary and Poisson crossover trials. Statistics and Computing, 31 (5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10029-3.

Wang:2021:BNM

[1651] Junyang Wang, Jon Cockayne, T. J.

Sullivan, and Chris. J. Oates. Bayesian numerical methods for nonlinear partial differential equations. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-021-10030-
w.

Fanaskov:2021:UCP

[1652] Vladimir Fanaskov. Uncertainty calibration for probabilistic projection methods. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10031-9.

Luo:2021:BIC

[1653] Yu Luo and David A. Stephens.

Bayesian inference for continuoustime hidden Markov models with an unknown number of states. Statistics and Computing, 31(5): ??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10032-8.

Ahfock:2021:DFU

[1654] Daniel Ahfock, Saumyadipta Pyne, and Geoffrey J. McLachlan. Data fusion using factor analysis and low-rank matrix completion. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10033-7.

Houssineau:2021:UMC

[1655] Jeremie Houssineau, Jiajie Zeng, and Ajay Jasra. Uncertainty modelling and computational aspects of data association. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10039-1.

Palade:2021:FGG

[1656] D. I. Palade and M. Vlad. Fast generation of Gaussian random fields for direct numerical simulations of stochastic transport. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10035-5.

Karamanis:2021:ESS

[1657] Minas Karamanis and Florian Beutler. Ensemble slice sampling. Statistics and Computing, 31(5): ??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10038-2.

Bon:2021:ASM

[1658] Joshua J. Bon, Anthony Lee, and Christopher Drovandi. Accelerating sequential Monte Carlo with surrogate likelihoods. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10036-4.

Ayed:2021:NBN

[1659] Fadhel Ayed and François Caron. Nonnegative Bayesian nonparametric factor models with completely random measures. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10037-3.

Caballero:2021:QUD

[1660] Renzo Caballero, Ahmed Kebaier, Marco Scavino, and Raúl Tempone. Quantifying uncertainty with a derivative tracking SDE model and application to wind power forecast data. Statistics and Computing, 31 (5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10040-8.

Ahipasaoglu:2021:BBA

[1661] Selin Damla Ahipasaoglu. A branchand-bound algorithm for the exact optimal experimental design problem. Statistics and Computing, 31 (5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10043-5.

Schreurs:2021:ODN

[1662] Joachim Schreurs, Iwein Vranckx, Johan A. K. Suykens, and Peter J. Rousseeuw. Outlier detection in non-elliptical data by kernel MRCD. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10041-7.

Latz:2021:GPT

[1663] Jonas Latz, Juan P. Madrigal-Cianci, Fabio Nobile, and Raúl Tempone. Generalized parallel tempering on Bayesian inverse problems. Statistics and Computing, 31 (5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10042-6.

Goldman:2021:SBB

[1664] Jacob Vorstrup Goldman and Sumeetpal S. Singh. Spatiotemporal blocking of the bouncy particle sampler for efficient inference in state-space models. Statistics and Computing, 31(5):??, September 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10034-6.

BenRached:2021:EIS

[1665] Nadhir Ben Rached, Abdul-Lateef Haji-Ali, and Raúl Tempone. Efficient importance sampling for large sums of independent and identically distributed random variables. Statistics and Computing, 31(6): ??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10055-1.

Dang:2021:PCP

[1666] Xuan Dang, Shuai Huang, and Xiaoning Qian. Penalized Cox’s proportional hazards model for highdimensional survival data with grouped predictors. Statistics and Computing, 31(6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10052-4.

dAvigneau:2021:APT

[1667] Alix Marie d’Avigneau, Sumeetpal S. Singh, and Lawrence M. Murray. Anytime parallel tempering. Statistics and Computing, 31 (6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10048-0.

Elvira:2021:PPF

[1668] Víctor Elvira, Joaquín Miguez, and Petar M. Djurić. On the performance of particle filters with adaptive number of particles. Statistics and Computing, 31(6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10056-0.

Harman:2021:ODM

[1669] Radoslav Harman, Lenka Filová, and Samuel Rosa. Optimal design of multifactor experiments via grid exploration. Statistics and Computing, 31(6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10046-2.

Hooker:2021:UPF

[1670] Giles Hooker, Lucas Mentch, and Siyu Zhou. Unrestricted permutation forces extrapolation: variable importance requires at least one more model, or there is no free variable importance. Statistics and Computing, 31(6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10057-z.

Huang:2021:CME

[1671] Chaofan Huang, V. Roshan Joseph, and Douglas M. Ray. Constrained minimum energy designs. Statistics and Computing, 31(6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-021-10054-
2.

Koppel:2021:COG

[1672] Alec Koppel, Hrusikesha Pradhan, and Ketan Rajawat. Consistent online Gaussian process regression without the sample complexity bottleneck. Statistics and Computing, 31 (6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10051-5.

Lee:2021:TBC

[1673] Jeong Eun Lee and Geoff K. Nicholls. Tree based credible set estimation. Statistics and Computing, 31 (6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10045-3.

Mao:2021:DCS

[1674] Yinan Mao, Xueou Wang, and Michael Evans. Detecting conflicting summary statistics in likelihood-free inference. Statistics and Computing, 31(6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10053-3.

Moriarty:2021:MCS

[1675] John Moriarty, Jure Vogrinc, and Alessandro Zocca. A Metropolis-class sampler for targets with non-convex support. Statistics and Computing, 31(6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10044-4.

Suotsalo:2021:HDS

[1676] Kimmo Suotsalo, Yingying Xu, and Johan Pensar. High-dimensional structure learning of sparse vector autoregressive models using fractional marginal pseudo-likelihood. Statistics and Computing, 31(6): ??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10049-z.

Tran:2021:VBM

[1677] Minh-Ngoc Tran, Dang H. Nguyen, and Duy Nguyen. Variational Bayes on manifolds. Statistics and Computing, 31(6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10047-1.

Zhao:2021:DSS

[1678] Zheng Zhao, Muhammad Emzir, and Simo Särkkä. Deep state-space Gaussian processes. Statistics and Computing, 31(6):??, November 2021. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10050-6.

Amato:2022:WBR

[1679] Umberto Amato, Anestis Antoniadis, and Irène Gijbels. Wavelet-based robust estimation and variable selection in nonparametric additive models. Statistics and Computing, 32 (1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10065-z.

Belomestny:2022:VRA

[1680] D. Belomestny, E. Moulines, and S. Samsonov. Variance reduction for additive functionals of Markov chains via martingale representations. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10073-z.

Boisvert-Beaudry:2022:MAP

[1681] Gabriel Boisvert-Beaudry and Mylène Bédard. MALA with annealed proposals: a generalization of locally and globally balanced proposal distributions. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10063-1.

Denti:2022:CTS

[1682] Francesco Denti, Andrea Cappozzo, and Francesca Greselin. Correction to: A two-stage Bayesian semiparametricmodel for novelty detection with robust prior information. Statistics and Computing, 32 (1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10028-4.

Dutang:2022:ESP

[1683] Christophe Dutang and Quentin Guibert. An explicit split point procedure in model-based trees allowing for a quick fitting of GLM trees and GLM forests. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10059-x.

Frisch:2022:LMD

[1684] Gabriel Frisch, Jean-Benoist Leger, and Yves Grandvalet. Learning from missing data with the binary latent block model. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10058-y.

Garcia-Escudero:2022:CPM

[1685] Luis A. García-Escudero, Agustín Mayo-Iscar, and Marco Riani. Constrained parsimonious model-based clustering. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10061-3.

Godsill:2022:PPS

[1686] Simon Godsill and Yaman Kindap. Point process simulation of generalised inverse Gaussian processes and estimation of the Jaeger integral. Statistics and Computing, 32 (1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10072-0.

Golightly:2022:APM

[1687] Andrew Golightly and Chris Sherlock. Augmented pseudo-marginal Metropolis–Hastings for partially observed diffusion processes. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-022-10083-
                                        5.

Goncalves:2022:ECE

[1688] Flávio B. Gonçalves, Lívia M. Dutra, and Roger W. C. Silva. Exact and computationally efficient Bayesian inference for generalized Markov modulated Poisson processes. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10074-y.

Hu:2022:GMB

[1689] Yaofang Hu, Wanjie Wang, and Yi Yu. Graph matching beyond perfectly-overlapping Erdős–Rényi random graphs. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10079-1.

Jones:2022:DCR

[1690] Corinne Jones, Vincent Roulet, and Zaid Harchaoui. Discriminative clustering with representation learning with any ratio of labeled to unlabeled data. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10067-x.

Jurek:2022:HSC

[1691] Marcin Jurek and Matthias Katzfuss. Hierarchical sparse Cholesky decomposition with applications to high-dimensional spatio-temporal filtering. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10077-9.

Kim:2022:EMM

[1692] Hyotae Kim and Athanasios Kottas. Erlang mixture modeling for Poisson process intensities. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-021-10064-
0.

Kuntz:2022:PFE

[1693] Juan Kuntz, Francesca R. Crucinio, and Adam M. Johansen. Productform estimators: exploiting independence to scale up Monte Carlo. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10069-9.

Lambert:2022:RVG

[1694] Marc Lambert, Silvère Bonnabel, and Francis Bach. The recursive variational Gaussian approximation (R-VGA). Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10068-w.

Liu:2022:SAC

[1695] Yang Liu and Robert J. B. Goudie. Stochastic approximation cut algorithm for inference in modularized Bayesian models. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-021-10070-

Paun:2022:EAH

[1696] L. Mihaela Paun and Dirk Husmeier. Emulation-accelerated Hamiltonian Monte Carlo algorithms for parameter estimation and uncertainty quantification in differential equation models. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10060-4.

Sembach:2022:RNT

[1697] Lena Sembach, Jan Pablo Burgard, and Volker Schulz. A Riemannian Newton trust-region method for fitting Gaussian mixture models. Statistics and Computing, 32 (1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10071-1.

Tomasetti:2022:UVB

[1698] Nathaniel Tomasetti, Catherine Forbes, and Anastasios Panagiotelis. Updating Variational Bayes: fast sequential posterior inference. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10062-2.

Wu:2022:EKF

[1699] Jiangqi Wu, Linjie Wen, and Simon Maskell. Ensemble Kalman filter based sequential Monte Carlo sampler for sequential Bayesian inference. Statistics and Computing, 32(1):??, February 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10075-x.

Abdulle:2022:EME

[1700] Assyr Abdulle, Grigorios A. Pavliotis, and Andrea Zanoni. Eigenfunction martingale estimating functions and filtered data for drift estimation of discretely observed multiscale diffusions. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10081-7.

Eigel:2022:LRT

[1701] Martin Eigel, Robert Gruhlke, and Manuel Marschall. Low-rank tensor reconstruction of concentrated densities with application to Bayesian inversion. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10087-1.

Guan:2022:SFP

[1702] Tianyu Guan, Zhenhua Lin, and Jiguo Cao. Sparse functional partial least squares regression with a locally sparse slope function. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10066-y.

Hainy:2022:OBD

[1703] Markus Hainy, David J. Price, and Christopher Drovandi. Optimal Bayesian design for model discrimination via classification. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10078-
2.

Kugler:2022:FBI

[1704] Benoit Kugler, Florence Forbes, and Sylvain Douté. Fast Bayesian inversion for high dimensional inverse problems. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-021-10019-5.

Manderson:2022:NSA

[1705] Andrew A. Manderson and Robert J. B. Goudie. A numerically stable algorithm for integrating Bayesian models using Markov melding. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10086-
2.

Molkenthin:2022:GES

[1706] Christian Molkenthin, Christian Donner, and Manfred Opper. GP-ETAS: semiparametric Bayesian inference for the spatio-temporal epidemic type aftershock sequence model. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10085-
3.

Papastamoulis:2022:IBF

[1707] Panagiotis Papastamoulis and Ioannis Ntzoufras. On the identifiability of Bayesian factor analytic models. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10084-4.

Passino:2022:LSB

[1708] Francesco Sanna Passino and Nicholas A. Heard. Latent structure blockmodels for Bayesian spectral graph clustering. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10082-6.

Sailynoja:2022:GTD

[1709] Teemu Säilynoja, Paul-Christian Bürkner, and Aki Vehtari. Graphical test for discrete uniformity and its applications in goodness-of-fit evaluation and multiple sample comparison. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10090-6.

Schmon:2022:OSR

[1710] Sebastian M. Schmon and Philippe Gagnon. Optimal scaling of random walk Metropolis algorithms using Bayesian large-sample asymptotics. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10080-8.

Suuronen:2022:CMR

[1711] Jarkko Suuronen, Neil K. Chada, and Lassi Roininen. Cauchy Markov random field priors for Bayesian inversion. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10089-z.

Titsias:2022:SCD

[1712] Michalis K. Titsias, Jakub Sygnowski, and Yutian Chen. Sequential changepoint detection in neural networks with checkpoints. Statistics and Computing, 32(2):??, April 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10088-0.

Botsas:2022:RBB

[1713] Themistoklis Botsas, Lachlan R. Mason, and Indranil Pan. Rulebased Bayesian regression. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10100-
7.

Canale:2022:ICS

[1714] Antonio Canale, Riccardo Corradin, and Bernardo Nipoti. Importance conditional sampling for Pitman-Yor mixtures. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10096-0.

Drovandi:2022:CLF

[1715] Christopher Drovandi and David T. Frazier. A comparison of likelihoodfree methods with and without summary statistics. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10092-4.

Hooker:2022:SDF

[1716] Giles Hooker and Han Lin Shang. Selecting the derivative of a functional covariate in scalar-on-function regression. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10091-5.

Hughes:2022:SOG

[1717] John Hughes. Sklar’s Omega: a Gaussian copula-based framework for assessing agreement. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10105-2.

James:2022:OAB

[1718] Nick James and Max Menzies. Optimally adaptive Bayesian spectral density estimation for stationary and nonstationary processes. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10103-
4.

Kook:2022:DAR

[1719] Lucas Kook, Beate Sick, and Peter Bühlmann. Distributional anchor regression. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10097-z.

Lu:2022:CZS

[1720] Jianfeng Lu and Lihan Wang. Complexity of zigzag sampling algorithm for strongly log-concave distributions. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10109-y.

Marchello:2022:CCE

[1721] Giulia Marchello, Audrey Fresse, and Charles Bouveyron. Coclustering of evolving count matrices with the dynamic latent block model: application to pharmacovigilance. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10098-y.

Ruzayqat:2022:MEN

[1722] Hamza Ruzayqat, Neil K. Chada, and Ajay Jasra. Multilevel estimation of normalization constants using ensemble Kalman–Bucy filters. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10094-2.

Samanta:2022:GLB

[1723] Srijata Samanta, Kshitij Khare, and George Michailidis. A generalized likelihood-based Bayesian approach for scalable joint regression and covariance selection in high dimensions. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10102-5.

Thesingarajah:2022:NWP

[1724] Denishrouf Thesingarajah and Adam M. Johansen. The node-wise pseudomarginal method: model selection with spatial dependence on latent graphs. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10101-6.

vandenBoom:2022:UAP

[1725] Willem van den Boom, Ajay Jasra, and Johan G. Eriksson. Unbiased approximation of posteriors via coupled particle Markov chain Monte Carlo. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10093-3.

Zhong:2022:BSR

[1726] Yan Zhong and Jianhua Z. Huang.

Biclustering via structured regularized matrix decomposition. Statistics and Computing, 32(3):??, June 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10095-
1.

Meila:2022:RIM

[1727] Marina Meila, Annelise Wagner, and Christopher Meek. Recursive inversion models for permutations. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10111-4.

Fisch:2022:RTA

[1728] Alexander T. M. Fisch, Lawrence Bardwell, and Idris A. Eckley. Real time anomaly detection and categorisation. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10112-3.

Hafych:2022:PMS

[1729] Vasyl Hafych, Philipp Eller, Oliver Schulz, and Allen Caldwel. Parallelizing MCMC sampling via space partitioning. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10116-z.

Zhang:2022:IPE

[1730] Jingnan Zhang and Junhui Wang. Identifiability and parameter estimation of the overlapped stochastic coblock model. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10114-1.

Deng:2022:AWS

[1731] Wei Deng, Guang Lin, and Faming Liang. An adaptively weighted stochastic gradient MCMC algorithm for Monte Carlo simulation and global optimization. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10120-3.

Zhuang:2022:PTB

[1732] Haoxin Zhuang, Liqun Diao, and Grace Yi. Polya tree-based nearest neighborhood regression. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-021-10076-w.

Davis:2022:ROR

[1733] Andrew D. Davis, Youssef Marzouk, Aaron Smith, and Natesh Pillai. Rate-optimal refinement strategies for local approximation MCMC. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10123-0.

Merlo:2022:QHS

[1734] Luca Merlo, Antonello Maruotti, Lea Petrella, and Antonio Punzo. Quantile hidden semi-Markov models for multivariate time series. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10130-1.

Morales-Otero:2022:FDH

[1735] Mabel Morales-Otero, Virgilio Gómez-Rubio, and Vicente Núñez-Antón. Fitting double hierarchical models with the integrated nested Laplace approximation. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10122-1.

Xu:2022:CAG

[1736] Zuheng Xu and Trevor Campbell. The computational asymptotics of Gaussian variational inference and the Laplace approximation. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10125-y.

Girard:2022:ABR

[1737] Stéphane Girard, Gilles Stupfler, and Antoine Usseglio-Carleve. On automatic bias reduction for extreme expectile estimation. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10118-x.

Paulin:2022:VMF

[1738] Daniel Paulin, Ajay Jasra, Alexandros Beskos, and Dan Crisan. A 4D-Var method with flow-dependent background covariances for the shallowwater equations. Statistics and Computing, 32(4):??, August 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10119-w.

Alparslan:2022:SSM

[1739] Baris Alparslan and Sinan Yildirim. Statistic selection and MCMC for differentially private Bayesian estimation. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10129-8.

Rotiroti:2022:CML

[1740] Frank Rotiroti and Stephen G. Walker. Computing marginal likelihoods via the Fourier integral theorem and pointwise estimation of posterior densities. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10131-0.

Sharp:2022:JLF

[1741] Alex Sharp and Ryan Browne. A joint latent factor analyzer and functional subspace model for clustering multivariate functional data. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10128-
9.

Zhang:2022:JEA

[1742] Nan Zhang, Muye Nanshan, and Jiguo Cao. A joint estimation approach to sparse additive ordinary differential equations. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10117-y.

Kock:2022:VIS

[1743] Lucas Kock, Nadja Klein, and David J.

Nott. Variational inference and sparsity in high-dimensional deep Gaussian mixture models. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10132-

z. See correction [1807].

Sixta:2022:CRB

[1744] Sabrina Sixta and Jeffrey S. Rosenthal. Convergence rate bounds for iterative random functions using oneshot coupling. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10134-x.

Xing:2022:IBE

[1745] Hanwen Xing. Improving bridge estimators via f-GAN. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-022-10133-
                                        y.

Marques:2022:NSM

[1746] Isa Marques, Thomas Kneib, and Nadja Klein. A non-stationary model for spatially dependent circular response data based on wrapped Gaussian processes. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10136-9.

Majumder:2022:KGA

[1747] Suman Majumder, Yawen Guan, Brian J. Reich, and Arvind K. Saibaba. Kryging: geostatistical analysis of large-scale datasets using Krylov subspace methods. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10104-3.

Sherwood:2022:QRF

[1748] Ben Sherwood and Shaobo Li. Quantile regression feature selection and estimation with grouped variables using Huber approximation. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10135-w.

Lin:2022:DRP

[1749] Cunjie Lin, Nan Qiao, Wenli Zhang, Yang Li, and Shuangge Ma. Default risk prediction and feature extraction using a penalized deep neural network. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10140-z.

Scutari:2022:AFS

[1750] Marco Scutari, Francesca Panero, and Manuel Proissl. Achieving fairness with a simple ridge penalty. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10143-
w.

Zhang:2022:GII

[1751] Benjamin J. Zhang, Youssef M. Marzouk, and Konstantinos Spiliopoulos. Geometry-informed irreversible perturbations for accelerated convergence of Langevin dynamics. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10147-
6.

Constantin:2022:MMG

[1752] Alexandre Constantin, Mathieu Fauvel, and Stéphane Girard. Mixture of multivariate Gaussian processes for classification of irregularly sampled satellite image timeseries. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10145-8.

Olsson:2022:SSJ

[1753] Jimmy Olsson, Tatjana Pavlenko, and Felix L. Rios. Sequential sampling of junction trees for decomposable graphs. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10113-2.

Sabino:2022:ESN

[1754] Piergiacomo Sabino. Exact simulation of normal tempered stable processes of OU type with applications. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10153-8.

Englezou:2022:ALI

[1755] Yiolanda Englezou, Timothy W. Waite, and David C. Woods. Approximate Laplace importance sampling for the estimation of expected Shannon information gain in highdimensional Bayesian design for nonlinear models. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10159-2.

Chevalier:2022:SRI

[1756] Jérôme-Alexis Chevalier, Tuan-Binh Nguyen, Bertrand Thirion, and Joseph Salmon. Spatially relaxed inference on high-dimensional linear models. Statistics and Computing, 32 (5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10139-6.

Liang:2022:ARN

[1757] Xitong Liang, Samuel Livingstone, and Jim Griffin. Adaptive random neighbourhood informed Markov chain Monte Carlo for high-dimensional Bayesian variable selection. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10137-
8.

Forbes:2022:SSD

[1758] Florence Forbes, Hien Duy Nguyen, TrungTin Nguyen, and Julyan Arbel. Summary statistics and discrepancy measures for approximate Bayesian computation via surrogate posteriors. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10155-6.

Casas:2022:SHM

[1759] Fernando Casas, Jesús María Sanz-Serna, and Luke Shaw. Split Hamiltonian Monte Carlo revisited. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10149-4.

Cai:2022:PNS

[1760] Xiaohao Cai, Jason D. McEwen, and Marcelo Pereyra. Proximal nested sampling for high-dimensional Bayesian model selection. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-022-10152-
                                        9.

Papamarkou:2022:RPD

[1761] Theodore Papamarkou, Farzana Nasrin, Austin Lawson, Na Gong, Orlando Rios, and Vasileios Maroulas. A random persistence diagram generator. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10141-y.

Stringer:2022:ADB

[1762] Alex Stringer. Automatic differentiation of Box–Cox transformations with application to random effects models for continuous non-normal response. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10160-9.

Bucci:2022:CUP

[1763] Andrea Bucci, Luigi Ippoliti, and Pasquale Valentini. Comparing unconstrained parametrization methods for return covariance matrix prediction. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10157-4.

Wang:2022:GDN

[1764] Tiandong Wang, Jun Yan, Yelie Yuan, and Panpan Zhang. Generating directed networks with predetermined assortativity measures. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10161-
8.

Duarte:2022:ODD

[1765] Belmiro P. M. Duarte, Anthony C.

Atkinson, and Nuno M. C. Oliveira. Optimal designs for dose-escalation trials and individual allocations in cohorts. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10158-3.

Chee:2022:DDU

[1766] Chew-Seng Chee and Byungtae Seo.

Density deconvolution under a kmonotonicity constraint. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-022-10150-
                                        x.

Greengard:2022:ERR

[1767] Philip Greengard and Michael O’Neil. Efficient reduced-rank methods for Gaussian processes with eigenfunction expansions. Statistics and Computing, 32(5):??, October 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10124-z.

Yatracos:2022:LWM

[1768] Yannis G. Yatracos. Limitations of the Wasserstein MDE for univariate data. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10146-7.

Mishra:2022:PVS

[1769] Swapnil Mishra, Seth Flaxman, Tresnia Berah, Harrison Zhu, Mikko Pakkanen, and Samir Bhatt. π VAE: a stochastic process prior for Bayesian deep learning with MCMC. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10151-w.

Wen:2022:AMB

[1770] Linjie Wen and Jinglai Li. Affinemapping based variational ensemble Kalman filter. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10156-5.

Shlomovich:2022:PEM

[1771] Leigh Shlomovich, Edward A. K.

Cohen, and Niall Adams. A parameter estimation method for multivariate binned Hawkes processes. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10121-2.

Ma:2022:PTF

[1772] Xiaoyu Ma, Sylvain Sardy, Nick Hengartner, Nikolai Bobenko, and Yen Ting Lin. A phase transition for finding needles in nonlinear haystacks with LASSO artificial neural networks. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10169-0.

Castilla:2022:RAC

[1773] Elena Castilla, María Jaenada, Nirian Martín, and Leandro Pardo. Robust approach for comparing two dependent normal populations through Wald-type tests based on Rényi’s pseudodistance estimators. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10162-7.

Rendleman:2022:RRS

[1774] Michael C. Rendleman, Brian J. Smith, Guadalupe Canahuate, Terry A. Braun, John M. Buatti, and Thomas L. Casavant. Representative random sampling: an empirical evaluation of a novel bin stratification method for model performance estimation. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10138-7.

Brocker:2022:UCT

[1775] Jochen Bröcker. Uniform calibration tests for forecasting systems with small lead time. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10144-9.

Roikjer:2022:GBA

[1776] Tobias Røikjer, Asger Hobolth, and Kasper Munch. Graph-based algorithms for phase-type distributions. Statistics and Computing, 32 (6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10174-3.

Laurini:2022:LTR

[1777] Fabrizio Laurini, Paul Fearnhead, and Jonathan Tawn. Limit theory and robust evaluation methods for the extremal properties of GARCH(p, q) processes. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10164-5.

Liu:2022:DRB

[1778] Chunshan Liu, Daniel R. Kowal, and Marina Vannucci. Dynamic and robust Bayesian graphical models. Statistics and Computing, 32 (6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10177-0.

He:2022:SEF

[1779] Shiyuan He, Hanxuan Ye, and Kejun He. Spline estimation of functional principal components via manifold conjugate gradient algorithm. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10175-2.

Corbella:2022:AZZ

[1780] Alice Corbella, Simon E. F. Spencer, and Gareth O. Roberts. Automatic zig-zag sampling in practice. Statistics and Computing, 32 (6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10142-x.

Ameli:2022:ILD

[1781] Siavash Ameli and Shawn C. Shadden. Interpolating log-determinant and trace of the powers of matrix A + tB. Statistics and Computing, 32(6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10173-4.

Schoen:2022:SED

[1782] Eric D. Schoen, Pieter T. Eendebak, Alan R. Vazquez, and Peter Goos. Systematic enumeration of definitive screening designs. Statistics and Computing, 32 (6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10171-6.

Hallgren:2022:CDN

[1783] Karl L. Hallgren, Nicholas A. Heard, and Niall M. Adams. Changepoint detection in non-exchangeable data. Statistics and Computing, 32 (6):??, December 2022. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10176-1.

Li:2023:ASI

[1784] Zheyuan Li and Jiguo Cao. Automatic search intervals for the smoothing parameter in penalized splines. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10178-z.

Krueger:2023:RDC

[1785] Rico Krueger, Michel Bierlaire, Thomas Gasos, and Prateek Bansal. Robust discrete choice models with t-distributed kernel errors. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10182-3.

Vargas:2023:BLN

[1786] Francisco Vargas, Andrius Ovsianas, David Fernandes, Mark Girolami, Neil D. Lawrence, and Nikolas Nüsken. Bayesian learning via neural Schrödinger–Föllmer flows. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10172-
5.

Grabchak:2023:ESE

[1787] Michael Grabchak and Piergiacomo Sabino. Efficient simulation of ptempered α-stable OU processes. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https://link.

springer.com/article/10.1007/s11222-
022-10165-4.

Capo:2023:LSD

[1788] Marco Capó, Aritz Pérez, and José A. Lozano. LASSO for streaming data with adaptative filtering. Statistics and Computing, 33 (1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10181-4.

Alexopoulos:2023:VRM

[1789] Angelos Alexopoulos, Petros Dellaportas, and Michalis K. Titsias. Variance reduction for Metropolis–Hastings samplers. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10183-2.

Vazquez:2023:CTL

[1790] Alan R. Vazquez, Weng Kee Wong, and Peter Goos. Constructing twolevel Q_{B}-optimal screening designs using mixed-integer programming and heuristic algorithms. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-022-10168-
                                        1.

Bierkens:2023:SPS

[1791] Joris Bierkens, Sebastiano Grazzi, Frank van der Meulen, and Moritz Schauer. Sticky PDMP samplers for sparse and local inference problems. Statistics and Computing, 33 (1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10180-5.

Yi:2023:MFG

[1792] Si-Yu Yi and Yong-Dao Zhou. Modelfree global likelihood subsampling for massive data. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10185-0.

Vo-Thanh:2023:BOT

[1793] Nha Vo-Thanh and Hans-Peter Piepho.

Bayesian A-optimal two-phase designs with a single blocking factor in each phase. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10126-x.

Smith:2023:PSD

[1794] Anna L. Smith, Tian Zheng, and Andrew Gelman. Prediction scoring of data-driven discoveries for reproducible research. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10154-7.

Yang:2023:SCE

[1795] Zhirong Yang, Yuwei Chen, Denis Sedov, Samuel Kaski, and Jukka Corander. Stochastic cluster embedding. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10186-z.

Ge:2023:SMS

[1796] Shufei Ge, Shijia Wang, and Lloyd Elliott. Shape modeling with spline partitions. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10170-7.

Pronzato:2023:PAG

[1797] Luc Pronzato. Performance analysis of greedy algorithms for minimising a Maximum Mean Discrepancy. Statistics and Computing, 33 (1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10184-1.

Zhong:2023:SLF

[1798] Rou Zhong, Shishi Liu, Haocheng Li, and Jingxiao Zhang. Sparse logistic functional principal component analysis for binary data. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10190-
3.

Cironis:2023:AMT

[1799] Lukas Cironis, Jan Palczewski, and Georgios Aivaliotis. Automatic model training under restrictive time constraints. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10166-3.

Riutort-Mayol:2023:PHS

[1800] Gabriel Riutort-Mayol, Paul-Christian Bürkner, Michael R. Andersen, Arno Solin, and Aki Vehtari. Practical Hilbert space approximate Bayesian Gaussian processes for probabilistic programming. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10167-2.

Kejzlar:2023:VIV

[1801] Vojtech Kejzlar and Tapabrata Maiti.

Variational inference with vine copulas: an efficient approach for Bayesian computer model calibration. Statistics and Computing, 33 (1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10194-z.

Maama:2023:BPI

[1802] Mohamed Maama, Ajay Jasra, and Hernando Ombao. Bayesian parameter inference for partially observed stochastic differential equations driven by fractional Brownian motion. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10193-0.

Spuck:2023:FTS

[1803] Nikolai Spuck, Matthias Schmid, Nils Heim, Ute Klarmann-Schulz, Achim Hörauf, and Moritz Berger. Flexible tree-structured regression models for discrete event times. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10196-
x.

Freguglia:2023:DRS

[1804] Victor Freguglia and Nancy L. Garcia. Detecting renewal states in chains of variable length via intrinsic Bayes factors. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10191-2.

Raim:2023:DSS

[1805] Andrew M. Raim. Direct sampling with a step function. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10188-x.

Pentland:2023:GTP

[1806] Kamran Pentland, Massimiliano Tamborrino, T. J. Sullivan, James Buchanan, and L. C. Appel. GParareal: a time-parallel ODE solver using Gaussian process emulation. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-022-10195-
                                        y.

Kock:2023:CVI

[1807] Lucas Kock, Nadja Klein, and David J. Nott. Correction to: Variational inference and sparsity in highdimensional deep Gaussian mixture models. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.

1007/s11222-022-10179-y.  See

[1743].

Gaedke-Merzhauser:2023:PIN

[1808] Lisa Gaedke-Merzhäuser, Janet van Niekerk, Olaf Schenk, and Håvard Rue. Parallelized integrated nested Laplace approximations for fast Bayesian inference. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10192-1.

Korhonen:2023:FUE

[1809] Pekka Korhonen, Francis K. C.

Hui, Jenni Niku, and Sara Taskinen. Fast and universal estimation of latent variable models using extended variational approximations. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10189-w.

Lee:2023:VSU

[1810] Yeonjoo Lee, Natalia Rojas-Perilla, Marina Runge, and Timo Schmid. Variable selection using conditional AIC for linear mixed models with data-driven transformations. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-022-10198-
9.

Rusch:2023:SBH

[1811] Thomas Rusch, Patrick Mair, and Kurt Hornik. Structure-based hyperparameter selection with Bayesian optimization in multidimensional scaling. Statistics and Computing, 33 (1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10197-w.

Poinas:2023:PVS

[1812] Arnaud Poinas and Rémi Bardenet. On proportional volume sampling for experimental design in general spaces. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10115-0.

Zhang:2023:TPD

[1813] Junyi Zhang and Angelos Dassios. Truncated Poisson–Dirichlet approximation for Dirichlet process hierarchical models. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10201-3.

Yamashita:2023:EH

[1814] Hiroshi Yamashita, Hideyuki Suzuki, and Kazuyuki Aihara. Entropic herding. Statistics and Computing, 33 (1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10199-8.

Valeriano:2023:MRN

[1815] Katherine A. L. Valeriano, Christian E. Galarza, and Larissa A. Matos. Moments and random number generation for the truncated elliptical family of distributions. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10200-4.

Chakraborty:2023:MBA

[1816] Atlanta Chakraborty, David J. Nott, Christopher C. Drovandi, David T. Frazier, and Scott A. Sisson. Modularized Bayesian analyses and cutting feedback in likelihood-free inference. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10207-5.

Ahfock:2023:RSA

[1817] Daniel Ahfock, William J. Astle, and Sylvia Richardson. On randomized sketching algorithms and the Tracy– Widom law. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10148-5.

Wakefield:2023:MBD

[1818] Bradley Wakefield, Yan-Xia Lin, Rathin Sarathy, and Krishnamurty Muralidhar. Moment-based density estimation of confidential microdata: a computational statistics approach. Statistics and Computing, 33(1):??, February 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-022-10203-1.

Pulido:2023:FEH

[1819] Belén Pulido, Alba M. Franco-Pereira, and Rosa E. Lillo. A fast epigraph and hypograph-based approach for clustering functional data. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10213-7.

Rugamer:2023:PTS

[1820] David Rügamer, Philipp F. M. Baumann, Thomas Kneib, and Torsten Hothorn. Probabilistic time series forecasts with autoregressive transformation models. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10212-8.

Chen:2023:NBM

[1821] Li-Pang Chen. De-noising boosting methods for variable selection and estimation subject to error-prone variables. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10209-3.

Yonekura:2023:ATP

[1822] Shouto Yonekura and Shonosuke Sugasawa. Adaptation of the tuning parameter in general Bayesian inference with robust divergence. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10205-7.

BenAmar:2023:SDI

[1823] Eya Ben Amar, Nadhir Ben Rached, Abdul-Lateef Haji-Ali, and Raúl Tempone. State-dependent importance sampling for estimating expectations of functionals of sums of independent random variables. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10202-2.

Duhamel:2023:VBC

[1824] Clément Duhamel, Céline Helbert, Miguel Munoz Zuniga, Clémentine Prieur, and Delphine Sinoquet. A SUR version of the bichon criterion for excursion set estimation. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10208-4.

Jarvenpaa:2023:PII

[1825] Marko Järvenpää and Jukka Corander.

On predictive inference for intractable models via approximate Bayesian computation. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10163-6.

Weng:2023:BHD

[1826] Huibin Weng and Olivier Parent. Beyond homophilic dyadic interactions: the impact of network formation on individual outcomes. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10215-5.

Paun:2023:SVI

[1827] Ionut Paun, Dirk Husmeier, and Colin J. Torney. Stochastic variational inference for scalable nonstationary Gaussian process regression. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10210-w.

Yang:2023:ASS

[1828] Zhuang Yang and Li Ma. Adaptive step size rules for stochastic optimization in large-scale learning. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10218-2.

Geenens:2023:SDA

[1829] Gery Geenens, Alicia Nieto-Reyes, and Giacomo Francisci. Statistical depth in abstract metric spaces. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11222-023-10216-
4.

Nezakati:2023:UDE

[1830] Ensiyeh Nezakati and Eugen Pircalabelu. Unbalanced distributed estimation and inference for the precision matrix in Gaussian graphical models. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10211-9.

Geoga:2023:FMS

[1831] Christopher J. Geoga, Oana Marin, Michel Schanen, and Michael L. Stein. Fitting Matérn smoothness parameters using automatic differentiation. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-022-10127-w.

Majka:2023:MIA

[1832] Mateusz B. Majka, Marc Sabate-Vidales, and Lukasz Szpruch. Multiindex antithetic stochastic gradient algorithm. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10220-8.

Gutierrez:2023:NFB

[1833] Iván Gutiérrez, Luis Gutiérrez, and Danilo Alvares. A new flexible Bayesian hypothesis test for multivariate data. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10214-6.

deFreitas:2023:SAE

[1834] João Victor B. de Freitas, Caio L. N. Azevedo, and Juvêncio S. Nobre. A stochastic approximation ECME algorithm to semi-parametric scale mixtures of centred skew normal regression models. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10223-5.

Al-Qadhi:2023:SNQ

[1835] Al-Fahad Al-Qadhi, Carey E. Priebe, Hayden S. Helm, and Vince Lyzinski. Subgraph nomination: query by example subgraph retrieval in networks. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10206-6.

Sterzinger:2023:MSP

[1836] Philipp Sterzinger and Ioannis Kosmidis. Maximum softly-penalized likelihood for mixed effects logistic regression. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10217-3.

Quiroz:2023:FBI

[1837] Zaida C. Quiroz, Marcos O. Prates, Dipak K. Dey, and H.åvard Rue.

Fast Bayesian inference of block nearest neighbor Gaussian models for large data. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10227-1.

Redivo:2023:QDF

[1838] Edoardo Redivo, Cinzia Viroli, and Alessio Farcomeni. Quantile-distribution functions and their use for classification, with application to naïve Bayes classifiers. Statistics and Computing, 33(2):??, April 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10224-4.

Kang:2023:CBS

[1839] Myeongjong Kang and Matthias Katzfuss. Correlation-based sparse inverse Cholesky factorization for fast Gaussian-process inference. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10231-
5.

Zhou:2023:FCH

[1840] Xiaoxiao Zhou and Xinyuan Song.

Functional concurrent hidden Markov model. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10226-2.

BenHammouda:2023:LBI

[1841] Chiheb Ben Hammouda, Nadhir Ben Rached, Raúl Tempone, and Sophia Wiechert. Learning-based importance sampling via stochastic optimal control for stochastic reaction networks. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10222-6.

Peng:2023:FPT

[1842] Chien-Yu Peng, Yi-Shian Dong, and Tsai-Hung Fan. The first-passage-time moments for the Hougaard process and its Birnbaum–Saunders approximation. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10235-1.

Antonazzo:2023:FGC

[1843] Filippo Antonazzo, Christophe Biernacki, and Christine Keribin. Frugal Gaussian clustering of huge imbalanced datasets through a bin-marginal approach. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10221-7.

Hawat:2023:ESF

[1844] Diala Hawat, Guillaume Gautier, Rémi Bardenet, and Raphaël Lachièze-Rey. On estimating the structure factor of a point process, with applications to hyperuniformity. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10219-1.

Liang:2023:OOA

[1845] Xin Liang. On the optimality of the Oja’s algorithm for online PCA. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10229-z.

Gryvill:2023:SMF

[1846] Håkon Gryvill and Håkon Tjelmeland. A sparse matrix formulation of model-based ensemble Kalman filter. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10228-0.

Basson:2023:VTG

[1847] Marno Basson, Tobias M. Louw, and Theresa R. Smith. Variational tobit Gaussian process regression. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10225-3.

Fort:2023:SVM

[1848] Gersende Fort and Eric Moulines.

Stochastic variable metric proximal gradient with variance reduction for non-convex composite optimization. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10230-6.

Coullon:2023:EGT

[1849] Jeremie Coullon, Leah South, and Christopher Nemeth. Efficient and generalizable tuning strategies for stochastic gradient MCMC. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10233-
3.

Xu:2023:BPM

[1850] Min Xu and Zhongfeng Qin. A Bayesian parametrized method for interval-valued regression models. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https://link.

springer.com/article/10.1007/s11222-
023-10234-2.

Pan:2023:GSI

[1851] Maolin Pan, Minggao Gu, Xianyi Wu, and Xiaodan Fan. Globally and symmetrically identified Bayesian multinomial probit model. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10232-4.

Makalic:2023:MLE

[1852] Enes Makalic and Daniel F. Schmidt.

Maximum likelihood estimation of the Weibull distribution with reduced bias. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10236-0.

Lambert:2023:LMR

[1853] Marc Lambert, Silvère Bonnabel, and Francis Bach. The limited-memory recursive variational Gaussian approximation (L-RVGA). Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10239-x.

ONeill:2023:VSU

[1854] Meadhbh O’Neill and Kevin Burke.

Variable selection using a smooth information criterion for distributional regression models. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10204-8.

Ding:2023:VSR

[1855] Yujia Ding, Qidi Peng, Zhengming Song, and Hansen Chen. Variable selection and regularization via arbitrary rectangle-range generalized elastic net. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10240-4.

Wang:2023:DSO

[1856] Kangning Wang and Shaomin Li. Distributed statistical optimization for non-randomly stored big data with application to penalized learning. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10247-x.

Liu:2023:GJE

[1857] Yiming Liu, Shaochen Wang, and Wang Zhou. General jackknife empirical likelihood and its applications. Statistics and Computing, 33(3):??, June 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10245-z.

Kang:2023:FDB

[1858] Ilsuk Kang, Hosik Choi, Young Joo Yoon, Junyoung Park, Soon-Sun Kwon, and Cheolwoo Park. Fréchet distance-based cluster analysis for multi-dimensional functional data. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https://link.

springer.com/article/10.1007/s11222-
023-10237-z.

Dixit:2023:PFA

[1859] Vaidehi Dixit and Ryan Martin. A PRticle filter algorithm for nonparametric estimation of multivariate mixing distributions. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10242-2.

Mastrototaro:2023:AOV

[1860] Alessandro Mastrototaro and Jimmy Olsson. Adaptive online variance estimation in particle filters: the AL-Var estimator. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10243-1.

Huang:2023:QUB

[1861] Chen Huang, Jieli Ding, and Yanqin Feng. A quadratic upper bound algorithm for regression analysis of credit risk under the proportional hazards model with casecohort data. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10248-w.

Yang:2023:CRM

[1862] Xiaodong Yang and Jun S. Liu. Convergence rate of multiple-try Metropolis independent sampler. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10241-
3.

Bettisworth:2023:PEC

[1863] Ben Bettisworth, Alexander I. Jordan, and Alexandros Stamatakis. Phylourny: efficiently calculating elimination tournament win probabilities via phylogenetic methods. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10246-
y.

Philipson:2023:FLM

[1864] Pete Philipson and Alan Huang. A fast look-up method for Bayesian meanparameterised Conway–Maxwell–Poisson regression models. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10244-0.

Botha:2023:AAN

[1865] Imke Botha, Robert Kohn, Leah South, and Christopher Drovandi. Automatically adapting the number of state particles in SMC^{2}. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10250-2.

Zhang:2023:SBB

[1866] Tianfang Zhang, Rasmus Bokrantz, and Jimmy Olsson. A similaritybased Bayesian mixture-of-experts model. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10238-y.

Beckman:2023:SCN

[1867] Paul G. Beckman, Christopher J.

Geoga, Michael L. Stein, and Mihai Anitescu. Scalable computations for nonstationary Gaussian processes. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10252-0.

Eftekhari:2023:FBE

[1868] Armin Eftekhari, Luis Vargas, and Konstantinos C. Zygalakis. The forward-backward envelope for sampling with the overdamped Langevin algorithm. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10254-y.

Zhang:2023:FMA

[1869] Haixiang Zhang and Xin Li. A framework for mediation analysis with massive data. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10255-x.

Palias:2023:EID

[1870] Efstratios Palias and Ata Kabán. The effect of intrinsic dimension on the Bayes-error of projected quadratic discriminant classification. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10251-
1.

Zhan:2023:PRI

[1871] Zishu Zhan, Xiangjie Li, and Jingxiao Zhang. Partial replacement imputation estimation for partially linear models with complex missing pattern covariates. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10258-8.

Lan:2023:BSM

[1872] Shiwei Lan, Shuyi Li, and Mirjeta Pasha. Bayesian spatiotemporal modeling for inverse problems. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10253-z.

Li:2023:DDM

[1873] Sijing Li, Cheng Zhang, Zhiwen Zhang, and Hongkai Zhao. A datadriven and model-based accelerated Hamiltonian Monte Carlo method for Bayesian elliptic inverse problems. Statistics and Computing, 33(4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10262-y.

Bonnet:2023:IME

[1874] Anna Bonnet, Miguel Martinez Herrera, and Maxime Sangnier. Inference of multivariate exponential Hawkes processes with inhibition and application to neuronal activity. Statistics and Computing, 33 (4):??, August 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10264-w.

Gilks:2023:WMC

[1875] Walter R. Gilks, Lukas Cironis, and Stuart Barber. Wavelet Monte Carlo: a principle for sampling from complex distributions. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10256-w.

Richter:2023:IBC

[1876] Robin Richter, Shankar Bhamidi, and Sach Mukherjee. Improved baselines for causal structure learning on interventional data. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10257-9.

Anastasiou:2023:GMC

[1877] Andreas Anastasiou and Angelos Papanastasiou. Generalized multiple change-point detection in the structure of multivariate, possibly highdimensional, data sequences. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10261-
z.

Boutin:2023:ETS

[1878] Rémi Boutin, Charles Bouveyron, and Pierre Latouche. Embedded topics in the stochastic block model. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10265-9.

Liu:2023:DAN

[1879] Ruiqi Liu, Ganggang Xu, and Zuofeng Shang. Distributed adaptive nearest neighbor classifier: algorithm and theory. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10267-7.

Liang:2023:RMI

[1880] Xinzhu Liang, Shangda Yang, Simon L. Cotter, and Kody J. H. Law. A randomized multi-index sequential Monte Carlo method. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10249-9.

Crispino:2023:EAI

[1881] Marta Crispino, Cristina Mollica, Valerio Astuti, and Luca Tardella. Efficient and accurate inference for mixtures of Mallows models with Spearman distance. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10266-8.

Karling:2023:GFT

[1882] Maicon J. Karling, Marc G. Genton, and Simos G. Meintanis. Goodnessof-fit tests for multivariate skewed distributions based on the characteristic function. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10260-0.

Forsyth:2023:ULM

[1883] Jessica E. Forsyth, Ali H. Al-Anbaki, Berenika Plusa, and Simon L. Cotter. Unlabelled landmark matching via Bayesian data selection, and application to cell matching across imaging modalities. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10259-7.

Li:2023:UGL

[1884] Xinmin Li, Haozhe Liang, Wolfgang Härdle, and Hua Liang. Use generalized linear models or generalized partially linear models? Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10278-
4.

Li:2023:LUS

[1885] Yanxin Li, Antonio Linero, and Stephen G. Walker. Latent uniform samplers on multivariate binary spaces. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10276-6.

Demange-Chryst:2023:EEM

[1886] Julien Demange-Chryst, François Bachoc, and Jérôme Morio. Efficient estimation of multiple expectations with the same sample by adaptive importance sampling and control variates. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10270-y.

Vicente:2023:HDO

[1887] Gonzalo Vicente, Aritz Adin, Tomás Goicoa, and María Dolores Ugarte. High-dimensional order-free multivariate spatial disease mapping. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10263-
x.

Hou-Liu:2023:GLM

[1888] Jason Hou-Liu and Ryan P. Browne. Generalized linear models for massive data via doubly-sketching. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10274-
8.

Tsagris:2023:FNP

[1889] Michail Tsagris, Abdulaziz Alenazi, and Connie Stewart. Flexible nonparametric regression models for compositional response data with zeros. Statistics and Computing, 33 (5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10277-5.

Garcia-Portugues:2023:TPD

[1890] Eduardo García-Portugués and Arturo Prieto-Tirado. Toroidal PCA via density ridges. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10273-9.

Palasciano:2023:TAA

[1891] Henry Antonio Palasciano and Guy P.

Nason. A test for the absence of aliasing or white noise in twodimensional locally stationary wavelet processes. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10269-5.

Mrkvicka:2023:FDR

[1892] Tomás Mrkvicka and Mari Myllymäki. False discovery rate envelopes. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10275-7.

Bodenham:2023:EEC

[1893] Dean A. Bodenham and Yoshinobu Kawahara. euMMD: efficiently computing the MMD twosample test statistic for univariate data. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10271-x.

Llewellyn:2023:PMP

[1894] Mary Llewellyn, Ruth King, Víctor Elvira, and Gordon Ross. A point mass proposal method for Bayesian state-space model fitting. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10268-
                                        6.

Watson:2023:ECO

[1895] Samuel I. Watson and Yi Pan. Evaluation of combinatorial optimisation algorithms for c-optimal experimental designs with correlated observations. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10280-w.

Li:2023:TCT

[1896] Yan Li, Na Han, Yuxiang Qin, Jing Zhang, and Jinxia Su. TranscGAN: transformer-unet-based generative adversarial networks for crossmodality magnetic resonance image synthesis. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10282-8.

Shi:2023:ODG

[1897] Yao Shi, Wanchunzi Yu, and John Stufken. Optimal designs for generalized linear mixed models based on the penalized quasi-likelihood method. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10279-3.

Alvares:2023:TSA

[1898] Danilo Alvares and Valeria Leiva-Yamaguchi. A two-stage approach for Bayesian joint models: reducing complexity while maintaining accuracy. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10281-9.

Merad:2023:RSL

[1899] Ibrahim Merad and Stéphane Gaïffas.

Robust supervised learning with coordinate gradient descent. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10283-

Zou:2023:MAS

[1900] Jiahui Zou, Chaoxia Yuan, Xinyu Zhang, Guohua Zou, and Alan T. K. Wan. Model averaging for support vector classifier by crossvalidation. Statistics and Computing, 33(5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10284-6.

Wu:2023:NNS

[1901] Sidi Wu, Cédric Beaulac, and Jiguo Cao. Neural networks for scalar input and functional output. Statistics and Computing, 33 (5):??, October 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10287-3.

Fang:2023:BAM

[1902] Junhan Fang and Grace Y. Yi. Bayesian analysis for matrix-variate logistic regression with/without response misclassification. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10286-
                                        4.

Pronello:2023:PMB

[1903] Nicola Pronello, Rosaria Ignaccolo, Luigi Ippoliti, and Sara Fontanella. Penalized model-based clustering of complex functional data. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10288-
2.

Tomarchio:2023:PMA

[1904] Salvatore D. Tomarchio, Antonio Punzo, and Luca Bagnato. Parsimonious mixtures for the analysis of tensor-variate data. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10291-
7.

Kwok:2023:LBB

[1905] Wai M. Kwok, George Streftaris, and Sarat C. Dass. Laplace based Bayesian inference for ordinary differential equation models using regularized artificial neural networks. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10289-1.

Lau:2023:EVM

[1906] Yuen Tsz Abby Lau, Tianying Wang, Jun Yan, and Xuebin Zhang. Extreme value modeling with errorsin-variables in detection and attribution of changes in climate extremes. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10290-8.

Viani:2023:CFH

[1907] Alessandro Viani, Adam M. Johansen, and Alberto Sorrentino. Cost free hyper-parameter selection/averaging for Bayesian inverse problems with vanilla and Rao–Blackwellized SMC samplers. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10294-4.

Rana:2023:BHM

[1908] Masud Rana, Justin Kosar, and Shaqil Peermohamed. Bayesian hierarchical models incorporating measurement error for interrupted time series design. Statistics and Computing, 33 (6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10295-3.

Lyu:2023:TSB

[1909] Guanjie Lyu and Mohamed Belalia. Testing symmetry for bivariate copulas using Bernstein polynomials. Statistics and Computing, 33 (6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10297-1.

Nascimento:2023:USP

[1910] Abraão D. C. Nascimento, Jodavid A. Ferreira, and Alejandro C. Frery. Unsupervised segmentation of Pol-SAR data with complex Wishart and }_{m}^{0} distributions and Shannon entropy. Statistics and Computing, 33 (6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10298-0.

Rao:2023:MNL

[1911] Aniruddha Rajendra Rao and Matthew Reimherr. Modern non-linear functionon-function regression. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10299-
z.

Chen:2023:ORF

[1912] Sigeng Chen, Jeffrey S. Rosenthal, Aki Dote, Hirotaka Tamura, and Ali Sheikholeslami. Optimization via rejection–free partial neighbor search. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10300-9.

Barabesi:2023:CFM

[1913] Lucio Barabesi, Andrea Cerioli, Luis Angel García-Escudero, and Agustín Mayo-Iscar. Consistency factor for the MCD estimator at the Student-t distribution. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10296-
2.

DeCanditiis:2023:LBU

[1914] Daniela De Canditiis. Learning binary undirected graph in low dimensional regime. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10321-4.

Golightly:2023:ABI

[1915] Andrew Golightly, Laura E. Wadkin, Sam A. Whitaker, Andrew W. Baggaley, Nick G. Parker, and Theodore Kypraios. Accelerating Bayesian inference for stochastic epidemic models using incidence data. Statistics and Computing, 33 (6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10311-6.

Rappl:2023:SJM

[1916] Anja Rappl, Thomas Kneib, Stefan Lang, and Elisabeth Bergherr. Spatial joint models through Bayesian structured piecewise additive joint modelling for longitudinal and time-to-event data. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10293-5.

Mai:2023:RRA

[1917] The Tien Mai. A reduced-rank approach to predicting multiple binary responses through machine learning. Statistics and Computing, 33 (6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10314-3.

Schorlepp:2023:SMC

[1918] Timo Schorlepp, Shanyin Tong, Tobias Grafke, and Georg Stadler. Scalable methods for computing sharp extreme event probabilities in infinite-dimensional stochastic systems. Statistics and Computing, 33 (6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10307-2.

Brouste:2023:OSC

[1919] Alexandre Brouste, Christophe Dutang, Lilit Hovsepyan, and Tom Rohmer. One-step closed-form estimator for generalized linear model with categorical explanatory variables. Statistics and Computing, 33 (6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10313-4.

Journel:2023:SDP

[1920] Lucas Journel and Pierre Monmarché.

Switched diffusion processes for nonconvex optimization and saddle points search. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10315-2.

Xu:2023:DET

[1921] Danli Xu and Yong Wang. Density estimation for toroidal data using semiparametric mixtures. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10305-
                                        4.

Chen:2023:CEE

[1922] Yan Chen, Ruipeng Dong, and Canhong Wen. Communication-efficient estimation for distributed subset selection. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10302-7.

Sun:2023:MCG

[1923] Yan Sun and Shihao Yang. Manifoldconstrained Gaussian process inference for time-varying parameters in dynamic systems. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10319-y.

Anonymous:2023:JN

[1924] Anonymous. Journal navigation. Statistics and Computing, 33(6):??, December 2023. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Cheung:2024:HDL

[1925] Kin Yap Cheung and Stephen M. S. Lee. High-dimensional local polynomial regression with variable selection and dimension reduction. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10308-
1.

Kim:2024:ISR

[1926] Kipoong Kim and Sungkyu Jung.

Integrative sparse reduced-rank regression via orthogonal rotation for analysis of high-dimensional multisource data. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10322-3.

Xue:2024:ELE

[1927] Liugen Xue. Empirical likelihood and estimation in single-index varying-coefficient models with censored data. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10310-7.

Li:2024:SEH

[1928] Xin Li and Dongya Wu. Sparse estimation in high-dimensional linear errors-in-variables regression via a covariate relaxation method. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10312-

Moindjie:2024:CMF

[1929] Issam-Ali Moindjié, Sophie Dabo-Niang, and Cristian Preda. Classification of multivariate functional data on different domains with Partial Least Squares approaches. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10324-
                                        1.

Ortega-Fernandez:2024:EGA

[1930] Ines Ortega-Fernandez, Marta Sestelo, and Nora M. Villanueva. Explainable generalized additive neural networks with independent neural network training. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10320-5. See cor-

rection [2077].

Lakra:2024:DAM

[1931] Arjun Lakra, Buddhananda Banerjee, and Arnab Kumar Laha. A data-adaptive method for outlier detection from functional data. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10301-
                                        8.

Jiang:2024:TMI

[1932] Jingjing Jiang, Chunjie Wang, Deng Pan, and Xinyuan Song. Transformation models with informative partly interval-censored data. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10306-
3.

Cavalheiro:2024:RFK

[1933] Lucca Portes Cavalheiro, Simon Bernard, Jean Paul Barddal, and Laurent Heutte. Random forest kernel for high-dimension low sample size classification. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10309-0.

Huang:2024:FRL

[1934] Hsin-Hsiung Huang, Feng Yu, Xing Fan, and Teng Zhang. A framework of regularized low-rank matrix models for regression and classification. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10318-z.

Perepolkin:2024:HEQ

[1935] Dmytro Perepolkin, Benjamin Goodrich, and Ullrika Sahlin. Hybrid elicitation and quantile-parametrized likelihood. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10325-0.

Allouche:2024:EEQ

[1936] Michaël Allouche, Stéphane Girard, and Emmanuel Gobet. Estimation of extreme quantiles from heavytailed distributions with neural networks. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10331-2.

Zanini:2024:VIB

[1937] Carlos Tadeu Pagani Zanini, Helio S. Migon, and Ronaldo Dias. Variational inference for Bayesian bridge regression. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10317-0.

Biswas:2024:CCI

[1938] Rahul Biswas and Somabha Mukherjee. Consistent causal inference from time series with PC algorithm and its time-aware extension. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10330-3.

Onizuka:2024:FLA

[1939] Takahiro Onizuka, Shintaro Hashimoto, and Shonosuke Sugasawa. Fast and locally adaptive Bayesian quantile smoothing using calibrated variational approximations. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10327-
y.

Ramos:2024:PLD

[1940] Pedro L. Ramos, Nixon Jerez-Lillo, Francisco A. Segovia, Osafu A. Egbon, and Francisco Louzada. Powerlaw distribution in pieces: a semiparametric approach with change point detection. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10336-x.

Bhattacharya:2024:CSV

[1941] Shrijita Bhattacharya, Zihuan Liu, and Tapabrata Maiti. Comprehensive study of variational Bayes classification for dense deep neural networks. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10338-9.

Alfo:2024:CFM

[1942] Marco Alfò, Nicola Salvati, and Ranalli M. Giovanna. Correction to: Finite mixtures of quantile and M-quantile regression models. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10316-1.  See

Prajapati:2024:FMM

[1943] Deepak Prajapati, Ayan Pal, and Debasis Kundu. A finite mixture model for multiple dependent competing risks with applications of automotive warranty claims data. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10326-z.

Rudolf:2024:DIS

[1944] Daniel Rudolf and Philip Schär. Dimension-independent spectral gap of polar slice sampling. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10335-
                                        y.

Zambon:2024:EPR

[1945] Lorenzo Zambon, Dario Azzimonti, and Giorgio Corani. Efficient probabilistic reconciliation of forecasts for real-valued and count time series. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10343-y.

Barroso:2024:FDM

[1946] M. Barroso, C. M. Alaíz, J. L. Torrecilla, and A. Fernández. Functional diffusion maps. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10332-1.

Liang:2024:MVH

[1947] Wanwan Liang, Ben Wu, and Bo Zhang. Modeling volatility for high-frequency data with rounding error: a nonparametric Bayesian approach. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10341-0.

Sun:2024:BTB

[1948] Rongqian Sun and Xinyuan Song.

Bayesian tree-based heterogeneous mediation analysis with a time-to-event outcome. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10340-1.

Nieto-Barajas:2024:MRM

[1949] Luis E. Nieto-Barajas. Multivariate and regression models for directional data based on projected Pólya trees. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10337-w.

Fayomi:2024:CRP

[1950] Aisha Fayomi, Yannis Pantazis, Michail Tsagris, and Andrew T. A. Wood. Cauchy robust principal component analysis with applications to high-dimensional data sets. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10328-x.

Aushev:2024:LFI

[1951] Alexander Aushev, Thong Tran, Henri Pesonen, Andrew Howes, and Samuel Kaski. Likelihood-free inference in state-space models with unknown dynamics. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10339-8.

Grabchak:2024:RSM

[1952] Michael Grabchak and Xingnan Zhang. Representation and simulation of multivariate Dickman distributions and Vervaat perpetuities. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10349-6.

Collins:2024:BPP

[1953] Gavin Collins, Devin Francom, and Kellin Rumsey. Bayesian projection pursuit regression. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10334-
                                        z.

Loria:2024:STB

[1954] Jorge Loría and Anindya Bhadra.

SURE-tuned bridge regression. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10350-
z.

Daniel:2024:PPL

[1955] Schalk Daniel, Bischl Bernd, and Rügamer David. Privacy-preserving and lossless distributed estimation of high-dimensional generalized additive mixed models. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10323-2.

Rebafka:2024:MBC

[1956] Tabea Rebafka. Model-based clustering of multiple networks with a hierarchical algorithm. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10329-w.

Kindap:2024:PPS

[1957] Yaman Kindap and Simon Godsill. Point process simulation of generalised hyperbolic Lévy processes. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10344-x.

Dlotko:2024:TDG

[1958] Paweł Dłotko, Niklas Hellmer, Łukasz Stettner, and Rafał Topolnicki. Topologydriven goodness-of-fit tests in arbitrary dimensions. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10333-0.

Fang:2024:OEC

[1959] Guanhua Fang, Owen G. Ward, and Tian Zheng. Online estimation and community detection of network point processes for event streams. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10342-z.

Park:2024:DAD

[1960] Yeonjoo Park, Hee-Seok Oh, and Yaeji Lim. A data-adaptive dimension reduction for functional data via penalized low-rank approximation. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10348-7.

Oechsler:2024:LLM

[1961] David Oechsler. Lévy Langevin Monte Carlo. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10345-w.

Singh:2024:PED

[1962] Rakhi Singh. Pareto-efficient designs for multi- and mixed-level supersaturated designs. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10354-9.

Xue:2024:DRE

[1963] Liugen Xue. Doubly robust estimation and robust empirical likelihood in generalized linear models with missing responses. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10347-8.

Liang:2024:HAC

[1964] Chunhui Liang and Wenqing Ma.

Heterogeneous analysis for clustered data using grouped finite mixture models. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10353-w.

Wang:2024:PET

[1965] Weiwei Wang, Yuqiang Li, and Xianyi Wu. Off-policy evaluation for tabular reinforcement learning with synthetic trajectories. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10351-y.

Alfo:2024:BMD

[1966] M. Alfó, M. F. Marino, and F. Martella. Biclustering multivariate discrete longitudinal data. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10292-6.

Chen:2024:RCQ

[1967] Yan Chen, Shuixin Fang, and Lu Lin. Renewable composite quantile method and algorithm for nonparametric models with streaming data. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10352-x.

Silverman:2024:BMS

[1968] Bernard W. Silverman, Lax Chan, and Kyle Vincent. Bootstrapping multiple systems estimates to account for model selection. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10346-9.

Xuan:2024:SVI

[1969] Hanwen Xuan, Luca Maestrini, Feng Chen, and Clara Grazian. Stochastic variational inference for GARCH models. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10356-7.

Vavra:2024:CMR

[1970] Jan Vávra, Arnošt Komárek, Bettina Grün, and Gertraud Malsiner-Walli. Clusterwise multivariate regression of mixed-type panel data. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10304-
5.

Shiu:2024:RSU

[1971] Shang-Ying Shiu, Yen-Shiu Chin, Szu-Han Lin, and Ting-Li Chen. Randomized self-updating process for clustering large-scale data. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10355-
8.

Whalley:2024:RTR

[1972] Peter A. Whalley, Daniel Paulin, and Benedict Leimkuhler. Randomized time Riemannian manifold Hamiltonian Monte Carlo. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10303-
6.

Shang:2024:GEM

[1973] Laixu Shang, Qian-Zhen Zheng, Ping-Feng Xu, Na Shan, and Man-Lai Tang. A generalized expectation model selection algorithm for latent variable selection in multidimensional item response theory models. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10360-x.

Warne:2024:GLP

[1974] David J. Warne, Oliver J. Maclaren, Elliot J. Carr, Matthew J. Simpson, and Christopher Drovandi. Generalised likelihood profiles for models with intractable likelihoods. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10361-
                                        w.

Nghiem:2024:LBS

[1975] Linh H. Nghiem, Francis K. C. Hui, Samuel Muller, and A. H. Welsh. Likelihood-based surrogate dimension reduction. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10357-6.

Pi:2024:EEC

[1976] Hung-Kai Pi, Mei-Hui Guo, Ray-Bing Chen, and Shih-Feng Huang. ECOP-ICA: empirical copula-based independent component analysis. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10368-
                                        3.

Naveau:2024:BHD

[1977] Marion Naveau, Guillaume Kon Kam King, Renaud Rincent, Laure Sansonnet, and Maud Delattre. Bayesian high-dimensional covariate selection in non-linear mixed-effects models using the SAEM algorithm. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10367-

4. See correction [2037].

David:2024:IDI

[1978] Étienne David, Jean Bellot, Sylvain Le Corff, and Luc Lehéricy. Identifiability of discrete input-output hidden Markov models with external signals. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10364-7.

Godichon-Baggioni:2024:RMB

[1979] Antoine Godichon-Baggioni and Stéphane Robin. A robust model-based clustering based on the geometric median and the median covariation matrix. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10362-9.

Jang:2024:MHT

[1980] Yuhyeong Jang, Raanju R. Sundararajan, and Wagner Barreto-Souza. A multivariate heavytailed integer-valued GARCH process with EM algorithm-based inference. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10372-7.

Kallioinen:2024:DDP

[1981] Noa Kallioinen, Topi Paananen, Paul-Christian Bürkner, and Aki Vehtari. Detecting and diagnosing prior and likelihood sensitivity with powerscaling. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10366-5.

Mosching:2024:ELR

[1982] Alexandre Mösching and Lutz Dümbgen. Estimation of a likelihood ratio ordered family of distributions. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-023-10370-
9.

Yang:2024:AGM

[1983] Kai Yang, Masoud Asgharian, and Sahir Bhatnagar. Accelerated gradient methods for sparse statistical learning with nonconvex penalties. Statistics and Computing, 34 (1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10371-8.

Nielsen:2024:DBD

[1984] Frank Nielsen and Kazuki Okamura.

On the f-divergences between densities of a multivariate location or scale family. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10373-6.

Pagani:2024:NNZ

[1985] Filippo Pagani, Augustin Chevallier, Sam Power, Thomas House, and Simon Cotter. NuZZ: Numerical Zig-Zag for general models. Statistics and Computing, 34(1):??, February 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10363-8.

Brusa:2024:MLE

[1986] Luca Brusa, Fulvia Pennoni, and Francesco Bartolucci. Maximum likelihood estimation for discrete latent variable models via evolutionary algorithms. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10358-5.

Prabpon:2024:NBO

[1987] Naruesorn Prabpon, Kitakorn Homsud, and Pat Vatiwutipong. Nonparametric Bayesian online change point detection using kernel density estimation with nonparametric hazard function. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10375-4.

Come:2024:BCC

[1988] Etienne Côme. Bayesian contiguity constrained clustering. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10376-
                                        3.

Fuh:2024:EET

[1989] Cheng-Der Fuh and Chuan-Ju Wang. Efficient exponential tilting with applications. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-023-10374-5.

Foroni:2024:EHM

[1990] Beatrice Foroni, Luca Merlo, and Lea Petrella. Expectile hidden Markov regression models for analyzing cryptocurrency returns. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-023-10377-
                                        2.

Xian:2024:VBA

[1991] Chengqian Xian, Camila P. E. de Souza, Wenqing He, Felipe F. Rodrigues, and Renfang Tian. Variational Bayesian analysis of survival data using a log-logistic accelerated failure time model. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10365-6.

Mosley:2024:SDF

[1992] Luke Mosley, Tak-Shing T. Chan, and Alex Gibberd. The sparse dynamic factor model: a regularised quasi-maximum likelihood approach. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10378-1.

Lu:2024:KFE

[1993] Shan Lu, Wenjing Wang, and Rong Guan. Kent feature embedding for classification of compositional data with zeros. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10382-z.

Wu:2024:GLS

[1994] Hongyu Wu and Jonathan R. Bradley.

Global-local shrinkage multivariate logit-beta priors for multiple responsetype data. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10380-1.

Echegoyen:2024:LSU

[1995] Carlos Echegoyen, Aritz Pérez, Guzmán Santafé, Unai Pérez-Goya, and María Dolores Ugarte. Large-scale unsupervised spatio-temporal semantic analysis of vast regions from satellite images sequences. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10383-y.

Levin:2024:FGE

[1996] Keith Levin and Brenda Betancourt. Fast generation of exchangeable sequences of clusters data. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10385-w.

Ameijeiras-Alonso:2024:RDB

[1997] Jose Ameijeiras-Alonso. A reliable data-based smoothing parameter selection method for circular kernel estimation. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10384-x.

Castro:2024:RQR

[1998] Matheus Castro, Caio Azevedo, and Juvêncio Nobre. A robust quantile regression for bounded variables based on the Kumaraswamy rectangular distribution. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10381-0.

Moka:2024:CBS

[1999] Sarat Moka, Benoit Liquet, Houying Zhu, and Samuel Muller. COMBSS: best subset selection via continuous optimization. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10387-8.

Mahdi:2024:NMP

[2000] Esam Mahdi. New mixed portmanteau tests for time series models. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10393-w.

Yang:2024:SAL

[2001] Zehan Yang, HaiYing Wang, and Jun Yan. Subsampling approach for least squares fitting of semi-parametric accelerated failure time models to massive survival data. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10391-y.

Jasra:2024:BPI

[2002] Ajay Jasra, Hamza Ruzayqat, and Amin Wu. Bayesian parameter inference for partially observed stochastic Volterra equations. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10389-6.

Xue:2024:EEN

[2003] Yushan Xue, Jie Ren, and Bin Yang. Enmsp: an elastic-net multistep screening procedure for highdimensional regression. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10394-
                                        9.

Tian:2024:MCD

[2004] Wan Tian and Zhongfeng Qin. The minimum covariance determinant estimator for interval-valued data. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10386-9.

Amato:2024:CLO

[2005] Francesco Amato, Julien Jacques, and Isabelle Prim-Allaz. Clustering longitudinal ordinal data via finite mixture of matrix-variate distributions. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10390-z.

Kissel:2024:FSM

[2006] Nicholas Kissel and Lucas Mentch. Forward stability and model path selection. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10395-8.

Yatracos:2024:DAS

[2007] Yannis G. Yatracos. Do applied statisticians prefer more randomness or less? Bootstrap or jackknife? Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10388-7.

Takazawa:2024:MLE

[2008] Yuki Takazawa and Tomonari Sei.

Maximum likelihood estimation of logconcave densities on tree space. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10400-
0.

Jiang:2024:ISS

[2009] Alex Ziyu Jiang and Abel Rodriguez. Improvements on scalable stochastic Bayesian inference methods for multivariate Hawkes process. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10392-x.

Chen:2024:HDS

[2010] Xin Chen, Chang Deng, Shuaida He, Runxiong Wu, and Jia Zhang. Highdimensional sparse single-index regression via Hilbert–Schmidt independence criterion. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10399-4.

Borgonovo:2024:TEC

[2011] Emanuele Borgonovo, Elmar Plischke, and Clémentine Prieur. Total effects with constrained features. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10398-5.

Lux:2024:ERS

[2012] Thomas Lux. Estimation of regimeswitching diffusions via Fourier transforms. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10397-6.

Hao:2024:MOQ

[2013] Ruiting Hao and Xiaorong Yang. Multiple-output quantile regression neural network. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10408-6.

Lbath:2024:LSC

[2014] Hanâ Lbath, Alexander Petersen, and Sophie Achard. Large-scale correlation screening under dependence for brain functional connectivity network inference. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10411-x.

Celani:2024:BVS

[2015] Alessandro Celani, Paolo Pagnottoni, and Galin Jones. Bayesian variable selection for matrix autoregressive models. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10402-y.

Zhang:2024:QGM

[2016] Xinyu Zhang, Hongwei Shi, Niwen Zhou, Falong Tan, and Xu Guo. Quantile generalized measures of correlation. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10414-8.

Scealy:2024:RSM

[2017] Janice L. Scealy, Kassel L. Hingee, John T. Kent, and Andrew T. A. Wood. Robust score matching for compositional data. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10412-w.

Farcomeni:2024:QRR

[2018] Alessio Farcomeni and Marco Geraci.

Quantile ratio regression. Statistics and Computing, 34(2):??, April 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10406-
                                        8.

Zhang:2024:MRH

[2019] Fengchuan Zhang, Sanguo Zhang, Shi-Ming Li, and Mingyang Ren. Matrix regression heterogeneity analysis. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10401-z.

Junwen:2024:DRE

[2020] Xia Junwen, Zhan Zishu, and Zhang Jingxiao. Doubly robust estimation of optimal treatment regimes for survival data using an instrumental variable. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10407-7.

Danilevicz:2024:EQR

[2021] Ian Meneghel Danilevicz, Valdério Anselmo Reisen, and Pascal Bondon. Expectile and M-quantile regression for panel data. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10396-7.

Chamroukhi:2024:FME

[2022] Faïcel Chamroukhi, Nhat Thien Pham, Van Hà Hoang, and Geoffrey J. McLachlan. Functional mixturesof-experts. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10379-0.

Ward:2024:CPI

[2023] Kes Ward, Gaetano Romano, Idris Eckley, and Paul Fearnhead. A constant-per-iteration likelihood ratio test for online changepoint detection for exponential family models. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10416-6.

Rodriguez-Vitores:2024:IMC

[2024] David Rodríguez-Vítores and Carlos Matrán. Improving model choice in classification: an approach based on clustering of covariance matrices. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10410-y.

Ruhl:2024:RBC

[2025] Jasmin Rühl and Sarah Friedrich.

Resampling-based confidence intervals and bands for the average treatment effect in observational studies with competing risks. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10420-w.

Shin:2024:SEV

[2026] Jungmin Shin, Seunghyun Gwak, Seung Jun Shin, and Sungwan Bang. Simultaneous estimation and variable selection for a non-crossing multiple quantile regression using deep neural networks. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10418-4.

Daouia:2024:ECC

[2027] Abdelaati Daouia, Gilles Stupfler, and Antoine Usseglio-Carleve. An expectile computation cookbook. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10403-
                                        x.

Leyder:2024:GSP

[2028] Sarah Leyder, Jakob Raymaekers, and Tim Verdonck. Generalized spherical principal component analysis. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10413-9.

Lin:2024:VSU

[2029] Youwu Lin, Xin Zeng, Pei Wang, Shuai Huang, and Kok Lay Teo. Variable selection using axis-aligned random projections for partial least-squares regression. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10417-5.

Kang:2024:NSM

[2030] Seungwoo Kang and Hee-Seok Oh.

Novel sampling method for the von Mises–Fisher distribution. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10419-
3.

Ouyang:2024:STT

[2031] Chen Ouyang, Chenkaixiang Lu, Xiong Zhao, Ruping Huang, Gonglin Yuan, and Yiyan Jiang. Stochastic threeterm conjugate gradient method with variance technique for non-convex learning. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10409-5.

Cavicchia:2024:PUG

[2032] Carlo Cavicchia, Maurizio Vichi, and Giorgia Zaccaria. Parsimonious ultrametric Gaussian mixture models. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10405-9.

James:2024:AGI

[2033] Jonathan James. Automated generation of initial points for adaptive rejection sampling of log-concave distributions. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10425-5.

Vu:2024:RVS

[2034] Bao Anh Vu, David Gunawan, and Andrew Zammit-Mangion. R-VGAL: a sequential variational Bayes algorithm for generalised linear mixed models. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10422-8. See cor-

rection [2093].

Rotiroti:2024:RPF

[2035] Frank Rotiroti and Stephen G. Walker. Reversed particle filtering for hidden Markov models. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10426-4.

Jiang:2024:STS

[2036] Xuejun Jiang, Yakun Liang, and Haofeng Wang. Screen then select: a strategy for correlated predictors in high-dimensional quantile regression. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10424-6.

Naveau:2024:CBH

[2037] Marion Naveau, Guillaume Kon Kam King, Renaud Rincent, Laure Sansonnet, and Maud Delattre. Correction to: Bayesian high-dimensional covariate selection in non-linear mixedeffects models using the SAEM algorithm. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10421-9.  See

[1977].

Bombelli:2024:PCH

[2038] Ilaria Bombelli and Maurizio Vichi.

Parsimonious consensus hierarchies, partitions and fuzzy partitioning of a set of hierarchies. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10415-7.

Yang:2024:CEO

[2039] Ziyang Yang, Idris A. Eckley, and Paul Fearnhead. A communicationefficient, online changepoint detection method for monitoring distributed sensor networks. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10428-2.

Martinez-Minaya:2024:FBT

[2040] Joaquín Martínez-Minaya and Haavard Rue. A flexible Bayesian tool for CoDa mixed models: logisticnormal distribution with Dirichlet covariance. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10427-3.

Choi:2024:GMC

[2041] Dongrak Choi, Woojung Bae, Jun Yan, and Sangwook Kang. A general model-checking procedure for semiparametric accelerated failure time models. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10431-7.

Ning:2024:SSB

[2042] Yu-Chien Bo Ning and Ning Ning. Spike and slab Bayesian sparse principal component analysis. Statistics and Computing, 34(3):??, June 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10430-
                                        8.

Cooper:2024:BCV

[2043] Alex Cooper, Aki Vehtari, Catherine Forbes, Dan Simpson, and Lauren Kennedy. Bayesian cross-validation by parallel Markov chain Monte Carlo. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10404-w.

Pastukhov:2024:FLN

[2044] Vladimir Pastukhov. Fused lasso nearly-isotonic signal approximation in general dimensions. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10432-
6.

Templ:2024:RMI

[2045] Matthias Templ. Robust multiple imputation with GAM. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10429-
1.

Tan:2024:GSS

[2046] Zuoxun Tan and Hu Yang. Group sparse structural smoothing recovery: model, statistical properties and algorithm. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10438-0.

ONeill:2024:TTB

[2047] Eoghan O’Neill. Type I Tobit Bayesian additive regression trees for censored outcome regression. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10434-4.

Ohishi:2024:GFL

[2048] Mineaki Ohishi. Generalized fused lasso for grouped data in generalized linear models. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10433-5.

Skhosana:2024:MET

[2049] Sphiwe B. Skhosana, Salomon M. Millard, and Frans H. J. Kanfer. A modified EM-type algorithm to estimate semi-parametric mixtures of non-parametric regressions. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10435-3.

Luciano:2024:IGS

[2050] Antoine Luciano, Christian P. Robert, and Robin J. Ryder. Insufficient Gibbs sampling. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10423-7.

Cubadda:2024:OGC

[2051] Gianluca Cubadda, Francesco Giancaterini, Alain Hecq, and Joann Jasiak. Optimization of the generalized covariance estimator in noncausal processes. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10437-1.

Khan:2024:TGF

[2052] Ruhul Ali Khan, Ayan Pal, and Debasis Kundu. Testing the goodnessof-fit of the stable distributions with applications to German stock index data and bitcoin cryptocurrency data. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10441-5.

Wiens:2024:JCS

[2053] Douglas P. Wiens. Jittering and clustering: strategies for the construction of robust designs. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10436-2.

Daouia:2024:BRV

[2054] Abdelaati Daouia, Gilles Stupfler, and Antoine Usseglio-Carleve. Biasreduced and variance-corrected asymptotic Gaussian inference about extreme expectiles. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-023-10359-4.

Chau:2024:EMS

[2055] H. Chau, J. L. Kirkby, D. H. Nguyen, D. Nguyen, N. Nguyen, and T. Nguyen. An efficient method to simulate diffusion bridges. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10439-z.

McLatchie:2024:EEC

[2056] Yann McLatchie and Aki Vehtari. Efficient estimation and correction of selection-induced bias with order statistics. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10442-4.

Li:2024:PSB

[2057] Xiaoyan Li, Xiaochao Xia, and Zhimin Zhang. Poisson subsampling-based estimation for growing-dimensional expectile regression in massive data. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https://link.

                                        springer.com/article/10.1007/s11222-
                                        024-10449-x.

Flock:2024:CCS

[2058] Rafael Flock, Yiqiu Dong, Felipe Uribe, and Olivier Zahm. Certified coordinate selection for highdimensional Bayesian inversion with Laplace prior. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10445-1.

Sportisse:2024:MBC

[2059] Aude Sportisse, Matthieu Marbac, Fabien Laporte, Gilles Celeux, Claire Boyer, Julie Josse, and Christophe Biernacki. Model-based clustering with missing not at random data. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10444-2.

Lehmann:2024:BMM

[2060] Brieuc Lehmann and Simon White. A Bayesian multilevel model for populations of networks using exponentialfamily random graphs. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10446-
0.

Vandeskog:2024:EWM

[2061] Silius M. Vandeskog, Sara Martino, and Raphaël Huser. An efficient workflow for modelling high-dimensional spatial extremes. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10448-y.

Zietkiewicz:2024:BMA

[2062] Patrick Zietkiewicz and Ioannis Kosmidis. Bounded-memory adjusted scores estimation in generalized linear models with large data sets. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10447-
z.

Bai:2024:GPB

[2063] Tianming Bai, Aretha L. Teckentrup, and Konstantinos C. Zygalakis. Gaussian processes for Bayesian inverse problems associated with linear partial differential equations. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10452-2.

Abergel:2024:RAR

[2064] Rémy Abergel, Olivier Bouaziz, and Grégory Nuel. A review on the Adaptive-Ridge Algorithm with several extensions. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10440-6.

Gallardo:2024:BSF

[2065] Diego I. Gallardo, Marcelo Bourguignon, and José S. Romeo. Birnbaum– Saunders frailty regression models for clustered survival data. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10458-
w.

Aselisewine:2024:ECR

[2066] Wisdom Aselisewine and Suvra Pal. Enhancing cure rate analysis through integration of machine learning models: a comparative study. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10456-
y.

Heaps:2024:SPD

[2067] Sarah Elizabeth Heaps and Ian Hyla Jermyn. Structured prior distributions for the covariance matrix in latent factor models. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10454-0.

Hu:2024:FBQ

[2068] Weitao Hu and Weiping Zhang.

Flexible Bayesian quantile regression based on the generalized asymmetric Huberised-type distribution. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10453-
1.

Cardoso:2024:GRS

[2069] Douglas O. Cardoso, João Domingos Gomes da Silva Junior, Carla Silva Oliveira, Celso Marques, and Laura Silva de Assis. Greedy recursive spectral bisection for modularity-bound hierarchical divisive community detection. Statistics and Computing, 34(4):??, August 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10451-3.

Latz:2024:CAS

[2070] Jonas Latz. Correction to: Analysis of stochastic gradient descent in continuous time. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10450-4.  See

[1635].

Ning:2024:SID

[2071] Ning Ning and Edward Ionides. Systemic infinitesimal over-dispersion on graphical dynamic models. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10443-
                                        3.

Jiang:2024:WCQ

[2072] Hu Jiang, Liu Yiming, and Zhou Wang. On weak convergence of quantilebased empirical likelihood process for ROC curves. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10457-x.

Bell:2024:ESP

[2073] Logan Bell, Nikhil Devanathan, and Stephen Boyd. Efficient Shapley performance attribution for least-squares regression. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10459-9.

Gurer:2024:LSR

[2074] Sude Gurer, Han Lin Shang, Abhijit Mandal, and Ufuk Beyaztas. Locally sparse and robust partial least squares in scalar-on-function regression. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10464-y.

Camattari:2024:CDF

[2075] Fabiana Camattari, Sabrina Guastavino, Francesco Marchetti, Michele Piana, and Emma Perracchione. Classifier-dependent feature selection via greedy methods. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10460-
2.

Au:2024:LFS

[2076] Siu-Kui Au. A limit formula and a series expansion for the bivariate normal tail probability. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10466-w.

Ortega-Fernandez:2024:CEG

[2077] Ines Ortega-Fernandez, Marta Sestelo, and Nora M. Villanueva. Correction to: Explainable generalized additive neural networks with independent neural network training. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10461-1.  See

[1930].

Tchomgui:2024:MER

[2078] Jean Steve Tamo Tchomgui, Julien Jacques, Guillaume Fraysse, Vincent Barriac, and Stéphane Chretien. A mixture of experts regression model for functional response with functional covariates. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10455-z.

Ohl:2024:SGA

[2079] Louis Ohl, Pierre-Alexandre Mattei, Charles Bouveyron, Mickaël Leclercq, Arnaud Droit, and Frédéric Precioso. Sparse and geometry-aware generalisation of the mutual information for joint discriminative clustering and feature selection. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10467-9.

Cui:2024:ODN

[2080] Elvis Han Cui, Zizhao Zhang, and Weng Kee Wong. Optimal designs for nonlinear mixed-effects models using competitive swarm optimizer with mutated agents. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10468-8.

Huang:2024:FNA

[2081] Yifei Huang, Keren Li, Abhyuday Mandal, and Jie Yang. ForLion: a new algorithm for D-optimal designs under general parametric statistical models with mixed factors. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10465-x.

Chen:2024:BRE

[2082] Canyi Chen and Zhengtian Zhu.

Byzantine-robust and efficient distributed sparsity learning: a surrogate composite quantile regression approach. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10470-0.

Zeng:2024:MTM

[2083] Jie Zeng, Guozhi Hu, and Weihu Cheng. A Mallows-type model averaging estimator for ridge regression with randomly right censored data. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10472-y.

Lee:2024:HDM

[2084] Yoonah Lee and Seongoh Park.

High-dimensional missing data imputation via undirected graphical model. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10475-9.

Li:2024:DSM

[2085] Xiaoyan Li, Xiaochao Xia, and Zhimin Zhang. Distributed subsampling for multiplicative regression. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10477-7.

Hollaway:2024:DSC

[2086] Michael J. Hollaway and Rebecca Killick. Detection of spatiotemporal changepoints: a generalised additive model approach. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10478-
                                        6.

Guo:2024:CCO

[2087] Guangbao Guo, Haoyue Song, and Lixing Zhu. The COR criterion for optimal subset selection in distributed estimation. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10471-z. See cor-

rection [2103].

Hirt:2024:LVA

[2088] Marcel Hirt, Vasileios Kreouzis, and Petros Dellaportas. Learning variational autoencoders via MCMC speed measures. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10481-x.

DeMagistris:2024:RRF

[2089] Anna De Magistris, Valentina De Simone, Elvira Romano, and Gerardo Toraldo. Roughness regularization for functional data analysis with free knots spline estimation. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10474-w.

Park:2024:AAA

[2090] Hyejoon Park and Hyunjoong Kim.

AR-ADASYN: angle radius-adaptive synthetic data generation approach for imbalanced learning. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10479-
5.

Poot:2024:BAM

[2091] Anne Poot, Pierre Kerfriden, Iuri Rocha, and Frans van der Meer. A Bayesian approach to modeling finite element discretization error. Statistics and Computing, 34 (5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10463-z.

Pfahler:2024:TNI

[2092] Simon Pfahler, Peter Georg, Rudolf Schill, Maren Klever, Lars Grasedyck, Rainer Spang, and Tilo Wettig. Taming numerical imprecision by adapting the KL divergence to negative probabilities. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10480-y.

Vu:2024:CRV

[2093] Bao Anh Vu, David Gunawan, and Andrew Zammit-Mangion. Correction to: R-VGAL: a sequential variational Bayes algorithm for generalised linear mixed models. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10469-7.  See

[2034].

Huan:2024:ICM

[2094] Cheng Huan, Xinyuan Song, and Hongwei Yuan. Individualized causal mediation analysis with continuous treatment using conditional generative adversarial networks. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10484-
8.

Beraha:2024:WPC

[2095] Mario Beraha and Matteo Pegoraro. Wasserstein principal component analysis for circular measures. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10473-x.

Zhou:2024:NMM

[2096] Bu Zhou, Zhi Peng Ong, and Jin-Ting Zhang. A new maximum mean discrepancy based two-sample test for equal distributions in separable metric spaces. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10483-9.

Helgoy:2024:SBL

[2097] Ingvild M. Helgøy, Hans J. Skaug, and Yushu Li. Sparse Bayesian learning using TMB (Template Model Builder). Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10476-8.

Stockman:2024:SEU

[2098] Samuel Stockman, Daniel J. Lawson, and Maximilian J. Werner. SB-ETAS: using simulation based inference for scalable, likelihood-free inference for the ETAS model of earthquake occurrences. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10486-6.

Liu:2024:CCG

[2099] Huiling Liu, Xinmin Li, Feifei Chen, Wolfgang Härdle, and Hua Liang. A comprehensive comparison of goodness-of-fit tests for logistic regression models. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10487-5.

Dai:2024:NFB

[2100] Shuang Dai, Ping Wu, and Zhou Yu. New forest-based approaches for sufficient dimension reduction. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10482-
w.

Peer:2024:OCI

[2101] Almog Peer and David Azriel. Optimal confidence interval for the difference between proportions. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10485-
7.

Roy:2024:RSV

[2102] Subhrajyoty Roy, Abhik Ghosh, and Ayanendranath Basu. Robust singular value decomposition with application to video surveillance background modelling. Statistics and Computing, 34(5):??, October 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10493-7.

Guo:2024:CCC

[2103] Guangbao Guo, Haoyue Song, and Lixing Zhu. Correction to: The COR criterion for optimal subset selection in distributed estimation. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10496-4.  See

[2087].

Wen:2024:NDS

[2104] Jiawei Wen, Songshan Yang, and Delin Zhao. Nonconvex Dantzig selector and its parallel computing algorithm. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10492-8.

Arbel:2024:SEP

[2105] Julyan Arbel, Stéphane Girard, and Hadrien Lorenzo. Shrinkage for extreme partial least-squares. Statistics and Computing, 34(6): ??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10490-w.

Neal:2024:HMM

[2106] Mackenzie R. Neal, Alexa A. Sochaniwsky, and Paul D. McNicholas. Hidden Markov models for multivariate panel data. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10462-0.

Chen:2024:AFT

[2107] Li-Pang Chen. Accelerated failure time models with error-prone response and nonlinear covariates. Statistics and Computing, 34(6): ??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10491-9.

Huan:2024:SMI

[2108] Yue Huan and Hai Xiang Lin. Sequential model identification with reversible jump ensemble data assimilation method. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10499-1.

Zhang:2024:NGH

[2109] Weiben Zhang, Michael Smith, Worapree Maneesoonthorn, and Rubén Loaiza-Maya. Natural gradient hybrid variational inference with application to deep mixed models. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10488-
4.

Xie:2024:ITP

[2110] Tianyu Xie, Musu Yuan, Minghua Deng, and Cheng Zhang. Improving tree probability estimation with stochastic optimization and variance reduction. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10498-2.

Yan:2024:DHT

[2111] Lu Yan, Jiang Hu, and Lixiu Wu. Distributed hypothesis testing for large dimensional two-sample mean vectors. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10489-3.

Wang:2024:SVM

[2112] Kangning Wang, Jin Liu, and Xiaofei Sun. Support vector machine in big data: smoothing strategy and adaptive distributed inference. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10506-5.

Wu:2024:IAE

[2113] Qiying Wu, Huiwen Wang, and Shan Lu. IDGM: an approach to estimate the graphical model of intervalvalued data. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10504-7.

Li:2024:CEL

[2114] Mengyu Li, Jun Yu, Tao Li, and Cheng Meng. Core-elements for large-scale least squares estimation. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10505-6.

Riccio:2024:SLE

[2115] Donato Riccio, Fabrizio Maturo, and Elvira Romano. Supervised learning via ensembles of diverse functional representations: the functional voting classifier. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10503-8.

Fan:2024:ITM

[2116] Guoliang Fan, Xilin Zhang, and Liping Zhu. Independence test via mutual information in the presence of measurement errors. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10502-9.

Yuan:2024:ASS

[2117] Zihao Yuan, Jiaqing Chen, Han Qiu, Houxiang Wang, and Yangxin Huang. Adaptive sufficient sparse clustering by controlling false discovery. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10507-4.

Li:2024:DNN

[2118] Jie Li, Yunquan Song, and Ling Jian. Deep neural networks for variable selection of higher-order nonparametric spatial autoregressive model. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10500-x.

Baya:2024:CVD

[2119] Ariel E. Bayá and Mónica G. Larese.

Clustering validation by distribution hypothesis learning. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10511-
8.

Tran:2024:TDS

[2120] Jimmy Huy Tran and Tore Selland Kleppe. Tuning diagonal scale matrices for HMC. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10494-6.

Rached:2024:DLI

[2121] Nadhir Ben Rached, Abdul-Lateef Haji-Ali, Shyam Mohan Subbiah Pillai, and Raúl Tempone. Doubleloop importance sampling for McKean– Vlasov stochastic differential equation. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10497-3.

Alvarez:2024:UMM

[2122] Miguel Alvarez, Ajay Jasra, and Hamza Ruzayqat. Unbiased and multilevel methods for a class of diffusions partially observed via marked point processes. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10518-1.

Bohyn:2024:OOD

[2123] Alexandre Bohyn, Eric D. Schoen, and Peter Goos. Optimal orthogonal designs for experiments with four-level and two-level factors. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10517-2.

Du:2024:EEE

[2124] Xinting Du and Hejin Wang. Efficient estimation of expected information gain in Bayesian experimental design with multi-index Monte Carlo. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10522-5.

Boente:2024:RVS

[2125] Graciela Boente and Alejandra Mercedes Martínez. Robust variable selection for partially linear additive models. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10520-7.

BenRached:2024:FAN

[2126] Nadhir Ben Rached, Håkon Hoel, and Johannes Vincent Meo. A fast and accurate numerical method for the left tail of sums of independent random variables. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10514-5.

Wu:2024:FAS

[2127] Sidi Wu, Cédric Beaulac, and Jiguo Cao. Functional autoencoder for smoothing and representation learning. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10501-w.

Li:2024:OPE

[2128] Zhengpin Li, Jian Wang, and Yanxi Hou. Online prediction of extreme conditional quantiles via B-spline interpolation. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10515-4.

Su:2024:SGC

[2129] Han Su, Qingyang Sun, Mengxi Yi, Gaorong Li, and Panxu Yuan. Stab-GKnock: controlled variable selection for partially linear models using generalized knockoffs. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10516-
3.

Kang:2024:FBM

[2130] Jizhou Kang and Athanasios Kottas. Flexible Bayesian modeling for longitudinal binary and ordinal responses. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10525-2.

James:2024:IAG

[2131] Zachary James and Joseph Guinness. Implementation and analysis of GPU algorithms for Vecchia Approximation. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10510-9.

Huang:2024:DAH

[2132] Peiyao Huang, Shuwei Li, and Xinyuan Song. Distributed additive hazards regression analysis of multi-site current status data without using individuallevel data. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10523-4.

Ranciati:2024:AGG

[2133] Saverio Ranciati and Alberto Roverato. On the application of Gaussian graphical models to paired data problems. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10513-6.

Jiang:2024:SPD

[2134] Haoyu Jiang and Jason Xu. The stochastic proximal distance algorithm. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10524-3.

Harman:2024:POA

[2135] Radoslav Harman, Lenka Filová, and Samuel Rosa. The polytope of optimal approximate designs: extending the selection of informative experiments. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10527-0.

deFreitas:2024:SMM

[2136] João Victor B. de Freitas, Pascal Bondon, Caio L. N. Azevedo, Valdério A. Reisen, and Juvêncio S. Nobre. Scale mixtures of multivariate centered skew-normal distributions. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10512-7.

Zhu:2024:CAH

[2137] Wanchuang Zhu and Ngoc Lan Chi Nguyen. Continuity approximation in hybrid Bayesian networks structure learning. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10531-4.

DUrso:2024:FCB

[2138] Pierpaolo D’Urso, Livia De Giovanni, Lorenzo Federico, and Vincenzina Vitale. Fuzzy clustering with Barber modularity regularization. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10495-5.

Liang:2024:BNA

[2139] Mingrui Liang, Matthew D. Koslovsky, Emily T. Hébert, Darla E. Kendzor, and Marina Vannucci. A Bayesian nonparametric approach for clustering functional trajectories over time. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10521-6.

Alvarez:2024:CUM

[2140] Miguel Alvarez, Ajay Jasra, and Hamza Ruzayqat. Correction: Unbiased and multilevel methods for a class of diffusions partially observed via marked point processes. Statistics and Computing, 34(6): ??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10530-5.

Beckman:2024:FAF

[2141] Paul G. Beckman and Christopher J.

Geoga. Fast adaptive Fourier integration for spectral densities of Gaussian processes. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10519-0.

Tzoumerkas:2024:SPR

[2142] G. Tzoumerkas and D. Fouskakis.

Shrinkage priors via random imaginary data. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10509-2.

Bahceci:2024:KLD

[2143] Ufuk Bahçeci. Kullback–Leibler divergence based multidimensional robust universal hypothesis testing. Statistics and Computing, 34 (6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10533-2.

Bibi:2024:PME

[2144] Abdelouahab Bibi and Fayçal Hamdi.

On periodic log GARCH model with empirical application. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10532-
3.

Yan:2024:NAP

[2145] Tianyu Yan, Kai-Tai Fang, and Hong Yin. A novel approach for parameter estimation of mixture of two Weibull distributions in failure data modeling. Statistics and Computing, 34(6):??, December 2024. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10534-1.

Rached:2025:MIS

[2146] Nadhir Ben Rached, Abdul-Lateef Haji-Ali, Shyam Mohan Subbiah Pillai, and Raúl Tempone. Multilevel importance sampling for rare events associated with the McKean–Vlasov equation. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10508-3.

Zhao:2025:BFP

[2147] Yawen Zhao, Mingzhe Zhang, Chenhao Zhang, Weitong Chen, Nan Ye, and Miao Xu. A boosting framework for positive-unlabeled learning. Statistics and Computing, 35 (1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10529-y.

Heng:2025:BEP

[2148] Qiang Heng and Kenneth Lange.

Bootstrap estimation of the proportion of outliers in robust regression. Statistics and Computing, 35 (1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10526-1.

Stringer:2025:EGE

[2149] Alex Stringer. Exact gradient evaluation for adaptive quadrature approximate marginal likelihood in mixed models for grouped data. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10536-
z.

Turinici:2025:HEM

[2150] Gabriel Turinici. Huber-energy measure quantization. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-024-10540-
3.

Bassett:2025:ESS

[2151] Robert L. Bassett and Micah Y.

Oh. Estimating a signal subspace in the presence of impulsive noise. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10528-z.

Maatouk:2025:LSC

[2152] Hassan Maatouk, Didier Rullière, and Xavier Bay. Large-scale constrained Gaussian processes for shaperestricted function estimation. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10541-
                                        2.

Carrington:2025:IPC

[2153] Rachel Carrington and Paul Fearnhead. Improving power by conditioning on less in post-selection inference for changepoints. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10542-1.

Robert:2025:SSM

[2154] Christian P. Robert and Julien Stoehr. Simulating signed mixtures. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10539-w.

Chen:2025:GMC

[2155] Yifan Chen, Bamdad Hosseini, Houman Owhadi, and Andrew M. Stuart. Gaussian measures conditioned on nonlinear observations: consistency, MAP estimators, and simulation. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10535-0.

Iorio:2025:FHA

[2156] Jacopo Di Iorio, Marzia A. Cremona, and Francesca Chiaromonte. funBIalign: a hierarchical algorithm for functional motif discovery based on mean squared residue scores. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10537-y.

Bartuska:2025:LBS

[2157] Arved Bartuska, Luis Espath, and Raúl Tempone. Laplace-based strategies for Bayesian optimal experimental design with nuisance uncertainty. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10544-z.

Cheng:2025:ORE

[2158] Guanghui Cheng, Wenjuan Hu, Ruitao Lin, and Chen Wang. Online robust estimation and bootstrap inference for function-on-scalar regression. Statistics and Computing, 35 (1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10538-x.

He:2025:RPS

[2159] Sijin He and Xiaochao Xia. Random perturbation subsampling for rank regression with massive data. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10548-
                                        9.

Li:2025:OSG

[2160] Lili Li, Bingfan Liu, Xiaodi Liu, Haolun Shi, and Jiguo Cao. Optimal subsampling for generalized additive models on large-scale datasets. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10546-
                                        x.

Altmann:2025:PTI

[2161] R. Altmann and A. Moradi. Probabilistic time integration for semi-explicit PDAEs. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10543-0.

Fattah:2025:IAB

[2162] Esmail Abdul Fattah, Janet Van Niekerk, and Håvard Rue. INLA^{+}: approximate Bayesian inference for nonsparse models using HPC. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10545-
                                        y.

Liang:2025:FCV

[2163] Zhongqi Liang, Li Cai, Suojin Wang, and Qihua Wang. Kfold cross-validation based frequentist model averaging for linear models with nonignorable missing responses. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10554-x.

Piancastelli:2025:MZI

[2164] Luiza S. C. Piancastelli and Rodrigo B. Silva. Multivariate zeroinflated INGARCH models: Bayesian inference and composite likelihood approach. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10549-8.

Au:2025:LFR

[2165] Siu-Kui Au. A limit formula and recursive algorithm for multivariate normal tail probability. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10552-z.

Piancastelli:2025:CMM

[2166] Luiza S. C. Piancastelli and Nial Friel. The clustered Mallows model. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10555-w.

Kim:2025:NNB

[2167] Minwoo Kim, Marc G. Genton, Raphaël Huser, and Stefano Castruccio. A neural network-based adaptive cut-off approach to normality testing for dependent data. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-024-10551-
                                        0.

Okazaki:2025:MTL

[2168] Akira Okazaki and Shuichi Kawano. Multi-task learning via robust regularized clustering with non-convex group penalties. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10550-1.

Qiao:2025:RES

[2169] Nan Qiao, Canyi Chen, and Zhengtian Zhu. Robust and efficient sparse learning over networks: a decentralized surrogate composite quantile regression approach. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10547-w.

Anonymous:2025:JNa

[2170] Anonymous. Journal navigation. Statistics and Computing, 35(1):??, February 2025. CODEN STACE3.

ISSN 0960-3174 (print), 1573-1375 (electronic).

Liu:2025:PEL

[2171] Yang Liu, Pengfei Li, and Yukun Liu. Penalized empirical likelihood estimation and EM algorithms for closed-population capture-recapture models. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10557-8.

Shao:2025:ODS

[2172] Yujing Shao, Lei Wang, and Heng Lian.

Optimal distributed subsampling under heterogeneity. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10558-7.

Tsagris:2025:CLS

[2173] Michail Tsagris. Constrained least squares simplicial-simplicial regression. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10560-z.

Gao:2025:PEP

[2174] Junzhuo Gao, Lei Wang, and Haiying Wang. Power enhancing probability subsampling using side information. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10556-9.

Hinz:2025:AMF

[2175] Florian B. Hinz, Amr H. Mahmoud, and Markus A. Lill. An analysis of the modality and flexibility of the inverse stereographic normal distribution. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10563-4.

Hong:2025:IIM

[2176] Zhimin Hong, Ruoxuan Wang, Zhiwen Wang, and Wala Du. Inference issue in multiscale geographically and temporally weighted regression. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10559-6.

Qin:2025:PUE

[2177] Shanshan Qin, Zhenni Tan, Weidong Wei, and Yuehua Wu. PCA-uCPD: an ensemble method for multiple changepoint detection in moderately highdimensional data. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-024-10553-y. See cor-

rection [2224].

Zhang:2025:EMQ

[2178] Ziang Zhang, Patrick Brown, and Jamie Stafford. Efficient modeling of quasi-periodic data with seasonal Gaussian process. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10565-2.

Chen:2025:GWQ

[2179] Vivian Yi-Ju Chen and Shi-Ting Wang.

Geographically weighted quantile regression for count data. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-025-10568-
z.

Hou:2025:PGL

[2180] Zhaoran Hou and Samuel W. K.

Wong. Particle Gibbs for likelihoodfree inference of stochastic volatility models. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10571-4.

Zhong:2025:SIG

[2181] Chen Zhong. Statistical inference and goodness-of-fit test in functional data via error distribution function. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10574-1.

Guglielmini:2025:APS

[2182] Sofia Guglielmini and Gerda Claeskens.

Asymptotic post-selection inference for regularized graphical models. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-025-10564-
3.

Wang:2025:RTI

[2183] Yidan Wang, Lingyun Zhang, and Yujie Gai. Real-time inference for smoothing quantile regression on streaming datasets with heterogeneity detection. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10572-3.

Yuan:2025:TCD

[2184] Mingao Yuan and Qianqian Yao. Testing common degree-correction parameters of multilayer networks. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-025-10576-
z.

Lachos:2025:EAF

[2185] Victor H. Lachos, Salvatore D. Tomarchio, Antonio Punzo, and Salvatore Ingrassia. An EM algorithm for fitting matrix-variate normal distributions on interval-censored and missing data. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10575-0.

Bian:2025:EIB

[2186] Yuan Bian, Grace Y. Yi, and Wenqing He. Empirical investigations of boosting with pseudooutcome imputation for missing responses. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-024-10561-y.

Lei:2025:LUC

[2187] Rayleigh Lei and Abel Rodriguez.

Logit unfolding choice models for binary data. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10570-5.

Jia:2025:FFR

[2188] Shifan Jia, Haolun Shi, and Tianyu Guan. Function-on-function regression models with nonlinear dynamic effect and linear concurrent effect. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10578-x.

Bach:2025:AMS

[2189] Paul Bach and Nadja Klein. Anisotropic multidimensional smoothing using Bayesian tensor product P-splines. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https://link.

                                        springer.com/article/10.1007/s11222-
                                        025-10569-y.

Shi:2025:EAD

[2190] Haosheng Shi and Wenlin Dai. Exploratory analysis of dynamic networks using latent functions. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-025-10573-
                                        2.

Wang:2025:OSB

[2191] Yanzhao Wang, Yaohua Zhang, Jian Zou, and Nalini Ravishanker. Online structural break detection in financial durations. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10577-y.

Zhang:2025:RBF

[2192] Jiarui Zhang, Jiguo Cao, and Liangliang Wang. Robust Bayesian functional principal component analysis. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10580-3.

Rodriguez-Alvarez:2025:DRD

[2193] María Xosé Rodríguez-Álvarez, Vanda Inácio, and Nadja Klein. Density regression via Dirichlet process mixtures of normal structured additive regression models. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10567-0.

DUrso:2025:FEC

[2194] Pierpaolo D’Urso, Livia De Giovanni, Lorenzo Federico, and Vincenzina Vitale. Fuzzy K-expectiles clustering. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10566-1.

Modok:2025:CCM

[2195] Bidhan Modok, Shovan Chowdhury, and Amarjit Kundu. Control charts for monitoring Weibull quantile under generalized hybrid and progressive censoring schemes. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10581-2.

Columbu:2025:MLC

[2196] S. Columbu, N. Piras, and J. K. Vermunt. Multilevel latent class models for cross-classified categorical data: model definition and estimation through stochastic EM. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10579-w.

Tsagris:2025:DDA

[2197] Michail Tsagris, Panagiotis Papastamoulis, and Shogo Kato. Directional data analysis: spherical Cauchy or Poisson kernel-based distribution? Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10583-0.

Albano:2025:IDP

[2198] Giuseppina Albano, Antonio Barrera, Virginia Giorno, and Francisco Torres-Ruiz. Inference on diffusion processes related to a general growth model. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10562-5.

Biziaev:2025:UPD

[2199] Timofei Biziaev, Karen Kopciuk, and Thierry Chekouo. Using prior-data conflict to tune Bayesian regularized regression models. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10582-1.

Anonymous:2025:JNb

[2200] Anonymous. Journal navigation. Statistics and Computing, 35(2):??, April 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Grabchak:2025:ASS

[2201] Michael Grabchak and Sina Saba. On approximations of subordinators in L^{p} and the simulation of tempered stable distributions. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10586-x.

Farcomeni:2025:ESI

[2202] Alessio Farcomeni. Exact score and information matrix for panel hidden semi-Markov models. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-025-10585-
y.

Dunbar:2025:HOR

[2203] Oliver R. A. Dunbar, Nicholas H.

Nelsen, and Maya Mutic. Hyperparameter optimization for randomized algorithms: a case study on random features. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10587-w.

Lv:2025:RPR

[2204] Jing Lv and Chaohui Guo. Robust ℓ_{2,0}-penalized rank regression for high-dimensional group selection. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10588-9.

Rimella:2025:SBC

[2205] Lorenzo Rimella, Chris Jewell, and Paul Fearnhead. Simulation based composite likelihood. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11222-025-10584-
                                        z.

Li:2025:FMA

[2206] Xinmin Li, Hua Liang, Huihang Liu, Tingting Tong, and Tian Xie. Frequentist model averaging under a linear exponential loss. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10589-8.

Yu:2025:GIH

[2207] Ke Yu, Xu Guo, and Shan Luo. Group inference for high-dimensional mediation models. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10591-0.

Liang:2025:MHS

[2208] Qiao Liang and Xinwei Deng. A multifacet hierarchical sentimenttopic model with application to multi-brand online review analysis. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10593-y.

Joudah:2025:AHA

[2209] Ibrahim Joudah, Samuel Muller, and Houying Zhu. Air-HOLP: adaptive regularized feature screening for high dimensional correlated data. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10599-6.

Sprungk:2025:MAI

[2210] Björn Sprungk, Simon Weissmann, and Jakob Zech. Metropolisadjusted interacting particle sampling. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10595-w.

Barnett:2025:MWA

[2211] Steven D. Barnett, Lauren J. Beesley, Annie S. Booth, Robert B. Gramacy, and Dave Osthus. Monotonic warpings for additive and deep Gaussian processes. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10598-7.

Reiser:2025:UQP

[2212] Philipp Reiser, Javier Enrique Aguilar, Anneli Guthke, and Paul-Christian Bürkner. Uncertainty quantification and propagation in surrogate-based Bayesian inference. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10597-8.

Zhan:2025:EMS

[2213] Tingting Zhan, Misung Yi, Amy R. Peck, Hallgeir Rui, and Inna Chervoneva. Estimation and model selection for finite mixtures of Tukey’s g- & h-distributions. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10596-9.

Wijayawardhana:2025:VBI

[2214] Anjana Wijayawardhana, David Gunawan, and Thomas Suesse. Variational Bayes inference for simultaneous autoregressive models with missing data. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10590-1.

Rilling:2025:NVB

[2215] Joseph Rilling and Cheng Yong Tang. A new p-value based multiple testing procedure for generalized linear models. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11222-025-10600-2.

Buchhorn:2025:BDS

[2216] Katie Buchhorn, Kerrie Mengersen, Edgar Santos-Fernandez, and James McGree. Bayesian design for sampling anomalous spatio-temporal data. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10594-x.

Liu:2025:BAD

[2217] Xiaoyu Liu and Liming Xiang.

Bayesian analysis of doubly semiparametric mixture cure models with interval-censored data. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-025-10601-
1.

Liu:2025:SDL

[2218] Jin Liu, Wei Ma, Lei Wang, and Heng Lian. Sparse and debiased lasso estimation and statistical inference for long time series via divideand-conquer. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10602-0.

Yang:2025:DTL

[2219] Jichen Yang, Lei Wang, and Heng Lian. Debiased transfer learning estimation and inference for multinomial regression. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10607-9.

Paroli:2025:FBV

[2220] Roberta Paroli, Dimitris Fouskakis, and Ioannis Ntzoufras. Fast Bayesian variable screening using correlation thresholds. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10608-8.

Corneck:2025:OBC

[2221] Joshua Corneck, Edward A. K. Cohen, James S. Martin, and Francesco Sanna Passino. Online Bayesian changepoint detection for network Poisson processes with community structure. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10606-w.

Guo:2025:SEE

[2222] Xu Guo, Hongwei Shi, Weichao Yang, Yiyuan Qian, and Niwen Zhou. Semiparametric efficient estimation of genetic relatedness with machine learning methods. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10605-x.

Long:2025:FGS

[2223] Xueying Long, Daniel F. Schmidt, Christoph Bergmeir, and Slawek Smyl. Fast Gibbs sampling for the localseasonal-global trend Bayesian exponential smoothing model. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL https://link.springer.com/
article/10.1007/s11222-025-10603-
z.

Qin:2025:CPU

[2224] Shanshan Qin, Zhenni Tan, Dongwei Wei, and Yuehua Wu. Correction: PCA-uCPD: an ensemble method for multiple changepoint detection in moderately highdimensional data. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10615-9.  See

[2177].

Cheng:2025:EEV

[2225] Ruonan Cheng and Xiuqing Zhou. Efficient estimation for varying coefficient modal regression. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10611-z.

Hurwitz:2025:PPC

[2226] Rebecca Hurwitz and Georg Hahn. Penalized principal component analysis using smoothing. Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL https:

//link.springer.com/article/10.
1007/s11222-025-10610-0.

Anonymous:2025:JNc

[2227] Anonymous. Journal navigation.

Statistics and Computing, 35(3):??, June 2025. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).